A Complete Bibliography of Publications in Mathematics of Computation, 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 October 2024 Version 2.44

Title word cross-reference

#2945 [Tho94]. #3506 [Köl97e]. #8373 [BW96].

(0, ∞) [MM95]. (2, 1) [BPv94]. (2, 3) [Sel99]. (4, 2) [Sel99]. (6, 1) [Sel99]. (div) [AFW97]. (i, 1) [Cos91]. (Mod2) [BBG91]. (t = 3, 5) [KM91]. −2 [BN92]. 0 [Zag93]. 1 [Cop90, CS99, Kvæ90, Lan96, YR92, Zag93]. 1.4778 ⋯ [WW97]. 10 [Sou94]. 10000 [Sho99]. 10^{10} [MM91]. 10^{10,000,000} [SW94]. 10^{12} [Jae90]. 10^{14} [Bre97]. 10^{15} [Pin93a]. 10^{20} [Sao98]. 10^{6} [HC98]. 125 000 [EM91]. 150 000 [EM91]. 16 [San92]. 19 [CMS92a, DD93]. 1^{k} + ⋯ + (x − 1)^{k} = x^{k} [MTU94]. 2 [BH94b, BC98, Che94b, Göt94, HS92, KS93a, Lab98, LS94, Lou95, Pea96, Sch93c, SW90, vW99]. 22 [KS93a]. 25 [Dui94]. 26 [Dui96b]. 27 [CMM92b]. 2^{β} [Fis90]. 2 · 3 · 5 ⋯ p ± 1 [Cal95]. 2 × 1 [Dui94, Dui96b]. 3 [CHW99, Coh98, Dui96a, Dur99, His94, vv93b].

3x + 1 [AL95a, AL95b, Oli99]. 4 [RI96]. 43 [ESA94]. 471, 600, 000 [KW90]. 49 [CD98]. 4 · 10^{11} [Sin93]. 5 [Gol91b, VZ95, Wag96a]. 6 [Boy96b, Wag96a]. 7 [Wag96a]. 8 [LJ96a, Sou94]. 84 [GGM98]. 9 [Fuj93, Tak94]. [−1, 1] [Pin99]. _{3}F_{2} [LGRA94]. a^{p−1} ≡ 1 (mod p^{2}) [Mon93]. abc [BB94c]. B [DMS92, Li94b]. B_{2} [Lin98, Zha93b, Zha94]. β ≃ 32 [Fis90]. β ≃ 48 [Fis90]. mod2 [BBG90]. modp [Rob93]. BV (Ω) [Coc91]. C^{1} [CH96, CH90b]. C^{2} [FT91]. C^{n} [LW96a]. C^{r} [DM98]. d [EL96]. D > −10^{6} [FW90b, FW94]. d^{(1)} [Sid95]. d^{3}/dx^{3} [GASS91]. d_{2} [GL98a]. E_{6}^{∗} [Gec93]. ϵ [FMOS98]. F_{2^{m}} [MVZ93]. F_{6} [Wil93]. G(2) [Phi95b]. g(k) [DD90, SW92]. G_{2}(q) [HS92]. Γ [BF92]. Γ(2) [Phi95a]. Γ_{2}(2, 4, 8) [vv93b]. GF (2) [KM91, Cop94]. GL(3, Z) [AM92]. h [HST98]. h^{+} [Jak98]. H_{0}^{1} [HJ95]. ˆ3M_{22} [LS94]. h · 2^{k} ± 1 [Bos93a, Bos93b]. hp [SS96a, Sur97]. K [Tak94, CLT93a, CLT93b, Tez94, Tra95]. k(ln k + ln ln k − 1) [Dus99]. K3 [CS96a]. k^{th} [Dus99]. k ≥ 2 [Dus99]. L [AT99, EM92c, SSW99, Sil99a, FS95, Jak98, Sch98b, EM92a, EM94, Lou92]. L^{1} [GW96, LX93]. L^{2} [BS90b, HJ95, BX91, Coc91, Xu91]. L^{r}(0, T; L^{ρ}(Ω)) [Ver98]. L^{ρ} [WL99]. L_{2} [TB97, Bai94c, BL95, Hei96]. L_{2}(ℝ^{d}) [RS98]. λ [FS95]. lcm [Mor89, Mor90]. ≤ 2 [Lou95]. ≤ 44497 [KM91]. LU [BTW90]. m [BH94a, CM94a, FR94, Kui96, Rob93]. m − a_{m}^{1/2} [Zha94]. ℚ(\sqrt{3}) [Hun91]. F_{2} [Roe99]. n! ± 1 [Cal95]. n [Hor97, MS95]. n = fu^{2} + gv^{2} [HMW90]. nk = 2ab [Goe98]. O(Δx)L^{1} [TT95]. p [BS92a, BL93, BSW90, EM92a, EM92c, EM94, EM97b, FS95, HST98, HGK98, Jep96, SS96a, Sem98a, Ste94b, Xu94a]. P − 1 [MS90]. ϕ [Ct98]. π [AW97]. π(x) [DR96]. ψ(x) [DR98a]. P SL(2, Z) [HA93b, HA93a]. q [GS90, Jep96, WH93]. Q(\sqrt{3}) [CD92]. Q(ζ_{p}) [Jha95]. Q_{1} [CO98]. QR [Ste92, BD92a]. qx + 1 [FP95]. r [Pap97, JS92, Joe93]. R(1, 6) [Mai91]. R^{3} [CPS92]. S [Sma99]. S(5, 6, v) [GGM98]. S_{n} [Gra95]. ∑_{i=1}^{n}(−1)^{n−i}i! [Živ99b]. ∑ g_{n}/2^{g_{n}} [BL90]. Sz(8) [Sch93c]. t [Bai94b, KM91]. τ [Kei96]. u [HMW90]. v [HMW90, BP93, Vas92]. v = 72 [GGM98]. ε [NTS98]. w [Pea96]. x, k [MTU94]. x^{3} + y^{3} + z^{3} = k [HBLT93]. x^{3} + y^{3} + z^{3} = n [KTS97]. x^{4} + 1 = Dy^{2} [Coh97c]. x_{1}^{6} + x_{2}^{6} + x_{3}^{6} = y_{1}^{6} + y_{2}^{6} + y_{3}^{6} [Del92b]. x^{p} − y^{q} = 1 [AI91]. x_{1}^{k} + ⋯ + x_{s}^{k} ≡ N( mod p^{n}) [SK99]. ξ [CSV94, Kei92]. Z_{3} [Tay96]. ζ_{K}(2) [Tak94]. |ax^{n} − by^{n}| = 1 [Bd98]. |ζ(12 + it)| [It91].

-Adic [BSW90, EM92c, EM94, FS95, Xu94a, EM92a]. -algorithm [NTS98]. -amicable [Ct98]. -approximation [GW96, BS90b]. -basic [Pea96]. -binomial [GS90]. -blocks [HS92]. -bridge [KS93a]. -Complete [EL96]. -conjecture [BB94c]. -connected [Dui96a]. -convergence [Li94b]. -cycle [BP93, Vas92]. -d [Dur99, BH94b]. -dimensional [CLT93a, CLT93b, Lab98, Tez94]. -Dirichlet [Kei96]. -distribution [Bai94b]. -divisibility [EM97b, FR94]. -error [TT95, HJ95, Ver98]. -extension [FS95]. -function [BF92, CSV94]. -Functions [EM92a, EM92c, EM94, Lou92, SSW99, AT99]. -groups [Pea96]. -invariant [FS95]. -invariants [Tay96]. -Laplacian [BL93]. -manifolds [CHW99]. -modular [Gol91b, HS92, His94]. -Nomials [KM91]. -operator [Tra95]. -patches [DMS92]. -power [Lou95]. -powerful [Coh98]. -Projection [Coc91]. -projections [HJ95]. -rank [Pap97]. -Rotational [YR92]. -sequence [Zha93b]. -sequences [Zha94, Lin98]. -series [Sil99a]. -sets [CM94a]. -simplex [MS95]. -simplices [Hor97]. -spline [DMS92, DM98]. -stability [LX93]. -star [BH94a, Kui96]. -subtetrahedron [LJ96a]. -th [Sch98b]. -torsion [Sem98a]. -transformation [GL98a, Sid95, GL98a]. -triangles [RI96]. -uniform [FMOS98]. -unit [Sma99]. -version [BS92a, HGK98].

1 [Köl97g, Hig96]. 15 [Köl97h]. 1943-1993 [Len97]. 1995 [Tho97].

2 [Krö91, Mon94a, Fro97]. 2-adic [SSW99]. 2-bridge [KS93b]. 2-D [Krö91, Mon94a]. 21 [Tho94]. 25 [BW96].

37 [MPR95]. 38 [Tho94].

43 [Köl97e]. 46 [BW96].

50th [Gau94]. 52 [BO94, Mor90]. 53 [CH90a]. 55 [FW94]. 57 [Ano92, GC93]. 599pp [Köl90].

61 [LLMP95]. 62 [Woz97b]. 64 [JMS97].

82k [MPR95]. 868a [Köl97h]. 88f [Köl97g]. 89k [BO94].

90c [CH90a]. 90m [FW94]. 91j [Ano92]. 92h [GC93]. 93c [Köl97f]. 93k [LLMP95]. 94g [Fik98, Köl95, Lam96, Lam97, Woz97b]. 95m [JMS97]. 9th [Smy91]. 9th-order [Smy91].

A. [Cai96, Fal98, Köl97c, Köl97a, Köl97b, Tho97]. Åke [Col97]. abelian [AK99, BJT97, Far94, Pil90, Sho99, Yam94]. Abell [Mea96, Wri96]. Abramowitz [RW92]. absolute [Fla96, Odl93, RS95, Str92b]. Absorbing [Fen99, Krö91, She93]. Abstract [LT91b, Loz97, GP99, LT91a]. Academic [Fik98, Köl95, Lam96, Lam97]. Accelerated [AKL98, Mül97]. Acceleration [Van94, Göt94, LP95]. Accuracy [Van97, GS95, MS97b, Sül92]. Accurate [Ber90, Eck93, Eck95, HHM94a, HHM94b, Mac97, Noe96, Sun95, Zho97]. acoustic [BS94b]. acoustics [Gil98]. Acta [Fal98, Tho97]. Adams [Wri97]. Adapted [Chu96, Sch91]. Adaptive [BGLT95, DY90, EJ93, NPV91b, DR98b, NPV91a, WG92]. add [LF93]. addenda [Tem98]. Addendum [LLMP95, MPR95]. addition [AS93a, AS93b]. additional [CMV92]. additive [Bre96b, HST98, Ste96b]. Adelaida [Shi96]. ADI [LM94b]. Adic [BSW90, EM92c, EM94, FS95, SSW99, Xu94a, EM92a]. adjoint [EG90]. adjustable [Bai94c]. admit [GM96]. advanced [Kin99a, Loz98a]. Advances [BV99, War96]. advection [LT99]. advection-diffusion [LT99]. affine [RS98]. Afternotes [Kin99a, Ser97a]. Ainsworth [Kin97]. Airy [Tem94b]. Airy-type [Tem94b]. Alain [Bar96]. Alan [War98]. alcoves [Sut98]. Alexander [Qua96a]. algebra [Gec93, Hig94, Rud98, Bos98, Ste99c, Wat98]. Algebraic [Gal92, Jus90, LT91b, Zag93, Beu96, BC98, Dre98, FJP97, Fla96, Fuj93, Gra98, GL98b, KSS98, Kvæ90, LT91a, Mat97, Mon94b, RR94, Sch93a, Sch91, Sch96a, Tak94, Vas92]. Algebras [Den90, HW90, San92]. Algorithm [HMW90, HTS90, LZ92, Ada95, AD93, Bai94a, BD92a, Bel97, CFS95, CGT97a, CGT97b, Cop94, DDD91, DM98, Dew98, EG99, FBA99, GW96, Göt94, HT95, Huc95, Kal95, LLS96, LH98, LN96, MS90, NTS98, NY97, NG95, Oli98, Pih93, Roe99, Sem98b, SW93, Smi96, Swa93b, Tei98, Tes98, TB97, TW92, WH93, Woź94, Woz97b, DH99a]. Algorithmic [Wag98]. Algorithms [Bai94c, Ben98b, BB94a, BP92, Gol94b, Grö92, Hol97, Sho90, Tod96, Bos98, BPX91b, BPWX91a, BP93, BLP94, BPV98, Che94a, CEL96, Che96d, Coh96, EB99, EST90, Hei96, Hig96, LT99, LG97, MES98, Nie96b, PT92, PJV94, PST98, RI96, ST91, Val91, vSC92, Ye91, Ye94, Van97, Wag98]. Aliquot [MM91, BV99, GG94, MM93b]. allowing [CP97]. Almost [CHL99]. along [LM94b]. alternating [BF93b]. alternating-direction [BF93b]. Ambiguous [Mol93]. Amicable [Cos91, Ct98, MM93b, tR94]. Among [Jus90]. Analysis [BF93a, BWW97, BG94, BP92, BPV98, Cal92, Cho97, CPS92, FBA99, Fis90, FK93, GHS91, GL98b, Kal95, KS92, KLL96, LM90, Mad90, Mel98, MSS97, NW96, Ste92, TB97, WT90, AM94, AN92, Bai94a, BS94b, BDR98, BPX91b, BLP94, Bra94, Cal91, Chu96, DVZ94, EHE96, ELV91, Gol91a, GS97, HDJ94, HL94, Isa96b, Kin99b, LT99, LS91b, LG93a, LG93b, Mel97, Néd94, NW98, Rum90, Sid95, SW93, Ste90, Ste99b, SO91, Tem91, Ye91, Zhe94, dB98, Bra96a, Dav96, Kin99a, Lar97, Ser97a]. Analytic [LT91b, Odl94, BR94, GS95, LT91a, LP95, OW97, Sch97]. and/or [DH90]. Anderson [Cai96]. Andreas [Isa96b, Ste98b]. Andrew [Wah98a]. angle [Žen95]. Anne [Sil99b]. Anniversary [Gau94]. annular [CM95]. Anti [Lau96]. Anti-Gaussian [Lau96]. antidiffusive [CL91c]. appendix [FS95]. Application [HW92, BS94b, Bra95, CQS98, KX94, Lan96, LSW94, Sid95, Hal96]. Applications [CD96, LT91b, Loz98a, Sch94, Abe94, AFW97, ABZ95, BBG93, BPWX91b, GvP98, GH97, Hob98, LSS99, LT91a, Ste98b, XZ94, Bra96b, Che96d, Sch96b, Ste99a, Tem97]. Applied [Wat98, AN92, Kvæ90, Zha98a]. Applying [Jac99]. Approach [CSW92, BPX91a, BLP97, CG96b, CL91c, GANT99, Gol94a, LS95, Lew94, LG97, NRS95, ST91, Auz98]. approximants [CD96]. Approximate [Che94a, Abr95, Cha94a, GANT99, Nes96, Nii92, Tho94, Ziv95]. Approximated [Dub92]. Approximating [HS97b, HS95a, KP92, OW97]. Approximation [Bos90, BR94, DY90, DHNL^{+}99, Ell96, FK91, HW92, Jia98, Sho98, AKL98, BO94, BL93, BB99a, BS90b, BS94b, BF93a, CTW92, CLM94, CH96, CL91b, Cop99, Dem90, Dem95, DR98b, Fas97, FK93, GT94, Grö99, GM92b, GHS91, GMP91, HB92, HR94, Hua91, IK98, Kha99, Lan96, LSS99, Leh90, Lev92, Lev98, LL98, Lin94, LO95, Mül97, Pao97, Pes96, Qua97, RAG91, SS94, SBS95, TB97, Vv96, VC94, Zhu92, GW96]. Approximations [LT91b, Sch92, BL95, BCM92, CDT93, Cot91, DDP95, Du98, EG99, FG91c, GT91, KM90, KB97, LT91a, LNS93, LST96, Mak93, MM94a, Pal94, Pal95, RS96a, RVW96, Tad93, Tho94, Wal92, WL99, Wu95, Zha93a]. arbitrary [LM94b]. architecture [Loz98a, Roj94]. architectures [LF93, Hou96b]. arcs [AS91b, He94]. areas [Bea97]. arguments [Bru94, CK99]. Arieh [Kin99b]. arising [AY94, CG90b, Mit90, Mon94a, Was91]. Arithmetic [Beu96, Abe94, CF94, DN97, Kno93, PMT95, RR96]. Arithmetical [MT93, Dil92a, Dil92b]. Arjeh [Rud98]. Arkadii [Tod96]. Arnoldi [MS97a, Mor96]. arrowhead [Oli98]. art [Loz98c]. Artificial [HW92]. Artin [Pap97]. Aspects [dR92, Che96a, Shu96, Tem94b]. Assembling [Ber90]. assisted [MM98]. associated [AT99, Bea97, Mas94, MPR95]. assumptions [BPWX91a, SW96]. Astfalk [Loz98a]. Asymptotic [BP96, Fal96, Fly98, Hag91, HHM94b, Tem92, CJT97, Dom96, HH99a, Sch98a, Tem94b, Til97, VC94, HHM94a]. Asymptotically [Bra93, Ser97b]. asynchronous [MES98]. Atkin [Dew98]. atlas [Lee99]. attractors [HS95a]. Augmented [LS95]. August [Gau94]. Aurifeuillian [Wag96c]. Automatic [EB99, HT91]. automorphism [CD98, CMS92b, CMS92a]. automorphisms [AK99]. Average [DLP93, EHL97, NRW95, EHE96]. Average-case [NRW95, EHE96]. Avner [Isa96a]. avoid [EH91]. AWC [CL94]. AWC/SWB [CL94]. Axelsson [Elm96].

B [Arn97b, BW96, Gil98, Köl90, Shi96, Peñ97]. B-splines [Peñ97]. Bach [Wag98]. Backward [AM94, AL91]. Bad [EdT90]. Bai [Cai96]. Bailey [Gau96]. Baiocchi [Bra96b]. Bal [Nie96b]. Balancing [CMW95, MB96]. Banach [Wan99]. band [Ser97b]. band-Toeplitz [Ser97b]. Bank [Ano95m]. Banks [Auc96]. Barrett [Ano95q]. barrier [CGT97a, CGT97b]. barriers [MO93, OZ91]. Base [KKS90]. Based [HTS90, Ad97, BLP97, Dem96, FP97, LS98, LJ96a, Mar90, Shy94, Str92a, VV93a]. Bases [Dör92a, DP96, DMS92, DS99, Dör92b, FvS99, Gaá96, GP97, Jae93, LSW94, Nie94, Zhe94]. basic [Pea96]. basis [DM98, FG96, GV95, Hob98, JKN91, Sch99, Ste96b, Wen99]. Battle [Lai94]. Bau [Ste99c]. be [AH99]. before [WS94]. begets [DMS92]. behave [CLR97]. Behavior [Ber90, Che91, Cot91]. behaviour [Mor99]. Bell [Wag94, Wag96c]. Below [MM91, SW94, Sho99]. BEM [BWW97, Car96, Li99]. Bengt [Shu97]. Berg [Ste98b]. Bernard [Hou96a]. Bernstein [FN90, FG96, Peh93]. Berry [Ano95q]. Bessel [Ben98a, FS92, FO93, GF97, GR95]. Best [BS90b, BS99, Dem90]. beta [Boy96a, Boy96b, Boy96c, Dom96]. better [DMS92, ST96]. Between [Len91b, BBN99, Bra95, CD99, EM91, Mar92, NHA97, Sil90]. Beyerm [BW96]. Bézier [FN90]. bifurcation [ABZ95, BGJ99, WBL94]. billion [CD97]. Binary [DJ91, Bea97, EK97, Živ94a, Živ94b]. Bini [Hig96]. binomial [ELS93, FG91a, GS90, SdW99]. biographical [Kel95]. biquadrates [DD93]. Birkhoff [BN90]. Bischof [Cai96, Ste98b]. bisection [Hor97, LJ94, NRW95]. Bivariate [CH90b, CH90a, Shp93, Xu92]. Björck [Col97]. Bjørstad [Gau96]. BKK [LW96a]. blessing [AH99]. block [CV95a, Cop94, Gol91b, Kal95, PJV94, Til97, vS96]. block-size [CV95a]. blocks [HS92, Til97]. Blossoming [DMS92]. bodies [NRS95]. body [MP96]. Boltzmann [Léc91]. Book [Ach99, Ano95g, Ano95h, Ano95m, Ano95q, Ano95u, Ano95r, Ano95j, Ano95t, Ano95i, Ano95b, Ano95n, Ano95v, Ano95y, Ano95c, Ano95a, Ano95l, Ano95x, Ano95k, Ano95z, Ano95d, Ano95f, Ano95o, Ano95p, Ano95w, Ano95e, Ano95s, Tod96, Arn97b, Auc96, Auz98, Bai96, Bar96, Ben98b, Beu96, Bos98, Bra96a, Bra96b, Bre96a, Bre97, Bre96c, Cai96, Che96a, Che96b, Che96c, Che96d, Che97, Chu96, Coh97b, Coo98, Dav96, Ell96, Elm96, Elm97, Fal96, Fal98, Fro97, Gau96, Gil98, Hig96, Hol96, Hou96a, Hou96b, Isa96a, Isa96b, Isa96c, Kar97, Kea98, Kea97, Kin97, Kin99a, Kin99b, Kli96, Kok96, Köl90, Lar97, Len97, Loz96, Loz97, Loz98a, Loz98b, Loz98c, Mea96, Mey96, N.93, Nie96b, Nie96c, Nie96a, Noc96, Qua96a, Qua97, Res97, Rud98, Sch96a, Sch96b]. Book [SS96b, Ser97a, Sha96, Shi96, Shu96, Sil99b, Ste96a, Ste99a, Ste98b, Ste99b, Ste99c, Tem97, Tem98, Tho97, Tod98, Wag96b, Wag98, Wah97a, Wah98b, Wah98a, War96, War98, Wat98, Wim98, Wri96, Zha98a, dB98]. Books [Ano95-27, Ano95-28, Ano95-29, Ano97, Col97, Hol97, Mat97, Shu97, Ste97a, Ste97b, Van97, Wah97b, Wri97]. Boolean [Del90]. Boston [Fik98, Köl95, Lam96, Lam97]. both [Zag93]. Bouleau [Ano95h]. bound [DGNP93, Hic98, NY97, PW98, Str92b, Ziv99a]. boundaries [BK94, Lev92]. Boundary [AC90, Bra96b, CPS92, DCNZ96, HW92, LM90, Sch92, Ste99a, Yan90, AY94, AD91, BL95, BS94b, BG99a, BS99, BLP93, BV92, BCMR95, Car94, CS95, CH98, DR98b, EG99, FMOS98, Fen99, FK93, FG91c, GT91, HDJ94, Hag91, HHM94a, HHM94b, Han90, HST98, Hon99, Hun92, Hun98b, Krö91, Mon94a, SS91, SS94, SS96a, SW99, SZ90, She93, Shi96, SO91, vBv92, Wu95, XZ96, Yin98, Zhu92]. boundary-element [BS94b]. bounded [Yin98]. Boundedness [ELR90]. Bounds [DS91, HGK98, Joe93, Pow97, AL95a, AL95b, Bac90, BS94a, BS96a, Bra93, BCT91, CGT97a, CGT97b, DM99, Ehr94, EHN92, EH94a, Fly98, Léc96, Nie90, Nii92, Sch99, Ste98a, Ziv95]. Boussinesq [DOSS91]. box [CQS98, Neu94]. BPX [Osw99]. BPX-preconditioner [Osw99]. brackets [Sei99]. branching [Coh93]. Braselton [Mea96, Wri96]. Brass [Ano95g]. Brauer [HW90]. Breach [Köl97g, Köl97f]. breakdown [Ye94]. breakdown-free [Ye94]. breaking [BGJ99]. Brenan [Mat97]. Brenner [Bre96c]. Brian [Loz98c, Sha96]. bridge [KS93a, KS93b]. brief [Isa94]. Briggs [Nie96c]. British [Gau94]. Brown [Isa96c]. Brun [Bre97]. Brychkov [Köl97g, Köl97f, Köl97a, Köl97b]. bubble [ZZ97]. building [Kli96]. built [DMS92]. Bulirsch [Ano95u]. Burgers [BT97, BH94b]. Buschman [Tem98, Wah97b]. Butuzov [Shi96]. Byrd [Köl97e, Köl97d]. Byrnes [Boy97].

C [Bre96c, FS95, Isa96c, Kok96, WH93]. C. [Bra96b, Cai96, Coo98, Elm97, För94, War96]. Cahn [EL92b, BB99a, EL92a, Sun95]. calculate [Iya95]. Calculation [AT99, Ber90, Bri91, Lau97, TF93, Ura96a, Ura96b, KSS98]. calculations [Mas94]. calculus [MW94a]. Calvo [Ano95p]. Campanato [Dol99]. Campbell [Mat97]. can [AH99]. Canas [Ste99a]. Canonical [Hip99, OS92, Bun97, CS96a, LJ96b, Sil90, Sil97, Sil99a]. capitulation [FK96]. Capturing [JSH90]. Cardinal [CSW92, PT94]. Carlo [Nie96a, BGLT95, HMN93, HM93, HM97, Léc91, Léc96, TW92]. Carmichael [BN97, Jae90, LN96, Pin93a, SW94]. carry [CL97b]. Cartesian [CG97]. case [AC90, CHS90, CT92a, CT92b, CT94, Cop90, DLP93, DS95, EHE96, FP90, GP99, HDJ94, NRW95, Yan96b, Yan96a]. Catalan [AI91, Ste98a]. Cauchy [Bia90, Die94, HT91, Mon94a, Rab90a, Rab90b, Rab91]. cavities [Lev92]. cell [CFS95, ELV91, GS97, JS94, MSS97, Sül92]. cell-centered [ELV91]. cell-vertex [GS97, MSS97]. census [CHW99]. centenary [Isa96c]. centered [ELV91]. Central [Den90]. centralizer [HW90]. centrally [Cha92]. century [Gau94, Len97]. Certain [BBG90, BBG91, WT90, Al 94, CL94, Gau91b, Kid99, Lew94, LP95, Mac97, MT93, Mit90, MW94b]. CFL [GT91, HR93]. chains [AS93a, AS93b, For99]. Chaitin [Ste97b]. Chaitin-Chatelin [Ste97b]. Chan [Ano95q]. Change [Mar92, Mar90]. Chaos [MM98]. characteristic [GLW99, Göt94, LS94, Sch93c, Sch93b, Sem98a]. characterization [Dem90, Xu94b]. characterizations [MO93]. characters [HS92, His94]. Charles [Che96a, Che96d]. chasing [Oli98]. Chatelin [Ste97b]. Chebyshev [BBN99, BE96, Fly98, För94, HS97a, HTS90, Kui95, Mac93, Mac94b, Not92, RO91]. Chebyshev-type [Kui95]. Checking [Sao98, Sin93]. Chevalley [HS92]. chi [BD92b, CK99]. chi-function [BD92b]. Choice [DGPR92]. Christian [Ste98b]. Christoph [Fro97]. Chu [Isa96c]. Chui [Che96d]. Ciarlet [Bra96a, Fal96]. circle [MPV98]. Circulant [CY92, YC93, FG91a, ST96]. Circular [GHO92a, GHO92b, He94]. Class [BSW90, MW92b, Sol92, Ste98a, Ste94b, Wag96a, BPv94, CL90, CM94b, CDO98, CS99, CMM90, EH93, FR94, Gau91a, GN96, GLW99, Gir93a, Gir93b, HMN93, Jac99, JMS95, JMS97, Jak98, Jha95, KLL96, Lou92, Lou95, Lou97, MPR95, MES98, Pv93, PW98, Ros95a, Ros95b, Rüc99, SH98, Sch98b, SSW99, Sho99, Sri98, Vv96, XZ96, Yam94]. classes [BS94a, BS96a, But94, JM92, LT91c, Mol93, Phi95a, Phi95b, Wal99, Wal21]. classical [RS92, Wim98]. Classification [Mai91, SH98].

Clenshaw [HST93]. climate [Res97]. close [Zag93]. Closed [CK95, SV98]. Closed-form [CK95]. Clothoid [MW92a]. clusters [For99]. CM [Lou97, van99]. CM-fields [Lou97]. Code [Mai91]. Codes [DJ91, vSv92]. coefficient [ELS93]. coefficients [BBBP98, BTT94a, BTT94b, Boy97, CHT96, FO94, Fly98, GS90, GH94, GZ93, Han96, HA93a, HA93b, HWC99, KT90, LSEG91, MB96, Osw99, lX96]. Cohen [Rud98, Sch96a, Shu96]. Cohomology [AM92, Pea96]. Collocating [Ste95]. Collocation [CPS92, Yan90, Arn94, BF93b, BS99, Bru94, BPV99, CDL97, CRV92, HS94, Hu98, KB97, Mak94, Rob94, Yin94]. Columbia [Gau94]. combinations [Peh90a]. combined [CG93, CLT93a, CLT93b, LT91c, L’E96, L’E99b, Tez94]. Comments [Bac97]. Comp [Ano92, BO94, BW96, CH90a, FW94, GC93, JMS97, LLMP95, MPR95, Tho94, Woz97b]. compact [Far94]. Compactly [RS98, CD96]. companion [EM95]. Comparison [BN90, AY94, ACEO98, PJV94, WH99]. Complete [EL96, MPR96, Pet91, Gra98]. complex [FW90b, FW94, Hig98, Pv93, Wei94, vW99]. Complexity [Gol92, HM93, Tei90]. component [LS93b]. componentwise [Hig94]. Composite [DS99, ADS98, Arn95, CL96, CDNY95, Vv96]. Composition [KL97]. Compound [EHE96, EHL97, Emm96]. compressible [BDR98, CL91a]. Comput [Mor90]. Computation [Abe94, BSW90, DD90, Dre94, Ell96, EM92a, EM92c, EM94, Law91, Len97, Lou97, Loz98b, MM97, Shi95, Tay96, BBP97, Bea97, BMO90, Bre99b, FW90b, FW94, Gau94, Hal96, Isa94, Jac99, Jha95, Lai94, Mas98, Oli92, Pv93, RR94, Tes98, Isa96b, Loz96, Qua97, Köl90]. Computational [BH97, Coo98, Gil98, Gol92, Nie96b, Ste98b, Zha98a, dR92, Auz98, BJT97, Gau94, Laz94, Len97, Oli99, Sch96a, Tei90, War96, War98, dB98]. Computations [DJ91, Pan90, Sko92, Sri98, BWW97, BS96b, Hig96, Odl94, Rum93a, Rum93b, Ste97b]. compute [Bel97, MP96]. Computer [Fro97, Rud98, Bos98, HBLT93, MM98, Roj94, SW90, Wag96b]. Computers [Köl90, BB94b, vSC92, WS94, Loz97, Loz98a]. Computing [AK99, BR90, BC94, CS96a, CDO98, Cre95, CS99, DR96, DR98a, DST97, Gaá96, HW90, JS92, McK94, MST99, OW94, OS97, Sil97, Sil99a, Ste96a, Wei94, Wei95, Xu94a, Bre93, But94, Hei96, Hou96b, Loz98c, MS93, Swa93b, Gau96, Hou96a]. concentration [BB99a]. concerning [Boy97, För94, FP95, FR94, Rum93a, Rum93b, Xu91, Zha93c]. Condition [Ede92, DN90, FN90, GT91, HR93, HGK98, LS91a, Žen95]. Conditionally [MN90, NW94]. conditioned [AY94, Ser97b]. Conditions [HW92, AC90, BG99a, BV92, EG99, Fen99, HDJ94, Hag91, HHM94a, HHM94b, KP92, Krö91, LL99, SZ90, SK99, She93, Sul93, vBv92]. conducting [HR94]. conductor [Sho99]. conductors [CDO98]. Conference [Wah98a, Isa96c]. configuration [Mas94]. Configurations [BG92b, BG92c]. conformal [LS98]. conforming [DDP95, DM99]. Congruence [AM92, MST99, SK99]. congruences [AI91, SSW99]. Congruential [BB94b, Fis90, BW96, CL94, EH91, EHN92, EH93, EHI94, EH94b, EH94a, EHE96, EHL97, EHN97, KWY96a, L’E99a, Nie90]. conjecture [AS93a, AS93b, BB94c, Coh98, Cus93, FP95, FK96, Fuk96, iHT99, KW90, Sao98, SW94, Sin93, Zha95a, Pap97]. conjectured [VC94]. Conjugacy [Phi95a, Phi95b, But94]. conjugate [Gra98, Wri97]. conjugates [Str92b]. conjugation [LS98]. connected [Dui96a, Sma95]. consecutive [DN97, Odl93]. consequences [DD93]. Conservation [JSH90, Cha97, Cha99, CDT93, CL97a, Che91, CHS90, CCF94, CG96b, CG97, CL91c, CFN95a, CFN95b, Fan98, HR93, LW96b, LW97, TT94, ThT97, vBv92]. conservative [CV99, DD92]. consistencies [WBL94]. Consistency [BS94b]. Consistent [MS95]. constant [BBC97, Bai94c, Bre94d, Leh90, LSEG91, Bre97]. Constants [Bri91, BBP97, Dil92a, Dil92b, KL97]. constrained [CQS98, Ked98, LR94, NY97, Sei99]. constraint [Car94, Sur97]. constraints [CGT97a, CGT97b]. Constructing [BS96c, PP93, CV99, DM98, LN96, Mac94b]. Construction [CH96, iN97, CMX99, Hip99, Lan96, LLS96]. constructive [BBG98]. contact [Cop99, Gau91b, Mac97]. continuation [Car94]. Continued [LS93a, Lor94, Shi95]. Continuity [Coc91, BC97]. Continuous [KB97, VZ95, Bai94c, Ber94, CY92, FP96, MO93, OZ91, Sho98, WBL94, YC93]. continuous-time [WBL94]. Contractivity [Len91a]. Contractivity-preserving [Len91a]. Control [Auc96, GHS91, HR94, Nie96b]. controlled [DR98b]. controls [GHS91]. Convection [Ric90, EJ93, GS97, LNS93, MM93a, MSS97, XZ99]. Convection-Diffusion [Ric90, EJ93, GS97, LNS93, MM93a, MSS97, XZ99]. converge [HL94]. Convergence [BG92a, BPWX91a, BPWX91b, Bre99a, CMP98, CH98, Cha99, CL93, CL97a, CL90, CT92a, CT92b, CL91c, CMO90, CMM90, FG91c, GS91, Gla92, HWC99, JSH90, LT99, Lu98, Mar92, MM95, Mor99, Nes96, NW98, Sid95, Sur90, Via91, Wal92, Wan99, BBN99, BD92a, BTT94a, BTT94b, BS99, BR94, CHL99, CFS95, Cha97, CQS98, Cho97, CMV92, DMT93, Dol99, FS92, GL96, Hon99, Hun92, Lab98, LP95, Li94b, Mar90, Mel97, NW95, NW96, OR92, Rab91, RS96a, Sul93, Van94, Ye91, DH99b]. convergent [BS90b, Che94a, CGT97a, CGT97b, Gau91a, Gau91b, Lew94, OS91, Sid95, SS96c, Van94]. Converting [Ziv95]. convex [Tod96, Dem95, LW97]. Convolution [Cal92, EL91, LO93, Pes96]. convolutions [Ste95]. Coombes [Sha96]. Cooper [Ano95r]. coordinates [Hal96]. Coppersmith [Kal95]. Coprime [HMW90]. Corliss [Ste98b]. Cornelius [Isa96c]. Corner [Bre99b]. corrected [CL91c, HHM94a, HHM94b]. correction [BPX91a, BCMR95, Hu98, Yin98]. Corrigenda [Ano92, BW96, CH90a, Mor90]. Corrigendum [Ano96, BO94, GC93, JMS97, Wal21, Woz97a]. Cosets [Mai91, Pow97]. Cosine [Dub92, ST91]. Cosine-Series [Dub92]. count [LW96a]. counterexample [Spi94]. Counterexamples [Xu91]. Counting [MVZ93, GN92]. coupled [Nes96]. Coupling [BDM90, BWW97, Li99]. Courant [Bre96a]. course [Isa96a, Isa96b, Kin99b, Sch96a]. Covering [DJ91, His94]. covers [MW94b]. covolume [Cho97, CV99, NW98]. Cox [Kli96]. Crandall [FP95]. Crank [Hun92]. crimes [BF93a]. criteria [Bos93a, Bos93b]. Criterion [JS92, Cik94, Cop90, DS95, Joe93]. critical [Kei96, Tol97]. crossings [KS93a, KS93b]. Croz [Cai96]. cryptographical [GvP98]. crystalline [KLL96, LL98]. cubature [CS96b, LC94, Xu92]. cube [SW98]. Cubes [Vas91, BRZ99]. cubic [Bel97, Cla96, DST97, FT91, Fuj93, FW90b, FW94, GS93, Oli92]. Cullen [KN95, Kel95]. Cunningham [For99]. Curtis [HST93]. curvature [Pao97]. curve [AM93b, Cre95, Sem98a, SW93, Smy91, Tei90, vW99]. curved [BK94, Lyn94]. Curves [Bar96, EdT90, AT99, AM93a, AM93b, BB93, Che96b, CW94, CS99, DH94, FR94, GLW99, iHT99, Kid99, MVZ93, Mey96, iN97, Ros95a, Ros95b, Rüc99, SV98, Sil90, Sil99a, Tra95, War98, Was91, van99, Beu96]. cusp [CW94, vv93b]. cusped [CHW99]. Cuspidal [AM92, Kno93]. cycle [BP93, Vas92]. Cycles [MM91, MM93b]. Cyclic [DJ91, Phe90a, Phe90b, Acc96, AM94, Gaá96, GP97, Gir93a, Gir93b, GS93, Lou95, Was91]. cyclically [EG90, EG91]. Cyclotomic [EM91, EM92b, Fuk96, Bre93, BCEM93, Jak98, Ste94b]. Cyclotomy [Laz93].

D [Ach99, Ano92, Gau96, Köl97e, Köl97d, Dur99, BH94b, Che94b, Krö91, Mon94a]. D. [Cai96, Che96c, Coo98, Wah97a]. DAEs [Zha98a]. Daniel [Wil97]. Dario [Hig96]. data [BL95, DS93, EL92a, EL92b, FK93, JSW99, LST96, Sch99]. David [Isa96c, Gau96, Ste99c]. Davies [War98]. Davis [SW95b]. Davy [Loz97]. Deák [N.93]. debate [Roj94]. Debnath [Tem97]. Decay [LW97]. Deciphering [TG94]. decomposable [Sma95]. Decomposing [CD97]. decomposition [AV99, BPWX91b, BPV98, CEL96, CMW95, DDD91, DD92, Gec93, HP92, LS95, LT99, MB96, Ste96b, Swa93a]. Dedekind [Coh96, Tol97]. Defect [BCMR95]. deficient [LW91]. defined [van99]. Definite [Hun91, Leh92a, MN90, CMP98, Hig98, Kap95, Leh92b, NW94, RS96b, Ser97b]. Deformations [vKP95]. degeneracies [BGJ99]. Degenerate [KX91, LP92, NW95, Zha95b]. Degree [DY90, KM91, Abe94, Acc96, AS91a, BBG98, Boy96b, Boy96c, Bra93, CS96b, Cre95, DO95, DH90, Fuj93, Gra98, Jak98, Roe99, Sel99, Sou94, Tak94]. degrees [CJT97, Gir93a, Gir93b, HM98, Lou95]. Del [Ura96b, Ura96a]. delay [Bru94, CG90b]. Delhi [War96]. delta [Laz93]. Demmel [Ano95q, Cai96, Wat98]. Dense [Pan90]. densest [Cha92]. Density [AL95a, AL95b, BN97, BS96b, FW90a, HB92]. Dependent [Ric90, BB99a, Du98, EL95, GP99]. depth [HH99b]. derivation [vBv92]. Derivative [Zha98b]. Derivatives [CG90a]. Descriptions [Ano95-27, Ano95-28, Ano95-29, Ano97, Col97, Hol97, Mat97, Shu97, Ste97a, Ste97b, Van97, Wah97b, Wri97]. design [Bar96]. Designing [Mey96]. designs [GT93]. Desingularization [Hou92]. Details [MM98]. Detecting [Ber98]. Detection [GM92b, BGJ99, Pea96]. Determinant [LZ92, FG91a, Hel97]. Determination [Lou95, Yam94]. determine [Coh93]. Determining [Sma99]. Deterministic [HMW90, AN92, Cot91]. Deuflhard [Isa96b]. development [Bra94, Ste94a]. developments [Gol94b]. device [CT92a, CT92b, CT94, Hol96, Zhu92]. Deville [Kea98]. Dew [Loz97]. DFT [Nie96c]. diagonal [Ish95]. diagonally [vSC92]. Diameters [Gra98]. Diaz [Res97]. Difference [ELR90, Wad90, AD91, CGJ95a, CGJ95b, CL97a, Che91, CL91c, DDD91, FMOS98, GASS91, Gol94a, HT94, HT95, KT90, Liu94, MW94a, SA94, Sil90, Sun95, TT95, Wu95]. Differences [Mor89, Mor90, Odl93]. different [L’E99a, LSW94]. differentiable [Wal91]. Differential [Sha96, WT90, Ale91, Auz98, By99, BS99, Bra96b, BI93, BC98, CG90b, CTW92, Coo98, CHT96, DCNZ96, HM97, Hu98, KL97, Kin97, Kin99b, Kvæ90, LTW98, Liu94, Lu98, Mac94b, Mak94, Mat97, Nie94, OR92, RR94, Ren90, Sch91, SV90, Ste94a, Tem94a, Wah97a, Zha93a]. differential-algebraic [Kvæ90, Mat97, RR94, Sch91]. differentiation [AH99, EB99, Ste98b]. Diffusion [EJ93, JSH90, Ric90, Ric92, Che94b, Che97, CT92a, CT92b, CT94, CG93, FK91, GS97, LSS99, LM99, LT99, LNS93, MM93a, MSS97, Pes96, SS96c, WL99, XZ99, Zho97]. digital [AMN90, Nie96b]. Dihedral [May92a, May92b]. Dikshit [War96]. dimension [Buh95, CH90a, CLT94, Gla92]. Dimensional [GM92c, SZ92, Ach99, BLS94, BCM92, BL94, BV92, CLT93a, CLT93b, Lab98, LSW94, Mad90, Sun95, Tez94, Yin98]. dimensionality [AH99]. Dimensions [NPV91b, Bai94c, Cha94b, Chi95, CL91c, GS91, NW98, NPV91a, OS91]. diminishing [GS98]. Diophantine [Bd98, Coh97c, Del92b, HBLT93, KTS97, Lan96, Leh90, SdW99, Sty93]. Dirac [Sei99]. direct [Han90]. direction [BF93b]. directional [Neu94]. Dirichlet [BL95, BK94, Kei96]. discontinuities [LM94b]. discontinuity [ZZ97]. Discontinuous [Ric92, Bai94c, CHS90, Eck93, JS94, KM98, LTW98, TB97]. discrepancies [TF93]. Discrepancy [NS90, EHN92, EH94a, GT93, Hei96, Hic98, Nie90, WW97]. Discrete [CF94, Dem95, Du98, RAG91, AD93, BL95, BLP97, Dem90, DH99a, DH99b, Fas97, FR94, LL99, LX93, LW96b, LW97, MST99, Nie96c, Pao97, PST98, Rüc99, Sem98b, Sem98a, ST91, TG94, Tsu90, Zha93a]. discretization [BB99b, CFS95, GASS91]. discretizations [AD91, Osw96, Ver94, Ver98, WBL94, WH99]. discretized [BI93, CRV92]. discriminant [BFP93, Fuj93, FW90b, FW94, Sel99, Str92b]. Discriminants [May92a, CDO98, DO95, HW90, May92b]. Dispersive [ELR90]. displacement [CL91a, GKO95]. Dissections [Jep96]. dissipative [CJT97, Hag91]. distance [Dem95]. distances [BH97]. distributed [Auc96, GHS91]. Distribution [CL97b, Ede92, EHN97, HM98, Bai94b, CLT93a, CLT93b, GP91, GC93, Per92, Tez94]. Divergence [Lor94]. divergent [Sid95]. dividing [Gra95]. Divisibility [MTU94, EM97b, FR94, GN92, Jak98]. division [McK94, Smi96]. divisor [Cus93, FR94, Ian99, Rüc99]. Divisors [Coh90, Coh97a, Sch93a, Vou95]. Dixon [LGRA94]. Domain [CEL96, LS95, AV99, BPWX91b, BPV98, CMW95, DDD91, DD92, LT99, MB96, She93, Swa93a]. domains [BK94, Bra95, Coh96, FK91, HHM94a, HHM94b, Hal96, Mad90, Wan91, Yin98]. Donald [Isa96c]. Donato [Ano95q]. Dong [Nie96b]. Dongarra [Ano95q, Cai96, Hou96a]. Dormand [Auz98]. Dorothy [BW96]. Double [Pin99, BS94b, CP97, Del90, GL98a]. double-layer [BS94b]. double-well [CP97]. Doubly [AM92]. Douglas [Hun98a]. drift [Che94b, CT92a, CT92b, CT94]. drift-diffusion [Che94b, CT92a, CT92b, CT94]. Drinfeld [Hay91, DH94]. dual [NRS95, Tod98]. Dumka [Köl90]. Dummit [Ano92]. dynamical [BS97, WBL94]. dynamics [CL93, WBL94].

E. [Ach99, Cai96]. Earl [Kli96]. edited [Auc96, Bra96a, Bra96b, Bre96a, Che96d, Che97, Ell96, Fal98, Gau96, Hou96a, Hou96b, Loz97, Loz98a, Mey96, Res97, Rud98, SS96b, Shu96, Ste98b, Tho97, Wah98b, Wah98a, War96, Wri97]. edition [Fik98, Köl97e, Lam96, Lam97, Loz98c]. Editor [Ano95y, Len97]. Editors [RW92, Ano95g, Ano95u, Ano95r, Ano95j, Ano95n, Ano95d]. effect [BS92a]. effects [SS94, SBS95]. Efficiency [Car96, HH99a]. Efficient [BC98, Hei96, Hol97, PT92, She94, Wag98, CHT96, Eck93, Mas98, Nie92a, Nie92b, Tes98]. Efi [SS96b]. Eigenforms [GM92a]. eigenfrequencies [RS96a]. Eigenvalue [BN92, Lev92, LZ92, Lin94, Bai94a, EST90, Leb94, Mas94, Mor96, Vv96]. Eigenvalues [Bou99, BO94, EM95, HGK98, Swa93b]. eigenvectors [BO94, Leh97]. Eight [Mor89, Mor90, Sel99]. Eijkhout [Ano95q]. Eisenbud [Ano95j]. Elastic [HW92, Fal96, Mak93]. elasticity [BS92b, Bre94a, Bre94b, Fal91]. elasto [NRS95]. elasto-perfectly [NRS95]. electrically [HR94]. electromagnetic [Fen99]. Electronic [Köl90, Mas94]. Element [BDM90, Coc91, CPS92, JSH90, NPV91b, Ric90, Sur90, ACM98, Arn97b, AB97, BO94, BL95, BS92a, BL93, BB99a, BB99b, BS94b, BF93a, BDR98, BG94, BK94, BS92b, Bre96a, Bre96b, Bre96c, CS95, Car97b, Cha94b, CTW92, Che94b, CL90, CL91a, CHS90, CT92a, CT92b, CT94, CL91b, Cop99, CLT94, DDP95, DCNZ96, DM99, Dol99, DL92, EL92a, EL92b, EL95, EJ93, Fal91, FO94, FK93, FP96, Gol91a, Gra95, GHS91, GMP91, Han90, HST98, Hon99, HR94, HWC99, HGK98, KM98, KS92, LSS99, Leb94, LL98, Mak93, MP95, NRS95, NPV91a, OS91, Pes96, RS96a, RVW96, SW95a, Sch98a, SS94, SS96a, SZ90, Ste90, Ste99a, SO91, SBS95, Sur97, Ver94, Ver98, Wal92, Wan91, WH99, XZ99, Zha90a, Zha90b, Zha95b, Zha98b, ZF93, Zho97, Zhu92]. element [Bre96a]. element-Galerkin [BO94]. elementary [SW98]. elements [Al 94, CO98, CH90b, CH96, CMW95, Dur99, Gir95, GH97, Hip99, KLL96, Pih93, Val91, WH99, Žen95]. elimination [GKO95]. Elkies [Dew98]. Elliptic [AM93a, Car92, Dör92a, EdT90, GM92c, LM90, SdW99, Sur90, AT99, AC95, AM93b, BLP93, BB93, Bre99a, Car91, CFS95, CEL96, CV99, CW94, Cre95, CS99, CLT94, DCNZ96, Dör92b, DR98b, DH94, ELV91, GLW99, GP91, GC93, Hol97, HWC99, Kha99, Kid99, Köl97e, Köl97d, KT90, MVZ93, MP95, iN97, Nii92, Noc95, OS91, Ros95a, Ros95b, RW93, RVW96, Sem98a, She94, Sil90, SW93, Sil99a, Smy91, Swa93a, Tra95, Vv96, Ver94]. elliptic-type [Kha99]. Ellison [Isa96c]. Embedded [vSC92]. Emden [Nie96c]. Emma [Dar91]. enclosing [APS93]. encyclopedia [Bai96]. endpoint [MM94a]. energy [Gla92, HDJ94]. energy-preserving [Gla92]. engineering [Zha98a]. engineers [Köl97e, Köl97d, Kok96]. Engl [Che97]. English [Fik98, Lam97, Köl97g]. enhanced [BR94]. entire [Grö99, Mül97, RS94]. entropy [BBP96, JS94, KP94, LL99, Noe96, Yan96b, Yan96a]. Enumeration [Bre97, BFP93, YR92]. environment [Res97]. Environments [Hou96a]. Epstein [Kha99]. Equal [Ekl96, Ekl98]. Equation [Del92b, Td91, WT90, AI91, Ad97, AN92, Arn94, AC90, AL91, BL95, BB99a, Bd98, BTT94a, BTT94b, BT97, BH94b, BV92, BI93, Car97a, CG90b, CTW92, Coh97c, DDD91, De 95, EL92a, EL92b, FK93, FP96, Goe98, Guo95, HDJ94, HHM94a, HHM94b, HS94, HBLT93, Hon99, Hua91, KM98, KTS97, LM99, Léc91, Li99, LNS93, Lu98, LST96, Mon94a, NTS98, SW95b, She93, SH95, Ste98a, Sun95, TW94, Zha98b, DOSS91, KM90]. Equations [GM92c, HW92, KX91, LT91b, Ric90, Yan90, Acc96, Ach99, APS93, Ale91, AV99, Ang96, AS91b, Auz98, BB99b, By99, BS99, BG94, Bra96b, Bre90, Bru94, BPV99, BC98, Car96, CGJ95a, CGJ95b, Che91, CT92a, CT92b, CT94, Coh93, Coh95, Cop94, CFN95a, CFN95b, Cot91, CHT96, CRV92, CMM90, DS99, DCNZ96, DHNL^{+}99, EL91, EL93, EDK97, Fal91, FJP97, FK91, GANT99, GS91, Gla92, GHS91, GMP91, GL98b, GL99, HT94, HT95, HM93, HS95a, HM97, Hu98, Iya95, KL97, KX94, KB97, Ked98, KSS98, Krö91, Kvæ90, Lar97, LTW98, LSS99, LT91a, LM94b, LSEG91, Liu94, LO93, LO95, Mac94b, Mak94, Mel97, Mel98, MPR96, MM98, MS98, Néd94, NW96, NW98, Nie94, Nii92, OR92, Pal96, Pes96, Pet91, RR94, Ren90, Rob94]. equations [Rum90, SV98, SW99, Sha96, She96, SA94, Sma95, Sma99, SV90, Ste97a, Ste94a, SdW99, Sty93, Tem94a, Tra95, vBv92, Ver94, Ver98, Wal92, Wal99, Wal21, WL99, XZ94, XZ96, XZ99, Yin90, Yin98, Zha90a, Zha90b, Zha93a, Zha95b, ZF93, Zhu92, Coo98, Elm97, Kar97, Kin97, Kin99b, Mat97, Wah97a]. equidistributed [L’E96, L’E99b]. equidistribution [EHL97]. equivalence [JM92, JMS95, JMS97]. Equivalent [Wei97]. Eratosthenes [CG90b]. Erdélyi [Köl97h, Köl97c]. Erdos [BL90, Coh98, LSW97, Zha93c]. ERH [Rum93b, Rum93a]. Eric [Wag98]. Eriksson [Coo98]. Errata [RW92, FW94, Köl97c, Köl97a, Köl97b, Köl97d, BR05, BR11, Köl95, Köl97e, Köl97g, Köl97f, Köl97h, Lam96, Lam97]. Erratum [Tho94, Woz97b, Fik98]. Error [Bai94a, By99, Cal92, CT94, CPS92, DDP95, Die94, DS91, Dur99, Ehr94, EL92a, EL92b, It91, JSW99, Leb94, Léc96, LP92, LM90, Mit90, NPV91b, Ste90, Ste92, Al 94, AM94, Bra93, Cal91, CS95, Car96, Car97a, Car97b, Cha94b, CCF94, CG96b, CG97, CL91b, DLP93, DM99, DN90, EL91, Fan98, Gas93, GT94, HJ95, Hic98, HL94, LST96, Nie96b, Nii92, NPV91a, Noc95, Noe96, Pao97, Sch99, SW96, Sch98a, She96, Tad93, TT95, Ver94, Ver98, WL99, Wei94, WH99, Ziv95, Ziv99a, de 96, ELV91]. errors [vS96]. Escott [Mal97]. essentially [Ber98]. Estep [Coo98]. estimate [Car97a, Car97b, Cha94b, CCF94, Leb94, Pao97, TT95]. Estimates [BX91, ELS93, NPV91b, Al 94, BLSW96, BL95, BS94b, BS99, BPWX91a, BPWX91b, BP93, CS95, Car96, CT94, CG96b, CG97, CL91b, CLT94, DLP93, Die94, DN90, Dur99, EL91, EL92a, EL92b, Fan98, FP90, Gas93, GT94, HJ95, JSW99, Kny97, LP92, LST96, Mit90, NPV91a, Noc95, Noe96, SW95a, SW96, Sch98a, She96, Tad93, Tra95, Ver94, Ver98, WL99]. Estimating [CJT97]. estimation [Auc96, By99, Rum96]. estimators [DDP95, WH99]. Euclid [Tei98]. Euclidean [Cla96, GS93, RS96b]. Eugene [Loz96, Qua96a]. Euler [Dil92b, Bre94d, Dil92a, Krö91, SW94, Yin98]. Euvrard [Ano95t]. evaluating [SV90]. Evaluation [CG90a, Sem98a, Tak94, BD92b, Car98, Chi95, Cra98, LG97, LO94, McL94, Mon94a, Rab90b, Sem98b, lX96]. Even [Hun91]. Every [HC98]. evidence [BH93]. evolution [Car94, FK91, HS95a, LST96]. evolutionary [DVZ94, FSSC90]. Exact [HW92, Ziv95]. example [Wri96]. Examples [van99]. exceeds [HC98, Ian99]. Exceptional [BN92, NS98]. excursion [Oli99]. Exhaustive [Fis90]. Existence [Fan98, ESA94, FMOS98, GMP91, JKN91, Nie92a, Nie92b]. Expansion [BBG91, Rum96, Ada95, Boy96b, Boy96c, Dom96, Fly98, Sch98a, VC94]. Expansions [Kei92, BF92, Boy96a, Eck93, Eck95, HHM94a, HHM94b, Mac93, Tem94b]. experiences [BH97]. Explicit [Bac90, BS94a, BS96a, Bos93a, Bos93b, DD92, DM99, EHI94, GHO92a, GHO92b, OS92, ACM98, GF97, GM96, JZ92, OZ91, VZ95]. Explicit/implicit [DD92]. Exponent [KM91, WW97]. Exponential [BBN99, RS90, Ste92, Apo93, GS95, Grö99, Kui95, RS94, Sty93, Wag94, Wal91]. exponents [Lou95, Sut98]. Extended [CMO90, BS96c, CMV92, GL96, HST93]. Extending [KW90]. Extension [GH97, Apo93, Cik94, FS95, HT94, LLS96, MS90, Sch93a]. Extensions [Del92a, Bun97, DP96, JKN91, Lou92]. exterior [HHM94a, HHM94b, Li99]. extraction [SW99]. extrapolation [Lyn94, Sid95, Wan91, XZ96]. extrema [LL99]. extremal [Che96c].

F [Köl97e, Köl97d, Köl97h, Shi96, Wah97a]. F. [Bos98, Köl97c, Köl97h]. Faber [GHO92b, BH94a, CSW92, CM95, GHO92a, He94, Kui96]. factor [BE99, Car91, ELS93, HC98, Jha95]. factored [Wil93]. Factoring [LM94a, Nie94, Roe99, SW98, Wan90, WS94, KS98, McK99, MS90, SW93, Val91].

Factorization [Bre99c, Del92a, AM93b, Göt94, LLMP93, LLMP95, NG95, Shp93, Sil97, Wag96b]. factorizations [BTW90, CV95a, HP92, Wag96c]. Factorizing [Hig98]. Factors [BR98, BR05, BR11, Car92, DK95, BN97, Bre99b, Bre93, CD99, FG91a, Gos95, HS97b, You98]. fails [HB92]. fair [Mey96]. Families [GM92a, Ada95, Goe98, Noe96, Pet91, Zha98b]. family [GP97, MPR96, NS98, PP96, Van94, Was91]. Farin [Che96b]. Fast [CB93, Cra98, GKO95, Nie96b, PST98, SV98, lX96, Bel97, EG99, Lab98, LM99, MP96, ST91, Str92a, Val91, WL99, Cal91]. Faster [Jha95, Lin98]. Faulhaber [Knu93]. Feigenbaum [Bri91]. FEM [BWW97, Li99]. FEM/BEM [BWW97]. Fermat [BR05, LLMP95, BR98, BR11, Bre99c, Cop90, CDNY95, DS95, DK95, EM97b, Gos95, LLMP93, McK99, RB94, You98]. Festschrift [Ell96]. FFT [HTS90, MS90]. FFTs [LF93]. Fibonacci [Bre94c, DK99, Mus93, TF93]. field [GLW99, Lou92, Pom94, Ste94b, Str92b]. Fields [Den90, Dia91a, Dia91b, HM92a, Len91b, MW92b, Sho90, Sho92, Sol92, Acc96, AK99, AD93, Bel97, BMO90, BG99b, BFP93, BPv94, BS96c, BE99, Cla96, CDO99, CW94, CS99, DO95, DST97, FJP97, Fuj93, FK96, Fuk96, FW90b, FW94, Gaá96, Ge94, Gir93a, Gir93b, GS93, HM98, Han96, HM92b, Iya95, JM92, JMS95, JMS97, Jak98, KS98, Kid99, Lou92, Lou95, Lou97, MM97, Mol93, MW94b, MM94b, MM94c, MST99, Nie94, NG95, NS98, Oli92, OS97, PR99, Pil90, Pv93, PW98, Pow97, Sch93a, Sch96b, Sch98b, SPD94, Sel99, Sem98b, Sho99, Shp93, Sri98, Tak94, Tay96, Tei98, Wan90, Wan97, Was91, Yam94, vKP95]. Fifth [Fik98, Lam96, Lam97, PP96]. fifth-order [PP96]. Fifty [Bre96a]. Figures [AMN90]. filtering [CHL99]. Finding [AM93b, Len91b, Lin98, Sho90, Zha94, FBA99, Pil90, tR94]. Finite [AD91, BO94, BL93, BB99a, BLS94, BDR98, BDM90, Bre96a, CGJ95a, CGJ95b, CTW92, Coc91, Cop99, Dub92, HM92a, HR94, JSH90, Len91b, MM93a, Mak93, NPV91b, Pes96, Ric90, SZ90, Sho90, Sho92, Sur90, Wad90, Zhu92, AD93, ACM98, Al 94, Arn97b, AB97, BR90, BL95, BS92a, BF93a, BG94, BK94, BS92b, Bre96b, Bre96c, BS96c, BJT97, Car97b, Cha94b, Che94b, Che91, CL91a, CH90b, CHS90, CT92a, CT92b, CT94, CCF94, CL91b, CL91c, CMW95, CLT94, DDP95, DDD91, DM99, Dol99, DL92, Dur99, Eck95, EL92a, EL92b, EL95, EJ93, Fal91, FO94, FMOS98, FK93, FP96, GASS91, Gir95, Gol91a, GHS91, GMP91, GS97, HM98, Han96, HM92b, Hip99, HWC99, HGK98, Iya95, KS98, KM98, KS92, KLL96, LSS99, Leb94, LL98, Lin98, Liu94]. finite [MW94a, MP95, MM94b, MM94c, MSS97, Mül97, NRS95, Nie94, NG95, NPV91a, Noe96, OS91, OS97, Pil90, RS96a, RVW96, SW95a, Sch98a, Sch96b, SS94, SS96a, Sem98b, Shp93, Sou94, Ste90, Ste97b, SO91, Sül92, SBS95, Sur97, Sut98, Tei98, Ver98, Wal92, Wan90, Wan97, Wan91, WH99, XZ99, Žen95, Zha90a, Zha90b, Zha94, Zha95b, Zha98b, ZF93, Zho97, Živ99b, Ver94]. Finn [Gil98]. First [BDM90, Yan90, AS91b, BS99, BLP97, Car96, Car97a, DS95, Isa96b, Jha95, Kap95, KB97, Kin99b, Nic99, Wal92]. First-Kind [Yan90, AS91b, Car96, Car97a]. first-order [BS99, Wal92]. fits [Bai94c]. Fiume [Loz96]. five [DD93]. Flannery [Loz98c]. flow [BF93a, BWW97, MP96]. flows [CIZ97, CH98, HR94]. fluid [BDR98, HR94, RS96a]. fluid-solid [BDR98]. fluid-structure [RS96a]. Flux [CG97, CL91c]. Flux-splitting [CG97]. Form [BL90, Cos91, Coh96, CK95, MPR96, Oli98, Sma95]. Forms [Hun91, Leh92a, Sko92, Bea97, BPX91b, CLR97, CW94, EK97, HB92, Ish95, Kap95, Leh92b, Leh97, LJ96b, vv93b]. Formula [AW97, DLD90, FO93, GASS91, GF97, GR95, Neu94, RS94, de 96]. Formulae [BN90, RS90, Ben98a, CS96b, Ehr94, EG99, GN96, Not92, Wei97, Xu94b]. Formulas [vSv92, CMM90, GM96, Hab93, Lau96, MM94a, Peh90a, Peh93, VV93a, Xu94c]. formulation [SS91]. formulations [EST90]. Fornberg [Shu97]. Fortan [Loz98c]. Fortran [Loz98c]. forward [AL91]. forward-backward [AL91]. Foufoula [SS96b]. Foufoula-Georgiou [SS96b]. four [BCEM93, CD97, Ekl96, GS93]. Fourier [Cal91, HA93b, LG93b, Nie96c, CB93, DGPR92, Eck95, HA93a, LG93a, Mas98, ST91]. Fourier-cosine [ST91]. Fourier-sine [ST91]. fourth [BCM92]. fourth-order [BCM92]. FPGA [BB94b]. fractional [HR93, OR92]. fractions [LS93a, Lor94, Shi95]. frames [RS98]. framework [CV95a, CV99, MS97b]. Francoise [Ste97b]. Frank [Qua97]. Frayssé [Ste97b]. Frederico [Ste99a]. Fredholm [HM93, Hu98, KX94, Lin94]. free [AB97, GN92, HS97b, SS91, Ye94]. free-boundary [SS91]. freedom [CJT97]. frequency [AV99, Ste96b]. Friedman [Isa96a, Köl97e, Köl97d]. Friedrichs [BL94, Che91]. Frobenius [Pil90]. Front [Shy94]. full [Bre99a, LT99]. fully [BL95, Kvæ90, Zha93a]. Function [Den90, Kei92, Alz93, Alz97b, BD92b, BF92, Car98, CHT96, CSV94, Cus93, Dem95, Dom96, Fly98, GLW99, GS95, GN92, LSW97, MM97, MST99, PR99, SW92, SW94, Shi96, Wah97b, Wan99, Wei94, XZ97, Ziv99a]. functional [Lar97, Tho94, dB98]. Functions [Ber90, BBG90, EM92a, Köl90, MN90, RS90, Tem92, AT99, Alz97a, Bai94c, Ben98a, Ber94, BR94, CM94a, CY92, CK99, DVZ94, Dim96, Eck93, Eck95, Eck98, EM92c, EM94, FS92, FO93, FK91, GS90, GN96, GLW99, GF97, GM96, Grö99, GR95, LSW94, Lou92, LO94, Mac93, MST96, MM94a, MM95, MS98, Mül97, NW94, OW97, Peh93, RS94, RS96b, Rum96, Sch99, Sch97, SZ90, Str92a, Tol97, Wal91, WG92, Wen99, Wim98, YC93, YZ97, ZZ97, SSW99, Köl97c, Loz98b]. Fundamental [Hig96, BBG93, SSW99]. fundamentals [Ste99a]. Further [Bur96, LSW97]. fused [LF93]. fuzzy [Kli96].

G [Bra96a, Fal96, Kel95, Köl90, Köl97c, Köl97h, Tem98, Wah97b]. G. [Che96c, Kin99a, Ser97a]. Galerkin [Arn97b, AL91, BO94, BS99, CHS90, DD92, DMT93, FSSC90, GANT99, GM92c, HST98, JS94, KM98, LTW98, Ric92, SW96, She94, Tem91, Tra95, WL99, Wen99, Zha93a, ZF93]. Galois [CM94a, Cla96, MM97]. Gamma [Tem92, Alz93, Alz97a, Alz97b]. Gander [Ste96a]. Gaps [CD99, Nic99]. Garrett [Sha96]. Gary [Hou96b, Shu96]. gas [CL93]. gauge [Du98]. Gaunt [lX96]. Gauss [Ehr94, FvS99, GvP98, GM96, KX94, Lau97, Not92, Peh90b, ZH90]. Gauss-type [KX94]. Gaussian [EG99, FP90, För94, Gau91a, GKO95, GZ93, Lau96, Peh93, Sch97, Xu92]. Gautschi [Len97, Ste99b, Ell96]. GBDF [Sul93].

Gegenbauer [GS95]. General [ZH90, CQS98, CV99, CG96b, CGT97a, CGT97b, FvS99, Fly98, LG93a, LG93b, MS97b, Hag91]. generalised [LS98]. generalization [LS91a]. Generalizations [LGRA94]. Generalized [Dil92a, Dil92b, Dub93, HTS90, LL99, Mus93, NW94, Rab90a, RB94, BR98, BR05, BR11, Bre94c, CGJ95a, CGJ95b, Cho97, DK95, EH94b, Far94, Hic98, Hor97, Kha99, Sid95, Wal91]. generated [BH94a, JKN91, Kui96]. generating [BH93, CY92, FG91b, LH98, Peh90a, TB97, YC93]. Generation [Val91, Bai94b, CM96, Emm96, Nie95]. generator [KWY96a, KWY96b, TW92]. Generators [Fis90, N.93, Wan97, CL94, CL96, CL97b, Deá90, GV95, LT91c, L’E96, L’E99a, L’E99b, WC93]. Genus [Den90, Sko92, iHT99, Kap95, MST99, van99, vW99]. geometric [Bar96, BG96]. geometrical [Che96a]. geometries [Lee99]. geometry [Che96b, War98]. geophysics [SS96b]. George [Ste98b]. Georgiou [SS96b]. Gerald [Che96b]. Germain [Dub96, IJ99]. GFSR [TF93, Tez94]. Gibbs [Gel97, GS95]. Gilbert [Gau96]. Ginzburg [Du98]. given [BBBP98, Bra93]. Glen [Wah98b]. Global [CQS98, Den90, Kea98, Sch98a]. globally [CGT97a, CGT97b, OS91]. GMRES [XC92]. Godunov [Fan98]. Goede [Ach99]. goes [Kin99a]. going [Ada95]. Goldbach [DGNP93, Sao98, Sin93]. Good [DS91, HMN93, CM94b, L’E99a, DOSS91]. Gordon [Köl97f, Köl97g]. graded [Noc95]. gradient [Wri97]. gradient-related [Wri97]. Gradshteın [Köl95, Lam97]. Gradshteyn [Fik98, Lam96]. graduate [Kin99a]. Graham [BL90]. graphical [Loz96]. Graphs [BN92, Dui96a]. Gray [Ano95i]. great [BB93]. greater [Dus99]. Green [XZ97]. Greenbaum [Cai96, Sil99b]. Greenberg [FK96, Fuk96]. Greg [Loz98a]. Grid [Yan95, LSEG91, TV93]. grids [Ad97, BLS94, CG97, ELV91, GASS91, LS95, Noe96, Sch98a]. Griewank [Ste98b]. group [CMS92b, CMS92a, FR94, Gol91b, His94, Mas98, Rüc99, San92, Sch93c, Sch93b, Sem98a, Tes98]. Groups [Sol92, BR90, BC94, BPv94, BJT97, But94, CM94a, CD98, CB93, CM94b, CDO98, Far94, HS92, Jac99, Kno93, Lee99, Leh90, Lou95, MM97, OS97, Pea96, Pv93, PW98, San92, Sch98b, Sou94, Sut98, SW98, Xu94a]. growth [Mül97]. guide [Kli96, Shu97, Cai96].

H [Gau96, Kea97, Loz98c, AFW97]. H. [Auc96, Kok96, War96, WH93]. Hackbusch [Ano95v, Ste97a]. Half [Coh95, EL91, Gau94, Len97]. half-century [Gau94, Len97]. half-line [EL91]. Half-step [Coh95]. Hamilton [CFN95b, CFN95a]. Hamiltonian [ASS93, Hua91, LR94, OS92, Sei99, vKP95]. Hammarling [Cai96]. Hammerlin [Ano95g]. Hammerstein [KX91]. Handbook [Bra96a, Köl97d, Qua96a, Kli96, Köl90, Mea96, Köl97e]. Hans [Wag96b]. Hansbo [Coo98]. Harald [Sch96b]. Hari [Nie96b]. Harmonic [Spi90, AV99, Coh97a, Zhe94]. harmonics [Gel97]. having [RS95, Rob93]. heat [AL91, DDD91, HS94, Hon99, LM94b]. heating [EL95]. Hecke [Gec93, Kno93, Leh90]. height [BB93, CS96a, Sil90]. heights [Dre98, Sil97, Sil99a]. Heinz [Che97]. Helmholtz [Ad97, Li99]. Hemker [Ano95n]. Henri [Sch96a]. Henrik [Gil98]. Henson [Nie96c]. Hentenryck [Kea98]. Hereditary [Den90]. Hermite [By99, Coh96, FN95, FK91, NW94, PT92, PT94, Qua96b, Žen95]. Hermitian [CMP98, EK97, Til97]. Hessenberg [vS96]. Heuristics [CM94b]. Hierarchical [Dör92a, Dör92b, Ste96b]. High [Hou96b, MS98, Wah98a, AS91a, CL97a, Eck98, FW90a, HJ95, KM90, LSW94, MS97b, iN97, Roe99, Shy94, Yan96b, Yan96a]. high-accuracy [MS97b]. high-degree [Roe99]. high-dimensional [LSW94]. high-order [HJ95, KM90]. high-rank [iN97]. high-resolution [Shy94, Yan96b, Yan96a]. Higham [Van97]. Higher [Köl97c, SZ92, Sur90, Gur90, Noe96, SA94]. Higher-Dimensional [SZ92]. Higher-Order [Sur90, Gur90, Noe96, SA94]. highly [Loz97, Mac97, Noc95]. Hilbert [Wei95]. Hill [Köl97h]. Hilliard [EL92b, BB99a, EL92a, Sun95]. history [Isa94]. hodograph [FN95]. Hohmann [Isa96b]. Hölder [GP99]. holders [Oli99]. holomorphic [Pal94, Pal95]. Holt [Kel95]. homogeneous [Cop94, HDJ94]. homotopies [LW91, LW93]. Homotopy [LZ92]. Homotopy-Determinant [LZ92]. honor [Ell96]. Hopf [ABZ95, EL93]. Horst [Gau96]. Hoschek [Ano95b]. Hrebícek [Ste96a]. Hubbell [Kha99]. Hunt [Sha96]. Hunter [Bia90]. Hurwitz [CK99, Leh90]. Hybrid [Vas92, NRW95]. Hyperbolic [JSH90, CHW99, Cha97, Cha99, Che91, FG91c, GS91, GT91, LSEG91, Ren90, Wal92, Yan95]. hyperelliptic [PR99]. hypergeometric [LP95, Lor94]. hypotenuse [Ziv99a]. hypothesis [BS96c].

I. [Kea97, Köl95, Lam97, Res97]. IBM [Roe99]. Ideal [Del92a, Smy91, Lou95]. ideas [Ste94a]. Identities [Spi90, BBG98]. Identity [BBG91, BBG93]. II [AL95b, BL96, CG97, EG91, EM94, MN90, MM98, NW96, Ols95b, PP93, SW95a, Zha90b, Živ94b]. III [Ste99c, Zha95b]. Ilin [Wah98a]. ill [AY94, ACEO98, Car94, Ser97b]. ill-conditioned [AY94, Ser97b]. ill-posed [ACEO98, Car94]. ILU [CV95a]. Ilya [Nie96a]. Image [CG93]. imaginary [BF92, BG99b, CW94, Gir93a, Gir93b, Hig98, Lou92, Lou95, PR99, Sho99, Yam94]. immersed [AY94]. impasse [RR94]. implementation [AC95, CB93, Coo91, Ste94a]. Implications [Laz94]. Implicit [ACM98, WT90, Coo91, DD92, Kvæ90, Len91a, MO93, vSC92]. Implicit-explicit [ACM98]. Implicitly [MS97a]. Imprimitive [DO95]. improper [XZ94]. Improved [BG92a, BCT91, EH94a, GH94, Sch99]. Improvement [Str92b]. including [BP93, DR98b]. Incomplete [Tem92, Alz97b, BTW90, Dom96]. incompressible [BDR98, GMP91, Kar97]. indecomposable [LS94, Sch93c, Sch93b]. indefinite [BLP93, Hab93, XC92]. independence [EH93]. independent [SA94]. index [BC98, Kvæ90, MPR96]. index- [BC98]. India [War96]. inductive [But94]. Industrial [Isa96a]. industry [Rud98]. inequalities [Alz93, Alz97a, Alz97b, AL95b, BBP96, CQS98, Che96c, KT90, Noe96, Sch98a, Yan96b, Yan96a]. inequality [CGT97a, CGT97b, JS94]. inertial [GANT99]. inexact [BPV98, Mar90, Mor99]. inf [DN90]. inf-sup [DN90]. Infinitary [Coh90]. Infinite [BBG90, Goe98, GL98a]. Infinitely [Wal91]. infinity [Ada95, LW91]. influence [Zhu92]. information [KP92, OW97, WG92]. inhomogeneous [HT95, Lev92]. Initial [Ste94a, AD91, CDL97, DS93, GT91, Hon99, Hun92, Hun98b, LS91b, Mat97, Wu95, Zhu92]. initial-boundary [GT91, Hun92, Hun98b, Wu95, Zhu92]. initial-Neumann [Hon99]. initial-value [Mat97]. inner [BLP97]. input [MS98, Shi95]. Integer [Coh90, Jus90, LSK91, Bai96, BE96, CF94, FBA99, HS97a, Val91]. Integers [HMW90, BGN93, CD97, CD99, De 95, EL96, Fla96, Gra98, HP99, HS97b, MTU94, Wag94, WS94]. Integral [Pea96, Ste97a, Tem97, Yan90, AC90, AS91b, Bru94, Car96, Car97a, CH98, Coh97a, CRV92, CMM90, DP96, Dre94, EL93, Fly98, Gaá96, GP97, HM93, JKN91, KX94, Kha99, Köl97h, Lin94, Mak94, Mon94a, MS98, Néd94, Pal96, PP93, SV98, SH98, SW99, Sou94, XZ94, XZ96]. Integrals [Car92, Köl97a, Köl97b, Köl97g, Rab90a, Bia90, Car91, Chi95, Die94, Fik98, GT94, GL98a, HT91, KX94, Köl95, Köl97e, Köl97d, Lam96, Lam97, Rab90b, Rab91, VC94, Wei94, XZ94, Köl97f]. integrands [Cro93, Rab90b]. Integration [Dim96, LSK91, NS90, Del90, EB99, Gol91a, GT93, Hab93, HMN93, LT94a, LSW94, Léc96, LR94, MM94a, MM95, Nie92a, Nie92b, Pal96, Sei99, SW97, Kea97]. integrators [BC98]. Integro [WT90, CTW92, Hu98, LTW98, Mak94, Zha93a]. Integro-Differential [WT90, CTW92, Hu98, LTW98, Mak94, Zha93a]. intensity [Bre99b]. interaction [Mas94, RS96a]. interface [AD91, LM94b]. Interfaces [LM90]. interfacial [CH98]. Interior [Tod96, RVW96, SW95a, Gas93, PJV94, Tod98]. Interior-point [Tod96, PJV94, Tod98]. Intermediate [BG92a, Zha98b]. International [Wah98a, Isa96c]. interpolant [BBN99]. interpolating [CMX99, FT91]. Interpolation [Bra95, CSW92, CMO90, HTS90, Hu98, MN90, dR92, Ber94, BC97, CD96, CMV92, EM97a, FN95, GL96, JSW99, MS97b, NW94, PT92, PT94, Qua96b, SX95, Sch99, SZ90, VV93a, Xu92, de 96]. interpolatory [CD96, Dem96]. interval [Abe94, Ber94]. intervals [FG91b, HP99, MES98]. intractability [SW97]. Introduction [Sch96b, Che94c, Loz96, War98, Wim98, dB98, Ste99b]. invariant [CMX99, Du98, FS95, LW91, MS95, Ura96a, Ura96b, Xu94a]. Invariants [EM92b, BCEM93, CM96, EM91, KS93a, KS93b, Tay96]. invented [Roj94]. Inverse [Che97, Buf90, Che94c, RS92, Wan99]. Inversion [DGPR92, Tem92, CV95b, MST96]. Inversive [EH91, EHN97, EHN92, EH93, EHI94, EH94b, EH94a, EHE96, Emm96, Nie90]. invert [MS97a]. investigations [Bur96]. Involving [Spi90, FO93]. Irreducible [Sho90, Sou94, BR90, BC94, BH97, HM98, OS97, Wan97]. Irregular [BCS92, BCEM93, Grö92, Grö99, BLS94, CG97, Noe96, Sch98a]. isentropic [CL93]. Iserles [Fal98, Kin99b, Tho97]. Isomorphisms [Len91b]. isoparametric [Leb94]. isospectral [CIZ97]. István [N.93]. Iterated [Bru94, EDK97, Odl93]. iteration [BS90a, DCNZ96, DJLZ97]. Iterative [BLP93, EG90, EG91, Elm96, Elm97, GM92c, GL99, MS93, Sil99b, BPWX91b, CV95b, MP96, MES98, Woź94, Woz97b]. IV [CHS90, GS95]. Iwasawa [FS95].

J [Bai96, CMM92a, Gau96, Hou96a, Isa96c, Kin97, Leh92b, Res97, Sha96, Tod98, Van97, Wri97]. J. [Bra96a, Bra96b, Cai96, Dav96, Hol96, Isa96c, Res97]. J.-I. [Res97]. J.-L [Res97]. J.-L. [Bra96b]. Jack [Hou96a]. Jacobi [CFN95b, Mac94a, CFN95a]. James [Mea96, Kel95, Wat98, Wri96]. Jean [Bar96]. Jeffrey [Wag98]. Jenn [Nie96b]. Jenn-Dong [Nie96b]. Jensen [Gil98]. Jianming [Loz98b]. Jin [Loz98b]. Jírí [Ste96a]. Joe [Kea97]. Johann [Knu93]. John [Auz98, Gau96, Loz97, Sha96]. Johnson [Coo98]. Jose [Ste99a]. Joule [EL95]. Juan [Fal96]. jumps [MB96, Osw99]. justification [Zha98b].

K. [Coo98, Mat97, Wah97a]. Kalachev [Shi96]. Kátai [Pih93]. Kato [IK98]. keep [LW91]. Keller [Kel95]. Kelley [Elm97]. Kernel [KX91, AS91b, CTW92]. kernels [BG99b, Pal96]. Kevin [Sha96]. Khintchine [BBC97]. Kiev [Köl90]. Kind [Yan90, AS91b, Car96, Car97a, KX94, KB97, Pal96]. kinetic [KP94]. Kirkman [CMM92a, MPR95, CMM92b]. knots [Buh95, CMV92, GH94, KS93a, KS93b]. Known [PSZ90, IJ96, IJ99]. Kolmogorov [BTT94b, BTT94a]. Korteweg [Hua91, KM90]. Kowalski [Qua97]. Kraft [Ano95u]. Krasikov [AL95b]. Kreiss [LS91a]. Krishna [Nie96b]. Krízek [Bre96a]. Kronrod [Ehr94, Lau97, Not92, Peh90b]. Krszysztof [Qua97]. Krylov [Sch92]. Kumar [SS96b]. Kuo [Nie96b]. Kuo-Yu [Nie96b]. Kuperman [Gil98]. Kutta [ASS93, BC98, CHS90, Coo91, CL99, GP99, GS98, JZ92, Kvæ90, Li94b, LO93, LO95, MO93, OR92, OZ91, PP96, Ren90, TV93, vSC92, VZ95, WT90]. Kuznetsov [TT95].

L [Bar96, Bra96a, Mat97, Mea96, Nie96c, Noc96, Res97, Sha96, Wri96, Wri97]. L. [Bra96b, Bre96c, Kar97, Tem97, Wah98a]. Lagarias [DR96]. Lagrange [Al 94, BC97, CMO90, CMV92, EM97a, GL96, SX95, Swa93a]. Lagrangian [CGT97b, CGT97a, LS95]. Laguerre [DJLZ97, Str92a]. Lanczos [Isa96c, Bun97, Tei98, Bai94a, Huc95, Ye91, Ye94]. Landau [Du98]. language [Kea98]. LAPACK [Cai96]. Laplace [Tem98, AC90, BF93b, CV95b, SH95, Str92a]. Laplace-modified [BF93b]. Laplacian [BL93]. Large [Che91, Cot91, Dub96, Phe90b, You98, CF94, For97, HS97b, LM94a, LN96, MB96, Mas94, Mor96, MST99, NY97, Phe90a, SH98, Wan90]. large-integer [CF94]. large-scale [NY97]. Large-time [Che91, Cot91]. larger [Zha93b, Zha94]. Largest [IJ96, IJ99, PSZ90, Gra95, Ian99]. Larry [Bar96]. Lars [Arn97b]. Lasser [Ano95b]. last [Cop90, DS95]. Lattice [DS91, JS92, Kea97, LS93b, NS90, SL90, BW96, Cha92, CL94, Hob98, Joe93, L’E99a, Lan96, LJ96b, Nie92a, Nie92b, SW90, Sut98]. lattices [CL96, PP93, SH98]. Lau [Kok96]. Laura [Che96d]. Laurent [Bar96, Kea98]. Lawrence [FS95]. Laws [JSH90, Cha97, Cha99, CDT93, CL97a, Che91, CHS90, CCF94, CG96b, CG97, CL91c, CFN95a, CFN95b, Fan98, HR93, LW96b, LW97, TT94, ThT97, vBv92]. Lax [BL94, Che91]. layer [BS94b, HS94, SS94]. layers [SS96a]. Layne [Gau96]. Least [Mar92, BG94, BLP97, BS98, Cha94b, Col97, ELS93, Fas97, Lev98, Mar90, PJV94, RAG91, ZF93]. Least-Change [Mar92]. least-squares [BLP97, Fas97, Lev98]. Lebedev [dB98]. Lectures [Ste97b, Kin99a]. Legendre [BD92b, CK99]. Lehman [Leh92b]. Lehmer [CSV94, Dar91, DR96, LL93, Ray94, Vou95]. Lehmers [Apo93]. Lemarié [Lai94]. Lemoine [Pih93]. length [Ada95, ESA94]. Leonid [Shi96]. Lepingle [Ano95h]. less [ESA94]. Letac [Smy91]. level [Bre96b]. Levels [Leh92a, Leh92b]. Levesley [Kin97]. LFSR [L’E99b]. Liapunov [ABZ95]. library [Kok96]. Lidl [Sch96b]. Light [Kin97]. like [CLR97, Ekl98, Yin94]. limited [NY97]. limiting [OW94]. Lin [Nie96b]. line [EL91, Wei95]. Linear [BD92a, Ber94, BS92b, HW92, KT90, Peh90a, Wri97, ZH90, Bai94c, BBN99, Ben98b, Ber98, Bre94a, Bre94b, Cop99, Cop94, CL94, CL96, EDK97, EG90, EG91, Elm97, Fal91, Gur90, HT94, HT95, Hig94, HS95a, KP92, Kal95, L’E99a, Len91a, LS93a, LO95, Nii92, PJV94, RR94, RI96, Rum90, Sil99b, SV90, Ste99c, Wal92, Wat98]. linear-quadratic [Ben98b]. linearization [Sho98]. linearized [Krö91, Sun95]. lines [BF92]. Links [DH91]. Lions [Bra96a, Bra96b, Res97]. Liouville [Abr95, Bou99, Cha94a, RS92]. Lipschitz [Bra95]. Lipsman [Sha96]. little [Sil97]. Littman [Isa96a]. Lloyd [Ste99c]. Local [ELV91, Mar90, Mar92, WBL94, CD96, CH96, CHS90, Gir95, Leh90, LJ96a, Sho98, TV93, Tra95]. localized [CLM94]. locally [Far94, KS92, Ray94, TB97]. Locking [AB97, SS94, SBS95]. Locking-free [AB97]. LOD [Hun92]. logarithm [FR94, Rüc99]. logarithmic [AS91b, Cro93, Sid95, Wal91]. logarithmic-kernel [AS91b]. logarithmically [Van94]. logarithms [AD93, MST99, Sem98b, Sem98a]. Lommel [FS92]. Long [TW94, HHM94a, HHM94b]. Long-time [TW94, HHM94a, HHM94b]. longest [RI96]. longest-side [RI96]. Lorenz [MM98]. Low [Huc95, AV99, Dre98, TF93]. low-frequency [AV99]. Low-rank [Huc95]. Lower [EHN92, Nie90, BCT91, EH94a]. Loyce [Wri97]. Lucas [GC93, Arn97a, DK99, GP91, Mus93, Vou95]. Luigia [Che96d]. Luís [Fal96]. lumped [EG99].

M [Che96c, Ell96, Fal96, Fik98, Kin97, Köl95, Köl97e, Lam96, Lam97, Loz97, Nie96a, Rud98, Wim98]. M. [Bre96a, Kin97, Köl97d]. MA [Fik98, Lam96, Lam97]. Maass [HA93b, HA93a]. MAC [NW96]. machine [Loz97]. magnetic [Zhu92]. magnetohydrodynamics [GMP91]. Magnus [Köl97c, Köl97h]. Mahler [Fla96, Mos98, RS95, Td91]. maintaining [Kli96]. Mangasarian [Noc96]. manifolds [CHW99, GANT99]. Manin [Ura96b, Ura96a]. manual [Nie96c]. MAPLE [Ste96a, Mea96, Wri96]. mapping [LS98]. maps [Pil90]. Marchuk [Ano95y]. Marek [Qua97]. Marián [Hol97]. Marichev [Köl97a, Köl97b, Köl97g, Köl97f]. Markov [MS93, ZH90]. Marletta [Kin97]. Martha [Mea96, Wri96]. Martin [SW95b, Ste98b]. Mascagni [Gau96]. mass [EG99]. matching [LS95]. Math [Ano92, BO94, BW96, CH90a, FW94, GC93, JMS97, LLMP95, MPR95, Mor90, Tho94, Woz97b]. Mathematica [Sha96]. Mathematical [Che96a, Shu96, Bre96c, Leh93, Wim98]. Mathematics [Isa94, Len97, Res97, Gau94, Isa96a, Laz94, Wah98b, War96, Zha98a, Gau94, Len97, dB98]. Mathon [CMM92a]. MATLAB [Ste96a]. Matrices [Pan90, BS90a, CMP98, CY92, GKO95, Grö99, Hig98, Mac94a, Oli98, Til97, YC93]. Matrix [Sch92, BL96, Buf90, EM95, GL99, Hig96, LLS96, LS91a, Li94a, NW94, Nie95, Sou94, Swa93b]. matrix-valued [NW94]. matters [Pin99]. maximal [Gra95, Nic99]. Maximally [L’E96, L’E99b]. Maximum [KP94, Oli99, Žen95, SW95a]. Maximum-angle [Žen95]. maximum-norm [SW95a]. Maxwell [AV99, BB99b, NW98]. Mayers [Wah97a]. McGraw [Köl97h]. McGraw-Hill [Köl97h]. McKean [BT97]. McKenney [Cai96]. Mean [It91, Ad97, Coh97a]. measure [Fla96, Mos98, RS95]. media [CL91a]. Mediants [Bos90]. medium [WL99]. Méhauté [Bar96]. Meissel [DR96]. Melas [vSv92]. memory [LST96, NY97, TW94]. merit [AMN90]. Mersenne [CW91, KM91]. meshes [BCMR95, Car96, FMOS98, LJ96a, Neu94, Noc95, Sül92, TB97, Zha90a, Zha90b, Zha95b]. Meshless [Wen99]. metallic [Lev92]. Method [BBG91, CPS92, DGPR92, DS91, KX91, NPV91b, Ric90, Ric92, Yan90, Abr95, AY94, Ale91, AL95a, AN92, Arn94, ABZ95, AM93b, AL91, BS92a, By99, BWW97, BTT94a, BTT94b, BS99, BT97, BBP96, BK94, BS98, Bre90, Bre94a, Bre94b, BPV99, BR94, CHL99, CDL97, Car97b, CH98, Cha94b, CGJ95a, CGJ95b, Che94b, CQS98, Cho97, CHS90, CT92a, CT92b, CT94, CL99, CHT96, CV95b, DR96, DM99, Dol99, Eck98, EDK97, EL92a, EL92b, Emm96, FSSC90, FP96, GANT99, Guo95, GS97, GL98b, HS94, Han90, HT94, HST98, HR93, HWC99, HGK98, HS95b, Hun98a, Iya95, KM98, Kha99, Lab98, LTW98, Léc91, Lev92, LP95, Liu94, Lu98, Mac94a, MP96, McK99, MP95, Mit90, Mor96, MSS97, Nie95, NPV91a, NRW95, OS91, Sch98a, SW92, SS96a]. method [Shi96, Sho98, SV90, Ste99a, Ste96b, SS96c, Sur97, Tei98, Tem91, Tra95, Vv96, Wan99, Woź94, Woz97b, XC92, XZ96, Yin94, Yin98, ZZ97, ZF93, Zho97, DOSS91, tR94, Bun97, Nie96a]. Methods [BDM90, GM92c, Hou92, JSH90, LM90, Mar92, Mas94, OS92, SZ92, Sur90, WT90, ASS93, Ach99, ACM98, AC95, AB97, AC90, Auz98, BLS94, BF93b, BG94, BPWX91b, Bra94, BS92b, Bre96a, Bre96b, Bre99a, Bre99b, Bru94, CS95, CMP98, CEL96, CO98, CL91a, CCF94, CG96b, CG97, Col97, Coo91, CRV92, DVZ94, Del90, DMT93, DL92, DS93, EL91, EL93, EG90, EG91, Elm97, EJ93, Fal91, FO94, Fal96, FMOS98, Fas97, Gal92, Gol91a, GP99, HDJ94, Hal96, HS95a, Hol96, Hon99, Hun92, JZ92, JS94, Kea97, KS92, Kvæ90, LT94a, Len91a, LR90, LM94b, Li94b, LT94b, LG93a, LG93b, LO93, Mak94, MS93, Mar90, Mel97, Mel98, Mon94a, MS98, Mor99, Mus93, OR92, OZ91, Ren90, SW95a, SW96, Sch91, Ser99, She94, She96, Shy94, Sil99b]. methods [Ste90, Sul93, Sül92, TT94, ThT97, Tem94a, TG94, TV93, VZ95, Wag96b, Wan91, Wen99, Zha90a, Zha90b, Zha95b, Wah98a, Arn97b, Bre96c, Elm96, Shu97, Tod98, Wri97]. Micchelli [Che96a, War96]. Michael [Gil98, Gau96]. Michel [Kea98]. microfiche [BR98, BN92, DO95, DH91, DJ91, DH94, Sou94, YR92]. microstructure [CP97, LL98]. microstructures [KLL96]. midpoint [Hun98b]. Miller [Arn95, DR96]. Million [BCS92, EM92b, BCEM93]. Milovanovic [Che96c]. Mindlin [BS98, DL92, DHNL^{+}99, Gas93, SS94]. Minimal [CS96b, GZ93, Pih93, SW96]. minimum [Fuj93, HH99b, PP93]. Minus [Sch98b, BLP94, BLP97]. miscible [CL91a]. Mitrinovic [Che96c]. Mixed [CL91a, Sur90, AC95, Arn94, Bre94a, Bre94b, Car97b, CEL96, CV99, CMW95, DL92, EG99, FO94, KS92, Lev92, MP95, RVW96, Wan91, WH99]. Mladen [Chu96]. mobility [BB99a]. Model [Ber90, ZH90, Che94b, Du98, DL92, EL95, Gas93, SS91, SS94]. modeling [Che96a]. Modelled [LT91b, LT91a]. modelling [Kea98]. models [Loz97, SBS95]. modes [DHNL^{+}99]. modification [GL98b, Huc95, Yan95]. Modified [LF93, Ale91, Gal92, Mac93, BF93b]. Modular [BBG90, CLR97, Gol92, GM92a, Sko92, BBG93, Coh93, Coh95, CM96, Cre95, Gol91b, HS92, His94, San92, Val91, Was91, YZ97]. Modular-Part [BBG90]. Modules [Hay91, DM98, DH94, LS94, Sch93c, Sch93b]. Moduli [CD92, ADS98, CL96]. modulo [Per92, Wag94, Wag96c]. Modulus [Fis90, EHN92, EHI94]. molecule [NTS98]. Molenaar [Hol96]. Mollin [Coh97b]. moment [BS90b]. Monier [Arn97a]. mono [MO93]. mono-implicit [MO93]. monotone [CG97, DCNZ96, LL99, Nii92, TT95, XZ99]. monotonic [Dem90, Dem95]. Monte [Nie96a, BGLT95, HMN93, HM93, HM97, Léc91, Léc96, TW92]. Montefusco [Che96d]. Moody [Isa96c]. Morgan [Leh93]. Morton [Wah97a]. most [Ros95a, Ros95b, WW97]. moving [Lev98]. MR [Ano92, BO94, BW96, CH90a, Fik98, FW94, GC93, Köl95, Köl97e, Köl97g, Köl97f, Köl97h, Lam96, Lam97, LLMP95, MPR95, Tho94, JMS97, Woz97b]. Muller [Mai91]. Mullineux [Wah98b].

Multiamicable [CGH95]. Multiconstrained [EST90]. multiderivative [LR90]. multidimensional [CDT93, CHS90, CCF94, Nie92a, Nie92b, SW90]. multigrades [Smy91]. Multigrid [BP92, Bre99b, CO98, Gol91a, Hol96, SZ92, BPX91b, BPWX91a, BLP94, Bra94, Bra95, Bre90, Bre94a, Bre94b, Bre99a, LG93a, LG93b, Zha90a, Zha90b, Zha95b, ZZ97]. Multilevel [BPX90, HST98, BPX91a, BP93, CO98, Ste96b, Vas92]. Multiple [NS90, Cra98, JRZ98, Kea97, Nie95, SW97, Smi96, SS96c]. multiple-precision [Smi96]. multiple-recursive [Nie95]. multiplication [vW99]. Multiplicative [Fis90, BW96, Pow97]. Multiplicities [May92a, May92b]. multiplicity [KSS98]. multipliers [Swa93a]. multiply [CL97b, LF93]. multiply-add [LF93]. multiply-with-carry [CL97b]. multipole [Lab98]. Multiquadric [Buh95, Mak94]. Multiresolution [Tem94a]. multiscale [HWC99]. multisplitting [CMP98]. multistep [ACM98, AMN90, HS95a, Len91a, LR90, Li94b, Pal95]. Multivariate [CG90a, MN90, AS91a, BC97, Jia98, LG97, SX95, Sch94, Wan90, Xu94c, de 96]. MUSCL [BBP96].

N [Ste99c]. N. [Bai96]. Nakamula [Str92b].

narrow [Al 94, Dur99]. natural [Hob98]. nature [GS93]. Naukova [Köl90]. Navier [Kar97, Lar97, Cot91, GANT99, GHS91, NW96, She96, Yin90, ZF93]. Nazareth [Wri97]. necessary [Sul93]. Negative [BL95, BS98]. negative-norm [BS98]. Neittaanmäki [Bre96a]. Nematic [Ber90]. Nemirovskii [Tod96]. Nesterov [Tod96]. nets [HH99a, Léc96]. networks [CLM94]. Neumann [BH94b, GHS91, Hon99]. Neural [CLM94]. Newton [Ale91, CQS98, GL98b, Mar90, Mor99, NY97, Wan99]. Nicely [Bre97]. Nicholas [Van97]. Nickolas [Mey96]. Nico [Wim98]. Nicolson [Hun92]. Niebuhr [Kel95]. Niederreiter [Göt94, Nie92b, Roe99, Sch96b]. ninth [DO95, LLMP93, LLMP95]. ninth-degree [DO95]. no [Ano92, BO94, CH90a, FW94, GC93, JMS97, LLMP95, MPR95, Mor90, Oli92, Sil97, Wal99, Wal21, Woz97b]. node [HMN93]. nodes [Bai94c, Ben98a, GR95, PT94]. noisy [KP92].

Nomials [KM91]. Non [BG99a, Sel99, BPV98, CMP98, CH98, DDP95, EG90, FMOS98, LS95, Mor96, Zha90b, Cla96]. non-conforming [DDP95]. non-existence [FMOS98]. Non-Galois [Cla96]. non-matching [LS95]. non-overlapping [BPV98]. Non-primitive [Sel99]. non-quasi-uniform [Zha90b]. Non-reflecting [BG99a]. non-self-adjoint [EG90]. non-stationary [CMP98]. non-stiff [CH98]. non-symmetric [Mor96]. Nonconforming [Fal91, LL98, AC95, Bre90, Bre94a, Bre94b, Bre96b, Bre99a, CO98, CL90, KLL96, Osw96, Wal92]. nonconservative [HL94]. nonconvex [CL91b, LW96b]. nondominant [HT95]. nonequally [Buh95]. nonequidistant [BCMR95]. Nonexistence [SK99]. nonhomogeneous [BL95, BK94, TT94]. noninherited [BPX91b]. Noninterpolatory [Rab90a]. Nonisomorphic [YR92]. Nonlinear [GM92c, LR90, LW96b, Noc96, WT90, ACM98, APS93, AC90, By99, BGJ99, BPV99, CL90, CJT97, CG93, DMT93, EL91, EHL97, Elm97, EST90, KM98, KWY96a, KWY96b, Lar97, LP92, Mak93, MP95, Nes96, NY97, Rob94, Rum90, Rum96, Tem91, Ver94, Ver98, Wri97]. Nonnegative [Buf90, PJV94]. Nonnested [SZ92, BPX91b, Zha90a, Zha90b, Zha95b]. nonprime [Sem98b]. nonquadratic [Lou95]. nonresidues [BS96c, Per92]. nonselfadjoint [EG91]. Nonsmooth [LST96, DS93, EL92a, EL92b, MS98, SZ90]. nonstandard [Dav96]. Nonsymmetric [LZ92, Bai94a, BS90a, BLP93, XC92, Ye91, Ye94]. nontrivial [CMS92b, CMS92a, Kap95, iN97]. nonuniform [GASS91]. norm [Acc96, BL95, BS98, Cla96, FJP97, SW95a]. norm-Euclidean [Cla96]. Normal [FvS99, Coh96, Dre94, GV95, JKN91, MM94b, MM94c, Nie94]. normals [OW94]. Note [Arn91, Dar91, Hun98a, Noe96, Tsu90]. Notes [LSK91]. Nu [BG92b, BG92c]. Nu-Configurations [BG92b, BG92c]. Number [Del92a, Dia91a, Dia91b, Ede92, Fis90, MW92b, N.93, Wah97b, Acc96, AK99, Bre99c, BG99b, BE99, CJT97, CL96, CL97b, CDNY95, CS99, Deá90, DO95, Dui96a, FJP97, Fuj93, Gaá96, Ge94, HC98, HS97b, Ian99, Ish95, Iya95, JM92, Jak98, Jha95, KWY96a, KWY96b, Kid99, LT91c, LT94a, LLMP93, LLMP95, Lou95, Nie96b, NS98, Odl94, Per92, Pom94, SH98, SW92, Sch93a, Sch96a, SPD94, Sel99, Sho99, Ste98a, Ste94b, Str92b, Sut98, Tak94, Wag96a, Wag98, WC93, Yam94, Živ99b]. Number-theoretic [Wah97b, LT94a, SW92]. Numbers [BL90, BSW90, Coh97a, Fla91, Jae90, Jus90, Rob93, Spi90, AMN90, Arn95, BN97, BR98, BR05, BR11, Boy96a, Boy96b, Boy96c, BCT91, CD97, CGH95, Coh98, Dre98, DK95, EH91, EHN92, EH93, EHI94, EH94b, EH94a, EHE96, EHL97, EHN97, Gec93, Gir93a, Gir93b, Gos95, HGK98, JKN91, JMS95, JMS97, LN96, Lou92, Lou97, Mon94b, Nie90, Pin93a, RB94, SSW99, Sri98, Wag96c, Wag96b, Zag93]. Numerica [Kea98, Tho97, Fal98]. Numerical [Ach99, AN92, Auz98, BGJ99, CIZ97, Car98, CP97, Chi95, Col97, CFN95a, CFN95b, DD93, Isa96b, Kar97, KSS98, LTW98, LT91a, LT91b, LM99, LS98, LO94, Loz98c, Mat97, McL94, Rab90b, Rum93a, Rum93b, Sei99, Ste99b, Ste99c, Tem94b, Wah97a, WT90, AH99, Ad97, ABZ95, AS91b, BS92a, BS94b, BF93a, BD92b, BBP96, Bra96a, CHL99, Cha97, CG96b, CG97, CRV92, CV95b, DDD91, DVZ94, EB99, Eck98, FN90, Fas97, GT94, Gol94a, Gol91a, GT93, Hab93, Hig94, Kin99a, Kin99b, Kok96, LT94a, LSW94, LS91b, LC94, Mak94, MST96, Mon94a, NW95, Nie92a, Nie92b, Ser97a, Shu96, SV90, Ste97a, TW94, Van97, WBL94, Wat98]. numerics [DH99a, Kin97]. Numerik [Fro97]. NURB [Che96b].

O [Köl97g, Köl97f]. O. [Köl97a, Köl97b].

Oberhettinger [Köl97c, Köl97h]. object [BS94b]. obtaining [Abr95, Cha94a]. occurrences [Nic99]. occurring [Gau91b, Mac97]. ocean [Gil98]. octahedra [Cha92]. octic [CDO99]. odd [BCT91, HC98, Ian99, Sao98]. ODEs [Zha98a]. Odlyzko [DR96]. off [NHA97]. Olver [HT94]. Olvi [Noc96]. One [BCS92, EdT90, EM92b, GM92c, Hay91, Bai94c, BLS94, BLP94, BLP97, Buh95, CRV92, CLT94, DH94, Gla92, Mad90, Yam94, Car91]. One-Dimensional [GM92c, BLS94, Mad90]. one-step [CRV92]. ones [Ziv95]. Only [EdT90, DH90, SO91]. onto [HJ95]. open [AS91b]. Operational [EL93, EL91]. operator [Arn94, DS99, EDK97, Gir95, HS94, LS98, Tra95, XZ97]. Operators [Wad90, BLP94, CLT94, Lin94, MS93, MW94a]. Optimal [CL91b, Dol99, OW97, Ser97b, Zha90a, Zha90b, Zha95b, AV99, Bai94c, Bra93, CHL99, CH96, FG96, GV95, GHS91, HR94, KM90, Pao97, Peñ97, TB97, Woź94, Woz97b]. Optimal-order [CL91b, Zha90a, Zha90b, Zha95b, CH96]. optimality [NRW95]. optimally [CD96]. Optimization [Dav96, BGLT95, CGT97a, CGT97b, Gol94b, NY97, Ben98b, Kea98]. Orbits [CL96, Dre98]. Order [MO93, OZ91, Phe90b, Sur90, Wah98a, AC95, BCM92, BS99, BLP94, BLP97, CD98, CL93, CEL96, CH96, CMS92b, CMS92a, CMM92b, CL91b, Dui94, Dui96b, Eck98, FSSC90, Fly98, GS91, Gra95, Gur90, HJ95, HT94, HT95, KM90, KP94, LJ96b, MPR95, Mel97, Mel98, MES98, MP95, MS98, Mül97, Noe96, PP96, Phe90a, Pow97, RS96a, San92, She93, She96, SA94, Smy91, SV90, Sun95, SW98, VZ95, Wal92, Yan95, Zha90a, Zha90b, Zha95b]. Orders [Den90, BB94a, Gaá96, GV95, GvP98, KL97, OR92]. ordinary [Ale91, BS99, CHT96, KL97, Kin97, Ste94a]. Oriented [DH91]. origin [MM95]. O’Rourke [Ano95c]. Orthogonal [BF93b, Rob94, FG91b, FP97, Gal92, Peh90a, Xu94c]. orthogonalization [Yin94]. Ortiz [Bun97]. Osborn [Sha96]. oscillating [HWC99]. Ostrouchov [Cai96]. overlapping [BPV98, LT99]. Owe [Elm96]. owner [Nie96c].

P [Köl97e, Köl97a, Köl97b, Köl97g, Köl97f, Loz98c, Mea96, War96, Wri96]. P. [Bra96a, Bre96a, Coo98, Köl97d]. packing [Cha92]. pair [For97]. Pairs [Cos91, Ct98, CSV94, MM93b, PP96, tR94]. Pan [Hig96]. pantograph [BI93]. Parabolic [LNS93, ACM98, BF93b, CTW92, CL90, DD92, FK93, Gol94a, GP99, GS97, LTW98, LP92, LO93, LO95, TW94, Ver98, Zha93a]. Parallel [BPX90, BPX91a, Gau96, CMP98, Hol97, Hou96a, Hou96b, Kal95, Loz97, Loz98c, Oli98, Swa93b, vSC92].

Parameter [DGPR92, Auc96]. parameterized [Ray94]. parameters [HMN93]. parametric [GP97, MS97b]. parametrization [Cre95]. Paris [Ste99a]. Parity [BBG90, JMS95, JMS97, Ste94b]. Part [BBG90, HDJ94, NPV91a, NPV91b, SW95a, CG96b, CG97, MM98, Sch98a]. Partial [Fis90, By99, Bra96b, DCNZ96, GKO95, GS95, Kin97, Kui95, Lu98, OR92, Ren90, Tem94a, Wah97a, Zha93a]. particle [AN92, BTT94a, BTT94b, BT97, Cot91]. Partition [BBG90, GS90, RI96]. Partitioned [ASS93]. Partitions [Dub92]. parts [Hig98, Ols95a, Ols95b, Ols95c]. Pascal [Kea98]. pass [Arn95]. patch [Zha96]. patches [DMS92]. paths [LL99]. PCG [Ser99]. PDEs [LG93b, IK98, LG93a, Zha98a]. Pell [Wal21, Ang96, Ked98, Wal99]. penalty [FG91c, RVW96]. penalty-type [FG91c]. pencils [BL96, Li94a]. penetrable [BS94b]. perfect [Ber98, BCT91, Dui94, Dui96b, HC98, Ian99]. perfectly [NRS95]. performance [Hou96b]. period [Ada95, Wag94, Wag96c]. Periodic [SA94, Bou99, DS93, OW97, PT92]. Periods [CW94, Bre94c, FvS99, GvP98, GZ93]. permutation [But94]. Perrin [Arn91]. perturbation [Hig94, LT94b, Shi96]. perturbations [BL96, Buf90, Li94a, LW97]. perturbed [BS99, OS91, Pao97, SO91, SS96c]. Perturbing [MPV98]. Peter [Isa96b, Loz97]. Petrov [GM92c, ZF93]. Petter [Gau96]. Petzold [Mat97]. Pezzo [Ura96b, Ura96a]. Phase [NPV91b, NPV91a]. phenomenon [Gel97, GS95]. Philip [War98]. Philippe [Fal96]. philosophy [Laz94]. physicists [Köl97d]. physics [Wim98]. picturesque [Apo93]. piecewise [Bai94c, BPV99, Dem95, GW96, GS95, HJ95, LSEG91, Nii92, ThT97, TB97, YC93]. Pierre [Bar96]. Pierre-Jean [Bar96]. Pisot [Boy96a]. pivoting [GKO95, PJV94]. Plana [DLD90]. planar [BS92b, Bre94a, Bre94b, Dui96a]. plane [LH98, MST96, Tei90, vKP95]. planes [CD98, EH91]. plastic [NRS95]. plate [DL92, DHNL^{+}99, Gas93, Gau91b, GH97, Mac97, SS94, SBS95]. plateau [DH99b, DH99a, Tsu90]. plates [BS98]. Plemmons [Isa96c]. Plis [Qua96a]. Ploufee [Bai96]. Pohst [Ano95x]. Poincaré [XZ97]. point [Tod96, BCMR95, CHL99, FMOS98, iN97, PJV94, RW93, SO91, Tod98, Via91]. Points [DS91, BBN99, BGJ99, BB93, Dem96, Fla96, MVZ93, RR94, Sem98a, Sil99a, Sut98, Zha98b]. Pointwise [CMV92, Noc95, Sch98a, DN90]. Poisson [Arn94, Yin94, Zha98b]. Poisson-like [Yin94]. Polar [Bai94b, Hal96]. Polygonal [Yan90]. polyharmonic [Dim96]. polyhedra [DH90]. polyhedral [Dui96a, HS95b]. polylogarithmic [BBP97]. Polynomial [Bos98, Dar91, EM95, Hig96, KS93a, KS93b, MM94a, Vas91, dR92, Tod96, BPV99, DM98, FT91, GW96, HS95b, LW91, LW93, Mül97, PST98, Shp93, ST91, Xu92, de 96]. Polynomials [CG90a, GHO92a, HM92a, Mos98, Sho90, BH94a, BH97, Boy97, Bre93, Bun97, Che96c, CM95, EM97a, Fas97, FS92, FG91b, FP97, FW90a, Gal92, GHO92b, GN96, GZ93, HS97a, HM98, HJ95, Han96, HM92b, He94, HH99b, KS98, Kui96, LG97, Mal97, McK94, MM94b, MM94c, MPV98, Nie94, NG95, Peh90a, Peh90b, RAG91, RS95, Roe99, TF93, TB97, Wan90, Wan97, WC93, Xu94a, Xu94c, Živ94a, Živ94b]. Polytope [Law91]. polytopes [LH98]. Pomerance [GC93]. Popov [Köl90]. porous [CL91a, WL99]. Porter [Gil98]. posed [ACEO98, Car94]. Positive [Hun91, Leh92a, MN90, Buf90, CMP98, CY92, Coh97a, Fla96, Hig98, Kap95, Leh92b, Liu94, LST96, NW94, Peh90a, RS96b, Ser97b, Xu94b]. positive-type [LST96]. Positivity [HST93]. Posse [För94]. possible [BS99]. posteriori [CS95, Car96, Car97a, Car97b, DR98b, Nii92, Noc95, Ver94, Ver98, WH99]. postprocessing [GANT99]. potential [BS94b, Car98]. Power [BBBP98, GP97, Kei92, BF92, CG96a, EHN92, EHI94, Gaá96, Lev98, Lou95, LJ96b, Pin99, WG92]. powerful [Coh98]. powers [Ber98, Ekl96, Ekl98, Knu93, Pih93]. Pozo [Ano95q]. practical [Bea97, Che96b, Shu97, SW93]. practice [Rud98]. practitioner [Kli96]. Prathap [Ano95l]. Praveen [SS96b]. Precise [Bri91, SH95]. precision [Smi96, Ste97b]. preconditioned [XC92]. preconditioner [AV99, Osw99, Ste96b, ST96]. Preconditioners [BPX90, Osw96, BPX91a, Bre96b, CY92, Mar93, RW93, RVW96, Ser97b, Vas92, YC93].

Preconditioning [AFW97, XZ97]. prescribed [Pao97, PP93]. Presentations [San92]. preserving [FT91, Gla92, Len91a]. Press [Fik98, Köl95, Lam96, Lam97, Loz98c]. prewavelets [Buh95]. Primal [Tod98]. Primal-dual [Tod98]. primality [Arn95, AM93a, Bac90, Bos93a, Bos93b, Cal95]. Prime [CW91, Del92a, EdT90, For99, Wag96b, Acc96, BN97, Bur96, CD99, DLP93, Dus99, ELS93, FG91a, FW90a, Gra95, HC98, HS97b, Ian99, JKN91, Jak98, LJ96b, Nic99, Per92, Pow97, Sho99, SW98, Wag94, Wag96c]. prime-power [LJ96b]. primer [Nie96a]. Primes [BCS92, EM92b, Mor89, PSZ90, RR96, BS94a, BS96a, Bre97, BCEM93, CM94b, CDP97, Dub93, Dub96, DN97, DK99, EM91, For97, IJ96, IJ99, Kel95, KN95, LM94a, Mor90, Odl93, PMT95, You98, Živ99b]. Primitive [HM92a, HM92b, KM91, MM94b, MM94c, Sho92, Vou95, Bac97, Han96, Sch93a, Sel99, Zha93c, Živ94a, Živ94b]. Principal [AM92, Rab90a, Bia90, Die94, Gol91b, HT91, Mon94a, Rab90b, Rab91]. principle [BV92, KP94]. principles [vBv92]. priori [CG96b, CG97, WL99]. probabilities [BP96]. probability [MS93]. probable [Bur96, DLP93]. Problem [CPS92, DD90, Hou96b, Rud98, ACEO98, AL95a, AL95b, Bai94a, BE96, Boy97, Bre94a, Bre94b, CP97, Cha94a, Cha94b, CG96a, Cho97, CL91b, Cop99, DDP95, Dem90, DGNP93, DN90, DH99b, EL95, FSSC90, FP95, FK96, GS97, Hal96, Hob98, KS92, Leb94, LP92, Mal97, MP95, OS91, Oli99, Pao97, Ray94, SS91, Ste90, SO91, SS96c, Tsu90, Yin90, Zha93c, DH99a]. Problems [BG92a, BN92, Dör92a, LM90, LZ92, NPV91b, Sch92, Ste96a, Sur90, ASS93, Abr95, AY94, AD91, ACM98, AC95, BO94, Bac90, BCM92, BF93b, BS99, Bou99, BLP93, BK94, Bra96b, BJT97, BCMR95, Car94, CDL97, CFS95, CEL96, Che94c, Che96c, Che97, CL90, CV99, DD92, DVZ94, Dör92b, DR98b, DS93, EJ93, ELV91, EST90, FO94, FMOS98, Gau91b, GP99, GT91, GHS91, Han90, Hol97, Hon99, HWC99, Hun92, Hun98b, KP92, Lab98, LT99, LS91b, Li99, LT94b, Mac97, MM93a, MB96, Mas94, Mat97, Mor96, MSS97, NW95, NPV91a, Noc95, PJV94, RS96a, RS92, RW93, RVW96, SS96a, She94, Ste94a, Sur97, Swa93a, Vv96, Ver94, Ver98, Wu95, XC92, Yan95, Yin94, Zhu92, Col97, Isa96a, Shi96]. procedures [Bac97, DD92]. Proceedings [Isa96c, Wah98a]. process [Sid95]. processes [FN90, Sho98, Che97]. processing [CG93, Gau96, Nie96b]. Product [Pal96, BBG98, BPWX91b, BLP97, BGN93, CMM90, HP99, Mal97, MM94a, MM95]. Products [BBG90, GN92, Mon94b, Fik98, Köl95, Lam96, Lam97, Wah97b]. programming [Tod96, Noc96]. progression [DN97, Pin93b, PMT95]. progressions [BG96, Dil92a, Dil92b, RR96]. project [LL93]. Projection [BX91, Coc91, CHS90, Gir95, She96, SL90, Xu91]. projections [HJ95, Ols95a, Ols95b, Ols95c]. projective [Che96b, LS94, Sch93c, Sch93b]. projectors [Mad90]. Proof [BBG91, VC94, MM98]. propagation [Fen99, Shy94, Shu96]. properties [Coh93, CL97b, CMM90, EHL97, GN92, Jia98, LT91c, Li94b, MTU94]. property [BC97]. Prouhet [Mal97]. provably [Val91]. Proving [vW99, AM93a]. Prudnikov [Köl97g, Köl97f, Köl97a, Köl97b]. Pryce [Ano95a]. pseudo [BW96]. pseudo-random [BW96]. pseudodifferential [BLP94, SV98]. pseudoinverses [Wei97]. pseudopowers [BLSW96]. Pseudoprimes [Arn91, Gur90, Arn97a, GP91, GC93, Jae93]. Pseudorandom [Emm96, Nie95, AMN90, EH91, EHN92, EH93, EHI94, EH94b, EH94a, EHE96, EHL97, EHN97, KWY96a, KWY96b, Mit90, Nie90, TW92]. Pseudospectral [DOSS91, FG91c, Shu97]. pseudosquares [LPW96a, LPW96b]. psi [Alz97a]. PSLQ [FBA99]. Puccio [Che96d]. Pure [Mal97, Bre94a, Bre94b, CG96a]. purification [MS97a]. Pythagorean [FN95].

QM [iHT99]. QM-curves [iHT99]. QR [HP92]. Quadratic [BSW90, Car92, FW90a, Hun91, Leh92a, MW92b, MW94b, Sol92, Ben98b, BMO90, BB94a, BS96b, BPX91b, BG99b, Car91, CH90b, CW94, CS99, DP96, De 95, FK96, Fuk96, Iya95, Jac99, Kid99, Leh92b, Lou92, Mol93, MST99, Oli92, PR99, Per92, Pin93b, Sri98, Tay96]. quadratically [Che94a]. Quadratics [Coh97b]. Quadrature [Ben98a, BN90, Cro93, Lyn94, RS90, VV93a, BS92a, Bia90, Bra93, Die94, EL91, EL93, Ehr94, EG99, FP90, För94, FO93, Gau91a, GL96, GN96, GF97, GM96, GR95, HT91, HST93, Hic98, HH99a, Kui95, Lan96, Lau96, Lau97, LO93, MS95, Not92, Peh90a, Peh90b, Peh93, RS94, SW90, Xu94b]. Quadratures [XZ94, KX94, Sch97]. quadrilateral [Gir95, Sül92]. Quality [LJ96a]. Quartapelle [Kar97]. quartic [Bea97, BFP93, BPv94, CH96, CDO99, JM92, JMS95, JMS97, MPR96, NS98, Pet91, Pv93, SW95b, Was91]. Quasi [Dem96, BGLT95, Car96, CL99, HMN93, HM97, Léc91, Léc96, LO95, Mit90, Pao97, RR94, Ser97b, Zha90a, Zha90b, DJLZ97, NY97]. Quasi-interpolatory [Dem96]. quasi-Laguerre [DJLZ97]. quasi-linear [LO95, RR94]. quasi-Monte [BGLT95, HMN93, HM97, Léc91, Léc96]. quasi-Newton [NY97]. quasi-optimal [Pao97, Ser97b]. quasi-random [Mit90]. quasi-randomized [CL99]. quasi-uniform [Zha90a]. quaternary [SW95b]. quaternion [Mac94a]. quaternion-Jacobi [Mac94a]. Questions [BL90, MT93]. quick [Cha94a]. Quintic [Dia91a, Dia91b, PW98, SPD94]. Quintics [Ano92, Dum91, FN95, GP97]. quintuple [BBG98]. Quintuples [Wag95]. quotient [BS90a]. quotients [ADS98, EM97b].

R [Auz98, BW96, Bre97, Gau96, Loz97, Mat97, Sha96, Tho94]. R. [Bre96a, Ell96, Tem98, Wah97b]. Rabin [Arn95, Arn97a]. radial [Sch99, Wen99]. Radiation [BV92, vBv92]. radicals [HH99b]. Radius [DJ91, Gol94a]. raising [KL97]. Ramanujan [Bre94d, Kei96]. Random [Deá90, Fis90, Bai94b, BW96, CL96, CL97b, KT90, LT91c, LH98, Lu98, Mit90, WC93, N.93]. randomized [CL99, HH99a]. range [Rum96]. Rank [DH94, Hay91, HP92, CS99, Huc95, Lan96, iN97, Pap97, Ros95a, Ros95b, Sil99a, SW90]. rank- [Lan96, SW90]. Rank-One [Hay91, DH94]. Rank-revealing [HP92]. rapid [BBP97]. rapidly [HWC99]. Rassias [Che96c]. Rate [BTT94a, BTT94b, CFS95, DMT93, FS92, LW97, Nes96]. Rates [BG92a]. Rational [Leh97, BBN99, Ber94, CHT96, CK99, GLW99, JKN91, Pal94, Pal95, Sil99a, VV93a]. Rationals [EdT90, van99]. Raviart [WH99]. ray [BBBP98, CDO98]. Rayleigh [BS90a]. reaction [CG93, LSS99, SS96c]. reaction-diffusion [LSS99]. Real [BSW90, Dia91a, Dia91b, MW92b, PR99, BL96, BF92, BS96b, CS99, DST97, Fuj93, FK96, Fuk96, Gaá96, GP97, Hig98, Isa96a, Jak98, Kid99, LW93, Li94a, MW94b, MST99, OW94, Sri98, Str92b, Tak94, Tay96, Wei95]. real-imaginary [BF92]. real-world [Isa96a]. recipes [Loz98c]. reciprocal [Zha93b]. reciprocals [BGN93]. Reciprocation [BBG90]. Recognizing [Ge94]. Reconstruction [Grö92, RS92, Eck93]. reconstructions [Eck95]. record [GG94, Oli99]. record-holders [Oli99]. Recovering [GS95]. recovery [Zha96]. rectangles [BF93b, Dui94, Dui96b, Jep96]. rectangular [She93, Wan91]. Recurrence [Xu94c, Gur90, MT93]. recurrences [Bre94c, LS93a]. recursive [Iya95, Nie95]. reduced [EG90, EG91, Sur97]. reducible [WC93]. Reduction [EdT90, Hay91, Kid99, AM94, CV95a, Mül97]. Reed [Mai91]. refinement [BE99, ELV91, LSEG91, LJ96a, RI96, TV93, WH93]. reflecting [BG99a]. reflection [BH94b, Xu94a]. Regions [Yan90, LS91b]. regular [Lee99]. regularity [BPWX91a, Bre99a, Eck95, NHA97, SW96, SL90]. Regularization [ZH90, NHA97]. regularizations [ACEO98]. regulator [PW98]. Reissner [BS98, DL92, DHNL^{+}99, Gas93, SS94]. Related [BN92, AI91, AS93a, AS93b, Bac90, CL94, GN96, Lor94, Mac93, MT93, Pin99, Wri97]. Relation [Mar92, FBA99]. Relations [Jus90]. Relative [Sho99, DP96, FJP97, Gir93a, Gir93b, Lou97]. relaxation [BS97, Cha99]. relaxed [BTW90]. remainder [Peh93, Sch97]. remark [FR94]. Remarks [Dew98, FO94, BB94c]. representable [BGN93]. Representation [LSW94]. representations [BR90, BC94, Gol91b, OS97, PR99, SL90]. repunit [Dub93]. residue [BS94a, BS96a, MW94b]. residues [Per92]. resolution [CL97a, Gel97, Shy94, Tei90, Yan96b, Yan96a]. Resolvent [CLT94, LS91a]. resonant [Lev92]. respect [Osw99]. restarted [MS97a]. restarting [Mor96]. restricted [BBBP98, Boy97]. result [LSS99, SW97]. Results [BLSW96, BDM90, CRV92, DD93, DH99b, Ekl98, FG91c, LPW96a, LPW96b, Oli99, Rab91]. revealing [HP92]. Review [Köl90, N.93, Gol94b]. Reviews [Ach99, Ano95g, Ano95h, Ano95m, Ano95q, Ano95u, Ano95r, Ano95j, Ano95t, Ano95i, Ano95b, Ano95n, Ano95v, Ano95y, Ano95c, Ano95a, Ano95l, Ano95x, Ano95k, Ano95z, Ano95d, Ano95f, Ano95o, Ano95p, Ano95w, Ano95e, Ano95s, Ano95-27, Ano95-28, Ano95-29, Tod96, Ano97, Arn97b, Auc96, Auz98, Bai96, Bar96, Ben98b, Beu96, Bos98, Bra96a, Bra96b, Bre96a, Bre97, Bre96c, Cai96, Che96a, Che96b, Che96c, Che96d, Che97, Chu96, Coh97b, Col97, Coo98, Dav96, Ell96, Elm96, Elm97, Fal96, Fal98, Fro97, Gau96, Gil98, Hig96, Hol96, Hol97, Hou96a, Hou96b, Isa96a, Isa96b, Isa96c, Kar97, Kea98, Kea97, Kin97, Kin99a, Kin99b, Kli96, Kok96, Lar97, Len97, Loz96, Loz97, Loz98a, Loz98b, Loz98c, Mat97, Mea96, Mey96, Nie96b, Nie96c, Nie96a, Noc96]. Reviews [Qua96a, Qua97, Res97, Rud98, Sch96a, Sch96b, SS96b, Ser97a, Sha96, Shi96, Shu96, Shu97, Sil99b, Ste96a, Ste97a, Ste99a, Ste97b, Ste98b, Ste99b, Ste99c, Tem97, Tem98, Tho97, Tod98, Van97, Wag96b, Wag98, Wah97a, Wah97b, Wah98b, Wah98a, War96, War98, Wat98, Wim98, Wri96, Wri97, Zha98a, dB98]. revisited [BF92]. Riccati [GL98b, LT91a, LT91b]. Richard [Coh97b]. Richardson [Sid95]. Ridgeway [Wah98a]. Ridgway [Bre96c]. Riemann [BS96c, CSV94, Kei92]. Riesel [Wag96b]. rigidity [Phi95a, Phi95b]. Rigorous [Rum90]. Rings [Del92a, Vas91, DM98]. Ritz [Kny97, SW96]. Robbiano [Ano95j]. Robert [Gau96, Isa96c]. robust [BK94, FK93, Ste96b]. robustness [Osw99]. Roger [Lar97]. Rombaldi [Ano95k]. Romine [Ano95q]. Ronald [Sha96]. Roof [BW96]. root [BBBP98, LW96a, WH93]. rootfinding [Mus93]. Roots [Sch92, Sho92, APS93, Bac97, EM95, HH99b, Mon94b, MPV98, Pil90, Pin99, Rob93, Sch98b]. Rosenbrock [Sch91]. Rosenbrock-type [Sch91]. Rotational [YR92]. rough [FO94, FK93]. Rounding [vS96, GH94, Ziv99a]. Roundoff [Cal92, Cal91]. Rubio [Dav96]. Rudolf [Sch96b]. Rudvalis [His94]. Rule [JS92, Bia90, Die94, Joe93]. Rules [NS90, Rab90a, Bra93, HST93, Lan96, Lau97, LS93b, LJ96b, MS95, MM95, Nie92a, Nie92b, SW90, SL90]. Rundell [Che97]. Runge [ASS93, BC98, CHS90, Coo91, CL99, GP99, GS98, JZ92, Kvæ90, Li94b, LO93, LO95, MO93, OR92, OZ91, PP96, Ren90, TV93, vSC92, VZ95, WT90]. Russian [Köl90]. Ryzhik [Köl95, Fik98, Lam96, Lam97].

S [Ano92, Che96c, CMM92a, Fik98, Gau96, Köl90, Köl95, Lam96, Lam97, Mat97, Mey96]. S. [Cai96, Kea97]. Sabot [Hou96b]. saddle [RW93]. Salem [Boy96c, Boy96b]. same [CD99]. Sampling [Grö92, Bou99, Far94, Grö99]. Samuels [War98]. Sanz [Ano95p]. Sanz-Serna [Ano95p]. Sapidis [Mey96]. satisfying [BBP96, MTU94, SZ90]. Sato [iHT99]. Sauer [de 96]. Saul [Loz98c]. scalar [CP97, CDT93, CG96b, CG97, HR93, LW96b, LW97, ThT97]. scale [NY97]. Scaled [Ede92]. scaling [LS93b]. scattered [JSW99, Sch99]. scattering [BS94b, Lab98]. schema [But94]. scheme [Ad97, BL94, DCNZ96, GS91, Gla92, NW96, NW98, Sun95, Via91, XZ99]. Schemes [ELR90, BLP93, Cha97, Cha99, CL93, CL97a, CV99, CG97, CL91c, CFN95a, CFN95b, Fan98, Gol94a, GS98, HL94, JRZ98, KL97, KP94, LL99, LSEG91, MO93, Noe96, She96, TT95, Yan96b, Yan96a]. Schiesser [Zha98a]. Schmidt [ABZ95, Gil98]. Schoenberg [Dem96]. Scholz [AS93b, AS93a]. Schoof [Dew98]. school [Kin99a]. Schreiber [Gau96]. Schrödinger [KM98, Rob94]. Schumaker [Ano95d, Bar96]. Schwarz [Bre96b, HST98]. science [Zha98a]. Scientific [Ste96a, Gau96, Hou96a, Isa96b, Loz96, Loz98c]. scientists [Kok96, Köl97e]. Scott [Bre96c, Wah98a]. Screen [CPS92]. Search [MM91, Bac97, BGLT95, CDP97, Mit90, MM93b, SW90, AL95a]. Searching [Sho92, KTS97]. Secant [Mar92, Mar93, NRW95]. secant-bisection [NRW95]. Second [Köl97e, She93, Sur90, AC95, BBG98, CL93, CEL96, CDNY95, GS91, HT94, HT95, Ian99, KX94, KP94, Loz98c, Mel97, MP95, Pal96, SV90, Sun95, Yan95, She96]. Second-Order [Sur90, She93, AC95, CL93, CEL96, GS91, HT94, HT95, KP94, Mel97, MP95, SV90, Sun95, Yan95, She96]. sector [RS95]. sectorial [HS95a]. Sectors [GHO92a, CM95, GHO92b]. secular [Mel97, Mel98]. segment [GM92b]. Selected [Qua97]. Self [Ber90, EG90]. Self-Assembling [Ber90]. selfadjoint [BO94]. Selfridge [LSW97]. semi [MS93]. semi-iterative [MS93]. semiconductor [Che94b, CT92a, CT92b, CT94, Hol96, Zhu92]. semicontinuity [HS95a]. semidiscrete [Yan96b, Yan96a]. semigroup [Wu95]. Semigroups [LT91b, LT91a, Pal94, Pal95]. semilinear [DCNZ96, FMOS98, GS91, SS96c]. semismoothness [BP96]. sensitivity [Rum90]. separable [ASS93]. sequence [GG94, MT93, Pih93, Zha93b]. Sequences [KKS90, BV99, BG96, CLT93a, CLT93b, CL94, ESA94, EL96, Gur90, Lin98, Mit90, Mus93, Sid95, TF93, Tez94, Van94, Vou95, Zha93c, Zha94, Bai96]. sequential [Hol97]. serendipity [Zha98b]. Serguei [Beu96]. serial [KWY96b]. Series [CSW92, DGPR92, Dub92, Kei92, Köl97a, Köl97b, BS90b, BBBP98, BF92, CLR97, CK95, Eck93, Eck95, Fik98, Fly98, Gau91a, Gau91b, GL98a, Kei96, Köl95, Köl97g, Kui95, Lam96, Lam97, LT94a, LSW94, Lew94, LP95, Lor94, Mac97, Mac94b, McL94, Pin99, Sil99a, Köl97f]. Serna [Ano95p]. Sets [Phe90b, BH93, CM94a, CMX99, Gra98, HMN93, Phe90a]. Seven [DN97, BRZ99, DD93]. seventh [Ekl96]. Several [dR92, CL91c, FG91b, Jae93]. sextic [BMO90, Gaá96, Oli92, Str92b]. shadowing [LSS99]. Shallit [Wag98]. shallow [Ach99]. Shampine [Ano95o]. Shanjie [Loz98b]. Shanks [Wil97]. Shannon [Bou99]. Shape [FT91, LJ94]. Shape-preserving [FT91]. Sharp [Ziv99a]. sharpness [HJ95, TT95]. sheets [CHL99]. shell [Sur97]. shells [AB97]. shift [MS97a]. shift-invert [MS97a]. shifted [BD92a, PT94]. Shikin [Qua96a]. Shock [JSH90]. Shock-Capturing [JSH90]. shocks [LX93, LW96b, LW97]. short [HP99]. side [RI96]. Siegel [Sko92]. sieve [CG90b, Pom94]. Sieved [GS90]. sieves [BB94b]. sieving [Jac99]. sigma [Woz97b, Woź94]. sigma-SOR [Woz97b, Woź94]. signal [Nie96b]. signature [BPv94, Sel99]. Signorini [Han90]. Sikorski [Qua97]. Sima [Ben98b]. Simon [Bai96, Gau96]. Simple [AW97, Den90, Dui94, Dui96b, APS93, CGT97a, CGT97b, CLT93a, CLT93b, Gol91b, His94, Lew94, LH98]. simplex [MS95]. simplices [Hor97]. simulate [Hua91]. simulation [Deá90, HM97, N.93, Hol96]. Simultaneous [Ang96, CMV92, Wal99, Wal21]. Sinc [Yin94, CDL97, Bia90]. Sinc-collocation [Yin94, CDL97]. sine [ST91]. single [HS94]. single-layer [HS94]. singly [Coo91]. Singular [CD92, Bre99b, BPV99, CTW92, CMM90, HP92, KX94, LT94b, MS98, Pal96, Rab90b, Shi96, Yin94, YZ97]. singularities [Bre99b, Eck98, GM92b, MM94a, MM95, Tei90, TG94, WG92]. singularity [Cro93]. singularly [BS99, OS91, Pao97, SO91, SS96c]. size [CV95a]. sizes [L’E99a]. sketch [Kel95]. skew [Buf90]. skew-symmetric [Buf90]. Sloan [Kea97]. Sloane [Bai96]. Slow [Car94]. slowly [Gau91a, Gau91b, Lew94]. Small [Sty93, YR92, BH93, BFP93, Coh97a, DO95, Gra98, HM98, Lee99, Mos98, OS97, Rob93, Sel99, Sma99, Val91]. Smallest [BN92]. Smith [Ano95f, Coh96]. smooth [AC90, AS91b, EL92a, EL92b, ThT97, Tra95]. Smoothers [BP92]. smoothing [CQS98]. Sobol’ [Nie96a]. Sobolev [Ano95w, Bra95]. Sociable [Fla91]. software [Chu96]. solid [BDR98]. solitary [Hua91]. solution [Ale91, AM94, AS91b, CIZ97, CP97, Cha94a, CL91b, DDD91, Elm96, GL99, HHM94a, HHM94b, HT94, Kal95, Kar97, KSS98, LTW98, LM99, Mak94, Mat97, Rob94, SK99, SH95, Sma95, TW94, Wah97a]. solutions [Abr95, Ad97, Bre99b, Che91, EDK97, Goe98, GMP91, HT95, HL94, Hu98, Iya95, KB97, KTS97, Mac94b, MM93a, MPR96, Mon93, Nes96, NW95, Nii92, Pet91, SW95b, SA94, Sma99, SV90, SS96c, ThT97, Tsu90, Wal99, Wal21, Zha98b]. Solvability [Acc96]. Solvable [Ano92, Dum91]. solvers [SV98, Hol97]. solves [BPV98]. Solving [Ano92, Cop94, Dum91, GM92c, HMW90, HH99b, Ked98, LW91, LZ92, LW93, Ste96a, Td91, AY94, AC90, CFS95, FJP97, HBLT93, Hou96b, HS95b, Isa96a, Mas94, Rud98, Sil99b, Swa93a, Zha90a, Zha90b, Zha95b, vS96]. Some [Alz93, BCM92, BL90, BX91, BB94c, Bun97, CD92, DH90, GT94, GN92, LSEG91, LPW96a, LPW96b, SW96, Spi90, Spi94, Sul93, AD91, Alz97a, Alz97b, BBG93, BJT97, CM94b, CK95, EG99, FK91, Hun92, McL94, Sem98b, She94, Ste90, Tak94]. Sommeijer [Ano95z]. Sookne [HT94]. Sophie [Dub96, IJ99]. SOR [Woz97b, Woź94]. Sorensen [Cai96]. source [Cha97, Cha99]. SP2 [Roe99]. Space [NPV91b, CL91c, FP96, Gla92, KM98, NPV91a, Tes98, Wan99]. space-time [FP96, KM98]. spaced [Buh95]. Spaces [Coc91, AS91a, BPX91b, Bra95, CH90a, Dol99, Leh97, RS96b]. sparse [HS95b, Kal95]. Special [Sol92, Wim98, Cik94, LO94, Loz98b]. Specific [Td91]. spectra [BL96, Li94a]. Spectral [BB99b, BDM90, CDT93, DS93, Hal96, LT94b, Wah98a, AKL98, Arn94, By99, BCM92, CHT96, EG99, GT91, Guo95, Mad90, Phi95a, Phi95b, She94, Tad93]. spectral-collocation [Arn94]. spectrum [Fla96, Kno93, Til97]. Speeding [McK99]. spheres [JSW99, RS96b]. Spherical [GF97, GT93, Gel97]. Spira [Tho94]. Spirals [MW92a]. Spline [MW92a, Zhe94, BF93b, CH96, DMS92, DM98, GH94, HS94, PT92, PT94, Rob94, RS98]. splines [CH90b, Dem96, FT91, Neu94, Peñ97, Qua96a, Sch94]. Splitting [DVZ94, Arn94, CG97, Hun98a, TT94, Yin90]. splittings [Hun98b]. Spontaneous [CM96]. sporadic [His94]. Springer [Köl97e]. Square [BG92b, HP99, It91, Mon94b, Sch92, ABZ95, BG92c, Wag95]. squared [Dui94, Dui96b]. Squares [Pin93b, BG94, BLP97, BS98, Cha94b, Col97, Dui94, Dui96b, Fas97, Lev98, PJV94, RAG91, Val91, ZF93]. Stability [BI93, CRV92, GP99, LM90, NPV91b, Pal95, Sur90, Tem91, BTW90, FG96, HDJ94, Hun98a, LS91b, LR90, LX93, LW96b, NHA97, Ols95a, Ols95b, Ols95c, Pal94, Peñ97, Van97, Wei97, Wu95, NPV91a]. stabilized [FSSC90, KS92]. stabilizing [Car94]. stable [Gol94a]. star [BH94a, Kui96]. Stark [DST97]. stationary [Bre90, CMP98, DVZ94, EJ93, GHS91, GMP91, LX93, LW96b, LW97, MS93, ZF93]. Statistical [BH93, EH93]. Stefan [NPV91a, NPV91b]. Stegun [RW92]. Steiner [CMS92b, CMM92a, CMS92a, MPR95, CMM92b, GGM98, Phe90a, Phe90b]. Steklov [XZ97]. Stenberg [Bre96a]. Stenger [Qua97]. step [Coh95, CRV92, JZ92, Ren90]. Stepanov [Beu96]. Stephen [Tod98]. steps [HR93]. stepsize [JZ92, Pal94]. Stewart [Kin99a, Ser97a]. Stieltjes [EM97a, GN96, Peh90b]. Stiff [Sch92, Ale91, CH98, Cha97, Cha99]. Stochastic [Cal92, BTT94a, BTT94b, BT97, Cal91, HM97, Sho98]. Stokes [Kar97, Lar97, BG94, Bre90, Cha94b, Cho97, Cot91, DDP95, DN90, GANT99, GHS91, Hal96, KS92, NW96, She96, Ste90, Yin90, ZF93]. stopping [Oli99]. strategy [DR98b, Woź94, Woz97b]. Streamline [EJ93, JSH90, Zho97]. stress [Bre99b]. stresses [NRS95]. Strictly [RS96b, Dem95]. strip [Kei96, Tol97]. Strong [NW95, Bur96, DLP93, Jae93]. strongly [MP95, Tra95]. Structural [LT91c]. structure [AS91a, BW96, CL94, GKO95, L’E99a, LS94, Mas94, RS96a, Sch93c, Sch93b, Tes98]. Structured [Pan90, Mar90]. structures [MS95, Vv96, Fal96]. Struve [Mac93]. Stuck [Sha96]. study [MT93]. Sturm [Abr95, Bou99, Cha94a, RS92]. subdivision [FN90, JRZ98, LJ96a]. subexponential [AD93]. subfield [BMO90, CDO99, Oli92]. subfields [Fuj93]. subgroup [GN92]. Subgroups [AM92]. subinterval [GS95]. Sublattices [LSK91]. submerged [MP96]. Subquadratic [KS98]. Subquadratic-time [KS98]. subschemes [LW91]. subsets [SW98]. subspace [NY97]. subspaces [HJ95]. Substructure [RW93]. subtetrahedron [LJ96a]. suitable [AM93b]. sum [BGN93, CG96a, GS95, LGRA94, Zha93b]. Summation [DLD90, Ols95a, Ols95b, Ols95c, CK95, Gau91a, Lew94, Mac97]. Sums [Vas91, Apo93, BRZ99, Cra98, Cus93, DD93, Ekl96, Ekl98, Knu93, Kui95]. Sun [Nie96b]. sup [DN90]. supercomputers [Ach99]. superconductivity [Du98]. Superconvergence [Arn97b, Wan91]. superconvergent [Zha98b]. Superlinear [Ser99, CQS98, Ser97b]. supersolvable [BC94, OS97]. superspline [AS91a, CH90a]. Supplement [AS93b, BTT94b, Bos93b, Boy96c, BG92c, Bre94b, CGJ95b, CT92b, CMS92b, CMM92a, CGT97b, CFN95b, CLT93b, Dia91a, Dil92b, Dör92b, EL92b, EM92c, GHO92b, Gir93b, HHM94b, HM92b, HA93b, KS93b, KN95, LT91b, Leh92b, LG93b, LPW96b, May92b, MM94c, Nie92b, NPV91b, Ols95c, Phe90b, Phi95b, Ros95b, Rum93b, Sch93b, Ura96b, Yan96b, BR98, BN92, DO95, DH91, DJ91, DH94, Kel95, Sou94, YR92]. supported [CD96, RS98]. supremum [Wei97]. surface [CH98, Chi95, GT94, Guo95, VC94]. surfaces [Bar96, CS96a, Che96b, Lyn94, OW94, Ura96a, Ura96b, Mey96, Wah98b, War98]. survey [Hig94, LC94]. Susanne [Bre96c]. Suzuki [Sch93b, Sch93c]. SWB [CL94]. sweep [LH98]. sweep-plane [LH98]. symbolic [Loz96]. Symbols [Gol92]. Symmetric [CM94a, Buf90, Car98, Cha92, CB93, Hig98, Liu94, Mac94a, Mas98, Mor96, Sch97, Ser99, Swa93b]. Symmetrizable [Wad90]. symmetry [BGJ99]. Symplectic [LR94]. Symposium [Gau94]. System [GM92c, BS97]. Systems [Ber90, LT91b, OS92, Phe90b, RO91, YR92, AM94, BDR98, BS99, BLP97, Che91, CJT97, CMS92b, CMM92a, CMS92a, CMM92b, EG90, EG91, FG91c, Gol94a, GGM98, HS95b, Kal95, Kli96, LT91a, LR94, LW91, LW93, LG93a, LG93b, MPR95, Nes96, Phe90a, Rum90, Sch91, Sei99, Ser97b, Ser99, WBL94, vS96, Auc96, Kli96, Sil99b]. Sz [Sch93b]. Szego [Peh93].

T [Auc96, Elm97, Gau96, Isa96c, Kok96, Loz98c]. T. [Kel95]. Table [RW92, BR05, BR11, Car92, Dia91a, Dia91b, Fik98, Köl95, Köl97c, Köl97e, Köl97a, Köl97b, Köl97g, Köl97f, Köl97d, Köl97h, Lam96, Lam97, Živ94b, Car91, FW90b, FW94, SPD94, Živ94a, Fik98]. Tables [Ano95-27, Ano95-28, Ano95-29, Ano97, CDO99, Col97, Hol97, Köl97h, L’E99a, L’E99b, Mat97, PW98, Shu97, Ste97a, Ste97b, Tem98, Van97, Wah97b, Wri97]. tabulating [SW92]. Tabulation [DH91, LSW97]. Tame [BG99b]. Tarry [Mal97]. Tate [iHT99]. Tau [Kha99, Bun97, EDK97]. Tausworthe [CLT93b, CLT93a, L’E96]. technique [SW99, Zha96]. techniques [AY94, BCT91, ELV91, RS92, Ste94a, Ste98b, SO91, Nie96b]. Temam [Lar97]. Temme [Wim98]. ten [Ian99]. tension [CH98]. tenth [Bre99c]. Term [It91, LST96, Sch97]. Termination [KP92]. terms [Cha97, Cha99, Pes96]. Ternary [Leh92a, vSv92, Kap95, Leh92b, Leh97, MT93]. Tesler [Köl90]. test [Arn95, Bur96, DLP93, KWY96b, MM93a]. testing [Bac90]. tetrahedra [LJ94]. tetrahedral [CD97, LJ96a]. Teukolsky [Loz98c]. th [Rob93, Sch98b, Ste94b, WH93]. Th. [Che96c]. Their [CG90a, Bra94, BGN93, CD98, Gau91a, KX94, MPV98, Sch96b, Tem97, WBL94, XZ94]. theorem [Arn97a, Bou99, Cop90, Dem90, DS95, Far94, För94, IK98, Lab98, LGRA94, LS91a, Sch93a, Wan99]. theorems [GH97]. theoretic [LT94a, SW92, Wah97b]. theoretical [Zha98b]. Theories [Mar92]. Theory [Kin97, Ste97a, BBG98, Bra95, Bre96c, Bun97, Che96d, Chu96, CB93, Hag91, Hig94, Mar90, Mar93, Nie96b, Odl94, Wag98, Sch96a]. thermoelasticity [Cop99]. theta [CLR97]. thin [Lee99]. Third [Wah98a, FSSC90, Mel98]. third-order [FSSC90, Mel98]. Thomas [Bre97, WH99]. thousand [Ian99]. Three [BB93, Ach99, BN97, BGN93, Cha94b, Chi95, HP99, Neu94, NW98]. three-directional [Neu94]. Thron [Ano95r]. Thue [De 95, MPR96, Pet91, Td91]. tight [RS98]. Tikhonov [NHA97]. Tiling [BG92b, BG92c]. Time [Ric90, TT94, AV99, Ber98, Che91, Cot91, Du98, EL95, FP96, GP99, HHM94a, HHM94b, KS98, KM90, KM98, Oli99, Sho98, TW94, WBL94]. Time-Dependent [Ric90, EL95, GP99]. time-harmonic [AV99]. Time-splitting [TT94]. Titchmarsh [Spi94]. Tits [Gol91b]. Toda [NTS98]. Toeplitz [CY92, Grö99, Ser97b, Ser99, Til97, YC93]. too [CM94b]. tools [Hou96a, Ste98b]. Topics [Che96c, Qua97]. topological [Abe94]. Torczon [Gau96]. torsion [iN97, Sem98a]. Total [GS98, Tad93]. Totally [Dia91a, Dia91b, DST97, Fla96, Fuj93, Gaá96, GP97, Pv93, Tak94]. Tourancheau [Hou96a]. Trabucho [Fal96]. tracking [Shy94]. traction [Bre94a, Bre94b]. trade [NHA97]. trade-off [NHA97]. transcendental [Köl97c]. transform [Str92a, Wei95, Cal91, Nie96c]. transformation [GL98a, MS97a, Sid95, GL98a]. transformed [BBN99]. transforming [Oli98]. transforms [CB93, CF94, CV95b, Köl97h, Mas98, PST98, ST91, Tem97, Tem98]. Transition [MW92a]. Transitive [CMM92a, CMM92b]. translation [CD98, Fik98, Köl97g, Lam97]. transonic [BF93a, BWW97]. Trapezoidal [Hun98b]. traveling [Fan98]. treatment [CRV92, FG91c, Ste97a]. Treatments [ZZ97]. Tree [AL95a]. Tree-search [AL95a]. Trefethen [Ste99c]. trends [Néd94]. triangles [LC94, RI96]. Triangular [LJ96b, Ad97, Žen95]. triangularly [Sma95]. Tricomi [Köl97c, Köl97h]. tridiagonal [AM94, Oli98, Swa93b]. Trigonometric [Qua96b, CS96b, CK95, Fas97, McL94, RAG91]. Triple [Phe90b, YR92, BBG98, CMS92b, CMM92a, CMS92a, CMM92b, MPR95, Phe90a]. triples [EHN97]. triplets [Wag95]. trivariate [Dre94]. Trotter [IK98]. truncated [Eck93, Eck95]. tuples [LH98]. Turán [CG96a, GM96]. Turyn [ESA94, KKS90]. Twenty [PMT95, CDNY95]. twenty-second [CDNY95]. Twenty-two [PMT95]. Twin [PSZ90, Bre97, For97, IJ96, IJ99]. Two [Bre96b, Car92, Fla96, Hab93, Mar92, MW92b, NPV91b, Ren90, Sko92, Ada95, Bai94c, BCM92, BL94, BV92, BCMR95, EHN92, EHI94, FMOS98, GLW99, GL99, iHT99, JZ92, LT91c, LM94a, NPV91a, OS91, PMT95, Ros95a, SA94, Str92b, Sty93, SO91, Sun95, Via91, Wal99, Wal21, Yin98, van99]. two-dimensional [BCM92, BL94, BV92, Sun95, Yin98]. Two-level [Bre96b]. Two-Phase [NPV91b, NPV91a]. two-point [BCMR95, FMOS98, SO91, Via91]. Two-step [Ren90, JZ92]. two-variable [Sty93]. Twofold [YR92]. Type [BN90, RS90, DVZ94, EST90, EG99, För94, FG91c, Gec93, Grö99, Kha99, LST96, Nii92, Not92, RS94, Tem94b, Žen95, Zha93a, KX94, Kui95, Sch91].

Überhuber [Fro97]. ULP [Ziv99a].

Ultraconvergence [Zha96]. ultraspherical [FP90]. unbounded [FK91, Hal96, Yin90]. Unconditional [Hun92]. unconstrained [Gol94b]. unconventional [KL97]. Unicity [GW96]. unified [LG97]. Uniform [Al 94, EL91, Gas93, Rab91, Car96, FMOS98, Tem94b, TV93, Zha90a, Zha90b]. uniformly [OS91, SS96c]. unifying [BPX91a, Mel97]. Unimodular [Hun91]. unique [SL90]. uniqueness [Fan98, GMP91]. unit [BPv94, JKN91, MPV98, Pv93, PW98, San92, Sma99]. units [BS96b, DST97, Fuk96, Ge94, Laz93, NS98, SSW99]. unity [HH99b, Pil90, Sch98b]. Universal [EK97]. unsymmetric [Huc95]. Update [Mar92]. updates [Mar90]. Updating [Ste92]. Upper [HS95a, DGNP93, Fly98, Str92b]. upstream [Via91]. use [LS91b, Néd94, Swa93a, WC93, Che96b]. Used [Ber90]. User [Cai96, Qua96a]. Using [Ste96a, Abe94, Ben98a, CHL99, CV95a, GM92b, Kli96, Nie94, Sei99, SO91, Wen99, XZ97]. utilization [Woź94, Woz97b].

V [Che96c, Ell96, Gau96, Mea96, Qua96a, Shi96, Wah98a, Wri96]. V. [dB98]. Vajtersic [Hol97]. Valentin [Shi96]. Valérie [Ste97b]. valid [SW94]. Value [Rab90a, Sch92, AD91, Ad97, Bia90, BS99, BLP93, Bra96b, BCMR95, CDL97, Die94, DS93, FMOS98, GT91, HT91, HP92, Hun92, Hun98b, LS91b, Mat97, Mon94a, Rab90b, Rab91, Ste94a, SO91, Wu95, Zhu92]. valued [NW94]. Values [CK99, AT99, FW90a, Odl93, Shi95, YZ97]. Vancouver [Gau94]. Vandermonde [RO91]. Variable [JZ92, BTT94a, BTT94b, Pal94, Sty93]. Variable-stepsize [JZ92]. Variables [dR92, De 95, PJV94, SA94]. variates [Bai94b]. variation [GS98, Tad93, Ye94]. Variational [SS91, vBv92, BF93a, BV92, CQS98, CL91b, EST90, NW95, Vv96]. varieties [Pil90]. various [BBP97]. Vasile [Ben98b]. Vasil’eva [Shi96]. Vector [JRZ98, Emm96, HBLT93, Hou96b, Nie95, vKP95]. vectors [BW96, CLT93a, CLT93b, Kny97, MS93]. version [BS92a, BL94, HST98, HGK98, Ray94]. versions [SS96a]. vertex [CH90b, DH90, GS97, MSS97, Sül92]. Vetterling [Loz98c]. via [Cop94, CFN95a, CFN95b, FvS99, GW96, NW94, Sil99a]. Viaño [Fal96]. vibration [DHNL^{+}99]. Victor [Chu96, Hig96]. viewpoint [Roj94]. Virginia [Gau96]. Viscosity [ThT97, CDT93, Tad93]. Viscous [Yin90]. Vlasov [BT97]. Vol [Hig96, Köl97g, Köl97f, Köl97h]. Volterra [Bru94, BPV99, CRV92, KB97, Mak94, Pal96]. Volume [Köl97a, Köl97b, Law91, Loz98c, Wag98, BLS94, CCF94, GS97, MM93a, MSS97, Noe96, Sül92]. Voronoı [Ada95]. Voronoı-algorithm [Ada95]. Vorst [Ano95q]. Vortex [Hou92, Yin98, CHL99]. vorticity [Guo95]. Vries [KM90, Hua91].

W [BW96, Che97, Hou96b, Kin97, Kin99a, Ser97a, Wah97a, Wat98]. W. [Köl97c, Köl97h]. Wahlbin [Arn97b]. walk [Lu98]. Walsh [LT94a, LSW94]. Walter [Ell96, Isa96a, Len97, Ste96a, Ste99b]. Ward [Leh93]. Waring [DD90, KW90]. was [Wil93]. Washington [FS95]. water [Ach99]. Watson [Gau96]. wave [BL95, BV92, Fen99, FP96, HDJ94, HHM94a, HHM94b, She93, Shu96, Shy94, vBv92]. Waveform [BS97]. waveforms [HA93a, HA93b]. waveguides [BG99a]. Wavelet [MW94a, Chu96, DS99, LLS96, Zhe94]. Wavelets [Che96d, FP97, SS96b, CMX99, CD96, Jia98, JRZ98, Lai94, Qua96b]. waves [Fan98, Hag91, Hua91, Mak93]. wavewise [Yan96b, Yan96a]. way [Cha94a]. weak [HB92, MM95]. weakly [BPV99, CTW92, CMM90, KX94, MS98, Nes96, Pal96]. Webb [Ano95d]. Weber [YZ97]. Weight [vSv92, GN96, GM96, Peh93, vv93b]. Weighted [BX91, DN90, CF94, Wei97, Xu91]. weighting [Via91]. weights [FP90, HST93]. Weil [Sil90]. well [CP97]. Wendt [FG91a, Hel97]. Wesseling [Ano95n]. Weyl [Sut98]. Which [BGN93, ST96, Arn95, CLR97, Cla96, FG91b, FW90a, GM96, HC98, HB92, LW91]. Whittaker [Bou99]. Who [Roj94]. Whose [KM91, Coh97a]. Wickerhauser [Chu96]. Wiedemann [Cop94, Kal95]. Wieferich [Cik94, Cop90, CDP97]. Wiener [EL93]. Wigner [AN92, Arn94]. wild [BG99b]. Wilfrid [Kel95]. William [Gil98, Loz98c, Che97, Loz98c, Nie96c, Zha98a]. Williams [WH93]. Williamson [BW96]. Wilson [ADS98, CL90, CDP97]. Windowing [Ste92]. Winkler [Bos98]. without [BPWX91a, Bre99a, BG96, HR93, Shi95]. Witt [JMS97, JM92, JMS95]. Wolfgang [Kel95, Ste97a]. work [Leh93]. world [Isa96a]. WR [DVZ94]. WR-type [DVZ94]. Wright [Tod98]. wrong [HL94].

Xu [de 96].

years [Bre96a]. Yokoi [Zha95a]. York [Köl97e, Köl97g, Köl97f, Köl97h]. Yu [Köl97g, Köl97f, Nie96b]. Yu. [Köl97a, Köl97b]. Yurii [Tod96]. Yves [Kea98].

Z. [Cai96]. Zahar [Ell96]. Zakharov [CGJ95b, CGJ95a, Gla92]. Zaremba [Cus93]. zero [BF92, CT92a, CT92b, CT94]. Zeros [EM92a, It91, SSW99, Tho94, Tol97, Ben98a, Che94a, CSV94, EM92c, EM94, FS92, FO93, GR95, HB92, Ish95, Kei96, Kui96, RS95, SV90, Spi94, Che96c]. Zeta [Den90, GLW99, Cra98, CK99, Tol97]. Zhang [Loz98b]. Zippel [Ano95s]. Zwillinger [Ano95e].

References

Arnold:1997:LFF

[AB97] Douglas N. Arnold and Franco Brezzi. Locking-free finite element methods for shells. Mathematics of Computation, 66(217):1– 14, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00785-
0&u=/mcom/1997-66-217/.

Aberth:1994:CTD

[Abe94] Oliver Aberth. Computation of topological degree using interval arithmetic, and applications. Mathematics of Computation, 62(205):171–178, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Abrate:1995:MOA

[Abr95] Serge Abrate. On a method for obtaining approximate solutions to Sturm– Liouville problems. Mathematics of Computation, 64(212):1511–1515, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ashwin:1995:NLM

[ABZ95] Peter Ashwin, Klaus Böhmer, and Mei Zhen. A numerical Liapunov–Schmidt method with applications to Hopf bifurcation on a square. Mathematics of Computation, 64(210):649–670, S19–S22, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Atkinson:1990:BIE

[AC90] Kendall E. Atkinson and Graeme Chandler. Boundary integral equation methods for solving Laplace’s equation with nonlinear boundary conditions: The smooth boundary case. Mathematics of Computation, 55(192):451– 472, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arbogast:1995:IMM

[AC95] Todd Arbogast and Zhangxin Chen. On the implementation of mixed methods as nonconforming methods for secondorder elliptic problems. Mathematics of Computation, 64(211):943–972, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Acciaro:1996:SNE

[Acc96]  Vincenzo Acciaro.  Solvability of

norm equations over cyclic number fields of prime degree. Mathematics of Computation, 65(216):1663– 1674, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00760-
0&u=/mcom/1996-65-216/.

Ames:1998:CRI

[ACEO98] Karen A. Ames, Gordon W. Clark, James F. Epperson, and Seth F. Oppenheimer. A comparison of regularizations for an ill-posed problem. Mathematics of Computation, 67(224):1451– 1471, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-01014-
X&u=/mcom/1998-67-224/.

Achdou:1999:BRN

[Ach99] Yves Achdou. Book reviews: Numerical methods for the three dimensional shallow water equations on supercomputers, by E. D. de Goede. Mathematics of Computation, 68(225): ??, January 1999. URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-00992-
8&u=/mcom/1999-68-225/.

Akrivis:1998:IEM

[ACM98] Georgios Akrivis, Michel Crouzeix, and Charalambos Makridakis. Implicitexplicit multistep finite element methods for nonlinear parabolic problems. Mathematics of Computation, 67(222): 457–477, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00930-
2&u=/mcom/1998-67-222/.

Akrivis:1991:FDD

[AD91] Georgios D. Akrivis and Vassilios A. Dougalis. Finite difference discretizations of some initial and boundary value problems with interface. Mathematics of Computation, 56(194):505–522, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Adleman:1993:SAD

[AD93] Leonard M. Adleman and Jonathan Demarrais. A subexponential algorithm for discrete logarithms over all finite fields. Mathematics of Computation, 61 (203):1–15, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Andrade:1997:NSB

[Ad97] M. G. Andrade and J. B. R. do Val. A numerical scheme based on mean value solutions for the Helmholtz equation on triangular grids. Mathematics of Computation, 66(218):477– 493, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00825-
                                              9&u=/mcom/1997-66-218/.

Adam:1995:VAE

[Ada95]  Brigitte Adam.  Voronoĭ-algorithm

expansion of two families with period length going to infinity. Mathematics of Computation, 64(212):1687– 1704, October 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Agoh:1998:WQC

[ADS98] Takashi Agoh, Karl Dilcher, and Ladislav Skula. Wilson quotients for composite moduli. Mathematics of Computation, 67(222):843– 861, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00951-
X&u=/mcom/1998-67-222/.

Arnold:1997:PA

[AFW97] Douglas N. Arnold, Richard S. Falk, and R. Winther. Preconditioning in h(div) and applications. Mathematics of Computation, 66(219):957– 984, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00826-
0&u=/mcom/1997-66-219/.

Anderssen:1999:NDD

[AH99] Robert S. Anderssen and Markus Hegland. For numerical differentiation, dimensionality can be a blessing! Mathematics of Computation, 68(227):1121– 1141, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01033-
9&u=/mcom/1999-68-227/.

Aaltonen:1991:CER

[AI91] M. Aaltonen and K. Inkeri. Catalan’s equation x^{p}−y^{q} = 1 and related congruences. Mathematics of Computation, 56 (193):359–370, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Acciaro:1999:CAA

[AK99] Vincenzo Acciaro and Jürgen Klüners.

Computing automorphisms of abelian number fields. Mathematics of Computation, 68(227):1179–1186, July 1999.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                              http://www.ams.org/jourcgi/jour-
                                              pbprocess?fn=110&arg1=S0025-5718-
                                              99-01084-4&u=/mcom/1999-68-227/.

Alam:1998:ASA

[AKL98] Rafikul Alam, Rekha P. Kulkarni, and Balmohan V. Limaye. Accelerated spectral approximation. Mathematics of Computation, 67(224):1401– 1422, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00980-
                                              6&u=/mcom/1998-67-224/.

AlShenk:1994:UEE

[Al 94] N. Al Shenk. Uniform error estimates for certain narrow Lagrange finite elements. Mathematics of Computation, 63(207):105–119, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Aziz:1991:GMF

[AL91] A. K. Aziz and J.-L. Liu. A Galerkin method for the forward-backward heat equation. Mathematics of Computation, 56(193):35–44, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Applegate:1995:DBPa

[AL95a] David Applegate and Jeffrey C. Lagarias. Density bounds for the 3x + 1 problem. I. Tree-search method. Mathematics of Computation, 64(209):411– 426, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Applegate:1995:DBPb

[AL95b] David Applegate and Jeffrey C. Lagarias. Density bounds for the 3x + 1 problem. II. Krasikov inequalities. Mathematics of Computation, 64(209): 427–438, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alexander:1991:MNM

[Ale91] Roger Alexander. The modified Newton method in the solution of stiff ordinary differential equations. Mathematics of Computation, 57(196):673– 701, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alzer:1993:SGF

[Alz93] Horst Alzer. Some gamma function inequalities. Mathematics of Computation, 60(201):337–346, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alzer:1997:SIG

[Alz97a] Horst Alzer. On some inequalities for the gamma and psi functions. Mathematics of Computation, 66(217):373– 389, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00807-
7&u=/mcom/1997-66-217/.

Alzer:1997:SII

[Alz97b] Horst Alzer. On some inequalities for the incomplete gamma function. Mathematics of Computation, 66(218): 771–778, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

                                              fn=110&arg1=S0025-5718-97-00814-
                                              4&u=/mcom/1997-66-218/.

Ash:1992:DCC

[AM92] Avner Ash and Mark McConnell. Doubly cuspidal cohomology for principal congruence subgroups of GL(3, ℤ). Mathematics of Computation, 59(200): 673–688, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Atkin:1993:ECP

[AM93a] A. O. L. Atkin and F. Morain. Elliptic curves and primality proving. Mathematics of Computation, 61(203): 29–68, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Atkin:1993:FSC

[AM93b] A. O. L. Atkin and F. Morain. Finding suitable curves for the elliptic curve method of factorization. Mathematics of Computation, 60(201):399– 405, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Amodio:1994:BEA

[AM94] Pierluigi Amodio and Francesca Mazzia. Backward error analysis of cyclic reduction for the solution of tridiagonal systems. Mathematics of Computation, 62(206):601–617, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Andre:1990:FMD

[AMN90] Debra A. André, Gary L. Mullen, and Harald Niederreiter. Figures of merit for digital multistep pseudorandom numbers. Mathematics of Computation, 54 (190):737–748, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arnold:1992:NAD

[AN92] Anton Arnold and Francis Nier. Numerical analysis of the deterministic particle method applied to the Wigner equation. Mathematics of Computation, 58 (198):645–669, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anglin:1996:SPE

[Ang96] W. S. Anglin. Simultaneous Pell equations. Mathematics of Computation, 65(213):355–359, January 1996. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
96-00687-4&u=/mcom/1996-65-213/.

Anonymous:1992:CSS

[Ano92] Anonymous. Corrigenda: “Solving Solvable Quintics” [Math. Comp. 57 (1991), no. 195, 387–401, MR 91j:12005] by D. S. Dummit. Mathematics of Computation, 59(199):309, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRPa

[Ano95a] Anonymous. Book reviews: 10. Pryce.

Mathematics of Computation, 64(210): ??, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRHa

[Ano95b] Anonymous. Book reviews: 11. Hoschek and Lasser. Mathematics of Computation, 64(210):??, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRO

[Ano95c] Anonymous. Book reviews: 12. O’Rourke. Mathematics of Computation, 64(210):??, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRSb

[Ano95d] Anonymous. Book reviews: 13. Schumaker and Webb, editors. Mathematics of Computation, 64(210):??, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRZa

[Ano95e] Anonymous. Book reviews: 14. Zwillinger. Mathematics of Computation, 64 (210):??, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRSc

[Ano95f] Anonymous. Book reviews: 15. Smith. Mathematics of Computation, 64(210): ??, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRBa

[Ano95g] Anonymous. Book reviews: 16. Brass and Hammerlin, editors. Mathematics of Computation, 64(210):??, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRBb

[Ano95h] Anonymous. Book reviews: 17. Bouleau and Lepingle. Mathematics of Computation, 64(210):??, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRG

[Ano95i] Anonymous. Book reviews: 18. Gray.

Mathematics of Computation, 64(210): ??, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BREa

[Ano95j] Anonymous. Book reviews: 19. Eisenbud and Robbiano, editors. Mathematics of Computation, 64(210):??, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRR

[Ano95k] Anonymous. Book reviews: 20. Rombaldi. Mathematics of Computation, 64 (211):??, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRPb

[Ano95l]  Anonymous.  Book reviews: 21.

Prathap. Mathematics of Computation, 64(211):??, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRBc

[Ano95m] Anonymous. Book reviews: 22.

Bank. Mathematics of Computation, 64 (211):??, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRHb

[Ano95n] Anonymous. Book reviews: 23. Hemker and Wesseling, editors. Mathematics of Computation, 64(211):??, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRSd

[Ano95o] Anonymous. Book reviews: 24. Shampine. Mathematics of Computation, 64(211):??, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRSe

[Ano95p] Anonymous. Book reviews: 25. Sanz-Serna and Calvo. Mathematics of Computation, 64(211):??, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRBd

[Ano95q] Anonymous. Book reviews: 26. Barrett, Berry, Chan, Demmel, Donato, Dongarra, Eijkhout, Pozo, Romine, and van der Vorst. Mathematics of Computation, 64(211):??, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRC

[Ano95r] Anonymous. Book reviews: 27. Cooper and Thron, editors. Mathematics of Computation, 64(211):??, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRZb

[Ano95s] Anonymous. Book reviews: 28. Zippel. Mathematics of Computation, 64 (211):??, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BREb

[Ano95t] Anonymous. Book reviews: 29. Euvrard. Mathematics of Computation, 64 (212):??, October 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRBe

[Ano95u] Anonymous. Book reviews: 30. Bulirsch and Kraft, editors. Mathematics of Computation, 64(212):??, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRHc

[Ano95v] Anonymous. Book reviews: 31. Hackbusch. Mathematics of Computation, 64 (212):??, October 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRSf

[Ano95w] Anonymous. Book reviews: 32.

Sobolev. Mathematics of Computation, 64(212):??, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRPc

[Ano95x] Anonymous. Book reviews: 33. Pohst.

Mathematics of Computation, 64(212): ??, October 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRM

[Ano95y] Anonymous. Book reviews: 8. Marchuk, editor. Mathematics of Computation, 64 (210):8–??, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:BRSa

[Ano95z] Anonymous. Book reviews: 9. Sommeijer. Mathematics of Computation, 64 (210):??, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:RDTa

[Ano95-27] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 64(210):891–906, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:RDTb

[Ano95-28] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 64(211):1341–1354, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1995:RDTc

[Ano95-29] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 64(212):1757– ??, October 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1996:C

[Ano96] Anonymous. Corrigendum. Mathematics of Computation, 65(213):445– 446, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00697-
7&u=/mcom/1996-65-213/.

Anonymous:1997:RDT

[Ano97] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 66(218):915–925, April 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Apostol:1993:ELP

[Apo93] Tom M. Apostol. An extension of the Lehmers’ picturesque exponential sums. Mathematics of Computation, 61 (203):25–28, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alefeld:1993:ESR

[APS93] G. Alefeld, F. A. Potra, and Yixun Shi.

On enclosing simple roots of nonlinear equations. Mathematics of Computation, 61(204):733–744, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arno:1991:NPP

[Arn91] Steven Arno. A note on Perrin pseudoprimes. Mathematics of Computation, 56(193):371–376, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arnold:1994:MSC

[Arn94] Anton Arnold. A mixed spectralcollocation and operator splitting method for the Wigner–Poisson equation. In Gautschi [Gau94], pages 249– 253. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Arnault:1995:RMP

[Arn95] F. Arnault. Rabin–Miller primality test:

composite numbers which pass it. Mathematics of Computation, 64(209):355– 361, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arnault:1997:RMT

[Arn97a] F. Arnault. The Rabin–Monier theorem for Lucas pseudoprimes. Mathematics of Computation, 66(218):869– 881, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00836-
                                              3&u=/mcom/1997-66-218/.

Arnold:1997:BRS

[Arn97b] D. N. Arnold. Book reviews: Superconvergence in Galerkin finite element methods, by Lars B. Wahlbin. Mathematics of Computation, 66(217): ??, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00793-
                                              X&u=/mcom/1997-66-217/.

Alfeld:1991:SMS

[AS91a] Peter Alfeld and Maritza Sirvent. The structure of multivariate superspline spaces of high degree. Mathematics of Computation, 57(195):299–308, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Atkinson:1991:NSF

[AS91b] Kendall E. Atkinson and Ian H. Sloan. The numerical solution of firstkind logarithmic-kernel integral equations on smooth open arcs. Mathematics of Computation, 56(193):119– 139, January 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Aiello:1993:CAC

[AS93a] Walter Aiello and M. V. Subbarao. A conjecture in addition chains related to Scholz’s conjecture. Mathematics of Computation, 61(203):17–23, S1–S6, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Aiello:1993:SCA

[AS93b] Walter Aiello and M. V. Subbarao. Supplement to “A conjecture in addition chains related to Scholz’s conjecture”. Mathematics of Computation, 61(203): S1–S6, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Abia:1993:PRK

[ASS93] L. Abia and J. M. Sanz-Serna. Partitioned Runge–Kutta methods for separable Hamiltonian problems. Mathematics of Computation, 60(202):617– 634, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Akiyama:1999:CVF

[AT99] Shigeki Akiyama and Yoshio Tanigawa.

Calculation of values of L-functions associated to elliptic curves. Mathematics of Computation, 68(227):1201– 1231, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01051-
0&u=/mcom/1999-68-227/.

Auchmuty:1996:BRC

[Auc96] Giles Auchmuty. Book reviews: Control and estimation in distributed parameter systems, edited by H. T. Banks. Mathematics of Computation, 65(213): ??, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Auzinger:1998:BRN

[Auz98] Winfried Auzinger. Book reviews:

Numerical methods for differential equations. A computational approach, by John R. Dormand. Mathematics of Computation, 67(221):??, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00891-
6&u=/mcom/1998-67-221/.

Alonso:1999:ODD

[AV99] Ana Alonso and Alberto Valli. An optimal domain decomposition preconditioner for low-frequency timeharmonic Maxwell equations. Mathematics of Computation, 68(226):607– 631, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01013-
3&u=/mcom/1999-68-226/.

Adamchik:1997:SF

[AW97] Victor Adamchik and Stan Wagon. A simple formula for π. American Mathematical Monthly, 104(9):852–855, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_nov97_toc.html.

The authors employ Mathematica to extend earlier work of Bailey, Borwein, and Plouffe, [BBP97], done in 1995, but only just published, that discovered an amazing formula for π as is a power series in 16^{−k}, enabling any base-16 digit of π to be computed without knowledge of any prior digits. In this paper, Mathematica is used to find several simpler formulas having powers of 4^{−k}. They also note that it has been proven that their methods cannot be used to exhibit similar formulas in powers of 10^{−k}.

Adams:1994:CTS

[AY94] Loyce Adams and Zhiyun Yang. A comparison of techniques for solving ill-conditioned problems arising from the immersed boundary method. In Gautschi [Gau94], pages 243–247. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Bach:1990:EBP

[Bac90] Eric Bach. Explicit bounds for primality testing and related problems. Mathematics of Computation, 55(191): 355–380, July 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bach:1997:CSP

[Bac97] Eric Bach. Comments on search procedures for primitive roots. Mathematics of Computation, 66(220):1719– 1727, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00890-
9&u=/mcom/1997-66-220/.

Bai:1994:EAL

[Bai94a] Zhao Jun Bai. Error analysis of the Lánczos algorithm for the nonsymmetric eigenvalue problem. Mathematics of Computation, 62(205):209– 226, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bailey:1994:PGR

[Bai94b] Ralph W. Bailey. Polar generation of random variates with the t-distribution. Mathematics of Computation, 62(206):

779–781, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baines:1994:AOD

[Bai94c] M. J. Baines. Algorithms for optimal discontinuous piecewise linear and constant L_{2} fits to continuous functions with adjustable nodes in one and two dimensions. Mathematics of Computation, 62(206):645–669, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bailey:1996:BRE

[Bai96] David H. Bailey. Book reviews: The encyclopedia of integer sequences, by N. J. A. Sloane and Simon Ploufee. Mathematics of Computation, 65(214): 894, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Barry:1996:BRC

[Bar96] Phillip Barry. Book reviews: Curves and surfaces in geometric design, by Pierre-Jean Laurent, Alain le Méhauté and Larry L. Schumaker. Mathematics of Computation, 65(213): 435, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Bremner:1993:TPG

[BB93] Andrew Bremner and Duncan A. Buell. Three points of great height on elliptic curves. Mathematics of Computation, 61(203):111–115, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Biehl:1994:AQO

[BB94a] Ingrid Biehl and Johannes Buchmann.

Algorithms for quadratic orders. In Gautschi [Gau94], pages 425–449. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Bronson:1994:CSF

[BB94b] Nathan D. Bronson and Duncan A.

Buell. Congruential sieves on FPGA computers. In Gautschi [Gau94], pages 547–551. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Browkin:1994:SRC

[BB94c] J. Browkin and J. Brzeziński. Some remarks on the abc-conjecture. Mathematics of Computation, 62(206):931– 939, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barrett:1999:FEA

[BB99a] John W. Barrett and James F. Blowey.

Finite element approximation of the Cahn–Hilliard equation with concentration dependent mobility. Mathematics of Computation, 68(226):487– 517, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01015-
7&u=/mcom/1999-68-226/.

Belgacem:1999:SED

[BB99b] F. Ben Belgacem and C. Bernardi.

Spectral element discretization of the Maxwell equations. Mathematics of Computation, 68(228):1497– 1520, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01086-
8&u=/mcom/1999-68-228/.

Beaucoup:1998:PSR

[BBBP98] Franck Beaucoup, Peter Borwein, David W. Boyd, and Christopher Pinner. Power series with restricted coefficients and a root on a given ray. Mathematics of Computation, 67(222): 715–736, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00960-
0&u=/mcom/1998-67-222/.

Bailey:1997:KC

[BBC97] David H. Bailey, Jonathan M. Borwein, and Richard E. Crandall. On the Khintchine constant. Mathematics of Computation, 66(217):417– 431, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00800-
4&u=/mcom/1997-66-217/.

Blecksmith:1990:MRI

[BBG90] Richard Blecksmith, John Brillhart, and Irving Gerst. On the mod2 reciprocation of infinite modular-part products and the parity of certain partition functions. Mathematics of Computation, 54(189):345–376, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blecksmith:1991:CIM

[BBG91] Richard Blecksmith, John Brillhart, and Irving Gerst. On a certain (Mod 2) identity and a method of proof by expansion. Mathematics of Computation, 56(194):775–794, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blecksmith:1993:FMI

[BBG93] Richard Blecksmith, John Brillhart, and Irving Gerst. A fundamental modular identity and some applications. Mathematics of Computation, 61(203): 83–95, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blecksmith:1998:CTT

[BBG98] Richard Blecksmith, John Brillhart, and Irving Gerst. A constructive theory of triple and quintuple product identities of the second degree. Mathematics of Computation, 67(222):797– 814, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00931-
4&u=/mcom/1998-67-222/.

Baltensperger:1999:ECL

[BBN99] Richard Baltensperger, Jean-Paul Berrut, and Benjamin Noël. Exponential convergence of a linear rational interpolant between transformed Chebyshev points. Mathematics of Computation, 68(227):1109– 1120, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01070-
4&u=/mcom/1999-68-227/.

Bouchut:1996:MMS

[BBP96] F. Bouchut, Ch. Bourdarias, and B. Perthame. A MUSCL method satisfying all the numerical entropy inequalities. Mathematics of Computation, 65(216):1439–1461, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00752-
1&u=/mcom/1996-65-216/.

Bailey:1997:RCV

[BBP97] David Bailey, Peter Borwein, and Simon Plouffe. On the rapid computation of various polylogarithmic constants. Mathematics of Computation, 66(218):903–913, April 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/journals/
mcom/1997-66-218/S0025-5718-97-
00856-9/S0025-5718-97-00856-9.pdf.

See [AW97].

Baum:1994:CIR

[BC94] Ulrich Baum and Michael Clausen. Computing irreducible representations of supersolvable groups. Mathematics of Computation, 63(207):351–359, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bloom:1997:CPM

[BC97] Thomas Bloom and Jean-Paul Calvi. A continuity property of multivariate Lagrange interpolation. Mathematics of Computation, 66(220):1561– 1577, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00858-
2&u=/mcom/1997-66-220/.

Butcher:1998:ERK

[BC98] J. C. Butcher and R. P. K. Chan.

Efficient Runge–Kutta integrators for index-2 differential algebraic equations. Mathematics of Computation, 67(223): 1001–1021, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00953-
3&u=/mcom/1998-67-223/.

Buhler:1993:IPC

[BCEM93] J. Buhler, R. Crandall, R. Ernvall, and T. Metsänkylä. Irregular primes and cyclotomic invariants to four million. Mathematics of Computation, 61(203): 151–153, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bernardi:1992:SSA

[BCM92] Christine Bernardi, Giuseppe Coppoletta, and Yvon Maday. Some spectral approximations of two-dimensional fourth-order problems. Mathematics of Computation, 59(199):63–76, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Butcher:1995:DCT

[BCMR95] J. C. Butcher, J. R. Cash, G. Moore, and R. D. Russell. Defect correction for two-point boundary value problems on nonequidistant meshes. Mathematics of Computation, 64(210):629–648, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buhler:1992:IPO

[BCS92] J. P. Buhler, R. E. Crandall, and R. W.

Sompolski. Irregular primes to one million. Mathematics of Computation, 59 (200):717–722, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1991:ITL

[BCT91] R. P. Brent, G. L. Cohen, and H. J. J. Te Riele. Improved techniques for lower bounds for odd perfect numbers. Mathematics of Computation, 57(196):857– 868, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Batterson:1992:LCS

[BD92a] Steve Batterson and David Day. Linear convergence in the shifted QR algorithm. Mathematics of Computation, 59(199):141–151, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boersma:1992:NEL

[BD92b] J. Boersma and J. P. Dempsey. On the numerical evaluation of Legendre’s chifunction. Mathematics of Computation, 59(199):157–163, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bennett:1998:DE

[Bd98] Michael A. Bennett and Benjamin M. M. de Weger. On the Diophantine equation |ax^{n} − by^{n}| = 1. Mathematics of Computation, 67(221):413– 438, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00900-
                                              4&u=/mcom/1998-67-221/.

Bernardi:1990:CFE

[BDM90] Christine Bernardi, Naïma Debit, and Yvon Maday. Coupling finite element and spectral methods: First results. Mathematics of Computation, 54(189):

21–39, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bermudez:1998:FEA

[BDR98] Alfredo Bermúdez, Ricardo Durán, and Rodolfo Rodríguez. Finite element analysis of compressible and incompressible fluid-solid systems. Mathematics of Computation, 67(221):111– 136, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00901-
6&u=/mcom/1998-67-221/.

Borwein:1996:ICP

[BE96] Peter Borwein and Tamás Erdélyi. The integer Chebyshev problem. Mathematics of Computation, 65(214):661– 681, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00702-
8&u=/mcom/1996-65-214/.

Buchmann:1999:FRN

[BE99] Johannes Buchmann and Friedrich Eisenbrand. On factor refinement in number fields. Mathematics of Computation, 68(225):345–350, January 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01023-
6&u=/mcom/1999-68-225/.

Bean:1997:PCA

[Bea97]  Michael A. Bean.  The practical

computation of areas associated with binary quartic forms. Mathematics of Computation, 66(219):1269– 1293, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00815-
                                              6&u=/mcom/1997-66-219/.

Belabas:1997:FAC

[Bel97] K. Belabas. A fast algorithm to compute cubic fields. Mathematics of Computation, 66(219):1213– 1237, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00846-
                                              6&u=/mcom/1997-66-219/.

BenGhanem:1998:QFU

[Ben98a]  Riadh Ben Ghanem.  Quadrature

formulae using zeros of Bessel functions as nodes. Mathematics of Computation, 67(221):323–336, January 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00882-
                                              5&u=/mcom/1998-67-221/.

Benner:1998:BRA

[Ben98b] Peter Benner. Book reviews: Algorithms for linear-quadratic optimization, by Vasile Sima. Mathematics of Computation, 67(221): ??, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00891-
                                              6&u=/mcom/1998-67-221/.

Berger:1990:ACF

[Ber90] Alan E. Berger. Accurate calculation of functions used in a model of the nematic behavior of self-assembling systems. Mathematics of Computation, 54 (189):313–330, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berrut:1994:LRI

[Ber94] Jean-Paul Berrut. Linear rational interpolation of continuous functions over an interval. In Gautschi [Gau94], pages 261–264. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Bernstein:1998:DPP

[Ber98] Daniel J. Bernstein. Detecting perfect powers in essentially linear time. Mathematics of Computation, 67(223): 1253–1283, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00952-
1&u=/mcom/1998-67-223/.

Beukers:1996:BRA

[Beu96] F. Beukers. Book reviews: Arithmetic of algebraic curves, by Serguei A. Stepanov. Mathematics of Computation, 65(216):??, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Bohman:1992:FRP

[BF92] Jan Bohman and Carl-Erik Fröberg.

The Γ-function revisited: power series expansions and real-imaginary zero lines. Mathematics of Computation, 58 (197):315–322, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berger:1993:AFE

[BF93a] Harald Berger and Miloslav Feistauer. Analysis of the finite element variational crimes in the numerical approximation of transonic flow. Mathematics of Computation, 61(204):493– 521, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bialecki:1993:OSC

[BF93b] Bernard Bialecki and Ryan I. Fernandes. Orthogonal spline collocation Laplace-modified and alternatingdirection methods for parabolic problems on rectangles. Mathematics of Computation, 60(202):545–573, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buchmann:1993:EQF

[BFP93] Johannes Buchmann, David Ford, and Michael Pohst. Enumeration of quartic fields of small discriminant. Mathematics of Computation, 61(204):873– 879, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beattie:1992:ICR

[BG92a] Christopher Beattie and W. M. Greenlee. Improved convergence rates for intermediate problems. Mathematics of Computation, 59(199):77–95, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bremner:1992:NCT

[BG92b] Andrew Bremner and Richard K. Guy. Nu-configurations in tiling the square. Mathematics of Computation, 59(199): 195–202, S1–S20, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bremner:1992:SNC

[BG92c] Andrew Bremner and Richard K.

Guy. Supplement to “Nu-configurations in tiling the square”. Mathematics of Computation, 59(199):S1–S20, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bochev:1994:ALS

[BG94] Pavel B. Bochev and Max D. Gunzburger. Analysis of least squares finite element methods for the Stokes equations. Mathematics of Computation, 63 (208):479–506, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brown:1996:SGP

[BG96] Brienne E. Brown and Daniel M.

Gordon. On sequences without geometric progressions. Mathematics of Computation, 65(216):1749– 1754, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00765-
X&u=/mcom/1996-65-216/.

Bendali:1999:NRB

[BG99a] A. Bendali and Ph. Guillaume.

Non-reflecting boundary conditions for waveguides. Mathematics of Computation, 68(225):123–144, January 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01016-
9&u=/mcom/1999-68-225/.

Browkin:1999:TWK

[BG99b] Jerzy Browkin and Herbert Gangl. Tame and wild kernels of quadratic imaginary number fields. Mathematics of Computation, 68(225):291– 305, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01000-
5&u=/mcom/1999-68-225/.

Bohmer:1999:NDS

[BGJ99] Klaus Böhmer, Willy Govaerts, and Vladimír Janovsky. Numerical detection of symmetry breaking bifurcation points with nonlinear degeneracies. Mathematics of Computation, 68(227): 1097–1108, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01052-
2&u=/mcom/1999-68-227/.

Biester:1995:ASQ

[BGLT95] Christian Biester, Peter J. Grabner, Gerhard Larcher, and Robert F. Tichy. Adaptive search in quasi-Monte Carlo optimization. Mathematics of Computation, 64(210):807–818, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bremner:1993:WIR

[BGN93] Andrew Bremner, Richard K. Guy, and Richard J. Nowakowski. Which integers are representable as the product of the sum of three integers with the sum of their reciprocals? Mathematics of Computation, 61(203):117–130, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bach:1993:SES

[BH93] Eric Bach and Lorenz Huelsbergen. Statistical evidence for small generating sets. Mathematics of Computation, 61 (203):69–82, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bartolomeo:1994:FPG

[BH94a] J. Bartolomeo and Matthew He. On Faber polynomials generated by an mstar. Mathematics of Computation, 62 (205):277–287, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brio:1994:NRD

[BH94b] M. Brio and J. K. Hunter. A von Neumann reflection for the 2-D Burgers equation. In Gautschi [Gau94], pages 265–268. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Berczes:1997:CED

[BH97] A. Bérczes and L. Hajdu. Computational experiences on the distances of polynomials to irreducible polynomials. Mathematics of Computation, 66(217): 391–398, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00801-
6&u=/mcom/1997-66-217/.

Buhmann:1993:SDP

[BI93] Martin Buhmann and Arieh Iserles.

Stability of the discretized pantograph differential equation. Mathematics of Computation, 60(202):575–589, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bialecki:1990:SHQ

[Bia90]  Bernard Bialecki.  A Sinc–Hunter

quadrature rule for Cauchy principal value integrals. Mathematics of Computation, 55(192):665–681, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buchmann:1997:SCP

[BJT97] Johannes Buchmann, Michael J. Jacobson, Jr., and Edlyn Teske. On some computational problems in finite abelian groups. Mathematics of Computation, 66(220):1663– 1687, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00880-
                                              6&u=/mcom/1997-66-220/.

Bramble:1994:RFE

[BK94] James H. Bramble and J. Thomas King. A robust finite element method for nonhomogeneous Dirichlet problems in domains with curved boundaries. Mathematics of Computation, 63(207): 1–17, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Borwein:1990:SQE

[BL90] Peter B. Borwein and Terry A. Loring. Some questions of Erdős and Graham on numbers of the form ∑ g_{n}/2^{g_{n}}. Mathematics of Computation, 54(189): 377–394, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barrett:1993:FEA

[BL93] John W. Barrett and W. B. Liu. Finite element approximation of the p-Laplacian. Mathematics of Computation, 61(204):523–537, October 1993.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boukadida:1994:NVT

[BL94] T. Boukadida and A. Y. LeRoux. A new version of the two-dimensional Lax– Friedrichs scheme. Mathematics of Computation, 63(208):541–553, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bales:1995:NNE

[BL95] L. Bales and I. Lasiecka. Negative norm estimates for fully discrete finite element approximations to the wave equation with nonhomogeneous L_{2} Dirichlet boundary data. Mathematics of Computation, 64(209):89–115, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bhatia:1996:PMP

[BL96] Rajendra Bhatia and Ren Cang Li.

On perturbations of matrix pencils with real spectra. II. Mathematics of Computation, 65(214):637– 645, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00699-
0&u=/mcom/1996-65-214/.

Bramble:1993:ISN

[BLP93] James H. Bramble, Zbigniew Leyk, and Joseph E. Pasciak. Iterative schemes for nonsymmetric and indefinite elliptic boundary value problems. Mathematics of Computation, 60(201):1–22, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1994:AMA

[BLP94] James H. Bramble, Zbigniew Leyk, and Joseph E. Pasciak. The analysis of multigrid algorithms for pseudodifferential operators of order minus one. Mathematics of Computation, 63(208):461– 478, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1997:LSA

[BLP97] James H. Bramble, Raytcho D. Lazarov, and Joseph E. Pasciak. A least-squares approach based on a discrete minus one inner product for first order systems. Mathematics of Computation, 66(219):935– 955, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00848-
X&u=/mcom/1997-66-219/.

Berger:1994:FVM

[BLS94] M. J. Berger, R. J. LeVeque, and L. G. Stern. Finite volume methods for irregular one-dimensional grids. In Gautschi [Gau94], pages 255–259. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Bach:1996:REP

[BLSW96] Eric Bach, Richard Lukes, Jeffrey Shallit, and H. C. Williams. Results and estimates on pseudopowers. Mathematics of Computation, 65(216):1737– 1747, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-96-00762-
4&u=/mcom/1996-65-216/.

Berge:1990:CSF

[BMO90] A.-M. Bergé, J. Martinet, and M. Olivier. The computation of sextic fields with a quadratic subfield. Mathematics of Computation, 54(190): 869–884, April 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bojanov:1990:CBT

[BN90] Borislav Bojanov and Geno Nikolov.

Comparison of Birkhoff type quadrature formulae. Mathematics of Computation, 54(190):627–648, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bussemaker:1992:EGS

[BN92] F. C. Bussemaker and A. Neumaier. Exceptional graphs with smallest eigenvalue −2 and related problems. with microfiche supplement. Mathematics of Computation, 59(200):583–608, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Balasubramanian:1997:DCN

[BN97] R. Balasubramanian and S. V. Nagaraj. Density of Carmichael numbers with three prime factors. Mathematics of Computation, 66(220):1705– 1708, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00857-
0&u=/mcom/1997-66-220/.

Babuska:1994:CFE

[BO94] I. Babuška and J. E. Osborn. Corrigendum: “Finite element-Galerkin approximation of the eigenvalues and eigenvectors of selfadjoint problems” [Math. Comp. 52 (1989), no. 186, 275–297, MR 89k:65132]. Mathematics of Computation, 63(208):831–832, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bosma:1990:AM

[Bos90] Wieb Bosma. Approximation by mediants. Mathematics of Computation, 54 (189):421–434, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bosma:1993:EPC

[Bos93a] Wieb Bosma. Explicit primality criteria for h · 2^{k} ± 1. Mathematics of Computation, 61(203):97–109, S7–S9, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bosma:1993:SEP

[Bos93b] Wieb Bosma. Supplement to “Explicit primality criteria for h · 2^{k} ± 1”. Mathematics of Computation, 61(203): S7–S9, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bosma:1998:BRP

[Bos98] Wieb Bosma. Book reviews: Polynomial algorithms in computer algebra, by F. Winkler. Mathematics of Computation, 67(224):??, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00989-
                                              2&u=/mcom/1998-67-224/.

Boumenir:1999:EPS

[Bou99] Amin Boumenir. Eigenvalues of periodic Sturm–Liouville problems by the Shannon–Whittaker sampling theorem. Mathematics of Computation, 68(227): 1057–1066, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01053-
4&u=/mcom/1999-68-227/.

Boyd:1996:BEP

[Boy96a] David W. Boyd. On beta expansions for Pisot numbers. Mathematics of Computation, 65(214):841– 860, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00693-
X&u=/mcom/1996-65-214/.

Boyd:1996:BES

[Boy96b] David W. Boyd. On the beta expansion for Salem numbers of degree 6. Mathematics of Computation, 65(214):861–875, S29–31, April 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Boyd:1996:SBE

[Boy96c] David W. Boyd. Supplement to “On the beta expansion for Salem numbers of degree 6”. Mathematics of Computation, 65(214):S29–S31, April 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boyd:1997:PBC

[Boy97] David W. Boyd. On a problem of Byrnes concerning polynomials with restricted coefficients. Mathematics of Computation, 66(220):1697– 1703, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00892-
                                              2&u=/mcom/1997-66-220/.

Bramble:1992:ASM

[BP92] James H. Bramble and Joseph E. Pasciak. The analysis of smoothers for multigrid algorithms. Mathematics of Computation, 58(198):467–488, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1993:NEM

[BP93] James H. Bramble and Joseph E. Pasciak. New estimates for multilevel algorithms including the V -cycle. Mathematics of Computation, 60(202):447– 471, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bach:1996:ASP

[BP96] Eric Bach and René Peralta. Asymptotic semismoothness probabilities. Mathematics of Computation, 65(216): 1701–1715, October 1996. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                              http://www.ams.org/jourcgi/jour-
                                              pbprocess?fn=110&arg1=S0025-5718-
                                              96-00775-2&u=/mcom/1996-65-216/.

Buchmann:1994:UGC

[BPv94] J. Buchmann, M. Pohst, and J. Graf von Schmettow. On unit groups and class groups of quartic fields of signature (2, 1). Mathematics of Computation, 62(205):387–390, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1998:ANO

[BPV98] James H. Bramble, Joseph E. Pasciak, and Apostol T. Vassilev. Analysis of non-overlapping domain decomposition algorithms with inexact solves. Mathematics of Computation, 67(221): 1–19, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00879-
5&u=/mcom/1998-67-221/.

Brunner:1999:PPC

[BPV99] Hermann Brunner, Arvet Pedas, and Gennadi Vainikko. The piecewise polynomial collocation method for nonlinear weakly singular Volterra equations. Mathematics of Computation, 68(227): 1079–1095, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01073-
X&u=/mcom/1999-68-227/.

Bramble:1991:CEM

[BPWX91a] James H. Bramble, Joseph E. Pasciak, Jun Ping Wang, and Jinchao Xu. Convergence estimates for multigrid algorithms without regularity assumptions. Mathematics of Computation, 57(195): 23–45, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1991:CEP

[BPWX91b] James H. Bramble, Joseph E. Pasciak, Jun Ping Wang, and Jinchao Xu. Convergence estimates for product iterative methods with applications to domain decomposition. Mathematics of Computation, 57(195):1–21, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1990:PMP

[BPX90] James H. Bramble, Joseph E. Pasciak, and Jin Chao Xu. Parallel multilevel preconditioners. Mathematics of Computation, 55(191):1–22, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1991:PMP

[BPX91a] J. H. Bramble, J. E. Pasciak, and J. Xu. Parallel multilevel preconditioners correction: A unifying approach. Mathematics of Computation, ??(??): ??, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1991:AMA

[BPX91b] James H. Bramble, Joseph E. Pasciak, and Jinchao Xu. The analysis of multigrid algorithms with nonnested spaces or noninherited quadratic forms. Mathematics of Computation, 56(193): 1–34, January 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Babai:1990:CIR

[BR90] László Babai and Lajos Rónyai. Computing irreducible representations of finite groups. Mathematics of Computation, 55(192):705–722, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bruno:1994:AAF

[BR94] Oscar P. Bruno and Fernando Reitich. Approximation of analytic functions: a method of enhanced convergence. Mathematics of Computation, 63(207): 195–213, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bjorn:1998:FGF

[BR98] Anders Björn and Hans Riesel. Factors of generalized Fermat numbers. With microfiche supplement. Mathematics of Computation, 67(221):441– 446, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00891-
6&u=/mcom/1998-67-221/. See errata

[BR05, BR11].

Bjorn:2005:TEF

[BR05] Anders Björn and Hans Riesel. Table errata to “Factors of generalized Fermat numbers”. Mathematics of Computation, 74(252):2099, October 2005. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/mcom/2005-74-
252/S0025-5718-05-01816-8/home.
html;   http://www.ams.org/mcom/
2005-74-252/S0025-5718-05-01816-
8/S0025-5718-05-01816-8.dvi;   http:
//www.ams.org/mcom/2005-74-252/
S0025-5718-05-01816-8/S0025-5718-
05-01816-8.pdf;   http://www.ams.
org/mcom/2005-74-252/S0025-5718-
05-01816-8/S0025-5718-05-01816-
8.ps;   http://www.ams.org/mcom/
2005-74-252/S0025-5718-05-01816-
8/S0025-5718-05-01816-8.tex. See

[BR11].

Bjorn:2011:TEF

[BR11] Anders Björn and Hans Riesel. Table errata 2 to “Factors of generalized Fermat numbers”. Mathematics of Computation, 80(275):1865–1866, July 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/journals/
mcom/2011-80-275/S0025-5718-10-
02371-9/home.html;   http://www.

                                              ams.org/journals/mcom/2011-80-275/
                                              S0025-5718-10-02371-9/S0025-5718-
                                              10-02371-9.pdf.

Brass:1993:AOE

[Bra93] H. Brass. Asymptotically optimal error bounds for quadrature rules of given degree. Mathematics of Computation, 61 (204):785–798, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1994:DMM

[Bra94] James H. Bramble. On the development of multigrid methods and their analysis. In Gautschi [Gau94], pages 5– 19. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Bramble:1995:IBS

[Bra95] James H. Bramble. Interpolation between Sobolev spaces in Lipschitz domains with an application to multigrid theory. Mathematics of Computation, 64(212):1359–1365, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Braess:1996:BRH

[Bra96a] Dietrich Braess. Book reviews: Handbook of numerical analysis, edited by P. G. Ciarlet and J. L. Lions. Mathematics of Computation, 65(213): ??, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00666-
                                              7&u=/mcom/1996-65-213/.

Bramble:1996:BRB

[Bra96b] James H. Bramble. Book reviews:

Boundary value problems for partial differential equations and applications, edited by J.-L. Lions and C. Baiocchi. Mathematics of Computation, 65 (213):??, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Brenner:1990:NMM

[Bre90] Susanne C. Brenner. A nonconforming multigrid method for the stationary Stokes equations. Mathematics of Computation, 55(192):411– 437, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1993:CFC

[Bre93] Richard P. Brent. On computing factors of cyclotomic polynomials. Mathematics of Computation, 61(203):131– 149, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brenner:1994:NMM

[Bre94a] Susanne C. Brenner. A nonconforming mixed multigrid method for the pure traction problem in planar linear elasticity. Mathematics of Computation, 63(208):435–460, S1–S5, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brenner:1994:SNM

[Bre94b] Susanne C. Brenner. Supplement to “A nonconforming mixed multigrid method for the pure traction problem in planar linear elasticity”. Mathematics of Computation, 63(208):S1–S5, October 1994.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1994:PGF

[Bre94c] Richard P. Brent. On the periods of generalized Fibonacci recurrences. Mathematics of Computation, 63(207): 389–401, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brent:1994:REC

[Bre94d] Richard P. Brent. Ramanujan and Euler’s constant. In Gautschi [Gau94], pages 541–545. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Brenner:1996:BRF

[Bre96a] Susanne C. Brenner. Book reviews: Finite element methods: Fifty years of the Courant element, edited by M. Krízek, P. Neittaanmäki and R. Stenberg. Mathematics of Computation, 65 (215):??, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00705-
3&u=/mcom/1996-65-215/.

Brenner:1996:TLA

[Bre96b] Susanne C. Brenner. Two-level additive Schwarz preconditioners for nonconforming finite element methods. Mathematics of Computation, 65(215): 897–921, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00746-
6&u=/mcom/1996-65-215/.

Brezzi:1996:BRM

[Bre96c] Franco Brezzi. Book reviews: The mathematical theory of finite element methods, by Susanne C. Brenner and L. Ridgway Scott. Mathematics of Computation, 65(214): ??, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Brent:1997:BRE

[Bre97] Richard P. Brent. Book reviews: Enumeration to 10^{14} of the twin primes and Brun’s constant, by Thomas R. Nicely. Mathematics of Computation, 66(218):??, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00856-
9&u=/mcom/1997-66-218/.

Brenner:1999:CNM

[Bre99a] Susanne C. Brenner. Convergence of nonconforming multigrid methods without full elliptic regularity. Mathematics of Computation, 68(225):25– 53, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01035-
2&u=/mcom/1999-68-225/.

Brenner:1999:MMC

[Bre99b] Susanne C. Brenner. Multigrid methods for the computation of singular solutions and stress intensity factors I: Corner singularities. Mathematics of Computation, 68(226):559– 583, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01017-
                                              0&u=/mcom/1999-68-226/.

Brent:1999:FTF

[Bre99c] Richard P. Brent. Factorization of the tenth Fermat number. Mathematics of Computation, 68(225):429– 451, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-00992-
                                              8&u=/mcom/1999-68-225/.

Briggs:1991:PCF

[Bri91] Keith Briggs. A precise calculation of the Feigenbaum constants. Mathematics of Computation, 57(195):435– 439, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brunner:1994:ICM

[Bru94] Hermann Brunner. Iterated collocation methods for Volterra integral equations with delay arguments. Mathematics of Computation, 62(206):581–599, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bertault:1999:SSC

[BRZ99] F. Bertault, O. Ramaré, and P. Zimmermann. On sums of seven cubes. Mathematics of Computation, 68(227): 1303–1310, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01071-
                                              6&u=/mcom/1999-68-227/.

Batterson:1990:RQI

[BS90a] Steve Batterson and John Smillie. Rayleigh quotient iteration for nonsymmetric matrices. Mathematics of Computation, 55(191):169–178, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baur:1990:BAC

[BS90b] Gerhard Baur and Bruce Shawyer. Best L^{2}-approximation of convergent moment series. Mathematics of Computation, 54(190):661–669, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Banerjee:1992:ENQ

[BS92a] Uday Banerjee and Manil Suri. The effect of numerical quadrature in the pversion of the finite element method. Mathematics of Computation, 59(199): 1–20, July 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brenner:1992:LFE

[BS92b] Susanne C. Brenner and Li-Yeng Y.

Sung. Linear finite element methods for planar linear elasticity. Mathematics of Computation, 59(200):321– 338, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bach:1994:EBP

[BS94a] Eric Bach and Jonathan Sorenson. Explicit bounds for primes in residue classes. In Gautschi [Gau94], pages 535– 539. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Bendali:1994:CED

[BS94b] A. Bendali and M. Souilah. Consistency estimates for a double-layer potential and application to the numerical analysis of the boundary-element approximation of acoustic scattering by a penetrable object. Mathematics of Computation, 62(205):65–91, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bach:1996:EBP

[BS96a] Eric Bach and Jonathan Sorenson. Explicit bounds for primes in residue classes. Mathematics of Computation, 65(216):1717–1735, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00763-
6&u=/mcom/1996-65-216/.

Bosma:1996:DCR

[BS96b] Wieb Bosma and Peter Stevenhagen. Density computations for real quadratic units. Mathematics of Computation, 65(215):1327– 1337, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00725-
9&u=/mcom/1996-65-215/.

Buchmann:1996:CNF

[BS96c] Johannes Buchmann and Victor Shoup. Constructing nonresidues in finite fields and the extended Riemann hypothesis. Mathematics of Computation, 65(215): 1311–1326, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00751-
X&u=/mcom/1996-65-215/.

Bjorhus:1997:WRD

[BS97] Morten Bjørhus and Andrew M.

Stuart. Waveform relaxation as a dynamical system. Mathematics of Computation, 66(219):1101– 1117, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00847-
8&u=/mcom/1997-66-219/.

Bramble:1998:NNL

[BS98] James H. Bramble and Tong Sun. A negative-norm least squares method for Reissner–Mindlin plates. Mathematics of Computation, 67(223):901– 916, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00972-
7&u=/mcom/1998-67-223/.

Blatov:1999:BPO

[BS99] I. A. Blatov and V. V. Strygin. On best possible order of convergence estimates in the collocation method and Galerkin’s method for singularly perturbed boundary value problems for systems of first-order ordinary differential equations. Mathematics of Computation, 68(226):683– 715, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01034-
0&u=/mcom/1999-68-226/.

Buchmann:1990:ACR

[BSW90] J. Buchmann, J. W. Sands, and H. C.

Williams. p-adic computation of real quadratic class numbers. Mathematics of Computation, 54(190):855–868, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bossy:1997:SPM

[BT97] Mireille Bossy and Denis Talay. A stochastic particle method for the McKean–Vlasov and the Burgers equation. Mathematics of Computation, 66(217):157–192, January 1997. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                              http://www.ams.org/jourcgi/jour-
                                              pbprocess?fn=110&arg1=S0025-5718-
                                              97-00776-X&u=/mcom/1997-66-217/.

Bernard:1994:RCS

[BTT94a] Pierre Bernard, Denis Talay, and Luciano Tubaro. Rate of convergence of a stochastic particle method for the Kolmogorov equation with variable coefficients. Mathematics of Computation, 63 (208):555–587, S11–S17, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bernard:1994:SRC

[BTT94b] Pierre Bernard, Denis Talay, and Luciano Tubaro. Supplement to “Rate of convergence of a stochastic particle method for the Kolmogorov equation with variable coefficients”. Mathematics of Computation, 63(208):S11– S17, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bruaset:1990:SRI

[BTW90] A. M. Bruaset, A. Tveito, and R. Winther. On the stability of relaxed incomplete LU factorizations. Mathematics of Computation, 54(190): 701–719, April 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buffoni:1990:NSS

[Buf90] Giuseppe Buffoni. Nonnegative and skew-symmetric perturbations of a matrix with positive inverse. Mathematics of Computation, 54(189):189– 194, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buhmann:1995:MPN

[Buh95] M. D. Buhmann. Multiquadric prewavelets on nonequally spaced knots in one dimension. Mathematics of Computation, 64(212):1611–1625, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bunchaft:1997:SEL

[Bun97] M. E. Froes Bunchaft. Some extensions of the Lanczos–Ortiz theory of canonical polynomials in the Tau Method. Mathematics of Computation, 66(218): 609–621, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00816-
8&u=/mcom/1997-66-218/.

Burthe:1996:FIS

[Bur96]  Ronald Joseph Burthe, Jr.  Fur-

ther investigations with the strong probable prime test. Mathematics of Computation, 65(213):373– 381, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00695-
3&u=/mcom/1996-65-213/.

Butler:1994:ISC

[But94] Greg Butler. An inductive schema for computing conjugacy classes in permutation groups. Mathematics of Computation, 62(205):363–383, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Broeze:1992:RBC

[BV92] Jan Broeze and Edwin F. G. Van Daalen. Radiation boundary conditions for the two-dimensional wave equation from a variational principle. Mathematics of Computation, 58(197):73– 82, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Benito:1999:AAS

[BV99] Manuel Benito and Juan L. Varona. Advances in aliquot sequences. Mathematics of Computation, 68(225):389– 393, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-00991-
6&u=/mcom/1999-68-225/.

Beyer:1996:CLS

[BW96] W. A. Beyer and W. W. Wood. Corrigenda: “The lattice structure of multiplicative congruential pseudo-random vectors” [Math. Comp. 25 (1971), 345– 363, MR 46 #8373] by W. A. Beyerm, R. B. Roof, and Dorothy Williamson. Mathematics of Computation, 65(213): 445–446, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berger:1997:AFB

[BWW97] H. Berger, G. Warnecke, and W. L. Wendland. Analysis of a FEM/ BEM coupling method for transonic flow computations. Mathematics of Computation, 66(220):1407– 1440, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00878-
8&u=/mcom/1997-66-220/.

Bramble:1991:SEW

[BX91] James H. Bramble and Jinchao Xu.

Some estimates for a weighted L^{2} projection. Mathematics of Computation, 56(194):463–476, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ben-yu:1999:EEH

[By99] Guo Ben-yu. Error estimation of Hermite spectral method for nonlinear partial differential equations. Mathematics of Computation, 68(227):1067– 1078, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01059-
5&u=/mcom/1999-68-227/.

Cai:1996:BRL

[Cai96] Xiao-Chuan Cai. Book reviews: LA-PACK Users’ guide, by E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. du Croz, A. Greenbaum, S. Hammarling, A. McKenney, S. Ostrouchov and D. Sorensen. Mathematics of Computation, 65(216): ??, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Calvetti:1991:SRE

[Cal91] Daniela Calvetti. A stochastic roundoff error analysis for the Fast Fourier Transform. Mathematics of Computation, 56(194):755–774, April 1991.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Calvetti:1992:SRE

[Cal92] Daniela Calvetti. A stochastic roundoff error analysis for the convolution. Mathematics of Computation, 59(200): 569–582, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Caldwell:1995:PP

[Cal95] Chris K. Caldwell. On the primality of n! ± 1 and 2 · 3 · 5 ⋯ p ± 1. Mathematics of Computation, 64(210):889–890, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1991:TEI

[Car91] B. C. Carlson. A table of elliptic integrals: One quadratic factor. Mathematics of Computation, 56(193):267– 280, January 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1992:TEI

[Car92] B. C. Carlson. A table of elliptic integrals: Two quadratic factors. Mathematics of Computation, 59(199): 165–180, July 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carasso:1994:SEB

[Car94] Alfred S. Carasso. Slow evolution from the boundary: a new stabilizing constraint in ill-posed continuation problems. In Gautschi [Gau94], pages 269– 273. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Carstensen:1996:EPB

[Car96] Carsten Carstensen. Efficiency of a posteriori BEM–error estimates for first-kind integral equations on quasi–uniform meshes. Mathematics of Computation, 65(213):69–84, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00671-
0&u=/mcom/1996-65-213/.

Carstensen:1997:PEEa

[Car97a]  Carsten Carstensen.  An a pos-

teriori error estimate for a firstkind integral equation. Mathematics of Computation, 66(217):139– 155, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00790-
4&u=/mcom/1997-66-217/.

Carstensen:1997:PEEb

[Car97b]  Carsten Carstensen.  A posteri-

ori error estimate for the mixed finite element method. Mathematics of Computation, 66(218):465– 476, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00837-
5&u=/mcom/1997-66-218/.

Carmack:1998:NES

[Car98] Lori A. Carmack. Numerical evaluation of a symmetric potential function. Mathematics of Computation, 67(222): 641–646, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

                                              fn=110&arg1=S0025-5718-98-00948-
                                              X&u=/mcom/1998-67-222/.

Clausen:1993:FFT

[CB93] Michael Clausen and Ulrich Baum. Fast Fourier transforms for symmetric groups: theory and implementation. Mathematics of Computation, 61(204): 833–847, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cockburn:1994:EEFb

[CCF94] Bernardo Cockburn, Frédéric Coquel, and Philippe Le Floch. An error estimate for finite volume methods for multidimensional conservation laws. Mathematics of Computation, 63(207): 77–103, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1992:SSM

[CD92] Harvey Cohn and Jesse Deutsch. Some singular moduli for Q(\sqrt{3}). Mathematics of Computation, 59(199):231– 247, July 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chui:1996:AOL

[CD96] Charles K. Chui and Johan M. De Villiers. Applications of optimally local interpolation to interpolatory approximants and compactly supported wavelets. Mathematics of Computation, 65(213):99– 114, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00672-
                                              2&u=/mcom/1996-65-213/.

Chou:1997:DBI

[CD97] Chung-Chiang Chou and Yuefan Deng.

Decomposing 40 billion integers by four tetrahedral numbers. Mathematics of Computation, 66(218):893– 901, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00818-
1&u=/mcom/1997-66-218/.

Charnes:1998:TPO

[CD98] C. Charnes and U. Dempwolff. The translation planes of order 49 and their automorphism groups. Mathematics of Computation, 67(223):1207– 1224, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00961-
2&u=/mcom/1998-67-223/.

Cochrane:1999:GBI

[CD99] Todd Cochrane and Robert E. Dressler.

Gaps between integers with the same prime factors. Mathematics of Computation, 68(225):395– 401, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01024-
8&u=/mcom/1999-68-225/.

Carlson:1997:SCM

[CDL97] Timothy S. Carlson, Jack Dockery, and John Lund. A sinc-collocation method for initial value problems. Mathematics of Computation, 66(217):215– 235, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

                                              fn=110&arg1=S0025-5718-97-00789-
                                              8&u=/mcom/1997-66-217/.

Crandall:1995:TSF

[CDNY95] R. Crandall, J. Doenias, C. Norrie, and J. Young. The twenty-second Fermat number is composite. Mathematics of Computation, 64(210):863–868, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1998:CRC

[CDO98] H. Cohen, F. Diaz y Diaz, and M. Olivier. Computing ray class groups, conductors and discriminants. Mathematics of Computation, 67(222):773– 795, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00912-
                                              0&u=/mcom/1998-67-222/.

Cohen:1999:TOF

[CDO99] H. Cohen, F. Diaz y Diaz, and M. Olivier. Tables of octic fields with a quartic subfield. Mathematics of Computation, 68(228):1701– 1716, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01074-
                                              1&u=/mcom/1999-68-228/.

Crandall:1997:SWW

[CDP97] Richard Crandall, Karl Dilcher, and Carl Pomerance. A search for Wieferich and Wilson primes. Mathematics of Computation, 66(217):433– 449, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00791-
                                              6&u=/mcom/1997-66-217/.

Chen:1993:SVA

[CDT93] Gui Qiang Chen, Qiang Du, and Eitan Tadmor. Spectral viscosity approximations to multidimensional scalar conservation laws. Mathematics of Computation, 61(204):629–643, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1996:DDA

[CEL96] Zhangxin Chen, Richard E. Ewing, and Raytcho Lazarov. Domain decomposition algorithms for mixed methods for second-order elliptic problems. Mathematics of Computation, 65(214): 467–490, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00703-
X&u=/mcom/1996-65-214/.

Crandall:1994:DWT

[CF94] Richard Crandall and Barry Fagin. Discrete weighted transforms and largeinteger arithmetic. Mathematics of Computation, 62(205):305–324, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Corrias:1995:NSC

[CFN95a] L. Corrias, M. Falcone, and R. Natalini. Numerical schemes for conservation laws via Hamilton–Jacobi equations. Mathematics of Computation, 64 (210):555–580, S13–18, S13–S18, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Corrias:1995:SNS

[CFN95b] L. Corrias, M. Falcone, and R. Natalini.

Supplement to “Numerical schemes for conservation laws via Hamilton–Jacobi equations”. Mathematics of Computation, 64(210):S13–S18, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cayco:1995:CRC

[CFS95] Maria Cayco, Leslie Foster, and Howard Swann. On the convergence rate of the cell discretization algorithm for solving elliptic problems. Mathematics of Computation, 64(212):1397–1419, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carnicer:1990:EMP

[CG90a] J. Carnicer and M. Gasca. Evaluation of multivariate polynomials and their derivatives. Mathematics of Computation, 54(189):231–243, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cheer:1990:DDE

[CG90b] A. Y. Cheer and D. A. Goldston. A differential delay equation arising from the sieve of Eratosthenes. Mathematics of Computation, 55(191):129–141, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cottet:1993:IPT

[CG93] G.-H. Cottet and L. Germain. Image processing through reaction combined with nonlinear diffusion. Mathematics of Computation, 61(204):659– 673, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cheer:1996:TPP

[CG96a] A. Y. Cheer and D. A. Goldston. Turán’s pure power sum problem. Mathematics of Computation, 65(215): 1349–1358, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00744-
2&u=/mcom/1996-65-215/.

Cockburn:1996:PEE

[CG96b] Bernardo Cockburn and Pierre-Alain Gremaud. A priori error estimates for numerical methods for scalar conservation laws. Part I: The general approach. Mathematics of Computation, 65(214): 533–573, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00701-
6&u=/mcom/1996-65-214/.

Cockburn:1997:PEE

[CG97] Bernardo Cockburn and Pierre-Alain Gremaud. A priori error estimates for numerical methods for scalar conservation laws. Part II: Fluxsplitting monotone schemes on irregular Cartesian grids. Mathematics of Computation, 66(218):547– 572, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00838-
7&u=/mcom/1997-66-218/.

Cohen:1995:MN

[CGH95] Graeme L. Cohen, Stephen F. Gretton, and Peter Hagis, Jr. Multiamicable numbers. Mathematics of Computation, 64(212):1743–1753, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chang:1995:FDM

[CGJ95a] Qian Shun Chang, Bo Ling Guo, and Hong Jiang. Finite difference method for generalized Zakharov equations.

Mathematics of Computation, 64(210): 537–553, S7–S11, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chang:1995:SFD

[CGJ95b] Qian Shun Chang, Bo Ling Guo, and Hong Jiang. Supplement to “Finite difference method for generalized Zakharov equations”. Mathematics of Computation, 64(210):S7–S12, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Conn:1997:GCL

[CGT97a] A. R. Conn, Nick Gould, and Ph. L. Toint. A globally convergent Lagrangian barrier algorithm for optimization with general inequality constraints and simple bounds. Mathematics of Computation, 66(217):261– 288, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00777-
                                              1&u=/mcom/1997-66-217/.

Conn:1997:SGC

[CGT97b] A. R. Conn, Nick Gould, and Ph. L. Toint. Supplement to “A globally convergent Lagrangian barrier algorithm for optimization with general inequality constraints and simple bounds”. Mathematics of Computation, 66(217):S1– S11, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chui:1990:CDB

[CH90a] Charles Chui and Tian Xiao He. Corrigenda: “On the dimension of bivariate superspline spaces” [Math. Comp. 53 (1989), no. 187, 219–234, MR 90c:41014]. Mathematics of Computation, 55(191):407–409, July 1990. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chui:1990:BQF

[CH90b] Charles K. Chui and Tian Xiao He. Bivariate C^{1} quadratic finite elements and vertex splines. Mathematics of Computation, 54(189):169–187, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chui:1996:CLQ

[CH96] Charles K. Chui and Dong Hong. Construction of local C^{1} quartic spline elements for optimal-order approximation. Mathematics of Computation, 65(213): 85–98, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00689-
8&u=/mcom/1996-65-213/.

Ceniceros:1998:CNS

[CH98] Héctor D. Ceniceros and Thomas Y.

Hou. Convergence of a non-stiff boundary integral method for interfacial flows with surface tension. Mathematics of Computation, 67(221):137– 182, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00911-
9&u=/mcom/1998-67-221/.

Chaladus:1992:DLP

[Cha92] Stefan Chaładus. On the densest lattice packing of centrally symmetric octahedra. Mathematics of Computation, 58 (197):341–345, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chambers:1994:QWO

[Cha94a] Ll. G. Chambers. A quick way of obtaining an approximate solution to a Sturm–Liouville problem. Mathematics of Computation, 62(206):577–579, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chang:1994:EEL

[Cha94b] Ching Lung Chang. An error estimate of the least squares finite element method for the Stokes problem in three dimensions. Mathematics of Computation, 63(207):41–50, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chalabi:1997:CNS

[Cha97] Abdallah Chalabi. On convergence of numerical schemes for hyperbolic conservation laws with stiff source terms. Mathematics of Computation, 66(218): 527–545, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00817-
                                              X&u=/mcom/1997-66-218/.

Chalabi:1999:CRS

[Cha99] A. Chalabi. Convergence of relaxation schemes for hyperbolic conservation laws with stiff source terms. Mathematics of Computation, 68(227): 955–970, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01089-
                                              3&u=/mcom/1999-68-227/.

Chern:1991:LTB

[Che91] I-Liang Chern. Large-time behavior of solutions of Lax–Friedrichs finite difference equations for hyperbolic systems of conservation laws. Mathematics of Computation, 56(193):107– 118, January 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1994:AZQ

[Che94a] Pengyuan Chen. Approximate zeros of quadratically convergent algorithms. Mathematics of Computation, 63(207): 247–270, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1994:FEM

[Che94b] Zhangxin Chen. A finite element method for the 2D drift-diffusion semiconductor model. In Gautschi [Gau94], pages 275–279. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Cheney:1994:IIP

[Che94c] Margaret Cheney. An introduction to inverse problems. In Gautschi [Gau94], pages 21–32. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Cheney:1996:BRM

[Che96a] E. W. Cheney. Book reviews: Mathematical aspects of geometrical modeling, by Charles A. Micchelli. Mathematics of Computation, 65(214):889– 890, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

                                              fn=110&arg1=S0025-5718-96-00700-
                                              4&u=/mcom/1996-65-214/.

Cheney:1996:BRN

[Che96b] E. W. Cheney. Book reviews: NURB curves and surfaces: from projective geometry to practical use, by Gerald E. Farin. Mathematics of Computation, 65(214):891– 892, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00700-
                                              4&u=/mcom/1996-65-214/.

Cheney:1996:BRT

[Che96c] E. W. Cheney. Book reviews: Topics in polynomials: extremal problems, inequalities, zeros, by G. V. Milovanovic, D. S. Mitrinovic and Th. M. Rassias. Mathematics of Computation, 65(213): ??, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00666-
                                              7&u=/mcom/1996-65-213/.

Cheney:1996:BRW

[Che96d] E. W. Cheney. Book reviews: Wavelets: theory, algorithms, and applications, edited by Charles K. Chui, Laura Montefusco and Luigia Puccio. Mathematics of Computation, 65(214):890– 891, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00700-
                                              4&u=/mcom/1996-65-214/.

Cheney:1997:BRI

[Che97] M. Cheney. Book reviews: Inverse problems in diffusion processes, edited by Heinz W. Engl and William Rundell. Mathematics of Computation, 66 (217):??, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00793-
X&u=/mcom/1997-66-217/.

Chien:1995:NES

[Chi95] David Chien. Numerical evaluation of surface integrals in three dimensions. Mathematics of Computation, 64(210): 727–743, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Caflisch:1999:AOC

[CHL99] Russel E. Caflisch, Thomas Y. Hou, and John Lowengrub. Almost optimal convergence of the point vortex method for vortex sheets using numerical filtering. Mathematics of Computation, 68(228):1465– 1496, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01108-
4&u=/mcom/1999-68-228/.

Chou:1997:ACC

[Cho97] S. H. Chou. Analysis and convergence of a covolume method for the generalized Stokes problem. Mathematics of Computation, 66(217):85– 104, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00792-
8&u=/mcom/1997-66-217/.

Cockburn:1990:RKL

[CHS90] Bernardo Cockburn, Suchung Hou, and Chi-Wang Shu. The Runge–Kutta local projection discontinuous Galerkin finite element method for conservation laws. IV. the multidimensional case. Mathematics of Computation, 54(190): 545–581, April 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Coutsias:1996:ESM

[CHT96] Evangelos A. Coutsias, Thomas Hagstrom, and David Torres. An efficient spectral method for ordinary differential equations with rational function coefficients. Mathematics of Computation, 65(214):611– 635, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00704-
1&u=/mcom/1996-65-214/.

Chui:1996:BRA

[Chu96] Charles Chui. Book reviews: Adapted wavelet analysis from theory to software, by Mladen Victor Wickerhauser. Mathematics of Computation, 65(213): ??, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Callahan:1999:CCH

[CHW99] Patrick J. Callahan, Martin V. Hildebrand, and Jeffrey R. Weeks. A census of cusped hyperbolic 3-manifolds. Mathematics of Computation, 68(225): 321–332, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01036-
4&u=/mcom/1999-68-225/. With mi-

crofiche supplement.

Cikanek:1994:SEW

[Cik94] Petr Cikánek. A special extension of Wieferich’s criterion. Mathematics of Computation, 62(206):923–930, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Calvo:1997:NSI

[CIZ97] Mari Paz Calvo, Arieh Iserles, and Antonella Zanna. Numerical solution of isospectral flows. Mathematics of Computation, 66(220):1461– 1486, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00902-
2&u=/mcom/1997-66-220/.

Cockburn:1997:ENA

[CJT97] Bernardo Cockburn, Don A. Jones, and Edriss S. Titi. Estimating the number of asymptotic degrees of freedom for nonlinear dissipative systems. Mathematics of Computation, 66(219): 1073–1087, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00850-
8&u=/mcom/1997-66-219/.

Cvijovic:1995:CFS

[CK95] Djurdje Cvijović and Jacek Klinowski. Closed-form summation of some trigonometric series. Mathematics of Computation, 64(209):205– 210, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cvijovic:1999:VLC

[CK99] Djurdje Cvijovic and Jacek Klinowski. Values of the Legendre chi and Hurwitz zeta functions at rational arguments. Mathematics of Computation, 68(228):1623–1630, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01091-
1&u=/mcom/1999-68-228/.

Chou:1990:CNW

[CL90] S. H. Chou and Q. Li. Convergence of the nonconforming Wilson element for a class of nonlinear parabolic problems. Mathematics of Computation, 54(190): 509–524, April 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chou:1991:MFE

[CL91a] So-Hsiang Chou and Qian Li. Mixed finite element methods for compressible miscible displacement in porous media. Mathematics of Computation, 57(196): 507–527, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Collins:1991:OOE

[CL91b] Charles Collins and Mitchell Luskin. Optimal-order error estimates for the finite element approximation of the solution of a nonconvex variational problem. Mathematics of Computation, 57 (196):621–637, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Coquel:1991:CFD

[CL91c] Frédéric Coquel and Philippe Le Floch. Convergence of finite difference schemes for conservation laws in several space dimensions: the corrected antidiffusive flux approach. Mathematics of Computation, 57(195):169–210, July 1991.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1993:CSO

[CL93] Gui-Qiang Chen and Jian-Guo Liu.

Convergence of second-order schemes for isentropic gas dynamics. Mathematics of Computation, 61(204):607– 627, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Couture:1994:LSC

[CL94] Raymond Couture and Pierre L’Ecuyer.

On the lattice structure of certain linear congruential sequences related to AWC/SWB generators. Mathematics of Computation, 62(206):799–808, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Couture:1996:OLL

[CL96] Raymond Couture and Pierre L’Ecuyer.

Orbits and lattices for linear random number generators with composite moduli. Mathematics of Computation, 65(213):189–201, January 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00673-
4&u=/mcom/1996-65-213/;   http:/
/www.jstor.org/stable/2153839.

Chen:1997:CDS

[CL97a] Gui-Qiang Chen and Jian-Guo Liu.

Convergence of difference schemes with high resolution for conservation laws. Mathematics of Computation, 66(219): 1027–1053, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-97-00859-
4&u=/mcom/1997-66-219/.

Couture:1997:DPM

[CL97b] Raymond Couture and Pierre L’Ecuyer. Distribution properties of multiplywith-carry random number generators. Mathematics of Computation, 66(218): 591–607, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00827-
2&u=/mcom/1997-66-218/.

Coulibaly:1999:QRR

[CL99] Ibrahim Coulibaly and Christian Lécot. A quasi-randomized Runge–Kutta method. Mathematics of Computation, 68(226):651–659, April 1999. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
99-01056-X&u=/mcom/1999-68-226/.

Clark:1996:NGC

[Cla96]  David A. Clark.  Non-Galois cu-

bic fields which are Euclidean but not norm-Euclidean. Mathematics of Computation, 65(216):1675– 1679, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00764-
8&u=/mcom/1996-65-216/.

Chui:1994:NNL

[CLM94] C. K. Chui, Xin Li, and H. N. Mhaskar. Neural networks for localized approximation. Mathematics of Computation, 63(208):607–623, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chakraborty:1997:MFW

[CLR97] K. Chakraborty, A. K. Lal, and B. Ramakrishnan. Modular forms which behave like theta series. Mathematics of Computation, 66(219):1169– 1183, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00872-
7&u=/mcom/1997-66-219/.

Couture:1993:DDV

[CLT93a] Raymond Couture, Pierre L’Ecuyer, and Shu Tezuka. On the distribution of k-dimensional vectors for simple and combined Tausworthe sequences. Mathematics of Computation, 60(202):749– 761, S11–S16, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Couture:1993:SDK

[CLT93b] Raymond Couture, Pierre L’Ecuyer, and Shu Tezuka. Supplement to “On the distribution of k-dimensional vectors for simple and combined Tausworthe sequences”. Mathematics of Computation, 60(202):S11–S16, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crouzeix:1994:REE

[CLT94] M. Crouzeix, S. Larsson, and V. Thomée.

Resolvent estimates for elliptic finite element operators in one dimension. Mathematics of Computation, 63(207): 121–140, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Casperson:1994:SFS

[CM94a] David Casperson and John McKay.

Symmetric functions, m-sets, and Galois groups. Mathematics of Computation, 63(208):749–757, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1994:HCG

[CM94b] Henri Cohen and Jacques Martinet. Heuristics on class groups: some good primes are not too good. Mathematics of Computation, 63(207):329–334, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Coleman:1995:FPA

[CM95] John P. Coleman and Nick J. Myers. The Faber polynomials for annular sectors. Mathematics of Computation, 64 (209):181–203, S1–S6, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1996:SGM

[CM96] Harvey Cohn and John McKay. Spontaneous generation of modular invariants. Mathematics of Computation, 65(215): 1295–1309, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00726-
                                              0&u=/mcom/1996-65-215/.

Criscuolo:1990:CPC

[CMM90] Giuliana Criscuolo, Giuseppe Mastroianni, and Giovanni Monegato. Convergence properties of a class of product formulas for weakly singular integral equations. Mathematics of Computation, 55(191):213–230, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Colbourn:1992:STSa

[CMM92a] Charles J. Colbourn, Spyros S. Magliveras, and Rudolf A. Mathon. Supplement to “Transitive Steiner and Kirkman triple systems” by Charles J. Colbourn, Spyros S. Magliveras, and Rudolf A. Mathon. Mathematics of Computation, 58(197):S23– S27, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Colbourn:1992:TSK

[CMM92b] Charles J. Colbourn, Spyros S. Magliveras, and Rudolf A. Mathon. Transitive Steiner and Kirkman triple systems of order 27. Mathematics of Computation, 58(197):441–449, S23– S27, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Criscuolo:1990:CEL

[CMO90] Giuliana Criscuolo, Giuseppe Mastroianni, and Donatella Occorsio. Convergence of extended Lagrange interpolation. Mathematics of Computation, 55(191):197–212, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Castel:1998:CNS

[CMP98] M. Jesús Castel, Violeta Migallón, and José Penadés. Convergence of non-stationary parallel multisplitting methods for Hermitian positive definite matrices. Mathematics of Computation, 67(221):209– 220, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00893-
X&u=/mcom/1998-67-221/.

Colbourn:1992:STSb

[CMS92a] Charles J. Colbourn, Spyros S. Magliveras, and D. R. Stinson. Steiner triple systems of order 19 with nontrivial automorphism group. Mathematics of Computation, 59(199):283–295, S25–S27, July 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Colbourn:1992:SST

[CMS92b] Charles J. Colbourn, Spyros S. Magliveras, and D. R. Stinson. Supplement to “Steiner triple systems of order 19 with nontrivial automorphism group”. Mathematics of Computation, 59(199): S25–S27, July 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Criscuolo:1992:PSC

[CMV92] Giuliana Criscuolo, Giuseppe Mastroianni, and Péter Vértesi. Pointwise simultaneous convergence of extended Lagrange interpolation with additional knots. Mathematics of Computation, 59 (200):515–531, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cowsar:1995:BDD

[CMW95] Lawrence C. Cowsar, Jan Mandel, and Mary F. Wheeler. Balancing domain decomposition for mixed finite elements. Mathematics of Computation, 64(211): 989–1015, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1999:CIW

[CMX99] Zhongying Chen, Charles A. Micchelli, and Yuesheng Xu. A construction of interpolating wavelets on invariant sets. Mathematics of Computation, 68(228):1569–1587, October 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01110-
2&u=/mcom/1999-68-228/.

Chen:1998:MMM

[CO98] Zhangxin Chen and Peter Oswald.

Multigrid and multilevel methods for nonconforming Q_{1} elements. Mathematics of Computation, 67(222):667– 693, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00920-
X&u=/mcom/1998-67-222/.

Cockburn:1991:CPF

[Coc91] Bernardo Cockburn. On the continuity in BV (Ω) of the L^{2}-projection into finite element spaces. Mathematics of Computation, 57(196):551– 561, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1990:IID

[Coh90]  Graeme L. Cohen.  On an inte-

ger’s infinitary divisors. Mathematics of Computation, 54(189):395–411, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1993:HBP

[Coh93] Harvey Cohn. How branching properties determine modular equations. Mathematics of Computation, 61(203): 155–170, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1995:HSM

[Coh95] Harvey Cohn. Half-step modular equations. Mathematics of Computation, 64 (211):1267–1285, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1996:HSN

[Coh96] Henri Cohen. Hermite and Smith normal form algorithms over Dedekind domains. Mathematics of Computation, 65(216):1681–1699, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00766-
1&u=/mcom/1996-65-216/.

Cohen:1997:NWP

[Coh97a] G. L. Cohen. Numbers whose positive divisors have small integral harmonic mean. Mathematics of Computation, 66(218):883– 891, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00819-
3&u=/mcom/1997-66-218/.

Cohn:1997:BRQ

[Coh97b] Harvey Cohn. Book reviews: Quadratics, by Richard A Mollin. Mathematics of Computation, 66(218): ??, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00856-
9&u=/mcom/1997-66-218/.

Cohn:1997:DE

[Coh97c] J. H. E. Cohn. The Diophantine equation x^{4} + 1 = Dy^{2}. Mathematics of Computation, 66(219):1347– 1351, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00851-
X&u=/mcom/1997-66-219/.

Cohn:1998:CEP

[Coh98] J. H. E. Cohn. A conjecture of Erdős on 3-powerful numbers. Mathematics of Computation, 67(221):439– 440, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00881-
3&u=/mcom/1998-67-221/.

Coleman:1997:RDT

[Col97] Thomas F. Coleman. Reviews and descriptions of tables and books: 26. Numerical methods for least squares problems, by Åke Björck. Mathematics of Computation, 66(220): ??, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00843-
0&u=/mcom/1997-66-220/.

Cooper:1991:ISI

[Coo91] G. J. Cooper. On the implementation of singly implicit Runge–Kutta methods. Mathematics of Computation, 57(196): 663–672, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cooper:1998:BRC

[Coo98] Jeffery Cooper. Book reviews: Computational differential equations, by K. Eriksson, D. Estep, P. Hansbo, and C. Johnson. Mathematics of Computation, 67(224):??, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00989-
                                              2&u=/mcom/1998-67-224/.

Coppersmith:1990:FLT

[Cop90] Don Coppersmith. Fermat’s last theorem (case 1) and the Wieferich criterion. Mathematics of Computation, 54(190): 895–902, April 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Coppersmith:1994:SHL

[Cop94] Don Coppersmith. Solving homogeneous linear equations over gf(2) via block Wiedemann algorithm. Mathematics of Computation, 62(205):333– 350, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Copetti:1999:FEA

[Cop99] M. I. M. Copetti. Finite element approximation to a contact problem in linear thermoelasticity. Mathematics of Computation, 68(227):1013– 1024, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01054-
                                              6&u=/mcom/1999-68-227/.

Costello:1991:APF

[Cos91] Patrick Costello. Amicable pairs of the form (i, 1). Mathematics of Computation, 56(194):859–865, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cottet:1991:LTB

[Cot91] Georges-Henri Cottet. Large-time behavior of deterministic particle approximations to the Navier–Stokes equations. Mathematics of Computation, 56 (193):45–59, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carstensen:1997:NSS

[CP97] Carsten Carstensen and Petr Plecháč.

Numerical solution of the scalar doublewell problem allowing microstructure. Mathematics of Computation, 66(219): 997–1026, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00849-
1&u=/mcom/1997-66-219/.

Costabel:1992:EAB

[CPS92] M. Costabel, F. Penzel, and R. Schneider. Error analysis of a boundary element collocation method for a screen problem in ℝ^{3}. Mathematics of Computation, 58(198):575–586, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1998:GSC

[CQS98] X. Chen, L. Qi, and D. Sun. Global and superlinear convergence of the smoothing Newton method and its application to general box constrained variational inequalities. Mathematics of Computation, 67(222):519– 540, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00932-
6&u=/mcom/1998-67-222/.

Crandall:1998:FEM

[Cra98] Richard E. Crandall. Fast evaluation of multiple zeta sums. Mathematics of Computation, 67(223):1163– 1172, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00950-
                                              8&u=/mcom/1998-67-223/.

Cremona:1995:CDM

[Cre95] J. E. Cremona. Computing the degree of the modular parametrization of a modular elliptic curve. Mathematics of Computation, 64(211):1235–1250, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crow:1993:QIL

[Cro93] John A. Crow. Quadrature of integrands with a logarithmic singularity. Mathematics of Computation, 60(201): 297–301, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Crisci:1992:SRO

[CRV92] M. R. Crisci, E. Russo, and A. Vecchio. Stability results for one-step discretized collocation methods in the numerical treatment of Volterra integral equations. Mathematics of Computation, 58(197):119–134, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carstensen:1995:PEE

[CS95] Carsten Carstensen and Ernst P. Stephan. A posteriori error estimates for boundary element methods. Mathematics of Computation, 64(210):483– 500, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Call:1996:CCH

[CS96a] Gregory S. Call and Joseph H. Silverman. Computing the canonical height on K3 surfaces. Mathematics of Computation, 65(213):259– 290, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00680-
1&u=/mcom/1996-65-213/.

Cools:1996:MCF

[CS96b] Ronald Cools and Ian H. Sloan. Minimal cubature formulae of trigonometric degree. Mathematics of Computation, 65(216):1583–1600, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00767-
3&u=/mcom/1996-65-216/.

Cremona:1999:CRE

[CS99] J. E. Cremona and P. Serf. Computing the rank of elliptic curves over real quadratic number fields of class number 1. Mathematics of Computation, 68(227):1187– 1200, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01055-
8&u=/mcom/1999-68-227/.

Csordas:1994:LPZ

[CSV94] George Csordas, Wayne Smith [Wayne Stewart Smith], and Richard S. Varga. Lehmer pairs of zeros and the Riemann ξ-function. In Gautschi [Gau94], pages 553–556. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Chui:1992:FSA

[CSW92] C. K. Chui, J. Stöckler, and J. D. Ward. A Faber series approach to cardinal interpolation. Mathematics of Computation, 58(197):255–273, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cockburn:1992:CFE

[CT92a] Bernardo Cockburn and Ioana Triandaf. Convergence of a finite element method for the drift-diffusion semiconductor device equations: the zero diffusion case. Mathematics of Computation, 59 (200):383–401, S29–S46, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cockburn:1992:SCF

[CT92b] Bernardo Cockburn and Ioana Triandaf. Supplement to “Convergence of a finite element method for the driftdiffusion semiconductor device equations: the zero diffusion case”. Mathematics of Computation, 59(200):S29– S46, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cockburn:1994:EEFa

[CT94] Bernardo Cockburn and Ioana Triandaf. Error estimates for a finite element method for the drift-diffusion semiconductor device equations: the zero diffusion case. Mathematics of Computation, 63(207):51–76, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1998:AP

[Ct98] Graeme L. Cohen and Herman J. J. te Riele. On ϕ-amicable pairs. Mathematics of Computation, 67(221):399– 411, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00933-
8&u=/mcom/1998-67-221/.

Chen:1992:FEA

[CTW92] C. Chen, V. Thomée, and L. B.

Wahlbin. Finite element approximation of a parabolic integro-differential equation with a weakly singular kernel. Mathematics of Computation, 58(198): 587–602, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cusick:1993:ZCS

[Cus93] T. W. Cusick. Zaremba’s conjecture and sums of the divisor function. Mathematics of Computation, 61(203): 171–176, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chan:1995:FBI

[CV95a] Tony F. Chan and Panayot S. Vassilevski. A framework for block ILU factorizations using block-size reduction. Mathematics of Computation, 64(209): 129–156, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cunha:1995:IMN

[CV95b] Cristina Cunha and Fermin Viloche.

An iterative method for the numerical inversion of Laplace transforms. Mathematics of Computation, 64(211): 1193–1198, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chou:1999:GMC

[CV99] So-Hsiang Chou and Panayot S. Vassilevski. A general mixed covolume framework for constructing conservative schemes for elliptic problems. Mathematics of Computation, 68(227):991– 1011, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01090-
                                              X&u=/mcom/1999-68-227/.

Colquitt:1991:NMP

[CW91] W. N. Colquitt and L. Welsh, Jr. A new Mersenne prime. Mathematics of Computation, 56(194):867–870, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cremona:1994:PCF

[CW94] J. E. Cremona and E. Whitley. Periods of cusp forms and elliptic curves over imaginary quadratic fields. Mathematics of Computation, 62(205):407– 429, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chan:1992:CPT

[CY92] Raymond H. Chan and Man-Chung Yeung. Circulant preconditioners for Toeplitz matrices with positive continuous generating functions. Mathematics of Computation, 58(197):233– 240, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Darmon:1991:NPE

[Dar91] Henri Darmon. Note on a polynomial of Emma Lehmer. Mathematics of Computation, 56(194):795–800, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Davidon:1996:BRO

[Dav96]  William C. Davidon.  Book re-

views: Optimization and nonstandard analysis, by J. E. Rubio. Mathematics of Computation, 65(213): ??, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

deBoor:1998:BRI

[dB98] Carl de Boor. Book reviews: An introduction to functional analysis and computational mathematics, by V. I. Lebedev. Mathematics of Computation, 67 (221):??, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00891-
6&u=/mcom/1998-67-221/.

Deng:1996:BEM

[DCNZ96] Yuanhua Deng, Goong Chen, Wei-Ming Ni, and Jianxin Zhou. Boundary element monotone iteration scheme for semilinear elliptic partial differential equations. Mathematics of Computation, 65(215):943– 982, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00743-
0&u=/mcom/1996-65-215/.

Delmer:1990:CWP

[DD90] Francine Delmer and Jean-Marc Deshouillers. On the computation of g(k) in Waring’s problem. Mathematics of Computation, 54(190):885–893, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dawson:1992:EIC

[DD92] Clint N. Dawson and Todd F. Dupont. Explicit/implicit conservative Galerkin domain decomposition procedures for parabolic problems. Mathematics of Computation, 58(197):21–34, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Deshouillers:1993:NRS

[DD93] J.-M. Deshouillers and F. Dress. Numerical results for sums of five and seven biquadrates and consequences for sums of 19 biquadrates. Mathematics of Computation, 61(203):195–207, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dawson:1991:FDD

[DDD91] Clint N. Dawson, Qiang Du, and Todd F. Dupont. A finite difference domain decomposition algorithm for numerical solution of the heat equation. Mathematics of Computation, 57(195): 63–71, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dari:1995:EEN

[DDP95] Enzo Dari, Ricardo Durán, and Claudio Padra. Error estimators for nonconforming finite element approximations of the Stokes problem. Mathematics of Computation, 64(211):1017– 1033, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DeWeger:1995:TEQ

[De 95] B. M. M. De Weger. A Thue equation with quadratic integers as variables. Mathematics of Computation, 64(210):855–861, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBoor:1996:SXF

[de 96] Carl de Boor. On the Sauer–Xu formula for the error in multivariate polynomial interpolation. Mathematics of Computation, 65(215):1231– 1234, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00727-
2&u=/mcom/1996-65-215/.

Deak:1990:RNG

[Deá90] István Deák. Random number generators and simulation, volume 4 of Mathematical methods of operations research. Akadémiai Kiadó, Budapest, Hungary, 1990. ISBN 963-05-5316-3. 341 pp. LCCN QA298 .D4313 1990.

Delvos:1990:BMD

[Del90] Franz-J. Delvos. Boolean methods for double integration. Mathematics of Computation, 55(192):683–692, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DelCorso:1992:FPI

[Del92a] Ilaria Del Corso. Factorization of prime ideal extensions in number rings. Mathematics of Computation, 58(198): 849–853, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Delorme:1992:DE

[Del92b] Jean-Joël Delorme. On the Diophantine equation x_{1}^{6} + x_{2}^{6} + x_{3}^{6} = y_{1}^{6} + y_{2}^{6} + y_{3}^{6}. Mathematics of Computation, 59(200): 703–715, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Demetriou:1990:CTD

[Dem90] I. C. Demetriou. A characterization theorem for the discrete best monotonic approximation problem. Mathematics of Computation, 55(191):191–195, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Demetriou:1995:DPM

[Dem95] I. C. Demetriou. Discrete piecewise monotonic approximation by a strictly convex distance function. Mathematics of Computation, 64(209):157– 180, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Demichelis:1996:QIS

[Dem96] V. Demichelis. Quasi-interpolatory splines based on Schoenberg points. Mathematics of Computation, 65(215): 1235–1247, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00728-
4&u=/mcom/1996-65-215/.

Denert:1990:GZF

[Den90] M. Denert. The genus zeta function of hereditary orders in central simple algebras over global fields. Mathematics of Computation, 54(189):449– 465, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dewaghe:1998:RSE

[Dew98] L. Dewaghe. Remarks on the Schoof– Elkies–Atkin algorithm. Mathematics of Computation, 67(223):1247– 1252, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00962-
4&u=/mcom/1998-67-223/.

Deshouillers:1993:UBG

[DGNP93] Jean-Marc Deshouillers, Andrew Granville, Władysław Narkiewicz, and Carl Pomerance. An upper bound in Goldbach’s problem. Mathematics of Computation, 61(203):209–213, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiNatale:1992:CPM

[DGPR92] M. Di Natale, L. Gotusso, R. Pavani, and D. Roux. On the choice of parameter in a method for the inversion of Fourier series. Mathematics of Computation, 58(198):737–744, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duijvestijn:1990:SNP

[DH90] A. J. W. Duijvestijn and B. Helthuis.

Some new polyhedra with vertex degree 4 and/or 5 only. Mathematics of Computation, 54(190):749–753, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Doll:1991:TOL

[DH91] Helmut Doll and Jim Hoste. A tabulation of oriented links. with microfiche supplement. Mathematics of Computation, 57(196):747–761, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dummit:1994:ROD

[DH94] D. S. Dummit and David Hayes. Rankone Drinfeld́ modules on elliptic curves. With microfiche supplement. Mathematics of Computation, 62(206):875– 883, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dziuk:1999:DPPa

[DH99a] Gerhard Dziuk and John E. Hutchinson. The discrete Plateau Problem: Algorithm and numerics. Mathematics of Computation, 68(225):1– 23, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01025-
X&u=/mcom/1999-68-225/.

Dziuk:1999:DPPb

[DH99b] Gerhard Dziuk and John E. Hutchinson. The discrete plateau problem: Convergence results. Mathematics of Computation, 68(226):519– 546, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01026-
1&u=/mcom/1999-68-226/.

Duran:1999:AVM

[DHNL^{+}99] R. G. Durán, L. Hervella-Nieto, E. Liberman, R. Rodríguez, and J. Solomin. Approximation of the vibration modes of a plate by Reissner– Mindlin equations. Mathematics of Computation, 68(228):1447–1463, October 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01094-
7&u=/mcom/1999-68-228/.

DiazyDiaz:1991:STT

[Dia91a] F. Diaz y Diaz. Supplement to A table of totally real quintic number fields. Mathematics of Computation, 56(194):

S1–S12, April 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiazyDiaz:1991:TTR

[Dia91b] F. Diaz y Diaz. A table of totally real quintic number fields. Mathematics of Computation, 56(194):801–808, S1–S12, April 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Diethelm:1994:EEQ

[Die94] Kai Diethelm. Error estimates for a quadrature rule for Cauchy principal value integrals. In Gautschi [Gau94], pages 287–291. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Dilcher:1992:GEC

[Dil92a] Karl Dilcher. Generalized Euler constants for arithmetical progressions. Mathematics of Computation, 59(199): 259–282, S21–S24, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dilcher:1992:SGE

[Dil92b] Karl Dilcher. Supplement to “Generalized Euler constants for arithmetical progressions”. Mathematics of Computation, 59(199):S21–S24, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dimitrov:1996:IPF

[Dim96] Dimitar K. Dimitrov. Integration of polyharmonic functions. Mathematics of Computation, 65(215):1269– 1281, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00747-
8&u=/mcom/1996-65-215/.

Dougherty:1991:CRC

[DJ91] Randall Dougherty and Heeralal Janwa. Covering radius computations for binary cyclic codes. with microfiche supplement. Mathematics of Computation, 57(195):415–434, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Du:1997:QLI

[DJLZ97] Qiang Du, Ming Jin, T. Y. Li, and Z. Zeng. The quasi-Laguerre iteration. Mathematics of Computation, 66(217): 345–361, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00786-
2&u=/mcom/1997-66-217/.

Dubner:1995:FGF

[DK95] Harvey Dubner and Wilfrid Keller. Factors of generalized Fermat numbers. Mathematics of Computation, 64(209): 397–405, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dubner:1999:NFL

[DK99] Harvey Dubner and Wilfrid Keller. New Fibonacci and Lucas primes. Mathematics of Computation, 68(225):417– 427, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-00981-
3&u=/mcom/1999-68-225/.

Duran:1992:MFE

[DL92] Ricardo Durán and Elsa Liberman. On mixed finite element methods for the Reissner–Mindlin plate model. Mathematics of Computation, 58(198):561– 573, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dempsey:1990:PSF

[DLD90] Kevin M. Dempsey, Dajin Liu, and John P. Dempsey. Plana’s summation formula for ∑_{m=1,3,...}^{∞} m^{−2} sin(mα), m^{−3} cos(mα), m^{−2}A^{m}, m^{−3}A^{m}. Mathematics of Computation, 55(192): 693–703, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Damgaard:1993:ACE

[DLP93] Ivan Damgård, Peter Landrock, and Carl Pomerance. Average case error estimates for the strong probable prime test. Mathematics of Computation, 61 (203):177–194, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Deo:1998:ACB

[DM98] Satya Deo and Lipika Mazumdar.

An algorithm for constructing a basis for C^{r}-spline modules over polynomial rings. Mathematics of Computation, 67(223):1107–1120, July 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00943-
0&u=/mcom/1998-67-223/.

Destuynder:1999:EEB

[DM99] Philippe Destuynder and Brigitte Métivet. Explicit error bounds in a conforming finite element method. Mathematics of Computation, 68(228):1379– 1396, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01093-
5&u=/mcom/1999-68-228/.

Dahmen:1992:BBS

[DMS92] Wolfgang Dahmen, Charles A. Micchelli, and Hans-Peter Seidel. Blossoming begets B-spline bases built better by B-patches. Mathematics of Computation, 59(199):97–115, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Devulder:1993:RCN

[DMT93] Christophe Devulder, Martine Marion, and Edriss S. Titi. On the rate of convergence of the nonlinear Galerkin methods. Mathematics of Computation, 60(202):495–514, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duran:1990:WIS

[DN90] Ricardo G. Durán and Ricardo H. Nochetto. Weighted inf-sup condition and pointwise error estimates for the Stokes problem. Mathematics of Computation, 54(189):63–79, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dubner:1997:SCP

[DN97] Harvey Dubner and Harry Nelson. Seven consecutive primes in arithmetic progression. Mathematics of Computation, 66(220):1743–1749, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-97-00875-
2&u=/mcom/1997-66-220/.

DiazyDiaz:1995:IND

[DO95] F. Diaz y Diaz and M. Olivier. Imprimitive ninth-degree number fields with small discriminants. with microfiche supplement. Mathematics of Computation, 64(209):305–321, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dolzmann:1999:OCF

[Dol99] Georg Dolzmann. Optimal convergence for the finite element method in Campanato spaces. Mathematics of Computation, 68(228):1397– 1427, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01175-
8&u=/mcom/1999-68-228/.

Doman:1996:AEI

[Dom96] B. G. S. Doman. An asymptotic expansion for the incomplete beta function. Mathematics of Computation, 65(215): 1283–1288, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00729-
6&u=/mcom/1996-65-215/.

Dorfler:1992:HBE

[Dör92a] W. Dörfler. Hierarchical bases for elliptic problems. Mathematics of Computation, 58(198):513–529, S29–S32, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dorfler:1992:SHB

[Dör92b] W. Dörfler. Supplement to “Hierarchical bases for elliptic problems”.

Mathematics of Computation, 58(198): S29–S32, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deFrutos:1991:PMG

[DOSS91] J. De Frutos, T. Ortega, and J. M. Sanz-Serna. Pseudospectral method for the “good” Boussinesq equation. Mathematics of Computation, 57(195): 109–122, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Daberkow:1996:IBR

[DP96] M. Daberkow and M. Pohst. On integral bases in relative quadratic extensions. Mathematics of Computation, 65(213): 319–329, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00686-
2&u=/mcom/1996-65-213/.

deBoor:1992:CAP

[dR92] Carl de Boor and Amos Ron. Computational aspects of polynomial interpolation in several variables. Mathematics of Computation, 58(198):705–727, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Deleglise:1996:CML

[DR96] M. Deléglise and J. Rivat. Computing π(x): the Meissel, Lehmer, Lagarias, Miller, Odlyzko method. Mathematics of Computation, 65(213):235– 245, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00674-
6&u=/mcom/1996-65-213/.

Deleglise:1998:C

[DR98a] Marc Deléglise and Joël Rivat. Computing ψ(x). Mathematics of Computation, 67(224):1691–1696, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00977-
6&u=/mcom/1998-67-224/.

Dorfler:1998:ASE

[DR98b]  W. Dörfler and M. Rumpf.  An

adaptive strategy for elliptic problems including a posteriori controlled boundary approximation. Mathematics of Computation, 67(224):1361– 1382, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00993-
4&u=/mcom/1998-67-224/.

Drezner:1994:CTN

[Dre94] Zvi Drezner. Computation of the trivariate normal integral. Mathematics of Computation, 62(205):289– 294, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dresden:1998:OAN

[Dre98] Gregory P. Dresden. Orbits of algebraic numbers with low heights. Mathematics of Computation, 67(222):815– 820, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00963-
6&u=/mcom/1998-67-222/.

Disney:1991:EBM

[DS91] Shaun Disney and Ian H. Sloan. Error bounds for the method of good lattice points. Mathematics of Computation, 56(193):257–266, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dutt:1993:SMP

[DS93] Pravir K. Dutt and A. K. Singh. Spectral methods for periodic initial value problems with nonsmooth data. Mathematics of Computation, 61(204):645– 658, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dilcher:1995:NCF

[DS95] Karl Dilcher and Ladislav Skula. A new criterion for the first case of Fermat’s last theorem. Mathematics of Computation, 64(209):363–392, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dahmen:1999:CWB

[DS99] Wolfgang Dahmen and Reinhold Schneider. Composite wavelet bases for operator equations. Mathematics of Computation, 68(228):1533– 1567, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01092-
                                              3&u=/mcom/1999-68-228/.

Dummit:1997:CSU

[DST97] David S. Dummit, Jonathan W. Sands, and Brett A. Tangedal. Computing Stark units for totally real cubic fields. Mathematics of Computation, 66(219): 1239–1267, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00852-
1&u=/mcom/1997-66-219/.

Du:1998:DGI

[Du98]  Qiang Du.  Discrete gauge in-

variant approximations of a time dependent Ginzburg–Landau model of superconductivity. Mathematics of Computation, 67(223):965– 986, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00954-
5&u=/mcom/1998-67-223/.

Dubner:1992:PAF

[Dub92] Harvey Dubner. Partitions approximated by finite cosine-series. Mathematics of Computation, 58(198):729– 736, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dubner:1993:GRP

[Dub93] Harvey Dubner. Generalized repunit primes. Mathematics of Computation, 61(204):927–930, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dubner:1996:LSG

[Dub96]  Harvey Dubner.  Large Sophie

Germain primes. Mathematics of Computation, 65(213):393–396, January 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00670-
9&u=/mcom/1996-65-213/.

Duijvestijn:1994:SPS

[Dui94] A. J. W. Duijvestijn. Simple perfect squared squares and 2 × 1 squared rectangles of order 25. Mathematics of Computation, 62(205):325–332, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duijvestijn:1996:NPC

[Dui96a] A. J. W. Duijvestijn. The number of polyhedral (3-connected planar) graphs. Mathematics of Computation, 65(215):1289–1293, July 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                              http://www.ams.org/jourcgi/jour-
                                              pbprocess?fn=110&arg1=S0025-5718-
                                              96-00749-1&u=/mcom/1996-65-215/.

Duijvestijn:1996:SPS

[Dui96b] A. J. W. Duijvestijn. Simple perfect squared squares and 2 × 1 squared rectangles of order 26. Mathematics of Computation, 65(215):1359– 1364, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00705-
                                              3&u=/mcom/1996-65-215/.

Dummit:1991:SSQ

[Dum91] D. S. Dummit. Solving solvable quintics. Mathematics of Computation, 57 (195):387–401, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duran:1999:EED

[Dur99] Ricardo G. Durán. Error estimates for 3-d narrow finite elements. Mathematics of Computation, 68(225):187– 199, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-00994-
1&u=/mcom/1999-68-225/.

Dusart:1999:PGT

[Dus99] Pierre Dusart. The k^{th} prime is greater than k(ln k + ln ln k − 1) for k ≥ 2. Mathematics of Computation, 68(225): 411–415, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01037-
6&u=/mcom/1999-68-225/.

DeMarchi:1994:SFN

[DVZ94] S. De Marchi, M. (Marco) Vianello, and R. Zanovello. Splitting functions and numerical analysis of WR-type methods for evolutionary and stationary problems. In Gautschi [Gau94], pages 281– 285. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

DeVore:1990:DAA

[DY90] Ronald A. DeVore and Xiang Ming Yu. Degree of adaptive approximation. Mathematics of Computation, 55(192): 625–635, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eberhard:1999:ADN

[EB99] Peter Eberhard and Christian Bischof.

Automatic differentiation of numerical integration algorithms. Mathematics of Computation, 68(226):717– 731, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01027-
                                              3&u=/mcom/1999-68-226/.

Eckhoff:1993:AER

[Eck93] Knut S. Eckhoff. Accurate and efficient reconstruction of discontinuous functions from truncated series expansions. Mathematics of Computation, 61 (204):745–763, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eckhoff:1995:ARF

[Eck95] Knut S. Eckhoff. Accurate reconstructions of functions of finite regularity from truncated Fourier series expansions. Mathematics of Computation, 64(210):671–690, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eckhoff:1998:HON

[Eck98]  Knut S. Eckhoff.  On a high

order numerical method for functions with singularities. Mathematics of Computation, 67(223):1063– 1087, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00949-
                                              1&u=/mcom/1998-67-223/.

Edelman:1992:DSC

[Ede92] Alan Edelman. On the distribution of a scaled condition number. Mathematics of Computation, 58(197):185– 190, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

El-Daou:1997:ISL

[EDK97] M. K. El-Daou and H. G. Khajah. Iterated solutions of linear operator equations with the Tau method. Mathematics of Computation, 66(217):207– 213, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00803-
X&u=/mcom/1997-66-217/.

Edixhoven:1990:ECR

[EdT90] Bas Edixhoven, Arnold de Groot, and Jaap Top. Elliptic curves over the rationals with bad reduction at only one prime. Mathematics of Computation, 54 (189):413–419, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elman:1990:IMC

[EG90] Howard C. Elman and Gene H. Golub.

Iterative methods for cyclically reduced non-self-adjoint linear systems. Mathematics of Computation, 54(190):671– 700, April 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elman:1991:IMC

[EG91] Howard C. Elman and Gene H. Golub.

Iterative methods for cyclically reduced nonselfadjoint linear systems. II. Mathematics of Computation, 56(193):215– 242, January 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ezzirani:1999:FAG

[EG99] Abdelkrim Ezzirani and Allal Guessab.

A fast algorithm for Gaussian type quadrature formulae with mixed boundary conditions and some lumped mass spectral approximations. Mathematics of Computation, 68(225):217– 248, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01001-
                                              7&u=/mcom/1999-68-225/.

Eichenauer-Herrmann:1991:ICP

[EH91] Jürgen Eichenauer-Herrmann. Inversive congruential pseudorandom numbers avoid the planes. Mathematics of Computation, 56(193):297–301, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.jstor.org/
                                              stable/2008543.

Eichenauer-Herrmann:1993:SIN

[EH93] Jürgen Eichenauer-Herrmann. Statistical independence of a new class of inversive congruential pseudorandom numbers. Mathematics of Computation, 60 (201):375–384, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

                                              www.jstor.org/stable/2153174.

Eichenauer-Herrmann:1994:ILB

[EH94a] Jürgen Eichenauer-Herrmann. Improved lower bounds for the discrepancy of inversive congruential pseudorandom numbers. Mathematics of Computation, 62(206):783–786, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eichenauer-Herrmann:1994:GIC

[EH94b] Jürgen Eichenauer-Herrmann. On generalized inversive congruential pseudorandom numbers. Mathematics of Computation, 63(207):293–299, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.jstor.org/stable/
                                              2153575.

Eichenauer-Herrmann:1996:CIC

[EHE96] Jürgen Eichenauer-Herrmann and Frank Emmerich. Compound inversive congruential pseudorandom numbers: an average-case analysis. Mathematics of Computation, 65(213):215– 225, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00675-
8&u=/mcom/1996-65-213/;   http:/
/www.jstor.org/stable/2153841.

Eichenauer-Herrmann:1994:EIC

[EHI94] Jürgen Eichenauer-Herrmann and Katja Ickstadt. Explicit inversive congruential pseudorandom numbers with power of two modulus. Mathematics of Computation, 62(206):787–797, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.jstor.org/stable/
2153539.

Eichenauer-Herrmann:1997:AEP

[EHL97] Jürgen Eichenauer-Herrmann and Gerhard Larcher. Average equidistribution properties of compound nonlinear congruential pseudorandom numbers. Mathematics of Computation, 66(217): 363–372, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00802-
8&u=/mcom/1997-66-217/.

Eichenauer-Herrmann:1992:LBD

[EHN92] Jürgen Eichenauer-Herrmann and Harald Niederreiter. Lower bounds for the discrepancy of inversive congruential pseudorandom numbers with power of two modulus. Mathematics of Computation, 58(198):775–779, April 1992.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eichenauer-Herrmann:1997:ICP

[EHN97] Jürgen Eichenauer-Herrmann and Harald Niederreiter. Inversive congruential pseudorandom numbers: Distribution of triples. Mathematics of Computation, 66(220):1629– 1644, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00867-
                                              3&u=/mcom/1997-66-220/.

Ehrich:1994:EBG

[Ehr94] Sven Ehrich. Error bounds for Gauss– Kronrod quadrature formulae. Mathematics of Computation, 62(205):295– 304, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eriksson:1993:ASD

[EJ93] Kenneth Eriksson and Claes Johnson. Adaptive streamline diffusion finite element methods for stationary convection-diffusion problems. Mathematics of Computation, 60(201):167– 188, S1–S2, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Earnest:1997:UBH

[EK97] A. G. Earnest and Azar Khosravani. Universal binary Hermitian forms. Mathematics of Computation, 66(219): 1161–1168, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00860-
                                              0&u=/mcom/1997-66-219/.

Ekl:1996:ESF

[Ekl96]  Randy L. Ekl.  Equal sums of

four seventh powers. Mathematics of Computation, 65(216):1755– 1756, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Ekl:1998:NRE

[Ekl98] Randy L. Ekl. New results in equal sums of like powers. Mathematics of Computation, 67(223):1309– 1315, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00979-
X&u=/mcom/1998-67-223/.

Eggermont:1991:UEE

[EL91] P. P. B. Eggermont and C. H. Lubich. Uniform error estimates of operational quadrature methods for nonlinear convolution equations on the half-line. Mathematics of Computation, 56(193): 149–176, January 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elliott:1992:EES

[EL92a] Charles M. Elliott and Stig Larsson. Error estimates with smooth and nonsmooth data for a finite element method for the Cahn–Hilliard equation. Mathematics of Computation, 58(198):603– 630, S33–S36, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elliott:1992:SEE

[EL92b] Charles M. Elliott and Stig Larsson.

Supplement to “Error estimates with smooth and nonsmooth data for a finite element method for the Cahn–Hilliard equation”. Mathematics of Computation, 58(198):S33–S36, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eggermont:1993:OQM

[EL93] P. P. B. Eggermont and Ch. Lubich. Operational quadrature methods for Wiener–Hopf integral equations. Mathematics of Computation, 60(202):699– 718, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elliott:1995:FEM

[EL95] Charles M. Elliott and Stig Larsson. A finite element model for the time-dependent Joule heating problem. Mathematics of Computation, 64(212): 1433–1453, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Erdos:1996:CSI

[EL96] P. Erdős and Mordechai Lewin. dcomplete sequences of integers. Mathematics of Computation, 65(214):837– 840, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00707-
7&u=/mcom/1996-65-214/.

Elliott:1996:BRA

[Ell96] David Elliott. Book reviews: Approximation and Computation: A Festschrift in honor of Walter Gautschi, edited by R. V. M. Zahar. Mathematics of Computation, 65(215): ??, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00705-
3&u=/mcom/1996-65-215/.

Elman:1996:BRI

[Elm96] Howard Elman. Book reviews: Iterative solution methods, by Owe Axelsson. Mathematics of Computation, 65(214): 886–888, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Elman:1997:BRI

[Elm97] Howard Elman. Book reviews: Iterative methods for linear and nonlinear equations, by C. T. Kelley. Mathematics of Computation, 66(218): ??, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00856-
9&u=/mcom/1997-66-218/.

Estep:1990:BDD

[ELR90] Donald Estep, Michael Loss, and Jeffrey Rauch. Boundedness of dispersive difference schemes. Mathematics of Computation, 55(191):55–87, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Erdos:1993:ELP

[ELS93] P. Erdős, C. B. Lacampagne, and J. L. Selfridge. Estimates of the least prime factor of a binomial coefficient. Mathematics of Computation, 61(203): 215–224, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ewing:1991:LRT

[ELV91] R. E. Ewing, R. D. Lazarov, and P. S. Vassilevski. Local refinement techniques for elliptic problems on cellcentered grids. I. Error analysis. Mathematics of Computation, 56(194):437– 461, April 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ernvall:1991:CIP

[EM91] R. Ernvall and T. Metsänkylä. Cyclotomic invariants for primes between 125 000 and 150 000. Mathematics of Computation, 56(194):851–858, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ernvall:1992:CZA

[EM92a] R. Ernvall and T. Metsänkylä. Computation of the zeros of p-ADIC Lfunctions. Mathematics of Computation, 58(198):815–830, S37–S53, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ernvall:1992:CIP

[EM92b] R. Ernvall and T. Metsänkylä. Cyclotomic invariants for primes to one million. Mathematics of Computation, 59(199):249–250, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ernvall:1992:SCZ

[EM92c] R. Ernvall and T. Metsankyla. Supplement to “Computation of the zeros of p-adic L-functions”. Mathematics of Computation, 58(198):S37–S53, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ernvall:1994:CZA

[EM94] R. Ernvall and T. Metsänkylä. Computation of the zeros of p-adic L-functions. II. Mathematics of Computation, 62 (205):391–406, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Edelman:1995:PRC

[EM95] Alan Edelman and H. Murakami. Polynomial roots from companion matrix eigenvalues. Mathematics of Computation, 64(210):763–776, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ehrich:1997:SPL

[EM97a] Sven Ehrich and Giuseppe Mastroianni. Stieltjes polynomials and Lagrange interpolation. Mathematics of Computation, 66(217):311– 331, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00808-
9&u=/mcom/1997-66-217/.

Ernvall:1997:DFQ

[EM97b] R. Ernvall and T. Metsänkylä. On the p-divisibility of Fermat quotients. Mathematics of Computation, 66(219): 1353–1365, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00843-
0&u=/mcom/1997-66-219/.

Emmerich:1996:PVG

[Emm96] Frank Emmerich. Pseudorandom vector generation by the compound inversive method. Mathematics of Computation, 65(214):749–760, April 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00706-
5&u=/mcom/1996-65-214/.

Edmondson:1994:ETS

[ESA94] Genet M. Edmondson, Jennifer Seberry, and Malcolm R. Anderson. On the existence of Turyn sequences of length less than 43. Mathematics of Computation, 62(205):351–362, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eydeland:1990:MVP

[EST90] Alexander Eydeland, Joel Spruck, and Bruce Turkington. Multiconstrained variational problems of nonlinear eigenvalue type: New formulations and algorithms. Mathematics of Computation, 55(192):509–535, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Falk:1991:NFE

[Fal91] Richard S. Falk. Nonconforming finite element methods for the equations of linear elasticity. Mathematics of Computation, 57(196):529– 550, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Falk:1996:BRA

[Fal96] Richard S. Falk. Book reviews: Asymptotic methods for elastic structures, by Philippe G. Ciarlet, Luís Trabucho and Juan M. Viaño. Mathematics of Computation, 65(216): ??, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Falk:1998:BRA

[Fal98]  Richard S. Falk.  Book reviews:

Acta numerica, edited by A. Iserles. Mathematics of Computation, 67(224): ??, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00989-
2&u=/mcom/1998-67-224/.

Fan:1998:EUT

[Fan98] Haitao Fan. Existence and uniqueness of traveling waves and error estimates for Godunov schemes of conservation laws. Mathematics of Computation, 67(221):87–109, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00921-
1&u=/mcom/1998-67-221/.

Faridani:1994:GST

[Far94] Adel Faridani. A generalized sampling theorem for locally compact Abelian groups. Mathematics of Computation, 63(207):307–327, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fassbender:1997:NMD

[Fas97] Heike Fassbender. On numerical methods for discrete least-squares approximation by trigonometric polynomials. Mathematics of Computation, 66(218): 719–741, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00845-
4&u=/mcom/1997-66-218/.

Ferguson:1999:API

[FBA99] Helaman R. P. Ferguson, David H. Bailey, and Steve Arno. Analysis of PSLQ, an integer relation finding algorithm. Mathematics of Computation, 68(225): 351–369, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-00995-
                                              3&u=/mcom/1999-68-225/.

Feng:1999:ABC

[Fen99] Xiaobing Feng. Absorbing boundary conditions for electromagnetic wave propagation. Mathematics of Computation, 68(225):145–168, January 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01028-
                                              5&u=/mcom/1999-68-225/.

Fee:1991:PFW

[FG91a] Greg Fee and Andrew Granville. The prime factors of Wendt’s binomial circulant determinant. Mathematics of Computation, 57(196):839–848, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fischer:1991:GPW

[FG91b] Bernd Fischer and Gene H. Golub. On generating polynomials which are orthogonal over several intervals. Mathematics of Computation, 56(194):711– 730, April 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Funaro:1991:CRP

[FG91c] Daniele Funaro and David Gottlieb. Convergence results for pseudospectral approximations of hyperbolic systems by a penalty-type boundary treatment. Mathematics of Computation, 57(196): 585–596, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Farouki:1996:OSB

[FG96] R. T. Farouki and T. N. T. Goodman. On the optimal stability of the Bernstein basis. Mathematics of Computation, 65(216):1553– 1566, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00759-
4&u=/mcom/1996-65-216/.

Fikioris:1998:TET

[Fik98] George Fikioris. Table erratum: Table of integrals, series, and products, [English translation, Fifth edition, Academic Press, Boston, MA, 1994; MR 94g:00008] by I. S. Gradshteyn and I. M. Ryzhik. Mathematics of Computation, 67(224):1753–1754, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fishman:1990:MCR

[Fis90] George S. Fishman. Multiplicative congruential random number generators with modulus 2^{β}: An exhaustive analysis for β ≃ 32 and a partial analysis for β ≃ 48. Mathematics of Computation, 54(189):331– 344, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

jstor.org/stable/2008698.

Fieker:1997:SRN

[FJP97] C. Fieker, A. Jurk, and M. Pohst.

On solving relative norm equations in algebraic number fields. Mathematics of Computation, 66(217):399– 410, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00761-
                                              8&u=/mcom/1997-66-217/.

Funaro:1991:ASD

[FK91] Daniele Funaro and Otared Kavian. Approximation of some diffusion evolution equations in unbounded domains by Hermite functions. Mathematics of Computation, 57(196):597– 619, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

French:1993:ARF

[FK93] Donald A. French and J. Thomas King. Analysis of a robust finite element approximation for a parabolic equation with rough boundary data. Mathematics of Computation, 60(201):79– 104, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fukuda:1996:CPG

[FK96] T. Fukuda and K. Komatsu. A capitulation problem and Greenberg’s conjecture on real quadratic fields. Mathematics of Computation, 65(213):313– 318, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00676-
                                              X&u=/mcom/1996-65-213/.

Flammenkamp:1991:NSN

[Fla91] Achim Flammenkamp. New sociable numbers. Mathematics of Computation, 56(194):871–873, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Flammang:1996:TNP

[Fla96]  V. Flammang.  Two new points

in the spectrum of the absolute Mahler measure of totally positive algebraic integers. Mathematics of Computation, 65(213):307–311, January 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00664-
3&u=/mcom/1996-65-213/.

Flyer:1998:AUB

[Fly98] Natasha Flyer. Asymptotic upper bounds for the coefficients in the Chebyshev series expansion for a general order integral of a function. Mathematics of Computation, 67(224):1601– 1616, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00976-
4&u=/mcom/1998-67-224/.

Farrell:1998:NEU

[FMOS98] Paul A. Farrell, John J. H. Miller, Eugene O’Riordan, and Grigorii I. Shishkin. On the non-existence of ϵuniform finite difference methods on uniform meshes for semilinear twopoint boundary value problems. Mathematics of Computation, 67(222):603– 617, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00922-
3&u=/mcom/1998-67-222/.

Farouki:1990:NCB

[FN90] R. T. Farouki and C. A. Neff. On the numerical condition of Bernstein– Bézier subdivision processes. Mathematics of Computation, 55(192):637– 647, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Farouki:1995:HIP

[FN95] R. T. Farouki and C. A. Neff. Hermite interpolation by Pythagorean hodograph quintics. Mathematics of Computation, 64(212):1589–1609, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Frappier:1993:QFI

[FO93] Clément Frappier and Patrick Olivier. A quadrature formula involving zeros of Bessel functions. Mathematics of Computation, 60(201):303–316, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Falk:1994:RMF

[FO94] Richard S. Falk and John E. Osborn. Remarks on mixed finite element methods for problems with rough coefficients. Mathematics of Computation, 62(205):1–19, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Forster:1994:TCP

[För94] Klaus-Jürgen Förster. On a theorem of C. Posse concerning Gaussian quadrature of Chebyshev type. Mathematics of Computation, 62(206):719–724, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Forbes:1997:LPT

[For97] Tony Forbes. A large pair of twin primes. Mathematics of Computation, 66(217):451–455, January 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
97-00793-X&u=/mcom/1997-66-217/.

Forbes:1999:PCC

[For99] Tony Forbes. Prime clusters and Cunningham chains. Mathematics of Computation, 68(228):1739– 1747, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01117-
5&u=/mcom/1999-68-228/.

Forster:1990:EWG

[FP90] Klaus-Jürgen Förster and Knut Petras.

On estimates for the weights in Gaussian quadrature in the ultraspherical case. Mathematics of Computation, 55(191):243–264, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Franco:1995:CCC

[FP95] Zachary Franco and Carl Pomerance.

On a conjecture of Crandall concerning the qx + 1 problem. Mathematics of Computation, 64(211):1333–1336, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

French:1996:CST

[FP96] Donald A. French and Todd E. Peterson. A continuous space-time finite element method for the wave equation. Mathematics of Computation, 65(214): 491–506, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00685-
0&u=/mcom/1996-65-214/.

Fischer:1997:WBO

[FP97] Bernd Fischer and Jürgen Prestin. Wavelets based on orthogonal polynomials. Mathematics of Computation, 66(220):1593–1618, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00876-
4&u=/mcom/1997-66-220/.

Frey:1994:RCD

[FR94] Gerhard Frey and Hans-Georg Rück. A remark concerning m-divisibility and the discrete logarithm in the divisor class group of curves. Mathematics of Computation, 62(206):865–874, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Frommer:1997:BRC

[Fro97] Andreas Frommer. Book reviews: Computer Numerik 1, 2, by Christoph Überhuber. Mathematics of Computation, 66(218):??, April 1997. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
97-00856-9&u=/mcom/1997-66-218/.

Feinsilver:1992:BFR

[FS92] P. Feinsilver and R. Schott. On Bessel functions and rate of convergence of zeros of Lommel polynomials. Mathematics of Computation, 59(199): 153–156, July 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Friedman:1995:AII

[FS95] Eduardo Friedman and Jonathan W.

Sands. On the l-adic Iwasawa λinvariant in a p-extension. with an appendix by Lawrence C. Washington. Mathematics of Computation, 64(212): 1659–1674, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fairweather:1990:SGM

[FSSC90] Graeme Fairweather, J. M. Sanz-Serna, and I. Christie. A stabilized Galerkin method for a third-order evolutionary problem. Mathematics of Computation, 55(192):497–507, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fiorot:1991:SPC

[FT91] J. C. Fiorot and J. Tabka. Shapepreserving C^{2} cubic polynomial interpolating splines. Mathematics of Computation, 57(195):291–298, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fujita:1993:MDT

[Fuj93] Hiroyuki Fujita. The minimum discriminant of totally real algebraic number fields of degree 9 with cubic subfields. Mathematics of Computation, 60(202): 801–810, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fukuda:1996:CUG

[Fuk96] Takashi Fukuda. Cyclotomic units and Greenberg’s conjecture for real quadratic fields. Mathematics of Computation, 65(215):1339–1348, July 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
96-00730-2&u=/mcom/1996-65-215/.

Feisel:1999:NBG

[FvS99] Sandra Feisel, Joachim von zur Gathen, and M. Amin Shokrollahi. Normal bases via general Gauss periods. Mathematics of Computation, 68(225): 271–290, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-00988-
6&u=/mcom/1999-68-225/.

Fung:1990:QPW

[FW90a] G. W. Fung and H. C. Williams. Quadratic polynomials which have a high density of prime values. Mathematics of Computation, 55(191):345– 353, July 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fung:1990:CTC

[FW90b] Gilbert W. Fung and H. C. Williams. On the computation of a table of complex cubic fields with discriminant D > −10^{6}. Mathematics of Computation, 55(191):313–325, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fung:1994:ECT

[FW94] G. W. Fung and H. C. Williams. Errata: “On the computation of a table of complex cubic fields with discriminant D > −10^{6}” [Math. Comp. 55 (1990), no. 191, 313–325, MR 90m:11155]. Mathematics of Computation, 63(207):433, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gaal:1996:CAP

[Gaá96] István Gaál. Computing all power integral bases in orders of totally real cyclic sextic number fields. Mathematics of Computation, 65(214):801– 822, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00708-
9&u=/mcom/1996-65-214/.

Galant:1992:AMM

[Gal92] David Galant. Algebraic methods for modified orthogonal polynomials. Mathematics of Computation, 59(200): 541–546, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Garcia-Archilla:1999:AIM

[GANT99] Bosco García-Archilla, Julia Novo, and Edriss S. Titi. An approximate inertial manifolds approach to postprocessing the Galerkin method for the Navier–Stokes equations. Mathematics of Computation, 68(227):893– 911, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01057-
1&u=/mcom/1999-68-227/.

Gastaldi:1993:UIE

[Gas93] Lucia Gastaldi. Uniform interior error estimates for the Reissner–Mindlin plate model. Mathematics of Computation, 61(204):539–567, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Garcia-Archilla:1991:FDF

[GASS91] B. García-Archilla and J. M. Sanz-Serna. A finite difference formula for the discretization of d^{3}/dx^{3} on nonuniform grids. Mathematics of Computation, 57(195):239–257, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1991:CSCa

[Gau91a] Walter Gautschi. A class of slowly convergent series and their summation by Gaussian quadrature. Mathematics of Computation, 57(195):309–324, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1991:CSCb

[Gau91b] Walter Gautschi. On certain slowly convergent series occurring in plate contact problems. Mathematics of Computation, 57(195):325–338, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943–1993: a halfcentury of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Gautschi:1996:BRP

[Gau96] Walter Gautschi. Book reviews: Parallel processing for scientific computing, edited by David H. Bailey, Petter E. Bjørstad, John R. Gilbert, Michael V. Mascagni, Robert S. Schreiber, Horst D. Simon, Virginia J. Torczon and Layne T. Watson. Mathematics of Computation, 65(213): ??, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Gordon:1993:CDL

[GC93] D. M. Gordon and C. Corrigendum:

“The distribution of Lucas Pomerance and elliptic pseudoprimes” [Math. Comp. 57 (1991), no. 196, 825–838, MR 92h:11081]. Mathematics of Computation, 60(202):877, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ge:1994:RUN

[Ge94] Guoqiang Ge. Recognizing units in number fields. Mathematics of Computation, 63(207):377–387, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Geck:1993:DNH

[Gec93] Meinolf Geck. The decomposition numbers of the Hecke algebra of type E_{6}^{∗}. Mathematics of Computation, 61(204): 889–899, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gelb:1997:RGP

[Gel97] Anne Gelb. The resolution of the Gibbs phenomenon for spherical harmonics. Mathematics of Computation, 66(218): 699–717, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00828-
4&u=/mcom/1997-66-218/.

Ghanem:1997:SBF

[GF97] Riadh Ben Ghanem and Clément Frappier. Spherical Bessel functions and explicit quadrature formula. Mathematics of Computation, 66(217):289– 296, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00794-
                                              1&u=/mcom/1997-66-217/.

Guy:1994:RAS

[GG94] Andrew W. P. Guy and Richard K. Guy. A record aliquot sequence. In Gautschi [Gau94], pages 557–559. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Grannell:1998:SS

[GGM98] M. J. Grannell, T. S. Griggs, and R. A. Mathon. Steiner systems S(5, 6, v) with v = 72 and 84. Mathematics of Computation, 67(221):357– 359, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00924-
                                              7&u=/mcom/1998-67-221/.

Grosse:1994:IRS

[GH94] Eric Grosse and John D. Hobby. Improved rounding for spline coefficients and knots. Mathematics of Computation, 63(207):175–194, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gu:1997:ETP

[GH97] Jinsheng Gu and Xiancheng Hu. Extension theorems for plate elements with applications. Mathematics of Computation, 66(220):1375– 1388, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00903-
4&u=/mcom/1997-66-220/.

Gatermann:1992:EFP

[GHO92a] Karin Gatermann, Christoph Hoffmann, and Gerhard Opfer. Explicit Faber polynomials on circular sectors. Mathematics of Computation, 58(197): 241–253, S1–S6, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gatermann:1992:SEF

[GHO92b] Karin Gatermann, Christoph Hoffmann, and Gerhard Opfer. Supplement to “Explicit Faber polynomials on circular sectors”. Mathematics of Computation, 58(197):S1–S6, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gunzburger:1991:AFE

[GHS91] M. D. Gunzburger, L. Hou, and T. P.

Svobodny. Analysis and finite element approximation of optimal control problems for the stationary Navier–Stokes equations with distributed and Neumann controls. Mathematics of Computation, 57(195):123–151, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gilbert:1998:BRC

[Gil98] R. P. Gilbert. Book reviews: Computational ocean acoustics, by Finn B. Jensen, William A. Kuperman, Michael B. Porter and Henrik Schmidt. Mathematics of Computation, 67(221):

??, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00891-
                                              6&u=/mcom/1998-67-221/.

Girstmair:1993:RCN

[Gir93a] Kurt Girstmair. The relative class numbers of imaginary cyclic fields of degrees 4, 6, 8, and 10. Mathematics of Computation, 61(204):881–887, S25– S27, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Girstmair:1993:SRC

[Gir93b] Kurt Girstmair. Supplement to “The relative class numbers of imaginary cyclic fields of degrees 4, 6, 8, and 10”. Mathematics of Computation, 61 (204):S25–S27, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Girault:1995:LPO

[Gir95] V. Girault. A local projection operator for quadrilateral finite elements. Mathematics of Computation, 64(212):1421– 1431, October 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gohberg:1995:FGE

[GKO95] I. Gohberg, T. Kailath, and V. Olshevsky. Fast Gaussian elimination with partial pivoting for matrices with displacement structure. Mathematics of Computation, 64(212):1557–1576, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1996:QCE

[GL96] Walter Gautschi and Shikang Li. On quadrature convergence of extended Lagrange interpolation. Mathematics of Computation, 65(215):1249– 1256, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00731-
4&u=/mcom/1996-65-215/.

Greif:1998:TID

[GL98a] Chen Greif and David Levin. The d_{2}-transformation for infinite double series and the D_{2}-transformation for infinite double integrals. Mathematics of Computation, 67(222):695– 714, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00955-
7&u=/mcom/1998-67-222/.

Guo:1998:AMN

[GL98b] Chun-Hua Guo and Peter Lancaster.

Analysis and modification of Newton’s method for algebraic Riccati equations. Mathematics of Computation, 67(223): 1089–1105, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00947-
8&u=/mcom/1998-67-223/.

Guo:1999:IST

[GL99] Chun-Hua Guo and Peter Lancaster. Iterative solution of two matrix equations. Mathematics of Computation, 68(228):1589–1603, October 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01122-
                                              9&u=/mcom/1999-68-228/.

Glassey:1992:CEP

[Gla92] R. T. Glassey. Convergence of an energy-preserving scheme for the Zakharov equations in one space dimension. Mathematics of Computation, 58 (197):83–102, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gekeler:1999:ZFC

[GLW99] Ernst-Ulrich Gekeler, Rita Leitl, and Bodo Wack. Zeta functions of a class of elliptic curves over a rational function field of characteristic two. Mathematics of Computation, 68(226):823– 833, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01058-
                                              3&u=/mcom/1999-68-226/.

Gouvea:1992:FME

[GM92a] F. Gouvêa and B. Mazur. Families of modular eigenforms. Mathematics of Computation, 58(198):793–805, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grothmann:1992:DSU

[GM92b] R. Grothmann and H. N. Mhaskar. Detection of singularities using segment approximation. Mathematics of Computation, 59(200):533–540, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Guo:1992:IPG

[GM92c] Ben Yu Guo and J. J. H. Miller. Iterative and Petrov–Galerkin methods for solving a system of one-dimensional nonlinear elliptic equations. Mathematics of Computation, 58(198):531– 547, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gori:1996:WFW

[GM96] Laura Gori and Charles A. Micchelli. On weight functions which admit explicit Gauss–Turán quadrature formulas. Mathematics of Computation, 65(216):1567–1581, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00769-
7&u=/mcom/1996-65-216/.

Gunzburger:1991:EUF

[GMP91] Max D. Gunzburger, Amnon J. Meir, and Janet S. Peterson. On the existence, uniqueness, and finite element approximation of solutions of the equations of stationary, incompressible magnetohydrodynamics. Mathematics of Computation, 56(194):523–563, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grady:1992:SDP

[GN92] Michael Grady and Morris Newman.

Some divisibility properties of the subgroup counting function for free products. Mathematics of Computation, 58 (197):347–353, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1996:SPR

[GN96] Walter Gautschi and Sotirios E. Notaris. Stieltjes polynomials and related quadrature formulae for a class of weight functions. Mathematics of Computation, 65(215):1257– 1268, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00732-
                                              6&u=/mcom/1996-65-215/.

Goetgheluck:1998:IFS

[Goe98] P. Goetgheluck. Infinite families of solutions of the equation \binom{n}{k} = 2\binom{a}{b}. Mathematics of Computation, 67(224):1727– 1733, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-01002-
                                              3&u=/mcom/1998-67-224/.

Goldstein:1991:MAF

[Gol91a] Charles I. Goldstein. Multigrid analysis of finite element methods with numerical integration. Mathematics of Computation, 56(194):409–436, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gollan:1991:MRT

[Gol91b] Holger W. Gollan. The 5-modular representations of the Tits simple group in the principal block. Mathematics of Computation, 57(195):369–386, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldfeld:1992:CCM

[Gol92] Dorian Goldfeld. On the computational complexity of modular symbols. Mathematics of Computation, 58(198):

807–814, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldberg:1994:NRA

[Gol94a] Moshe Goldberg. A numerical radius approach to stable difference schemes for parabolic systems. In Gautschi [Gau94], pages 293–296. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Goldfarb:1994:AUO

[Gol94b] Donald Goldfarb. Algorithms for unconstrained optimization: a review of recent developments. In Gautschi [Gau94], pages 33–48. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Gostin:1995:NFF

[Gos95] Gary B. Gostin. New factors of Fermat numbers. Mathematics of Computation, 64(209):393–395, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gottfert:1994:ANF

[Göt94] Rainer Göttfert. An acceleration of the Niederreiter factorization algorithm in characteristic 2. Mathematics of Computation, 62(206):831–839, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gordon:1991:DLE

[GP91] Daniel M. Gordon and Carl Pomerance.

The distribution of Lucas and elliptic pseudoprimes. Mathematics of Computation, 57(196):825–838, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gaal:1997:PIB

[GP97] István Gaál and Michael Pohst. Power integral bases in a parametric family of totally real cyclic quintics. Mathematics of Computation, 66(220):1689– 1696, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00868-
5&u=/mcom/1997-66-220/.

Gonzalez:1999:SRK

[GP99] C. González and C. Palencia. Stability of Runge–Kutta methods for abstract time-dependent parabolic problems: The Hölder case. Mathematics of Computation, 68(225):73– 89, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01018-
2&u=/mcom/1999-68-225/.

Grozev:1995:QFZ

[GR95] Georgi R. Grozev and Qazi I. Rahman. A quadrature formula with zeros of Bessel functions as nodes. Mathematics of Computation, 64(210):715–725, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grantham:1995:LPD

[Gra95] Jon Grantham. The largest prime dividing the maximal order of an element of S_{n}. Mathematics of Computation, 64(209):407–410, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grandcolas:1998:DCS

[Gra98]  Michel Grandcolas.  Diameters of

complete sets of conjugate algebraic integers of small degree. Mathematics of Computation, 67(222):821– 831, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00923-
5&u=/mcom/1998-67-222/.

Grochenig:1992:RAI

[Grö92] Karlheinz Gröchenig. Reconstruction algorithms in irregular sampling. Mathematics of Computation, 59(199):181– 194, July 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grochenig:1999:IST

[Grö99] Karlheinz Gröchenig. Irregular sampling, Toeplitz matrices, and the approximation of entire functions of exponential type. Mathematics of Computation, 68(226):749– 765, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01029-
7&u=/mcom/1999-68-226/.

Garvan:1990:SPF

[GS90] Frank Garvan and Dennis Stanton.

Sieved partition functions and qbinomial coefficients. Mathematics of Computation, 55(191):299–311, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glassey:1991:CSO

[GS91] Robert Glassey and Jack Schaeffer.

Convergence of a second-order scheme for semilinear hyperbolic equations in 2 + 1 dimensions. Mathematics of Computation, 56(193):87–106, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Godwin:1993:ENF

[GS93] H. J. Godwin and J. R. Smith. On the Euclidean nature of four cyclic cubic fields. Mathematics of Computation, 60 (201):421–423, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gottlieb:1995:GPI

[GS95] David Gottlieb and Chi-Wang Shu. On the Gibbs phenomenon. IV. Recovering exponential accuracy in a subinterval from a Gegenbauer partial sum of a piecewise analytic function. Mathematics of Computation, 64(211):1081– 1095, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Guo:1997:ACV

[GS97] Wen Guo and Martin Stynes. An analysis of a cell-vertex finite volume method for a parabolic convectiondiffusion problem. Mathematics of Computation, 66(217):105–124, January 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00795-
3&u=/mcom/1997-66-217/.

Gottlieb:1998:TVD

[GS98] Sigal Gottlieb and Chi-Wang Shu. Total variation diminishing Runge– Kutta schemes. Mathematics of Computation, 67(221):73–85, January 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00913-
2&u=/mcom/1998-67-221/.

Gottlieb:1991:CCS

[GT91] David Gottlieb and Eitan Tadmor. The CFL condition for spectral approximations to hyperbolic initial-boundary value problems. Mathematics of Computation, 56(194):565–588, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grabner:1993:SDD

[GT93] Peter J. Grabner and Robert F.

Tichy. Spherical designs, discrepancy and numerical integration. Mathematics of Computation, 60(201):327– 336, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Georg:1994:SEE

[GT94] Kurt Georg and Johannes Tausch.

Some error estimates for the numerical approximation of surface integrals. Mathematics of Computation, 62(206): 755–763, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Guo:1995:SMV

[Guo95] Ben Yu Guo. A spectral method for the vorticity equation on the surface. Mathematics of Computation, 64(211): 1067–1079, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gurak:1990:PHO

[Gur90] S. Gurak. Pseudoprimes for higherorder linear recurrence sequences. Mathematics of Computation, 55(192):

783–813, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gao:1995:OON

[GV95] Shuhong Gao and Scott A. Vanstone. On orders of optimal normal basis generators. Mathematics of Computation, 64(211):1227–1233, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gao:1998:GPO

[GvP98] Shuhong Gao, Joachim von zur Gathen, and Daniel Panario. Gauss periods: orders and cryptographical applications. Mathematics of Computation, 67(221): 343–352, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00935-
1&u=/mcom/1998-67-221/.

Gayle:1996:UPP

[GW96] R. C. Gayle and J. M. Wolfe. Unicity in piecewise polynomial L^{1}- approximation via an algorithm. Mathematics of Computation, 65(214):647– 660, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00709-
0&u=/mcom/1996-65-214/.

Gupta:1993:CMP

[GZ93] S. Gupta and D. Zagier. On the coefficients of the minimal polynomials of Gaussian periods. Mathematics of Computation, 60(201):385– 398, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hejhal:1993:FCM

[HA93a] D. A. Hejhal and S. Arno. On Fourier coefficients of Maass waveforms for psl(2, z). Mathematics of Computation, 61(203):245–267, S11–S16, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hejhal:1993:SFC

[HA93b] D. A. Hejhal and S. Arno. Supplement to “On Fourier coefficients of Maass waveforms for P SL(2, Z)”. Mathematics of Computation, 61(203): S11–S16, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Haber:1993:TFN

[Hab93] Seymour Haber. Two formulas for numerical indefinite integration. Mathematics of Computation, 60(201):279– 296, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagstrom:1991:ABC

[Hag91] Thomas Hagstrom. Asymptotic boundary conditions for dissipative waves: General theory. Mathematics of Computation, 56(194):589–606, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Halpern:1996:SMP

[Hal96] Laurence Halpern. Spectral methods in polar coordinates for the Stokes problem. Application to computation in unbounded domains. Mathematics of Computation, 65(214):507– 531, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00710-
7&u=/mcom/1996-65-214/.

Han:1990:DBE

[Han90] Hou De Han. A direct boundary element method for Signorini problems. Mathematics of Computation, 55(191): 115–128, July 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Han:1996:CPP

[Han96] Wen Bao Han. The coefficients of primitive polynomials over finite fields. Mathematics of Computation, 65(213):331– 340, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00663-
                                              1&u=/mcom/1996-65-213/.

Hayes:1991:RRO

[Hay91] David R. Hayes. On the reduction of rank-one Drinfeld modules. Mathematics of Computation, 57(195):339– 349, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Heath-Brown:1992:DZF

[HB92] D. R. Heath-Brown. The density of zeros of forms for which weak approximation fails. Mathematics of Computation, 59(200):613–623, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Heath-Brown:1993:SDE

[HBLT93] D. R. Heath-Brown, W. M. Lioen, and H. J. J. Te Riele. On solving the Diophantine equation x^{3} +y^{3} +z^{3} = k on a vector computer. Mathematics of Computation, 61(203):235–244, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1998:EOP

[HC98] Peter Hagis, Jr. and Graeme L. Cohen. Every odd perfect number has a prime factor which exceeds 10^{6}. Mathematics of Computation, 67(223):1323– 1330, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00982-
X&u=/mcom/1998-67-223/.

Ha-Duong:1994:SAB

[HDJ94] T. Ha-Duong and P. Joly. On the stability analysis of boundary conditions for the wave equation by energy methods. part I. the homogeneous case. Mathematics of Computation, 62(206): 539–563, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

He:1994:FPC

[He94] Matthew He. The Faber polynomials for circular arcs. In Gautschi [Gau94], pages 301–304. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Heinrich:1996:EAC

[Hei96] S. Heinrich. Efficient algorithms for computing the L_{2} discrepancy. Mathematics of Computation, 65(216):1621– 1633, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00756-
9&u=/mcom/1996-65-216/.

Helou:1997:WD

[Hel97] Charles Helou. On Wendt’s determinant. Mathematics of Computation, 66(219):1341–1346, July 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
97-00870-3&u=/mcom/1997-66-219/.

Hu:1998:BEC

[HGK98] Ning Hu, Xian-Zhong Guo, and I. Norman Katz. Bounds for eigenvalues and condition numbers in the p-version of the finite element method. Mathematics of Computation, 67(224):1423– 1450, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00983-
1&u=/mcom/1998-67-224/.

Hickernell:1999:AER

[HH99a] Fred J. Hickernell and Hee Sun Hong. The asymptotic efficiency of randomized nets for quadrature. Mathematics of Computation, 68(226):767– 791, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01019-
4&u=/mcom/1999-68-226/.

Horng:1999:SPR

[HH99b] Gwoboa Horng and Ming-Deh Huang. Solving polynomials by radicals with roots of unity in minimum depth. Mathematics of Computation, 68(226):881– 885, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-99-01060-
1&u=/mcom/1999-68-226/.

Hagstrom:1994:ALT

[HHM94a] Thomas Hagstrom, S. I. Hariharan, and R. C. MacCamy. On the accurate longtime solution of the wave equation in exterior domains: Asymptotic expansions and corrected boundary conditions. Mathematics of Computation, 63 (208):507–539, S7–S10, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagstrom:1994:SAL

[HHM94b] Thomas Hagstrom, S. I. Hariharan, and R. C. MacCamy. Supplement to “On the accurate long-time solution of the wave equation in exterior domains: Asymptotic expansions and corrected boundary conditions”. Mathematics of Computation, 63(208):S7–S10, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hickernell:1998:GDQ

[Hic98] Fred J. Hickernell. A generalized discrepancy and quadrature error bound. Mathematics of Computation, 67(221): 299–322, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00894-
1&u=/mcom/1998-67-221/.

Higham:1994:SCP

[Hig94] Nicholas J. Higham. A survey of componentwise perturbation theory in numerical linear algebra. In Gautschi [Gau94], pages 49–77. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Higham:1996:BRP

[Hig96] Nicholas J. Higham. Book reviews: Polynomial and matrix computations: Fundamental algorithms, Vol. 1, by Dario Bini and Victor Y. Pan. Mathematics of Computation, 65(214):888– 889, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00700-
                                              4&u=/mcom/1996-65-214/.

Higham:1998:FCS

[Hig98] Nicholas J. Higham. Factorizing complex symmetric matrices with positive definite real and imaginary parts. Mathematics of Computation, 67(224):1591– 1599, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00978-
                                              8&u=/mcom/1998-67-224/.

Hiptmair:1999:CCF

[Hip99] R. Hiptmair. Canonical construction of finite elements. Mathematics of Computation, 68(228):1325– 1346, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01166-
                                              7&u=/mcom/1999-68-228/.

Hiss:1994:MCR

[His94] Gerhard Hiss. The 3-modular characters of the Rudvalis sporadic simple group and its covering group. Mathematics of Computation, 62(206):851– 863, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Han:1995:SEE

[HJ95] Weimin Han and Søren Jensen. On the sharpness of L^{2}-error estimates of H_{0}^{1}- projections onto subspaces of piecewise, high-order polynomials. Mathematics of Computation, 64(209):51–70, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hou:1994:WNS

[HL94] Thomas Y. Hou and Philippe G. Le Floch. Why nonconservative schemes converge to wrong solutions: error analysis. Mathematics of Computation, 62 (206):497–530, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hansen:1992:PPF

[HM92a] Tom Hansen and Gary L. Mullen.

Primitive polynomials over finite fields. Mathematics of Computation, 59(200): 639–643, S47–S50, October 1992. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hansen:1992:SPP

[HM92b] Tom Hansen and Gary L. Mullen. Supplement to “Primitive polynomials over finite fields”. Mathematics of Computation, 59(200):S47–S50, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Heinrich:1993:MCC

[HM93] Stefan Heinrich and Peter Mathé. The Monte Carlo complexity of Fredholm integral equations. Mathematics of Computation, 60(201):257–278, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hofmann:1997:QMC

[HM97] Norbert Hofmann and Peter Mathé. On quasi-Monte Carlo simulation of stochastic differential equations. Mathematics of Computation, 66(218):573– 589, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00820-
X&u=/mcom/1997-66-218/.

Ham:1998:DIP

[HM98] Kie H. Ham and Gary L. Mullen. Distribution of irreducible polynomials of small degrees over finite fields. Mathematics of Computation, 67(221):337– 341, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00904-
1&u=/mcom/1998-67-221/.

Hansen:1993:GPC

[HMN93] Tom Hansen, Gary L. Mullen, and Harald Niederreiter. Good parameters for a class of node sets in quasi-Monte Carlo integration. Mathematics of Computation, 61(203):225–234, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hardy:1990:DAS

[HMW90] Kenneth Hardy, Joseph B. Muskat, and Kenneth S. Williams. A deterministic algorithm for solving n = fu^{2} + gv^{2} in coprime integers u and v. Mathematics of Computation, 55(191):327–343, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hobby:1998:NLB

[Hob98] John D. Hobby. A natural lattice basis problem with applications. Mathematics of Computation, 67(223):1149– 1161, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00936-
3&u=/mcom/1998-67-223/.

Holst:1996:BRM

[Hol96] Michael Holst. Book reviews: Multigrid methods for semiconductor device simulation, by J. Molenaar. Mathematics of Computation, 65(216): ??, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Holst:1997:RDT

[Hol97] Michael Holst. Reviews and descriptions of tables and books: 21. Algorithms for elliptic problems: Efficient sequential and parallel solvers, by Marián Vajteršic. Mathematics of Computation, 66(220):??, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00843-
0&u=/mcom/1997-66-220/.

Hongtao:1999:CBE

[Hon99]  Yang Hongtao.  On the conver-

gence of boundary element methods for initial-Neumann problems for the heat equation. Mathematics of Computation, 68(226):547– 557, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01022-
4&u=/mcom/1999-68-226/.

Horst:1997:GBS

[Hor97] Reiner Horst. On generalized bisection of n-simplices. Mathematics of Computation, 66(218):691– 698, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00809-
                                              0&u=/mcom/1997-66-218/.

Hou:1992:NDV

[Hou92] Thomas Y. Hou. A new desingularization for vortex methods. Mathematics of Computation, 58(197):103– 117, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Houstis:1996:BRE

[Hou96a] Elias N. Houstis. Book reviews: Environments and tools for parallel scientific computing, edited by Jack J. Dongarra and Bernard Tourancheau. Mathematics of Computation, 65(214): 879–880, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00700-
                                              4&u=/mcom/1996-65-214/.

Houstis:1996:BRH

[Hou96b] Elias N. Houstis. Book reviews: High performance computing: Problem solving with parallel and vector architectures, edited by Gary W. Sabot. Mathematics of Computation, 65(216): ??, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00768-
                                              5&u=/mcom/1996-65-216/.

Hong:1992:RRQ

[HP92] Y. P. Hong and C.-T. Pan. Rankrevealing QR factorizations and the singular value decomposition. Mathematics of Computation, 58(197):213– 232, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hajdu:1999:SPT

[HP99] L. Hajdu and Á. Pintér. Square product of three integers in short intervals. Mathematics of Computation, 68(227): 1299–1301, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01095-
9&u=/mcom/1999-68-227/.

Holden:1993:MFS

[HR93] Helge Holden and Nils Henrik Risebro.

A method of fractional steps for scalar conservation laws without the CFL condition. Mathematics of Computation, 60 (201):221–232, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hou:1994:FEA

[HR94] L. S. Hou and S. S. Ravindran. Finite element approximation for optimal control of electrically conducting fluid flows. In Gautschi [Gau94], pages 305– 309. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Hiss:1992:BMC

[HS92] Gerhard Hiss and Josephine Shamash.

2-blocks and 2-modular characters of the Chevalley groups G_{2}(q). Mathematics of Computation, 59(200):645– 672, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hamina:1994:SCM

[HS94] Martti Hamina and Jukka Saranen. On the spline collocation method for the single-layer heat operator equation. Mathematics of Computation, 62(205): 41–64, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hill:1995:USA

[HS95a] Adrian T. Hill and Endre Süli. Upper semicontinuity of attractors for linear multistep methods approximating sectorial evolution equations. Mathematics of Computation, 64(211):1097– 1122, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Huber:1995:PMS

[HS95b] Birkett Huber and Bernd Sturmfels. A polyhedral method for solving sparse polynomial systems. Mathematics of Computation, 64(212):1541–1555, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Habsieger:1997:ICP

[HS97a] Laurent Habsieger and Bruno Salvy. On integer Chebyshev polynomials. Mathematics of Computation, 66(218): 763–770, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00829-
6&u=/mcom/1997-66-218/.

Hunter:1997:ANI

[HS97b] Simon Hunter and Jonathan Sorenson.

Approximating the number of integers free of large prime factors. Mathematics of Computation, 66(220):1729– 1741, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00874-
0&u=/mcom/1997-66-220/.

Hasegawa:1993:PWE

[HST93] Takemitsu Hasegawa, Hirosi Sugiura, and Tatsuo Torii. Positivity of the weights of extended Clenshaw–Curtis quadrature rules. Mathematics of Computation, 60(202):719–734, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Heuer:1998:MAS

[HST98] Norbert Heuer, Ernst P. Stephan, and Thanh Tran. Multilevel additive Schwarz method for the h-p version of the Galerkin boundary element method. Mathematics of Computation, 67(222): 501–518, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00926-
0&u=/mcom/1998-67-222/.

Hasegawa:1991:AQC

[HT91] Takemitsu Hasegawa and Tatsuo Torii.

An automatic quadrature for Cauchy principal value integrals. Mathematics of Computation, 56(194):741–754, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hasegawa:1994:EOS

[HT94] Takemitsu Hasegawa and Tatsuo Torii.

An extension of the Olver–Sookne method for the solution of second-order linear difference equations. In Gautschi [Gau94], pages 297–300. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385

1993. URL http://math.nist.gov/

nesf/. See also SIAM Review, September 1995, 37(3), p. 483.

Hasegawa:1995:ANS

[HT95] Takemitsu Hasegawa and Tatsuo Torii. An algorithm for nondominant solutions of linear second-order inhomogeneous difference equations. Mathematics of Computation, 64(211):1199– 1214, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hasegawa:1990:ABF

[HTS90] Takemitsu Hasegawa, Tatsuo Torii, and Hiroshi Sugiura. An algorithm based on the FFT for a generalized Chebyshev interpolation. Mathematics of Computation, 54(189):195–210, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hu:1998:ICC

[Hu98] Qiya Hu. Interpolation correction for collocation solutions of Fredholm integro-differential equations. Mathematics of Computation, 67(223):987– 999, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00956-
9&u=/mcom/1998-67-223/.

Huang:1991:HAS

[Hua91] Ming You Huang. A Hamiltonian approximation to simulate solitary waves of the Korteweg–de Vries equation.

Mathematics of Computation, 56(194): 607–620, April 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Huckle:1995:LRM

[Huc95] Thomas Huckle. Low-rank modification of the unsymmetric Lánczos algorithm. Mathematics of Computation, 64(212):1577–1588, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hung:1991:EPD

[Hun91] David C. Hung. Even positive definite unimodular quadratic forms over ∥(\sqrt{3}). Mathematics of Computation, 57(195): 351–368, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hundsdorfer:1992:UCS

[Hun92] Willem Hundsdorfer. Unconditional convergence of some Crank–Nicolson LOD methods for initial-boundary value problems. Mathematics of Computation, 58(197):35–53, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hundsdorfer:1998:NSD

[Hun98a] Willem Hundsdorfer. A note on stability of the Douglas splitting method. Mathematics of Computation, 67(221): 183–190, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00914-
4&u=/mcom/1998-67-221/.

Hundsdorfer:1998:TMS

[Hun98b] Willem Hundsdorfer. Trapezoidal and midpoint splittings for initialboundary value problems. Mathematics of Computation, 67(223):1047– 1062, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00984-
3&u=/mcom/1998-67-223/.

Hanlon:1990:CDB

[HW90] Phil Hanlon and David Wales. Computing the discriminants of Brauer’s centralizer algebras. Mathematics of Computation, 54(190):771–796, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Han:1992:AEB

[HW92] Hou De Han and Xiao Nan Wu. The approximation of the exact boundary conditions at an artificial boundary for linear elastic equations and its application. Mathematics of Computation, 59 (199):21–37, July 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hou:1999:CMF

[HWC99] Thomas Y. Hou, Xiao-Hui Wu, and Zhiqiang Cai. Convergence of a multiscale finite element method for elliptic problems with rapidly oscillating coefficients. Mathematics of Computation, 68(227):913– 943, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01077-
7&u=/mcom/1999-68-227/.

Iannucci:1999:SLP

[Ian99] Douglas E. Iannucci. The second largest prime divisor of an odd perfect number exceeds ten thousand. Mathematics of Computation, 68(228):1749– 1760, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01126-
6&u=/mcom/1999-68-228/.

Hashimoto:1999:STC

[iHT99] Ki ichiro Hashimoto and Hiroshi Tsunogai. On the Sato–Tate conjecture for QM-curves of genus two. Mathematics of Computation, 68(228):1649– 1662, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01061-
3&u=/mcom/1999-68-228/.

Indlekofer:1996:LKT

[IJ96] Karl-Heinz Indlekofer and Antal Járai.

Largest known twin primes. Mathematics of Computation, 65(213):427– 428, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Indlekofer:1999:LKT

[IJ99] Karl-Heinz Indlekofer and Antal Járai.

Largest known twin primes and Sophie Germain primes. Mathematics of Computation, 68(227):1317– 1324, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01079-
0&u=/mcom/1999-68-227/.

Ito:1998:TKT

[IK98] Kazufumi Ito and Franz Kappel.

The Trotter–Kato theorem and approximation of PDEs. Mathematics of Computation, 67(221):21–44, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00915-
6&u=/mcom/1998-67-221/.

Nagao:1997:CHR

[iN97] Koh ichi Nagao. Construction of high-rank elliptic curves with a nontrivial torsion point. Mathematics of Computation, 66(217):411– 415, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00779-
5&u=/mcom/1997-66-217/.

Isaacson:1994:MCB

[Isa94] Eugene Isaacson. Mathematics of computation: a brief history. In Gautschi [Gau94], pages xvii–xix. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Isaacson:1996:BRI

[Isa96a] Eugene Isaacson. Book reviews: Industrial mathematics: A course in solving real-world problems, by Avner Friedman and Walter Littman. Mathematics of Computation, 65(214):878– 879, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Isaacson:1996:BRN

[Isa96b] Eugene Isaacson. Book reviews: Numerical analysis: A first course in scientific computation, by Peter Deuflhard and Andreas Hohmann. Mathematics of Computation, 65(216): ??, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Isaacson:1996:BRP

[Isa96c] Eugene Isaacson. Book reviews: Proceedings of the Cornelius Lanczos international centenary conference, by J. David Brown, Moody T. Chu, Donald C. Ellison and Robert J. Plemmons. Mathematics of Computation, 65(213): 423–435, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Ishikawa:1995:NZD

[Ish95] Tsuneo Ishikawa. On the number of zeros of diagonal forms. Mathematics of Computation, 64(210):841–854, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ivic:1991:ZET

[It91] A. Ivić and H. J. J. te Riele. On the zeros of the error term for the mean square of |ζ(\frac{1}{2} + it)|. Mathematics of Computation, 56(193):303–328, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Iyanaga:1995:RMC

[Iya95] Kenichi Iyanaga. A recursive method to calculate the number of solutions of quadratic equations over finite fields. Mathematics of Computation, 64(211): 1319–1331, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jacobson:1999:ASC

[Jac99] Michael J. Jacobson, Jr. Applying sieving to the computation of quadratic class groups. Mathematics of Computation, 68(226):859– 867, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01003-
0&u=/mcom/1999-68-226/.

Jaeschke:1990:CN

[Jae90] Gerhard Jaeschke. The Carmichael numbers to 10^{12}. Mathematics of Computation, 55(191):383–389, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jaeschke:1993:SPS

[Jae93] Gerhard Jaeschke. On strong pseudoprimes to several bases. Mathematics of Computation, 61(204):915– 926, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jakubec:1998:DCN

[Jak98] Stanislav Jakubec. On divisibility of the class number h^{+} of the real cyclotomic fields of prime degree l. Mathematics of Computation, 67(221):369– 398, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00916-
8&u=/mcom/1998-67-221/.

Jepsen:1996:DR

[Jep96]  Charles H. Jepsen.  Dissections

of p:q rectangles. Mathematics of Computation, 65(214):771–778, April 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00711-
9&u=/mcom/1996-65-214/.

Jha:1995:FCF

[Jha95] Vijay Jha. Faster computation of the first factor of the class number of ℚ(ζ_{p}). Mathematics of Computation, 64(212): 1705–1710, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jia:1998:APM

[Jia98] Rong-Qing Jia. Approximation properties of multivariate wavelets. Mathematics of Computation, 67(222):647– 665, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00925-
9&u=/mcom/1998-67-222/.

Jakubec:1991:EIN

[JKN91] Stanislav Jakubec, Juraj Kostra, and Karol Nemoga. On the existence of an integral normal basis generated by a unit in prime extensions of rational numbers. Mathematics of Computation, 56(194):809–815, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jakubec:1992:WEC

[JM92] Stanislav Jakubec and František Marko.

Witt equivalence classes of quartic number fields. Mathematics of Computation, 58(197):355–368, January 1992.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jakubec:1995:PCN

[JMS95] Stanislav Jakubec, František Marko, and Kazimierz Szymiczek. Parity of class numbers and Witt equivalence of quartic fields. Mathematics of Computation, 64(212):1711–1715, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jakubec:1997:CPC

[JMS97] Stanislav Jakubec, František Marko, and Kazimierz Szymiczek. Corrigendum: “Parity of class numbers and Witt equivalence of quartic fields” [math. comp. 64 (1995), no. 212, 1711–1715, MR 95m:11123]. Mathematics of Computation, 66(218): 927, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00842-
9&u=/mcom/1997-66-218/.

Joe:1993:BLR

[Joe93] Stephen Joe. Bounds on the lattice rule criterion R. Mathematics of Computation, 61(204):821–831, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jia:1998:VSS

[JRZ98] Rong-Qing Jia, S. D. Riemenschneider, and Ding-Xuan Zhou. Vector subdivision schemes and multiple wavelets. Mathematics of Computation, 67(224):1533–1563, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-98-00985-
5&u=/mcom/1998-67-224/.

Joe:1992:CLR

[JS92] Stephen Joe and Ian H. Sloan. On computing the lattice rule criterion R. Mathematics of Computation, 59(200): 557–568, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jiang:1994:CEI

[JS94] Guang Shan Jiang and Chi-Wang Shu.

On a cell entropy inequality for discontinuous Galerkin methods. Mathematics of Computation, 62(206):531– 538, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1990:CSC

[JSH90] Claes Johnson, Anders Szepessy, and Peter Hansbo. On the convergence of shock-capturing streamline diffusion finite element methods for hyperbolic conservation laws. Mathematics of Computation, 54(189):107– 129, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jetter:1999:EES

[JSW99] Kurt Jetter, Joachim Stöckler, and Joseph D. Ward. Error estimates for scattered data interpolation on spheres. Mathematics of Computation, 68(226): 733–747, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01080-
7&u=/mcom/1999-68-226/.

Just:1990:IRA

[Jus90] Bettina Just. Integer relations among algebraic numbers. Mathematics of Computation, 54(189):467–477, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jackiewicz:1992:VSE

[JZ92] Z. Jackiewicz and M. Zennaro. Variable-stepsize explicit two-step Runge–Kutta methods. Mathematics of Computation, 59(200):421–438, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kaltofen:1995:ACB

[Kal95] Erich Kaltofen. Analysis of Coppersmith’s block Wiedemann algorithm for the parallel solution of sparse linear systems. Mathematics of Computation, 64(210):777–806, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kaplansky:1995:FNG

[Kap95] Irving Kaplansky. The first nontrivial genus of positive definite ternary forms. Mathematics of Computation, 64(209): 341–345, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Karakashian:1997:BRN

[Kar97] Ohannes Karakashian. Book reviews: Numerical solution of the incompressible Navier–Stokes equations, by L. Quartapelle. Mathematics of Computation, 66(219):??, July 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kauthen:1997:CCA

[KB97] J.-P. Kauthen and H. Brunner. Continuous collocation approximations to solutions of first kind Volterra equations. Mathematics of Computation, 66(220):1441–1459, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00905-
8&u=/mcom/1997-66-220/.

Keast:1997:BRL

[Kea97] Patrick Keast. Book reviews: Lattice methods for multiple integration, by I. H. Sloan and S. Joe. Mathematics of Computation, 66(217): ??, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00793-
X&u=/mcom/1997-66-217/.

Kearfott:1998:BRN

[Kea98] R. Baker Kearfott. Book reviews: Numerica: A modelling language for global optimization, by Pascal Van Hentenryck, Laurent Michel, and Yves Deville. Mathematics of Computation, 67 (224):??, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00989-
2&u=/mcom/1998-67-224/.

Kedlaya:1998:SCP

[Ked98]  Kiran S. Kedlaya.  Solving con-

strained Pell equations. Mathematics of Computation, 67(222):833– 842, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00918-
                                              1&u=/mcom/1998-67-222/.

Keiper:1992:PSE

[Kei92] J. B. Keiper. Power series expansions of Riemann’s ξ function. Mathematics of Computation, 58(198):765–773, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keiper:1996:ZRD

[Kei96]  J. B. Keiper.  On the zeros of

the Ramanujan τ-Dirichlet series in the critical strip. Mathematics of Computation, 65(216):1613–1619, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00734-
                                              X&u=/mcom/1996-65-216/.

Keller:1995:NCP

[Kel95] Wilfrid Keller. New Cullen primes. with a biographical sketch of James Cullen by T. G. Holt and a supplement by Wilfrid Keller and Wolfgang Niebuhr. Mathematics of Computation, 64(212):1733–1741, 539–546, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Khajah:1999:TMA

[Kha99] H. G. Khajah. Tau method approximation of a generalized Epstein– Hubbell elliptic-type integral. Mathematics of Computation, 68(228):1615– 1621, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01128-
                                              X&u=/mcom/1999-68-228/.

Kida:1999:REC

[Kid99] Masanari Kida. Reduction of elliptic curves over certain real quadratic number fields. Mathematics of Computation, 68(228):1679–1685, October 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01129-
1&u=/mcom/1999-68-228/.

King:1997:BRT

[Kin97] J. Thomas King. Book reviews: Theory and numerics of ordinary and partial differential equations, by M. Ainsworth, J. Levesley, W. A. Light and M. Marletta. Mathematics of Computation, 66 (219):??, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

King:1999:BRA

[Kin99a] J. Thomas King. Book reviews: Afternotes goes to graduate school, lectures in advanced numerical analysis, by G. W. Stewart. Mathematics of Computation, 68(226):??, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01060-
1&u=/mcom/1999-68-226/.

King:1999:BRF

[Kin99b] Tom King. Book reviews: A first course in the numerical analysis of differential equations, by Arieh Iserles. Mathematics of Computation, 68(225): ??, January 1999. URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-00992-
8&u=/mcom/1999-68-225/.

Koukouvinos:1990:BTS

[KKS90] C. Koukouvinos, S. Kounias, and K. Sotirakoglou. On base and Turyn sequences. Mathematics of Computation, 55(192):825–837, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kahan:1997:CCR

[KL97] William Kahan and Ren-Cang Li. Composition constants for raising the orders of unconventional schemes for ordinary differential equations. Mathematics of Computation, 66(219):1089– 1099, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00873-
9&u=/mcom/1997-66-219/.

Klir:1996:BRF

[Kli96] George J. Klir. Book reviews: The fuzzy systems handbook: A practitioner’s guide to building, using, and maintaining fuzzy systems, by Earl Cox. Mathematics of Computation, 65(213): ??, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Kloucek:1996:ACN

[KLL96] Petr Klouček, Bo Li, and Mitchell Luskin. Analysis of a class of nonconforming finite elements for crystalline microstructures. Mathematics of Computation, 65(215):1111– 1135, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-96-00735-
1&u=/mcom/1996-65-215/.

Karakashian:1990:OHO

[KM90] Ohannes Karakashian and William McKinney. On optimal high-order in time approximations for the Korteweg– de Vries equation. Mathematics of Computation, 55(192):473–496, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kurita:1991:PNW

[KM91] Yoshiharu Kurita and Makoto Matsumoto. Primitive t-nomials (t = 3, 5) over GF (2) whose degree is a Mersenne exponent ≤ 44497. Mathematics of Computation, 56(194):817–821, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Karakashian:1998:STF

[KM98] Ohannes Karakashian and Charalambos Makridakis. A space-time finite element method for the nonlinear Schrödinger equation: the discontinuous Galerkin method. Mathematics of Computation, 67(222):479– 499, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00946-
6&u=/mcom/1998-67-222/.

Keller:1995:SNC

[KN95] Wilfrid Keller and Wolfgang Niebuhr.

Supplement to “New Cullen primes”. Mathematics of Computation, 64(212): S39–S46, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knopp:1993:CSA

[Kno93] Marvin I. Knopp. On the cuspidal spectrum of the arithmetic Hecke groups. Mathematics of Computation, 61(203): 269–275, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knuth:1993:JFS

[Knu93] Donald E. Knuth. Johann Faulhaber and sums of powers. Mathematics of Computation, 61(203):277–294, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knyazev:1997:NER

[Kny97]  Andrew V. Knyazev.  New esti-

mates for Ritz vectors. Mathematics of Computation, 66(219):985– 995, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00855-
                                              7&u=/mcom/1997-66-219/.

Kok:1996:BRN

[Kok96] Jan Kok. Book reviews: A numerical library in C for scientists and engineers, by H. T. Lau. Mathematics of Computation, 65(215): ??, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00705-
                                              3&u=/mcom/1996-65-215/.

Kolbig:1990:BRC

[Köl90] K. S. Kölbig. Book review: Computation of Functions on Electronic Computers — Handbook (in Russian). Naukova Dumka, Kiev, 194, 599pp, by B. A. Popov, G. S. Tesler. Mathematics of Computation, 55(191):395–397, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.jstor.org/stable/
2008818.

Kolbig:1995:TET

[Köl95] K. S. Kölbig. Table errata: Table of integrals, series, and products by I. S. Gradshteĭn, 1994 I. M. Ryzhik [Academic Press, Boston, MR 94g:00008]. Mathematics of Computation, 64(209): 449–460, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1997:TEIa

[Köl97a] K. S. Kölbig. Table errata: 629. integrals and series, volume 1, by A. P. Prudnikov, Yu. A. Brychkov, and O. I. Marichev. Mathematics of Computation, 66(220):??, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1997:TEIb

[Köl97b] K. S. Kölbig. Table errata: 630. integrals and series, volume 1, by A. P. Prudnikov, Yu. A. Brychkov, and O. I. Marichev. Mathematics of Computation, 66(220):1765–1767, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1997:TEE

[Köl97c] K. S. Kölbig. Table errata: 631. Higher transcendental functions, by A. Erdelyi, W. Magnus, F. Oberhettinger, and F. G. Tricomi. Mathematics of Computation, 66(220):1766–1777, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1997:TEP

[Köl97d] K. S. Kölbig. Table errata: 632. Handbook of elliptic integrals for engineers and physicists, by P. F. Byrd and M. D. Friedman. Mathematics of Computation, 66(220):1767, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1997:TEH

[Köl97e] K. S. Kölbig. Table errata: Handbook of elliptic integrals for engineers and scientists [Second edition, Springer, New York, 1971, MR 43 #3506] by P. F. Byrd and M. D. Friedman. Mathematics of Computation, 66(220):1767, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1997:TEId

[Köl97f] K. S. Kölbig. Table errata: Integrals and series. Vol. 4 [Gordon and Breach, New York, 1992 and MR 93c:44003] by A. P. Prudnikov, Yu. A. Brychkov and O. I. Marichev. Mathematics of Computation, 66(220):1766, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1997:TEIc

[Köl97g] K. S. Kölbig. Table errata: Integrals and series. Vol. 1 [English translation, Gordon & Breach, New York, 1986, MR 88f:00013 by A. P. Prudnikov, Yu. A. Brychkov and O. I. Marichev]. Mathematics of Computation, 66(220):1765– 1766, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1997:TET

[Köl97h] K. S. Kölbig. Table errata: Tables of integral transforms. Vol. I [McGraw-Hill, New York, 1954, MR 15, 868a] by A. Erdélyi, W. Magnus and F. Oberhettinger and F. G. Tricomi. Mathematics of Computation, 66(220):1766– 1767, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kacewicz:1992:TCA

[KP92] B. Z. Kacewicz and L. Plaskota. Termination conditions for approximating linear problems with noisy information. Mathematics of Computation, 59(200): 503–513, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Khobalatte:1994:MPE

[KP94] Brahim Khobalatte and Benoît Perthame. Maximum principle on the entropy and second-order kinetic schemes. Mathematics of Computation, 62(205): 119–131, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kroner:1991:ABC

[Krö91] Dietmar Kröner. Absorbing boundary conditions for the linearized Euler equations in 2-D. Mathematics of Computation, 57(195):153–167, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kechkar:1992:ALS

[KS92] Nasserdine Kechkar and David Silvester. Analysis of locally stabilized mixed finite element methods for the Stokes problem. Mathematics of Computation, 58(197):1–10, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kanenobu:1993:PIB

[KS93a] Taizo Kanenobu and Toshio Sumi.

Polynomial invariants of 2-bridge knots through 22 crossings. Mathematics of Computation, 60(202):771–778, S17– S28, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kanenobu:1993:SPI

[KS93b] Taizo Kanenobu and Toshio Sumi. Supplement to “Polynomial invariants of 2-bridge knots through 22 crossings”. Mathematics of Computation, 60(202): S17–S28, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kaltofen:1998:STF

[KS98] Erich Kaltofen and Victor Shoup. Subquadratic-time factoring of polynomials over finite fields. Mathematics of Computation, 67(223):1179– 1197, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00944-
2&u=/mcom/1998-67-223/.

Kobayashi:1998:NCM

[KSS98] Hidetsune Kobayashi, Hideo Suzuki, and Yoshihiko Sakai. Numerical calculation of the multiplicity of a solution to algebraic equations. Mathematics of Computation, 67(221):257– 270, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00906-
5&u=/mcom/1998-67-221/.

Kuo:1990:LED

[KT90] Hung-Ju Kuo and Neil S. Trudinger. Linear elliptic difference inequalities with random coefficients. Mathematics of Computation, 55(191):37–53, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Koyama:1997:SSD

[KTS97] Kenji Koyama, Yukio Tsuruoka, and Hiroshi Sekigawa. On searching for solutions of the Diophantine equation x^{3} + y^{3} + z^{3} = n. Mathematics of Computation, 66(218):841– 851, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00830-
2&u=/mcom/1997-66-218/.

Kuijlaars:1995:CTQ

[Kui95]  Arno Kuijlaars.  Chebyshev-type

quadrature and partial sums of the exponential series. Mathematics of Computation, 64(209):251–263, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kuijlaars:1996:ZFP

[Kui96] Arno B. J. Kuijlaars. The zeros of Faber polynomials generated by an m-star. Mathematics of Computation, 65(213): 151–156, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00684-
9&u=/mcom/1996-65-213/.

Kvaerno:1990:RKM

[Kvæ90] Anne Kværnø. Runge–Kutta methods applied to fully implicit differentialalgebraic equations of index 1. Mathematics of Computation, 54(190):583– 625, April 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kubina:1990:EWC

[KW90] Jeffrey M. Kubina and Marvin C. Wunderlich. Extending Waring’s conjecture to 471, 600, 000. Mathematics of Computation, 55(192):815–820, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kato:1996:NCP

[KWY96a] Takashi Kato, Li-Ming Wu, and Niro Yanagihara. On a nonlinear congruential pseudorandom number generator. Mathematics of Computation, 65 (213):227–233, January 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:/

/www.jstor.org/stable/2153842;
http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
96-00694-1&u=/mcom/1996-65-213/.

Kato:1996:STN

[KWY96b] Takashi Kato, Li-Ming Wu, and Niro Yanagihara. The serial test for a nonlinear pseudorandom number generator. Mathematics of Computation, 65 (214):761–769, April 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:/

/www.jstor.org/stable/2153612;
http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
96-00712-0&u=/mcom/1996-65-214/.

Kaneko:1991:DKM

[KX91] Hideaki Kaneko and Yuesheng Xu. Degenerate kernel method for Hammerstein equations. Mathematics of Computation, 56(193):141–148, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kaneko:1994:GTQ

[KX94] Hideaki Kaneko and Yuesheng Xu.

Gauss-type quadratures for weakly singular integrals and their application to Fredholm integral equations of the second kind. Mathematics of Computation, 62(206):739–753, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Labreuche:1998:CTF

[Lab98] Christophe Labreuche. A convergence theorem for the fast multipole method for 2-dimensional scattering problems. Mathematics of Computation, 67(222): 553–591, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00937-
5&u=/mcom/1998-67-222/.

Lai:1994:CBL

[Lai94] Ming Jun Lai. On the computation of Battle–Lemarié’s wavelets. Mathematics of Computation, 63(208):689– 699, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lambert:1996:TET

[Lam96] Adeline Lambert. Table errata: Table of integrals, series, and products [fifth edition, Academic Press, Boston, MA, 1994 and MR 94g:00008] by I. S. Gradshteyn and I. M. Ryzhik. Mathematics of Computation, 65(216):1769, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lambert:1997:TET

[Lam97] Adeline Lambert. Table errata: Table of integrals, series, and products [fifth edition, English translation, Academic Press, Boston, MA, 1994, MR 94g:00008] by I. S. Gradshteĭn and I. M. Ryzhik. Mathematics of Computation, 66(217):463, January 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Langtry:1996:ADA

[Lan96] T. N. Langtry. An application of Diophantine approximation to the construction of rank-1 lattice quadrature rules. Mathematics of Computation, 65(216):1635–1662, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00758-
2&u=/mcom/1996-65-216/.

Larsson:1997:BRN

[Lar97] Stig Larsson. Book reviews: Navier– Stokes equations and nonlinear functional analysis, by Roger Temam. Mathematics of Computation, 66(219): ??, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Laurie:1996:AGQ

[Lau96]  Dirk P. Laurie.  Anti-Gaussian

quadrature formulas. Mathematics of Computation, 65(214):739– 747, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00713-
2&u=/mcom/1996-65-214/.

Laurie:1997:CGK

[Lau97] Dirk P. Laurie. Calculation of Gauss– Kronrod quadrature rules. Mathematics of Computation, 66(219):1133– 1145, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00861-
2&u=/mcom/1997-66-219/.

Lawrence:1991:PVC

[Law91] Jim Lawrence. Polytope volume computation. Mathematics of Computation, 57(195):259–271, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lazarus:1993:CDU

[Laz93] Andrew J. Lazarus. Cyclotomy and delta units. Mathematics of Computation, 61(203):295–305, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lazarus:1994:ICM

[Laz94] Andrew J. Lazarus. Implications of computational mathematics for the philosophy of mathematics. In Gautschi [Gau94], pages 561–565. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Lyness:1994:SNC

[LC94] J. N. Lyness and Ronald Cools. A survey of numerical cubature over triangles. In Gautschi [Gau94], pages 127– 150. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

LEcuyer:1996:MEC

[L’E96] Pierre L’Ecuyer. Maximally equidistributed combined Tausworthe generators. Mathematics of Computation, 65(213):203–213, January 1996.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
96-00696-5&u=/mcom/1996-65-213/.

LEcuyer:1999:TLC

[L’E99a] Pierre L’Ecuyer. Tables of linear congruential generators of different sizes and good lattice structure. Mathematics of Computation, 68(225):249– 260, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-00996-
5&u=/mcom/1999-68-225/.

LEcuyer:1999:TME

[L’E99b] Pierre L’Ecuyer. Tables of maximally equidistributed combined LFSR generators. Mathematics of Computation, 68(225):261–269, January 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
99-01039-X&u=/mcom/1999-68-225/.

Lebaud:1994:EEI

[Leb94] M. P. Lebaud. Error estimate in an isoparametric finite element eigenvalue problem. Mathematics of Computation, 63(207):19–40, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lecot:1991:QMC

[Léc91] Christian Lécot. A quasi-Monte Carlo method for the Boltzmann equation. Mathematics of Computation, 56(194): 621–644, April 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lecot:1996:EBQ

[Léc96] Christian Lécot. Error bounds for quasi-Monte Carlo integration with nets. Mathematics of Computation, 65(213):179–187, January 1996. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
96-00690-4&u=/mcom/1996-65-213/.

Leemans:1999:ART

[Lee99] Dimitri Leemans. An atlas of regular thin geometries for small groups. Mathematics of Computation, 68(228):1631– 1647, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01130-
8&u=/mcom/1999-68-228/.

Lehner:1990:LHC

[Leh90] J. Lehner. The local Hurwitz constant and Diophantine approximation on Hecke groups. Mathematics of Computation, 55(192):765–781, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehman:1992:LPD

[Leh92a] J. Larry Lehman. Levels of positive definite ternary quadratic forms. Mathematics of Computation, 58(197):399– 417, S17–S22, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehman:1992:SLP

[Leh92b] J. Larry Lehman. Supplement to “Levels of positive definite ternary quadratic forms” by J. Larry Lehman. Mathematics of Computation, 58(197):S17– S22, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1993:MWM

[Leh93] D. H. Lehmer. The mathematical work of Morgan Ward. Mathematics of Computation, 61(203):307–311, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehman:1997:RES

[Leh97] Larry Lehman. Rational eigenvectors in spaces of ternary forms. Mathematics of Computation, 66(218):833– 839, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00821-
                                              1&u=/mcom/1997-66-218/.

Lenferink:1991:CPI

[Len91a] H. W. J. Lenferink. Contractivitypreserving implicit linear multistep methods. Mathematics of Computation, 56(193):177–199, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lenstra:1991:FIB

[Len91b] H. W. Lenstra, Jr. Finding isomorphisms between finite fields. Mathematics of Computation, 56(193):329– 347, January 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lenstra:1997:BRM

[Len97] H. W. Lenstra, Jr. Book reviews: Mathematics of Computation 1943-1993: A half-century of computational mathematics, by Walter Gautschi (Editor). Mathematics of Computation, 66 (219):??, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Levillain:1992:EAM

[Lev92] Vincent Levillain. Eigenvalue approximation by a mixed method for resonant inhomogeneous cavities with metallic boundaries. Mathematics of Computation, 58(197):11–20, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Levin:1998:APM

[Lev98]  David Levin.  The approximation

power of moving least-squares. Mathematics of Computation, 67(224):1517– 1531, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00974-
0&u=/mcom/1998-67-224/.

Lewanowicz:1994:SAS

[Lew94]  Stanisław Lewanowicz.  A simple

approach to the summation of certain slowly convergent series. Mathematics of Computation, 63(208):741– 745, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Linzer:1993:MFF

[LF93] Elliot Linzer and Ephraim Feig. Modified FFTs for fused multiply-add architectures. Mathematics of Computation, 60(201):347–361, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lotstedt:1993:FAM

[LG93a] Per Lötstedt and Bertil Gustafsson.

Fourier analysis of multigrid methods for general systems of PDEs. Mathematics of Computation, 60(202):473– 493, S3–S9, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lotstedt:1993:SFA

[LG93b] Per Lötstedt and Bertil Gustafsson. Supplement to “Fourier analysis of multigrid methods for general systems of PDEs”. Mathematics of Computation, 60(202):S3–S9, April 1993. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lodha:1997:UAE

[LG97] Suresh K. Lodha and Ron Goldman. A unified approach to evaluation algorithms for multivariate polynomials. Mathematics of Computation, 66(220):1521–1553, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00862-
4&u=/mcom/1997-66-220/.

Lavoie:1994:GDT

[LGRA94] J. L. Lavoie, F. Grondin, A. K. Rathie, and K. Arora. Generalizations of Dixon’s theorem on the sum of a _{3}F_{2}. Mathematics of Computation, 62(205): 267–276, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Leydold:1998:SPA

[LH98] Josef Leydold and Wolfgang Hörmann. A sweep-plane algorithm for generating random tuples in simple polytopes. Mathematics of Computation, 67(224):1617–1635, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-98-01004-
7&u=/mcom/1998-67-224/.

Li:1994:PMP

[Li94a] Ren Cang Li. On perturbations of matrix pencils with real spectra. Mathematics of Computation, 62(205):231– 265, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1994:BCP

[Li94b] Shou Fu Li. B-convergence properties of multistep Runge–Kutta methods. Mathematics of Computation, 62(206): 565–575, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1999:CBF

[Li99] Ruixia Li. On the coupling of BEM and FEM for exterior problems for the Helmholtz equation. Mathematics of Computation, 68(227):945– 953, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01064-
9&u=/mcom/1999-68-227/.

Lin:1994:EAF

[Lin94] E. B. Lin. Eigenvalue approximation of Fredholm integral operators. In Gautschi [Gau94], pages 317–321. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Lindstrom:1998:FFS

[Lin98]  Bernt Lindström.  Finding finite

B_{2}-sequences faster. Mathematics of Computation, 67(223):1173– 1178, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00986-
7&u=/mcom/1998-67-223/.

Liu:1994:FDM

[Liu94] Jinn Liang Liu. A finite difference method for symmetric positive differential equations. Mathematics of Computation, 62(205):105–118, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Liu:1994:STB

[LJ94] Anwei Liu and Barry Joe. On the shape of tetrahedra from bisection. Mathematics of Computation, 63(207): 141–154, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Liu:1996:QLR

[LJ96a] Anwei Liu and Barry Joe. Quality local refinement of tetrahedral meshes based on 8-subtetrahedron subdivision. Mathematics of Computation, 65(215): 1183–1200, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00748-
X&u=/mcom/1996-65-215/.

Lyness:1996:TCF

[LJ96b]  J. N. Lyness and S. Joe.  Tri-

angular canonical forms for lattice rules of prime-power order. Mathematics of Computation, 65(213):165– 178, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00691-
6&u=/mcom/1996-65-213/.

Lehmer:1993:LP

[LL93] D. H. Lehmer and Emma Lehmer. The Lehmer project. Mathematics of Computation, 61(203):313–317, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1998:NFE

[LL98] Bo Li and Mitchell Luskin. Nonconforming finite element approximation of crystalline microstructure. Mathematics of Computation, 67(223):917– 946, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00941-
7&u=/mcom/1998-67-223/.

LeFloch:1999:GMS

[LL99] Philippe G. LeFloch and Jian-Guo Liu. Generalized monotone schemes, discrete paths of extrema, and discrete entropy conditions. Mathematics of Computation, 68(227):1025– 1055, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01062-
5&u=/mcom/1999-68-227/.

Lenstra:1993:FNF

[LLMP93] A. K. Lenstra, H. W. Lenstra, Jr., M. S.

Manasse, and J. M. Pollard. The factorization of the ninth Fermat number. Mathematics of Computation, 61(203): 319–349, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lenstra:1995:AFN

[LLMP95] A. K. Lenstra, H. W. Lenstra, Jr., M. S. Manasse, and J. M. Pollard.

Addendum: “The factorization of the ninth Fermat number” [Math. Comp. 61 (1993), no. 203, 319–349, MR 93k:11116]. Mathematics of Computation, 64(211):1357, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lawton:1996:AME

[LLS96] W. Lawton, S. L. Lee, and Zuowei Shen. An algorithm for matrix extension and wavelet construction. Mathematics of Computation, 65(214):723– 737, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00714-
4&u=/mcom/1996-65-214/.

Li:1990:ESA

[LM90] Zi-Cai Li and Rudolf Mathon. Error and stability analysis of boundary methods for elliptic problems with interfaces. Mathematics of Computation, 54(189):41–61, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lenstra:1994:FTL

[LM94a] A. K. Lenstra and M. S. Manasse. Factoring with two large primes. Mathematics of Computation, 63(208):785– 798, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1994:AMH

[LM94b] Zhilin Li and Anita Mayo. ADI methods for heat equations with discontinuities along an arbitrary interface. In Gautschi [Gau94], pages 311–315. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

LeRoux:1999:NSF

[LM99] Marie-Noelle Le Roux and Paul-Emile Mainge. Numerical solution of a fast diffusion equation. Mathematics of Computation, 68(226):461– 485, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01063-
7&u=/mcom/1999-68-226/.

Loh:1996:NAC

[LN96] Günter Löh and Wolfgang Niebuhr.

A new algorithm for constructing large Carmichael numbers. Mathematics of Computation, 65(214):823– 836, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00692-
8&u=/mcom/1996-65-214/.

Loheac:1993:PAC

[LNS93] J.-P. Lohéac, F. Nataf, and M. Schatzman. Parabolic approximations of the convection-diffusion equation. Mathematics of Computation, 60(202):515– 530, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lubich:1993:RKM

[LO93] Ch. Lubich and A. Ostermann. Runge– Kutta methods for parabolic equations and convolution quadrature. Mathematics of Computation, 60(201):105– 131, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lozier:1994:NES

[LO94] D. W. Lozier and F. W. J. Olver. Numerical evaluation of special functions. In Gautschi [Gau94], pages 79– 125. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. URL http://math.

                                              nist.gov/mcsd/Reports/2001/nesf/.

See also SIAM Review, September 1995, 37(3), p. 483.

Lubich:1995:RKA

[LO95] Christian Lubich and Alexander Ostermann. Runge–Kutta approximation of quasi-linear parabolic equations. Mathematics of Computation, 64(210): 601–627, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lorentzen:1994:DCF

[Lor94] Lisa Lorentzen. Divergence of continued fractions related to hypergeometric series. Mathematics of Computation, 62(206):671–686, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Louboutin:1992:FCN

[Lou92] Stéphane Louboutin. L-functions and class numbers of imaginary quadratic fields and of quadratic extensions of an imaginary quadratic field. Mathematics of Computation, 59(199):213–230, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Louboutin:1995:DAN

[Lou95] Stéphane Louboutin. Determination of all nonquadratic imaginary cyclic number fields of 2-power degrees with ideal class groups of exponents ≤ 2. Mathematics of Computation, 64(209):323– 340, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Louboutin:1997:CRC

[Lou97] Stéphane Louboutin. Computation of relative class numbers of CM-fields. Mathematics of Computation, 66(219): 1185–1194, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00863-
6&u=/mcom/1997-66-219/.

Lozier:1996:BRI

[Loz96] Daniel W. Lozier. Book reviews: An introduction to scientific, symbolic, and graphical computation, by Eugene Fiume. Mathematics of Computation, 65 (216):??, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Lozier:1997:BRA

[Loz97] Daniel W. Lozier. Book reviews: Abstract machine models for highly parallel computers, edited by John R. Davy and Peter M. Dew. Mathematics of Computation, 66(218): ??, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00856-
9&u=/mcom/1997-66-218/.

Lozier:1998:BRA

[Loz98a] Daniel W. Lozier. Book reviews:

Applications on advanced architecture computers, edited by Greg Astfalk. Mathematics of Computation, 67(221):

??, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00891-
6&u=/mcom/1998-67-221/.

Lozier:1998:BRC

[Loz98b] Daniel W. Lozier. Book reviews: Computation of special functions, by Shanjie Zhang and Jianming Jin. Mathematics of Computation, 67(221): ??, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00891-
6&u=/mcom/1998-67-221/.

Lozier:1998:BRN

[Loz98c] Daniel W. Lozier. Book reviews: Numerical recipes in Fortran 90: The art of parallel scientific computing, Volume 2 of Fortan numerical recipes, second edition, by William H. Press, Saul A. Teukolsky, William T. Vetterling and Brian P. Flannery. Mathematics of Computation, 67(221): ??, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00891-
6&u=/mcom/1998-67-221/.

Lesaint:1992:EEN

[LP92] P. Lesaint and J. Pousin. Error estimates for a nonlinear degenerate parabolic problem. Mathematics of Computation, 59(200):339–358, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewanowicz:1995:AMC

[LP95] Stanisław Lewanowicz and Stefan Paszkowski. An analytic method for convergence acceleration of certain hypergeometric series. Mathematics of Computation, 64(210):691–713, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lukes:1996:SRP

[LPW96a] R. F. Lukes, C. D. Patterson, and H. C. Williams. Some results on pseudosquares. Mathematics of Computation, 65(213):361–372, S25–S27, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00678-
3&u=/mcom/1996-65-213/.

Lukes:1996:SSR

[LPW96b] R. F. Lukes, C. D. Patterson, and H. C. Williams. Supplement to “Some results on pseudosquares”. Mathematics of Computation, 65(213):S25– S27, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1990:NSM

[LR90] Shou Fu Li and Baogen Ruan. Nonlinear stability of multistep multiderivative methods. Mathematics of Computation, 55(192):581–589, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Leimkuhler:1994:SIC

[LR94] B. Leimkuhler and S. Reich. Symplectic integration of constrained Hamiltonian systems. Mathematics of Computation, 63(208):589–605, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lenferink:1991:GRC

[LS91a] H. W. J. Lenferink and M. N. Spijker. On a generalization of the resolvent condition in the Kreiss matrix theorem. Mathematics of Computation, 57(195): 211–220, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lenferink:1991:USR

[LS91b] H. W. J. Lenferink and M. N. Spijker. On the use of stability regions in the numerical analysis of initial value problems. Mathematics of Computation, 57(195):221–237, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lenstra:1993:CFL

[LS93a] H. W. Lenstra, Jr. and J. O. Shallit. Continued fractions and linear recurrences. Mathematics of Computation, 61(203):351–354, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1993:LRC

[LS93b] J. N. Lyness and T. Sørevik. Lattice rules by component scaling. Mathematics of Computation, 61(204):799– 820, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lempken:1994:SPI

[LS94] W. Lempken and R. Staszewski. The structure of the projective indecomposable modules of ˆ3M_{22} in characteristic 2. Mathematics of Computation, 62(206):841–850, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

LeTallec:1995:DDN

[LS95] Patrick Le Tallec and Taoufik Sassi. Domain decomposition with non-matching grids: Augmented Lagrangian approach. Mathematics of Computation, 64(212):1367–1396, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1998:NCM

[LS98] Bao Cheng Li and Stavros Syngellakis.

Numerical conformal mapping based on the generalised conjugation operator. Mathematics of Computation, 67(222): 619–639, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00957-
0&u=/mcom/1998-67-222/.

Lin:1991:SGR

[LSEG91] T. Lin, J. Sochacki, R. Ewing, and J. George. Some grid refinement schemes for hyperbolic equations with piecewise constant coefficients. Mathematics of Computation, 56(193):61– 86, January 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1991:NII

[LSK91] J. N. Lyness, T. Sørevik, and P. Keast.

Notes on integration and integer sublattices. Mathematics of Computation, 56 (193):243–255, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Larsson:1999:SRA

[LSS99] Stig Larsson and J.-M. Sanz-Serna.

A shadowing result with applications to finite element approximation of reaction-diffusion equations. Mathematics of Computation, 68(225):55– 72, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01038-
8&u=/mcom/1999-68-225/.

Lubich:1996:NDE

[LST96] Ch. Lubich, I. H. Sloan, and V. Thomée. Nonsmooth data error estimates for approximations of an evolution equation with a positive-type memory term. Mathematics of Computation, 65(213): 1–17, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00677-
1&u=/mcom/1996-65-213/.

Larcher:1994:RFW

[LSW94] Gerhard Larcher, Wolfgang Ch. Schmid, and Reinhard Wolf. Representation of functions as Walsh series to different bases and an application to the numerical integration of high-dimensional Walsh series. Mathematics of Computation, 63(208):701–716, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lukes:1997:FTE

[LSW97] Richard F. Lukes, Renate Scheidler, and Hugh C. Williams. Further tabulation of the Erdős–Selfridge function. Mathematics of Computation, 66(220):1709–1717, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00864-
8&u=/mcom/1997-66-220/.

Lasiecka:1991:NAA

[LT91a] I. Lasiecka and R. Triggiani. Numerical approximations of algebraic Riccati equations for abstract systems modelled by analytic semigroups, and applications. Mathematics of Computation, 57 (196):639–662, S13–S37, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lasiecka:1991:SNA

[LT91b] I. Lasiecka and R. Triggiani. Supplement to numerical approximations of algebraic Riccati equations for abstract systems modelled by analytic semigroups, and applications. Mathematics of Computation, 57(196):s13– s37, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

LEcuyer:1991:SPT

[LT91c] Pierre L’Ecuyer and Shu Tezuka. Structural properties for two classes of combined random number generators. Mathematics of Computation, 57(196): 735–746, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Larcher:1994:NIW

[LT94a] Gerhard Larcher and Claudia Traunfellner. On the numerical integration of Walsh series by number-theoretic methods. Mathematics of Computation, 63 (207):277–291, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Liu:1994:SMS

[LT94b] Wen Bin Liu and Tao Tang. Spectral methods for singular perturbation problems. In Gautschi [Gau94], pages 323– 326. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

LeTallec:1999:CAD

[LT99] P. Le Tallec and M. D. Tidriri. Convergence analysis of domain decomposition algorithms with full overlapping for the advection-diffusion problems. Mathematics of Computation, 68(226): 585–606, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01030-
3&u=/mcom/1999-68-226/.

Larsson:1998:NSP

[LTW98] Stig Larsson, Vidar Thomée, and Lars B. Wahlbin. Numerical solution of parabolic integro-differential equations by the discontinuous Galerkin method. Mathematics of Computation, 67(221): 45–71, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00883-
7&u=/mcom/1998-67-221/.

Lu:1998:CRW

[Lu98]  Weidong Lu.  Convergence of a

random walk method for a partial differential equation. Mathematics of Computation, 67(222):593– 602, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00917-
X&u=/mcom/1998-67-222/.

Li:1991:SDP

[LW91] T. Y. Li and Xiao Shen Wang. Solving deficient polynomial systems with homotopies which keep the subschemes at infinity invariant. Mathematics of Computation, 56(194):693–710, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1993:SRP

[LW93] T. Y. Li and Xiao Shen Wang. Solving real polynomial systems with real homotopies. Mathematics of Computation, 60(202):669–680, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1996:BRC

[LW96a] T. Y. Li and Xiaoshen Wang. The BKK root count in C^{n}. Mathematics of Computation, 65(216):1477– 1484, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00778-
8&u=/mcom/1996-65-216/.

Liu:1996:NSS

[LW96b] Hailiang Liu and Jinghua Wang.

Nonlinear stability of stationary discrete shocks for nonconvex scalar conservation laws. Mathematics of Computation, 65(215):1137–1153, July 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00733-
8&u=/mcom/1996-65-215/.

Liu:1997:DRP

[LW97] Hailiang Liu and Jinghua Wang. Decay rate for perturbations of stationary discrete shocks for convex scalar conservation laws. Mathematics of Computation, 66(217):69– 84, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00804-
                                              1&u=/mcom/1997-66-217/.

Liu:1993:SSD

[LX93] Jian Guo Liu and Zhou Ping Xin. L^{1}- stability of stationary discrete shocks. Mathematics of Computation, 60(201): 233–244, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Xu:1996:FEG

[lX96]  Yu lin Xu.  Fast evaluation of

the Gaunt coefficients. Mathematics of Computation, 65(216):1601– 1612, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00774-
                                              0&u=/mcom/1996-65-216/.

Lyness:1994:QCS

[Lyn94] J. N. Lyness. Quadrature over curved surfaces by extrapolation. Mathematics of Computation, 63(208):727– 740, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Li:1992:HDA

[LZ92] T. Y. Li and Zhonggang Zeng. Homotopy-determinant algorithm for solving nonsymmetric eigenvalue problems. Mathematics of Computation, 59 (200):483–502, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

MacLeod:1993:CEM

[Mac93] Allan J. MacLeod. Chebyshev expansions for modified Struve and related functions. Mathematics of Computation, 60(202):735–747, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mackey:1994:QJM

[Mac94a] Niloufer Mackey. A quaternion-jacobi method for symmetric matrices. In Gautschi [Gau94], pages 327–331. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

MacLeod:1994:CCS

[Mac94b] Allan J. MacLeod. On constructing Chebyshev series solutions of differential equations. In Gautschi [Gau94], pages 333–336. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

MacDonald:1997:HAS

[Mac97] D. A. MacDonald. On the highly accurate summation of certain series occurring in plate contact problems. Mathematics of Computation, 66(220):1619– 1627, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00869-
7&u=/mcom/1997-66-220/.

Maday:1990:ASP

[Mad90] Y. Maday. Analysis of spectral projectors in one-dimensional domains. Mathematics of Computation, 55(192):537– 562, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Maiorana:1991:CCR

[Mai91] James A. Maiorana. A classification of the cosets of the Reed–Muller code ℛ(1, 6). Mathematics of Computation, 57(195):403–414, July 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Makridakis:1993:FEA

[Mak93] Charalambos G. Makridakis. Finite element approximations of nonlinear elastic waves. Mathematics of Computation, 61(204):569–594, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Makroglou:1994:MCM

[Mak94] Athena Makroglou. Multiquadric collocation methods in the numerical solution of Volterra integral and integrodifferential equations. In Gautschi [Gau94], pages 337–341. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Maltby:1997:PPP

[Mal97] Roy Maltby. Pure product polynomials and the Prouhet–Tarry–Escott problem. Mathematics of Computation, 66(219):1323–1340, July 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                              http://www.ams.org/jourcgi/jour-
                                              pbprocess?fn=110&arg1=S0025-5718-
                                              97-00865-X&u=/mcom/1997-66-219/.

Martinez:1990:LCT

[Mar90] José Mario Martínez. Local convergence theory of inexact Newton methods based on structured least change updates. Mathematics of Computation, 55(191):143–167, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Martinez:1992:RBT

[Mar92] José Mario Martínez. On the relation between two local convergence theories of least-change secant update methods. Mathematics of Computation, 59(200): 457–481, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Martinez:1993:TSP

[Mar93] José Mario Martínez. A theory of secant preconditioners. Mathematics of Computation, 60(202):681–698, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Maschhoff:1994:MSL

[Mas94] Kristyn J. Maschhoff. Methods for solving large eigenvalue problems associated with configuration interaction electronic structure calculations. In Gautschi [Gau94], pages 343–347. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Maslen:1998:ECF

[Mas98] David K. Maslen. The efficient computation of Fourier transforms on the symmetric group. Mathematics of Computation, 67(223):1121– 1147, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00964-
8&u=/mcom/1998-67-223/.

Mattheij:1997:RDT

[Mat97] R. Mattheij. Reviews and descriptions of tables and books: 24. Numerical solution of initial-value problems in differential-algebraic equations, by K. E. Brenan, S. L. Campbell, and L. R. Petzold. Mathematics of Computation, 66 (220):??, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00843-
                                              0&u=/mcom/1997-66-220/.

Mayer:1992:MDD

[May92a] Daniel C. Mayer. Multiplicities of dihedral discriminants. Mathematics of Computation, 58(198):831–847, S55– S58, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mayer:1992:SMD

[May92b] Daniel C. Mayer. Supplement to “Multiplicities of dihedral discriminants”. Mathematics of Computation, 58(198): S55–S58, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mandel:1996:BDD

[MB96] Jan Mandel and Marian Brezina. Balancing domain decomposition for problems with large jumps in coefficients. Mathematics of Computation, 65(216):1387–1401, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00757-
                                              0&u=/mcom/1996-65-216/.

McKee:1994:CDP

[McK94] James McKee. Computing division polynomials. Mathematics of Computation, 63(208):767–771, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McKee:1999:SFF

[McK99] James McKee. Speeding Fermat’s factoring method. Mathematics of Computation, 68(228):1729–1737, October 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01133-
3&u=/mcom/1999-68-228/. This pa-

per present an O(N^{1/4+ϵ} integer factoring algorithm that never requires arithmetic on numbers larger than the one to be factored.

McLean:1994:NES

[McL94] W. McLean. Numerical evaluation of some trigonometric series. Mathematics of Computation, 63(207):271–275, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meade:1996:BRM

[Mea96] Douglas B. Meade. Book reviews:

The Maple V handbook, by Martha L. Abell and James P. Braselton. Mathematics of Computation, 65(216): ??, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Melman:1997:UCA

[Mel97]  A. Melman.  A unifying conver-

gence analysis of second-order methods for secular equations. Mathematics of Computation, 66(217):333– 344, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00787-
4&u=/mcom/1997-66-217/.

Melman:1998:ATO

[Mel98] A. Melman. Analysis of third-order methods for secular equations. Mathematics of Computation, 67(221):271– 286, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00884-
9&u=/mcom/1998-67-221/.

Miellou:1998:NCA

[MES98] J. C. Miellou, D. El Baz, and P. Spiteri. A new class of asynchronous iterative algorithms with order intervals. Mathematics of Computation, 67(221): 237–255, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00885-
0&u=/mcom/1998-67-221/.

Meyer:1996:BRD

[Mey96] Weston Meyer. Book reviews: Designing fair curves and surfaces, edited by Nickolas S. Sapidis. Mathematics of Computation, 65(214):892– 893, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Mitchell:1990:EEA

[Mit90]  Ricardo A. Mitchell.  Error esti-

mates arising from certain pseudorandom sequences in a quasi-random search method. Mathematics of Computation, 55(191):289–297, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moews:1991:SAC

[MM91] David Moews and Paul C. Moews. A search for aliquot cycles below 10^{10}. Mathematics of Computation, 57(196): 849–855, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mackenzie:1993:FVS

[MM93a] J. A. Mackenzie and K. W. Morton.

Finite volume solutions of convectiondiffusion test problems. Mathematics of Computation, 60(201):189–220, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moews:1993:SAC

[MM93b] David Moews and Paul C. Moews. A search for aliquot cycles and amicable pairs. Mathematics of Computation, 61 (204):935–938, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mastroianni:1994:PAF

[MM94a] Giuseppe Mastroianni and Giovanni Monegato. Polynomial approximations of functions with endpoint singularities and product integration formulas. Mathematics of Computation, 62(206): 725–738, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Morgan:1994:PNP

[MM94b] Ilene H. Morgan and Gary L. Mullen.

Primitive normal polynomials over finite fields. Mathematics of Computation, 63(208):759–765, S19–S23, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Morgan:1994:SPN

[MM94c] Ilene H. Morgan and Gary L. Mullen. Supplement to “Primitive normal polynomials over finite fields”. Mathematics of Computation, 63(208):S19– S23, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mastroianni:1995:CPI

[MM95] G. Mastroianni and G. Monegato. Convergence of product integration rules over (0, ∞) for functions with weak singularities at the origin. Mathematics of Computation, 64(209):237– 249, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mattman:1997:CGG

[MM97] Thomas Mattman and John McKay. Computation of Galois groups over function fields. Mathematics of Computation, 66(218):823–831, April 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
97-00831-4&u=/mcom/1997-66-218/.

Mischaikow:1998:CLE

[MM98] Konstantin Mischaikow and Marian Mrozek. Chaos in the Lorenz equations: A computer assisted proof. Part II: Details. Mathematics of Computation, 67(223):1023– 1046, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-98-00945-
4&u=/mcom/1998-67-223/.

Madych:1990:MIC

[MN90] W. R. Madych and S. A. Nelson. Multivariate interpolation and conditionally positive definite functions. II. Mathematics of Computation, 54(189):211– 230, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Muir:1993:OBC

[MO93] Paul Muir and Brynjulf Owren. Order barriers and characterizations for continuous mono-implicit Runge–Kutta schemes. Mathematics of Computation, 61(204):675–699, October 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mollin:1993:ACQ

[Mol93] R. A. Mollin. Ambiguous classes in quadratic fields. Mathematics of Computation, 61(203):355–360, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Montgomery:1993:NSP

[Mon93] Peter L. Montgomery. New solutions of a^{p−1} ≡ 1 (mod p^{2}). Mathematics of Computation, 61(203):361–363, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Monegato:1994:NEC

[Mon94a] Giovanni Monegato. The numerical evaluation of a 2-D Cauchy principal value integral arising in boundary integral equation methods. Mathematics of Computation, 62(206):765–777, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Montgomery:1994:SRP

[Mon94b] Peter L. Montgomery. Square roots of products of algebraic numbers. In Gautschi [Gau94], pages 567–571. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Morain:1989:DEP

[Mor89] François Morain. On the lcm of the differences of eight primes. Mathematics of Computation, 52(185):225– 229, January 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigenda [Mor90].

Morain:1990:CLD

[Mor90] François Morain. Corrigenda: “On the lcm of the differences of eight primes” [Math. Comput. 52 (1989), no. 185, 225–229]. Mathematics of Computation, 54(190):911, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Mor89].

Morgan:1996:RAM

[Mor96] Ronald B. Morgan. On restarting the Arnoldi method for large nonsymmetric eigenvalue problems. Mathematics of Computation, 65(215):1213– 1230, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00745-
4&u=/mcom/1996-65-215/.

Morini:1999:CBI

[Mor99] Benedetta Morini. Convergence behaviour of inexact Newton methods. Mathematics of Computation, 68 (228):1605–1613, October 1999. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
99-01135-7&u=/mcom/1999-68-228/.

Mossinghoff:1998:PSM

[Mos98] Michael J. Mossinghoff. Polynomials with small Mahler measure. Mathematics of Computation, 67(224):1697– 1705, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-01006-
0&u=/mcom/1998-67-224/.

Milner:1995:MFE

[MP95] F. A. Milner and E.-J. Park. A mixed finite element method for a strongly nonlinear second-order elliptic problem. Mathematics of Computation, 64(211): 973–988, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Malmliden:1996:FIM

[MP96] Johan F. Malmliden and N. Anders Petersson. A fast iterative method to compute the flow around a submerged body. Mathematics of Computation, 65(215): 1067–1083, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00715-
6&u=/mcom/1996-65-215/.

Mathon:1995:ACS

[MPR95] Rudolf A. Mathon, Kevin T. Phelps, and Alexander Rosa. Addendum: “A class of Steiner triple systems of order 21 and associated Kirkman systems” [Math. Comp. 37 (1981), no. 155, 209– 222, MR 82k:05022]. Mathematics of Computation, 64(211):1355–1356, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mignotte:1996:CSF

[MPR96] Maurice Mignotte, Attila Pethö, and Ralf Roth. Complete solutions of a family of quartic Thue and index form equations. Mathematics of Computation, 65(213):341– 354, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00662-
                                              X&u=/mcom/1996-65-213/.

Mossinghoff:1998:PPA

[MPV98] Michael J. Mossinghoff, Christopher G. Pinner, and Jeffrey D. Vaaler. Perturbing polynomials with all their roots on the unit circle. Mathematics of Computation, 67(224):1707– 1726, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-01007-
                                              2&u=/mcom/1998-67-224/.

Montgomery:1990:FEF

[MS90] Peter L. Montgomery and Robert D. Silverman. An FFT extension to the P −1 factoring algorithm. Mathematics of Computation, 54(190):839–854, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Marek:1993:ISI

[MS93] Ivo Marek and Daniel B. Szyld. Iterative and semi-iterative methods for computing stationary probability vectors of Markov operators. Mathematics of Computation, 61(204):719– 731, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Maeztu:1995:CSI

[MS95] J. I. Maeztu and E. Sáinz de la Maza. Consistent structures of invariant quadrature rules for the n-simplex. Mathematics of Computation, 64(211): 1171–1192, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meerbergen:1997:IRA

[MS97a] Karl Meerbergen and Alastair Spence.

Implicitly restarted Arnoldi with purification for the shift-invert transformation. Mathematics of Computation, 66(218):667–689, April 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
97-00844-2&u=/mcom/1997-66-218/.

Morken:1997:GFH

[MS97b] Knut Mørken and Karl Scherer. A general framework for high-accuracy parametric interpolation. Mathematics of Computation, 66(217):237– 260, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00796-
5&u=/mcom/1997-66-217/.

Monegato:1998:HOM

[MS98] G. Monegato and L. Scuderi. High order methods for weakly singular integral equations with nonsmooth input functions. Mathematics of Computation, 67(224):1493– 1515, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-01005-
9&u=/mcom/1998-67-224/.

Morton:1997:ACV

[MSS97] K. W. Morton, Martin Stynes, and Endre Süli. Analysis of a cell-vertex finite volume method for convectiondiffusion problems. Mathematics of Computation, 66(220):1389–1406, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00886-
7&u=/mcom/1997-66-220/.

Malta:1996:NIF

[MST96] Iaci Malta, Nicolau C. Saldanha, and Carlos Tomei. The numerical inversion of functions from the plane to the plane. Mathematics of Computation, 65(216):1531– 1552, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00770-
3&u=/mcom/1996-65-216/.

Muller:1999:CDL

[MST99] Volker Müller, Andreas Stein, and Christoph Thiel. Computing discrete logarithms in real quadratic congruence function fields of large genus. Mathematics of Computation, 68(226): 807–822, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01040-
6&u=/mcom/1999-68-226/.

Mignotte:1993:ASC

[MT93] M. Mignotte and N. Tzanakis. Arithmetical study of a certain ternary recurrence sequence and related questions. Mathematics of Computation, 61(204): 901–913, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moree:1994:DPI

[MTU94] P. Moree, H. J. J. Te Riele, and J. Urbanowicz. Divisibility properties of integers x, k satisfying 1^{k}+⋯+(x−1)^{k} = x^{k}. Mathematics of Computation, 63 (208):799–815, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Muller:1997:APA

[Mül97] Jürgen Müller. Accelerated polynomial approximation of finite order entire functions by growth reduction. Mathematics of Computation, 66(218): 743–761, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00832-
6&u=/mcom/1997-66-218/.

Muskat:1993:GFL

[Mus93] Joseph B. Muskat. Generalized Fibonacci and Lucas sequences and rootfinding methods. Mathematics of Computation, 61(203):365–372, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Menezes:1993:CPE

[MVZ93] Alfred J. Menezes, Scott A. Vanstone, and Robert J. Zuccherato. Counting points on elliptic curves over f_{2^{m}}. Mathematics of Computation, 60(201):407– 420, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meek:1992:CST

[MW92a] D. S. Meek and D. J. Walton. Clothoid spline transition spirals. Mathematics of Computation, 59(199):117–133, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mollin:1992:RQF

[MW92b] R. A. Mollin and H. C. Williams. On real quadratic fields of class number two. Mathematics of Computation, 59 (200):625–632, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCormick:1994:WCF

[MW94a] Kent McCormick and Raymond O. Wells, Jr. Wavelet calculus and finite difference operators. Mathematics of Computation, 63(207):155–173, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mollin:1994:QRC

[MW94b] R. A. Mollin and H. C. Williams. Quadratic residue covers for certain real quadratic fields. Mathematics of Computation, 62(206):885–897, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

N:1993:BRB

[N.93] H. N. Book review: Random Number Generators and Simulation, by István Deák. Mathematics of Computation, 60 (201):442, January 1993. URL http://

                                              www.jstor.org/stable/2153189. See

[Deá90].

Nedelec:1994:NTU

[Néd94] J.-C. Nédélec. New trends in the use and analysis of integral equations. In Gautschi [Gau94], pages 151–176. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt.

1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Nessyahu:1996:CRA

[Nes96] Haim Nessyahu. Convergence rate of approximate solutions to weakly coupled nonlinear systems. Mathematics of Computation, 65(214):575– 586, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00716-
8&u=/mcom/1996-65-214/.

Neuman:1994:NFB

[Neu94] Edward Neuman. A new formula for box splines on three-directional meshes. Mathematics of Computation, 62(205): 227–229, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1995:NFA

[NG95] Harald Niederreiter and Rainer Göttfert.

On a new factorization algorithm for polynomials over finite fields. Mathematics of Computation, 64(209):347– 353, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nair:1997:TBR

[NHA97] M. Thamban Nair, Markus Hegland, and Robert S. Anderssen. The tradeoff between regularity and stability in Tikhonov regularization. Mathematics of Computation, 66(217):193– 206, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

                                              fn=110&arg1=S0025-5718-97-00811-
                                              9&u=/mcom/1997-66-217/.

Nicely:1999:NMP

[Nic99] Thomas R. Nicely. New maximal prime gaps and first occurrences. Mathematics of Computation, 68(227):1311– 1315, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01065-
                                              0&u=/mcom/1999-68-227/.

Niederreiter:1990:LBD

[Nie90] Harald Niederreiter. Lower bounds for the discrepancy of inversive congruential pseudorandom numbers. Mathematics of Computation, 55(191):277– 287, July 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1992:EEL

[Nie92a] Harald Niederreiter. The existence of efficient lattice rules for multidimensional numerical integration. Mathematics of Computation, 58(197):305– 314, S7–S16, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1992:SEE

[Nie92b] Harald Niederreiter. Supplement to “The existence of efficient lattice rules for multidimensional numerical integration” by Harald Niederreiter. Mathematics of Computation, 58(197):S7– S16, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1994:FPF

[Nie94] Harald Niederreiter. Factoring polynomials over finite fields using differential equations and normal bases. Mathematics of Computation, 62(206): 819–830, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1995:PVG

[Nie95] Harald Niederreiter. Pseudorandom vector generation by the multiplerecursive matrix method. Mathematics of Computation, 64(209):279– 294, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1996:BRP

[Nie96a] Harald Niederreiter. Book reviews: A primer for the Monte Carlo method, by Ilya M. Sobol’. Mathematics of Computation, 65(216):??, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Niederreiter:1996:BRC

[Nie96b] Harald Niederreiter. Book reviews:

Computational number theory and digital signal processing: Fast algorithms and error control techniques, by Hari Krishna, Bal Krishna, Kuo-Yu Lin and Jenn-Dong Sun. Mathematics of Computation, 65(214):893– 894, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Niederreiter:1996:BRD

[Nie96c] Harald Niederreiter. Book reviews:

The DFT: An owner’s manual for the discrete Fourier transform, by William L. Briggs and Van Emden Henson. Mathematics of Computation, 65 (215):??, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00705-
                                              3&u=/mcom/1996-65-215/.

Niijima:1992:PEB

[Nii92] Koichi Niijima. A posteriori error bounds for piecewise linear approximate solutions of elliptic equations of monotone type. Mathematics of Computation, 58(198):549–560, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nochetto:1995:PPE

[Noc95] Ricardo H. Nochetto. Pointwise a posteriori error estimates for elliptic problems on highly graded meshes. Mathematics of Computation, 64(209): 1–22, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nocedal:1996:BRN

[Noc96] Jorge Nocedal. Book reviews: Nonlinear programming, by Olvi L. Mangasarian. Mathematics of Computation, 65 (214):884, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00700-
                                              4&u=/mcom/1996-65-214/.

Noelle:1996:NEI

[Noe96] Sebastian Noelle. A note on entropy inequalities and error estimates for higher-order accurate finite volume schemes on irregular families of grids. Mathematics of Computation, 65(215):

1155–1163, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00737-
5&u=/mcom/1996-65-215/.

Notaris:1992:GKQ

[Not92]  Sotirios E. Notaris.  On Gauss–

Kronrod quadrature formulae of Chebyshev type. Mathematics of Computation, 58(198):745–753, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nochetto:1991:AFE

[NPV91a] R. H. Nochetto, M. Paolini, and C. Verdi. An adaptive finite element method for two-phase Stefan problems in two space dimensions. part I: Stability and error estimates. Mathematics of Computation, 57(195):73–108, S1–S11, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nochetto:1991:SAF

[NPV91b] R. H. Nochetto, M. Paolini, and C. Verdi. Supplement to an adaptive finite element method for two-phase Stefan problems in two space dimensions. part I: Stability and error estimates. Mathematics of Computation, 57(195): s1–s11, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Neittaanmaki:1995:DFE

[NRS95] P. Neittaanmäki, V. Rivkind, and G. Serëgin. A dual finite element approach for stresses of elasto-perfectly plastic bodies. Mathematics of Computation, 64(212):1455–1462, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Novak:1995:ACO

[NRW95] Erich Novak, Klaus Ritter, and Henryk Woźniakowski. Average-case optimality of a hybrid secant-bisection method. Mathematics of Computation, 64(212): 1517–1539, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niederreiter:1990:LRM

[NS90] Harald Niederreiter and Ian H. Sloan. Lattice rules for multiple integration and discrepancy. Mathematics of Computation, 54(189):303–312, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Niklasch:1998:EUF

[NS98] G. Niklasch and N. P. Smart. Exceptional units in a family of quartic number fields. Mathematics of Computation, 67(222):759– 772, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00958-
                                              2&u=/mcom/1998-67-222/.

Nagai:1998:TME

[NTS98] Atsushi Nagai, Tetsuji Tokihiro, and Junkichi Satsuma. The Toda molecule equation and the ε-algorithm. Mathematics of Computation, 67(224):1565– 1575, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00987-
                                              9&u=/mcom/1998-67-224/.

Narcowich:1994:GHI

[NW94] Francis J. Narcowich and Joseph D. Ward. Generalized Hermite interpolation via matrix-valued conditionally positive definite functions. Mathematics of Computation, 63(208):661– 687, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nicolaides:1995:SCN

[NW95] R. A. Nicolaides and Noel J. Walkington. Strong convergence of numerical solutions to degenerate variational problems. Mathematics of Computation, 64(209):117–127, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nicolaides:1996:ACM

[NW96] R. A. Nicolaides and X. Wu. Analysis and convergence of the MAC scheme. II. Navier–Stokes equations. Mathematics of Computation, 65(213):29– 44, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00665-
5&u=/mcom/1996-65-213/.

Nicolaides:1998:CAC

[NW98] R. A. Nicolaides and D.-Q. Wang.

Convergence analysis of a covolume scheme for Maxwell’s equations in three dimensions. Mathematics of Computation, 67(223):947– 963, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00971-
5&u=/mcom/1998-67-223/.

Ni:1997:SLM

[NY97] Q. Ni and Y. Yuan. A subspace limited memory quasi-Newton algorithm for large-scale nonlinear bound constrained optimization. Mathematics of Computation, 66(220):1509– 1520, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00866-
                                              1&u=/mcom/1997-66-220/.

Odlyzko:1993:IAV

[Odl93] Andrew M. Odlyzko. Iterated absolute values of differences of consecutive primes. Mathematics of Computation, 61(203):373–380, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Odlyzko:1994:ACN

[Odl94] Andrew M. Odlyzko. Analytic computations in number theory. In Gautschi [Gau94], pages 451–463. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Olivier:1992:CSF

[Oli92] M. Olivier. The computation of sextic fields with a cubic subfield and no quadratic subfield. Mathematics of Computation, 58(197):419–432, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Oliveira:1998:NPC

[Oli98] Suely Oliveira. A new parallel chasing algorithm for transforming arrowhead matrices to tridiagonal form. Mathematics of Computation, 67(221):221– 235, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-98-00895-
3&u=/mcom/1998-67-221/.

OliveiraeSilva:1999:MES

[Oli99]  Tomás Oliveira e Silva.  Maxi-

mum excursion and stopping time record-holders for the 3x + 1 problem: computational results. Mathematics of Computation, 68(225):371– 384, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01031-
5&u=/mcom/1999-68-225/.

Olsson:1995:SPPa

[Ols95a] Pelle Olsson. Summation by parts, projections, and stability. I. Mathematics of Computation, 64(211):1035–1065, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olsson:1995:SPPb

[Ols95b] Pelle Olsson. Summation by parts, projections, and stability. II. Mathematics of Computation, 64(212):1473– 1493, October 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olsson:1995:SSP

[Ols95c] Pelle Olsson. Supplement to “Summation by parts, projections, and stability, I”. Mathematics of Computation, 64(211):S23–S26, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ostermann:1992:RKM

[OR92] A. Ostermann and M. Roche. Runge– Kutta methods for partial differential equations and fractional orders of convergence. Mathematics of Computation, 59(200):403–420, October 1992.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

ORiordan:1991:GUC

[OS91] Eugene O’Riordan and Martin Stynes. A globally uniformly convergent finite element method for a singularly perturbed elliptic problem in two dimensions. Mathematics of Computation, 57 (195):47–62, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Okunbor:1992:ECM

[OS92] Daniel Okunbor and Robert D. Skeel. Explicit canonical methods for Hamiltonian systems. Mathematics of Computation, 59(200):439–455, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Omrani:1997:CIR

[OS97] A. Omrani and A. Shokrollahi. Computing irreducible representations of supersolvable groups over small finite fields. Mathematics of Computation, 66(218):779– 786, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00839-
9&u=/mcom/1997-66-218/.

Oswald:1996:PND

[Osw96] Peter Oswald. Preconditioners for nonconforming discretizations. Mathematics of Computation, 65(215):923– 941, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00717-
X&u=/mcom/1996-65-215/.

Oswald:1999:RBP

[Osw99] Peter Oswald. On the robustness of the BPX-preconditioner with respect to jumps in the coefficients. Mathematics of Computation, 68(226):633– 650, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01041-
8&u=/mcom/1999-68-226/.

OShea:1994:CLN

[OW94] Donal O’Shea and Les Wilson. Computing limiting normals to real surfaces. In Gautschi [Gau94], pages 349– 353. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Osipenko:1997:OIA

[OW97] K. Yu. Osipenko and K. Wilderotter.

Optimal information for approximating periodic analytic functions. Mathematics of Computation, 66(220):1579– 1592, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00896-
X&u=/mcom/1997-66-220/.

Owren:1991:OBC

[OZ91] Brynjulf Owren and Marino Zennaro.

Order barriers for continuous explicit Runge–Kutta methods. Mathematics of Computation, 56(194):645–661, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Palencia:1994:SVS

[Pal94] C. Palencia. On the stability of variable stepsize rational approximations of holomorphic semigroups. Mathematics of Computation, 62(205):93– 103, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Palencia:1995:SRM

[Pal95] C. Palencia. Stability of rational multistep approximations of holomorphic semigroups. Mathematics of Computation, 64(210):591–599, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

PalamaraOrsi:1996:PIV

[Pal96] Annamaria Palamara Orsi. Product integration for Volterra integral equations of the second kind with weakly singular kernels. Mathematics of Computation, 65(215):1201– 1212, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00736-
3&u=/mcom/1996-65-215/.

Pan:1990:CDS

[Pan90] Victor Pan. On computations with dense structured matrices. Mathematics of Computation, 55(191):179–190, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Paolini:1997:QOE

[Pao97] Maurizio Paolini. A quasi-optimal error estimate for a discrete singularly perturbed approximation to the prescribed curvature problem. Mathematics of Computation, 66(217):45– 67, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-97-00771-
0&u=/mcom/1997-66-217/.

Pappalardi:1997:RAC

[Pap97] Francesco Pappalardi. On the rrank Artin Conjecture. Mathematics of Computation, 66(218):853– 868, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00805-
3&u=/mcom/1997-66-218/.

Pearson:1996:ICD

[Pea96] Kimberly Pearson. Integral cohomology and detection of w-basic 2-groups. Mathematics of Computation, 65(213): 291–306, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00679-
5&u=/mcom/1996-65-213/.

Peherstorfer:1990:LCO

[Peh90a] Franz Peherstorfer. Linear combinations of orthogonal polynomials generating positive quadrature formulas. Mathematics of Computation, 55(191): 231–241, July 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Peherstorfer:1990:SPG

[Peh90b] Franz Peherstorfer. On Stieltjes polynomials and Gauss–Kronrod quadrature. Mathematics of Computation, 55(192): 649–664, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Peherstorfer:1993:RGQ

[Peh93]  F. Peherstorfer.  On the remain-

der of Gaussian quadrature formulas for Bernstein–Szegő weight functions. Mathematics of Computation, 60(201): 317–325, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pena:1997:BSO

[Peñ97]  J. M. Peña.  B-splines and op-

timal stability. Mathematics of Computation, 66(220):1555–1560, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00897-
1&u=/mcom/1997-66-220/.

Peralta:1992:DQR

[Per92] René Peralta. On the distribution of quadratic residues and nonresidues modulo a prime number. Mathematics of Computation, 58(197):433– 440, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Peszynska:1996:FEA

[Pes96] Małgorzata Peszyńska. Finite element approximation of diffusion equations with convolution terms. Mathematics of Computation, 65(215):1019– 1037, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00738-
7&u=/mcom/1996-65-215/.

Petho:1991:CSF

[Pet91] Attila Pethö. Complete solutions to families of quartic Thue equations. Mathematics of Computation, 57(196): 777–798, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Phelps:1990:CLS

[Phe90a] Kevin T. Phelps. Cyclic large sets of Steiner triple systems of order 15. Mathematics of Computation, 55(192):821– 824, S1–S6, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Phelps:1990:SCL

[Phe90b] Kevin T. Phelps. Supplement to cyclic large sets of Steiner triple systems of order 15. Mathematics of Computation, 55(192):S1–S6, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Phillips:1995:CCS

[Phi95a] Ralph Phillips. Conjugacy classes of Γ(2) and spectral rigidity. Mathematics of Computation, 64(211):1287–1306, S35–S37, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Phillips:1995:SCC

[Phi95b] Ralph Phillips. Supplement to “Conjugacy classes of G(2) and spectral rigidity”. Mathematics of Computation, 64 (211):S35–S37, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pihko:1993:MES

[Pih93] Jukka Pihko. On the minimal elements for the sequence of all powers in the Lemoine–Kátai algorithm. Mathematics of Computation, 60(201):425– 430, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pila:1990:FMA

[Pil90] J. Pila. Frobenius maps of Abelian varieties and finding roots of unity in finite fields. Mathematics of Computation, 55(192):745–763, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pinch:1993:CN

[Pin93a] R. G. E. Pinch. The Carmichael numbers up to 10^{15}. Mathematics of Computation, 61(203):381–391, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pinch:1993:SQP

[Pin93b] R. G. E. Pinch. Squares in quadratic progression. Mathematics of Computation, 60(202):841–845, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pinner:1999:DRP

[Pin99] Christopher Pinner. Double roots of [−1, 1] power series and related matters. Mathematics of Computation, 68(227): 1149–1178, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01042-
X&u=/mcom/1999-68-227/.

Portugal:1994:CBP

[PJV94] Luís F. Portugal, Joaquím J. Júdice, and Luís N. Vicente. A comparison of block pivoting and interior-point algorithms for linear least squares problems with nonnegative variables. Mathematics of Computation, 63(208):625– 643, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pritchard:1995:TTP

[PMT95] Paul A. Pritchard, Andrew Moran, and Anthony Thyssen. Twenty-two primes in arithmetic progression. Mathematics of Computation, 64(211):1337– 1339, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pomerance:1994:NFS

[Pom94] Carl Pomerance. The number field sieve. In Gautschi [Gau94], pages 465– 480. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Powell:1997:BMC

[Pow97] Corey Powell. Bounds for multiplicative cosets over fields of prime order. Mathematics of Computation, 66(218): 807–822, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00797-
7&u=/mcom/1997-66-218/.

Plesken:1993:CIL

[PP93] W. Plesken and M. Pohst. Constructing integral lattices with prescribed minimum. II. Mathematics of Computation, 60(202):817–825, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Papakostas:1996:FFO

[PP96] S. N. Papakostas and G. Papageorgiou. A family of fifth-order Runge–Kutta pairs. Mathematics of Computation, 65(215):1165– 1181, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00718-
1&u=/mcom/1996-65-215/.

Paulus:1999:RIQ

[PR99] Sachar Paulus and Hans-Georg Rück. Real and imaginary quadratic representations of hyperelliptic function fields. Mathematics of Computation, 68(227): 1233–1241, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01066-
                                              2&u=/mcom/1999-68-227/.

Potts:1998:FAD

[PST98] Daniel Potts, Gabriele Steidl, and Manfred Tasche. Fast algorithms for discrete polynomial transforms. Mathematics of Computation, 67(224):1577– 1590, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00975-
                                              2&u=/mcom/1998-67-224/.

Parady:1990:LKT

[PSZ90] B. K. Parady, Joel F. Smith, and Sergio E. Zarantonello. Largest known twin primes. Mathematics of Computation, 55(191):381–382, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plonka:1992:EAP

[PT92] G. Plonka and M. Tasche. Efficient algorithms for periodic Hermite spline interpolation. Mathematics of Computation, 58(198):693–703, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Plonka:1994:CHS

[PT94] Gerlind Plonka and Manfred Tasche. Cardinal Hermite spline interpolation with shifted nodes. Mathematics of Computation, 63(208):645–659, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pohst:1993:CUG

[Pv93] M. Pohst and J. Graf von Schmettow.

On the computation of unit groups and class groups of totally complex quartic fields. Mathematics of Computation, 60(202):793–800, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pohst:1998:TUG

[PW98] M. Pohst and K. Wildanger. Tables of unit groups and class groups of quintic fields and a regulator bound. Mathematics of Computation, 67(221):361– 367, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00927-
2&u=/mcom/1998-67-221/.

Quak:1996:BRH

[Qua96a] Ewald Quak. Book reviews: Handbook on splines for the user, by Eugene V. Shikin and Alexander I. Plis. Mathematics of Computation, 65(216): ??, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00768-
5&u=/mcom/1996-65-216/.

Quak:1996:TWH

[Qua96b] Ewald Quak. Trigonometric wavelets for Hermite interpolation. Mathematics of Computation, 65(214):683– 722, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00719-
                                              3&u=/mcom/1996-65-214/.

Quak:1997:BRS

[Qua97] Ewald Quak. Book reviews: Selected topics in approximation and computation, by Marek A. Kowalski, Krszysztof A. Sikorski and Frank Stenger. Mathematics of Computation, 66(219):??, July 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1990:GNR

[Rab90a] Philip Rabinowitz. Generalized noninterpolatory rules for Cauchy principal value integrals. Mathematics of Computation, 54(189):271–279, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1990:NEC

[Rab90b] Philip Rabinowitz. Numerical evaluation of Cauchy principal value integrals with singular integrands. Mathematics of Computation, 55(191):265–276, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1991:UCR

[Rab91] Philip Rabinowitz. Uniform convergence results for Cauchy principal value integrals. Mathematics of Computation, 56(194):731–740, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Reichel:1991:DLS

[RAG91] L. Reichel, G. S. Ammar, and W. B. Gragg. Discrete least squares approximation by trigonometric polynomials. Mathematics of Computation, 57(195): 273–289, July 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ray:1994:LPV

[Ray94] Gary A. Ray. A locally parameterized version of Lehmer’s problem. In Gautschi [Gau94], pages 573–576. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Riesel:1994:GFN

[RB94] Hans Riesel and Anders Björn. Generalized Fermat numbers. In Gautschi [Gau94], pages 583–587. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Renaut:1990:TSR

[Ren90] R. A. Renaut. Two-step Runge– Kutta methods and hyperbolic partial differential equations. Mathematics of Computation, 55(192):563– 579, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Restrepo:1997:BRM

[Res97] Juan Mario Restrepo. Book reviews:

Mathematics, climate and environment, edited by J.-I. Diaz and J.-L. Lions. Mathematics of Computation, 66 (218):??, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00856-
9&u=/mcom/1997-66-218/.

Rivara:1996:TLS

[RI96] María-Cecilia Rivara and Gabriel Iribarren. The 4-triangles longestside partition of triangles and linear refinement algorithms. Mathematics of Computation, 65(216):1485– 1502, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00772-
                                              7&u=/mcom/1996-65-216/.

Richter:1990:FEM

[Ric90] Gerard R. Richter. A finite element method for time-dependent convectiondiffusion equations. Mathematics of Computation, 54(189):81–106, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Richter:1992:DGM

[Ric92] Gerard R. Richter. The discontinuous Galerkin method with diffusion. Mathematics of Computation, 58(198): 631–643, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Reichel:1991:CVS

[RO91] Lothar Reichel and Gerhard Opfer. Chebyshev–Vandermonde systems. Mathematics of Computation, 57(196):703– 721, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Robinson:1993:NHS

[Rob93] Raphael M. Robinson. Numbers having m small mth roots mod p. Mathematics of Computation, 61(203):393–413, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Robinson:1994:OSC

[Rob94]  Mark P. Robinson.  Orthogonal

spline collocation solution of nonlinear Schrödinger equations. In Gautschi [Gau94], pages 355–360. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Roelse:1999:FHD

[Roe99] Peter Roelse. Factoring high-degree polynomials over F_{2} with Niederreiter’s algorithm on the IBM SP2. Mathematics of Computation, 68(226):869– 880, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01008-
X&u=/mcom/1999-68-226/.

Rojas:1994:WIC

[Roj94] Raúl Rojas. Who invented the computer? The debate from the viewpoint of computer architecture. In Gautschi [Gau94], pages 361–365. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Rose:1995:CEC

[Ros95a] H. E. Rose. On a class of elliptic curves with rank at most two. Mathematics of Computation, 64(211):1251–1265, S27–S34, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rose:1995:SCE

[Ros95b] H. E. Rose. Supplement to “On a class of elliptic curves with rank at most 2”.

Mathematics of Computation, 64(211): S27–S34, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabier:1994:CIP

[RR94] Patrick J. Rabier and Werner C. Rheinboldt. On the computation of impasse points of quasi-linear differentialalgebraic equations. Mathematics of Computation, 62(205):133–154, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ramare:1996:PAP

[RR96] Olivier Ramaré and Robert Rumely. Primes in arithmetic progressions. Mathematics of Computation, 65(213): 397–425, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00669-
                                              2&u=/mcom/1996-65-213/.

Rahman:1990:QFF

[RS90] Qazi I. Rahman and Gerhard Schmeisser. Quadrature formulae and functions of exponential type. Mathematics of Computation, 54(189):245–270, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rundell:1992:RTC

[RS92] William Rundell and Paul E. Sacks. Reconstruction techniques for classical inverse Sturm–Liouville problems. Mathematics of Computation, 58(197):161– 183, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rahman:1994:QFE

[RS94] Qazi Ibadur Rahman and Gerhard Schmeisser. A quadrature formula for entire functions of exponential type. Mathematics of Computation, 63(207): 215–227, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rhin:1995:AMM

[RS95] Georges Rhin and Christopher Smyth.

On the absolute Mahler measure of polynomials having all zeros in a sector. Mathematics of Computation, 64(209): 295–304, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rodriguez:1996:OCE

[RS96a] Rodolfo Rodríguez and Jorge E.

Solomin. The order of convergence of eigenfrequencies in finite element approximations of fluid-structure interaction problems. Mathematics of Computation, 65(216):1463– 1475, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00739-
9&u=/mcom/1996-65-216/.

Ron:1996:SPD

[RS96b] Amos Ron and Xingping Sun. Strictly positive definite functions on spheres in Euclidean spaces. Mathematics of Computation, 65(216):1513– 1530, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00780-
6&u=/mcom/1996-65-216/.

Ron:1998:CST

[RS98] Amos Ron and Zuowei Shen. Compactly supported tight affine spline frames in L_{2}(ℝ^{d}). Mathematics of Computation, 67(221):191–207, January 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00898-
                                              9&u=/mcom/1998-67-221/.

Ruck:1999:DLD

[Rüc99]  Hans-Georg Rück.  On the dis-

crete logarithm in the divisor class group of curves. Mathematics of Computation, 68(226):805–806, April 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01043-
                                              1&u=/mcom/1999-68-226/.

Rudin:1998:BRC

[Rud98] Bernard D. Rudin. Book reviews: Computer algebra in industry: Problem solving in practice, edited by Arjeh M. Cohen. Mathematics of Computation, 67 (224):??, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00989-
                                              2&u=/mcom/1998-67-224/.

Rump:1990:RSA

[Rum90] Siegfried M. Rump. Rigorous sensitivity analysis for systems of linear and nonlinear equations. Mathematics of Computation, 54(190):721–736, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rumely:1993:NCC

[Rum93a] Robert Rumely. Numerical computations concerning the ERH. Mathematics of Computation, 61(203):415–440, 517–523, S17–S23, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rumely:1993:SCC

[Rum93b] Robert Rumely. Supplement to “Numerical computations concerning the ERH”. Mathematics of Computation, 61(203):S17–S23, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rump:1996:EER

[Rum96] S. M. Rump. Expansion and estimation of the range of nonlinear functions. Mathematics of Computation, 65(216):1503–1512, October 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00773-
9&u=/mcom/1996-65-216/.

Rusten:1996:IPP

[RVW96] Torgeir Rusten, Panayot S. Vassilevski, and Ragnar Winther. Interior penalty preconditioners for mixed finite element approximations of elliptic problems. Mathematics of Computation, 65(214): 447–466, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00720-
X&u=/mcom/1996-65-214/.

Anonymous:1992:TEA

[RW92] S. M. Rudolpher and P. C. Watson.

Table errata Abramowitz and Stegun, editors. Mathematics of Computation, 59(200):727, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rusten:1993:SPE

[RW93] Torgeir Rusten and Ragnar Winther. Substructure preconditioners for elliptic saddle point problems. Mathematics of Computation, 60(201):23–48, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sheng:1994:PSH

[SA94] Qin Sheng and Ravi P. Agarwal. Periodic solutions of higher-order difference equations in two independent variables. In Gautschi [Gau94], pages 377– 381. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Sandling:1992:PUG

[San92] Robert Sandling. Presentations for unit groups of modular group algebras of groups of order 16. Mathematics of Computation, 59(200):689– 701, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saouter:1998:COG

[Sao98] Yannick Saouter. Checking the odd Goldbach conjecture up to 10^{20}. Mathematics of Computation, 67(222):863– 866, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00928-
4&u=/mcom/1998-67-222/.

Suri:1995:LEF

[SBS95] Manil Suri, Ivo Babuška, and Christoph Schwab. Locking effects in the finite element approximation of plate models. Mathematics of Computation, 64(210): 461–482, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schneider:1991:RTM

[Sch91]  Claus Schneider.  Rosenbrock-type

methods adapted to differentialalgebraic systems. Mathematics of Computation, 56(193):201–213, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schmitt:1992:KAM

[Sch92] Bernhard A. Schmitt. Krylov approximations for matrix square roots in stiff boundary value problems. Mathematics of Computation, 58(197):191– 212, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schinzel:1993:ETP

[Sch93a] A. Schinzel. An extension of the theorem on primitive divisors in algebraic number fields. Mathematics of Computation, 61(203):441–444, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schneider:1993:SSP

[Sch93b] Gerhard A. Schneider. Supplement to “The structure of the projective indecomposable modules of the Suzuki group Sz(8) in characteristic 2”. Mathematics of Computation, 60(202):S29– S32, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schneider:1993:SPI

[Sch93c] Gerhard J. A. Schneider. The structure of the projective indecomposable modules of the Suzuki group sz(8) in characteristic 2. Mathematics of Computation, 60(202):779–786, S29–S32, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schumaker:1994:AMS

[Sch94] Larry L. Schumaker. Applications of multivariate splines. In Gautschi [Gau94], pages 177–203. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Schoof:1996:BRC

[Sch96a]  René Schoof.  Book reviews: A

course in computational algebraic number theory, by Henri Cohen. Mathematics of Computation, 65(213): ??, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Schoof:1996:BRI

[Sch96b] René Schoof. Book reviews: Introduction to finite fields and their applications, by Rudolf Lidl and Harald Niederreiter. Mathematics of Computation, 65(215):??, July 1996. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
96-00705-3&u=/mcom/1996-65-215/.

Schira:1997:RTA

[Sch97] Thomas Schira. The remainder term for analytic functions of symmetric Gaussian quadratures. Mathematics of Computation, 66(217):297– 310, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00798-
9&u=/mcom/1997-66-217/.

Schatz:1998:PEE

[Sch98a]  Alfred H. Schatz.  Pointwise er-

ror estimates and asymptotic error expansion inequalities for the finite element method on irregular grids: Part I. Global estimates. Mathematics of Computation, 67(223):877– 899, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00959-
4&u=/mcom/1998-67-223/.

Schoof:1998:MCG

[Sch98b] René Schoof. Minus class groups of the fields of the l-th roots of unity. Mathematics of Computation, 67(223): 1225–1245, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00939-
9&u=/mcom/1998-67-223/.

Schaback:1999:IEB

[Sch99] R. Schaback. Improved error bounds for scattered data interpolation by radial basis functions. Mathematics of Computation, 68(225):201– 216, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01009-
                                              1&u=/mcom/1999-68-225/.

Stroeker:1999:EBD

[SdW99] Roelof J. Stroeker and Benjamin M. M. de Weger. Elliptic binomial Diophantine equations. Mathematics of Computation, 68(227):1257– 1281, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01047-
                                              9&u=/mcom/1999-68-227/.

Seiler:1999:NIC

[Sei99] Werner M. Seiler. Numerical integration of constrained Hamiltonian systems using Dirac brackets. Mathematics of Computation, 68(226):661– 681, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01010-
                                              8&u=/mcom/1999-68-226/.

Selmane:1999:NPN

[Sel99] Schehrazad Selmane. Non-primitive number fields of degree eight and of signature (2, 3), (4, 2), and (6, 1) with small discriminant. Mathematics of Computation, 68(225):333– 344, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-00998-
                                              9&u=/mcom/1999-68-225/.

Semaev:1998:EDL

[Sem98a] I. A. Semaev. Evaluation of discrete logarithms in a group of p-torsion points of an elliptic curve in characteristic p.

Mathematics of Computation, 67(221): 353–356, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00887-
4&u=/mcom/1998-67-221/.

Semaev:1998:AED

[Sem98b] Igor A. Semaev. An algorithm for evaluation of discrete logarithms in some nonprime finite fields. Mathematics of Computation, 67(224):1679– 1689, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00969-
7&u=/mcom/1998-67-224/.

Serbin:1997:BRA

[Ser97a] Steven M. Serbin. Book reviews: Afternotes on numerical analysis, by G. W. Stewart. Mathematics of Computation, 66(219):??, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Serra:1997:OQO

[Ser97b] Stefano Serra. Optimal, quasi-optimal and superlinear band-Toeplitz preconditioners for asymptotically illconditioned positive definite Toeplitz systems. Mathematics of Computation, 66(218):651–665, April 1997. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
97-00833-8&u=/mcom/1997-66-218/.

Serra:1999:SPM

[Ser99] Stefano Serra. Superlinear PCG methods for symmetric Toeplitz systems. Mathematics of Computation, 68(226):

793–803, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01045-
                                              5&u=/mcom/1999-68-226/.

Shi:1995:PSL

[SH95] Zhixin Shi and Brian Hassard. Precise solution of Laplace’s equation. Mathematics of Computation, 64(210):515– 536, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scharlau:1998:CIL

[SH98] Rudolf Scharlau and Boris Hemkemeier. Classification of integral lattices with large class number. Mathematics of Computation, 67(222):737– 749, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00938-
                                              7&u=/mcom/1998-67-222/.

Shampine:1996:BRD

[Sha96] Lawrence F. Shampine. Book reviews: Differential equations with Mathematica, by Kevin R. Coombes, Brian R. Hunt, Ronald L. Lipsman, John E. Osborn and Garrett J. Stuck. Mathematics of Computation, 65(215): ??, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00705-
                                              3&u=/mcom/1996-65-215/.

Sheen:1993:SOA

[She93] Dongwoo Sheen. Second-order absorbing boundary conditions for the wave equation in a rectangular domain.

Mathematics of Computation, 61(204): 595–606, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shen:1994:ESG

[She94] Jie Shen. Efficient spectral Galerkin methods for some elliptic problems. In Gautschi [Gau94], pages 373–376. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Shen:1996:EEP

[She96] Jie Shen. On error estimates of the projection methods for the Navier–Stokes equations: Second-order schemes. Mathematics of Computation, 65(215): 1039–1065, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00750-
8&u=/mcom/1996-65-215/.

Shiu:1995:CCF

[Shi95] P. Shiu. Computation of continued fractions without input values. Mathematics of Computation, 64(211):1307– 1317, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shih:1996:BRB

[Shi96] Shagi-Di Shih. Book reviews: The boundary function method for singular perturbation problems, by Adelaida B. Vasil’eva, Valentin F. Butuzov and Leonid V. Kalachev. Mathematics of Computation, 65(215): ??, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00705-
                                              3&u=/mcom/1996-65-215/.

Shoup:1990:NAF

[Sho90] Victor Shoup. New algorithms for finding irreducible polynomials over finite fields. Mathematics of Computation, 54 (189):435–447, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shoup:1992:SPR

[Sho92] Victor Shoup. Searching for primitive roots in finite fields. Mathematics of Computation, 58(197):369– 380, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shoji:1998:ACT

[Sho98] Isao Shoji. Approximation of continuous time stochastic processes by a local linearization method. Mathematics of Computation, 67(221):287– 298, January 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00888-
                                              6&u=/mcom/1998-67-221/.

Shokrollahi:1999:RCN

[Sho99] M. A. Shokrollahi. Relative class number of imaginary Abelian fields of prime conductor below 10000. Mathematics of Computation, 68(228):1717– 1728, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01139-
                                              4&u=/mcom/1999-68-228/.

Shparlinski:1993:BPF

[Shp93] Igor E. Shparlinski. On bivariate polynomial factorization over finite fields. Mathematics of Computation, 60(202): 787–791, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shu:1996:BRM

[Shu96] Chi-Wang Shu. Book reviews: Mathematical and numerical aspects of wave propagation, edited by Gary Cohen. Mathematics of Computation, 65 (215):??, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00705-
3&u=/mcom/1996-65-215/.

Shu:1997:RDT

[Shu97] Chi-Wang Shu. Reviews and descriptions of tables and books: 23. A practical guide to pseudospectral methods, by Bengt Fornberg. Mathematics of Computation, 66(220): ??, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00843-
0&u=/mcom/1997-66-220/.

Shyue:1994:FTB

[Shy94] Keh-Ming Shyue. Front tracking based on high-resolution wave propagation methods. In Gautschi [Gau94], pages 383–387. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Sidi:1995:CAG

[Sid95] Avram Sidi. Convergence analysis for a generalized Richardson extrapolation process with an application to the d^{(1)}- transformation on convergent and divergent logarithmic sequences. Mathematics of Computation, 64(212):1627– 1657, October 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Silverman:1990:DBW

[Sil90] Joseph H. Silverman. The difference between the Weil height and the canonical height on elliptic curves. Mathematics of Computation, 55(192):723– 743, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Silverman:1997:CCH

[Sil97]  Joseph H. Silverman.  Comput-

ing canonical heights with little (or no) factorization. Mathematics of Computation, 66(218):787– 805, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00812-
0&u=/mcom/1997-66-218/.

Silverman:1999:CRP

[Sil99a] Joseph H. Silverman. Computing rational points on rank 1 elliptic curves via L-series and canonical heights. Mathematics of Computation, 68(226):835– 858, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01068-
6&u=/mcom/1999-68-226/.

Silvester:1999:BRI

[Sil99b] David Silvester. Book reviews: Iterative methods for solving linear systems, by Anne Greenbaum. Mathematics of Computation, 68(226): ??, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01060-
1&u=/mcom/1999-68-226/.

Sinisalo:1993:CGC

[Sin93] Matti K. Sinisalo. Checking the Goldbach conjecture up to 4 · 10^{11}. Mathematics of Computation, 61(204):931– 934, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sekigawa:1999:NCS

[SK99] Hiroshi Sekigawa and Kenji Koyama.

Nonexistence conditions of a solution for the congruence x_{1}^{k} + ⋯ + x_{s}^{k} ≡ N( mod p^{n}). Mathematics of Computation, 68(227):1283– 1297, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01067-
4&u=/mcom/1999-68-227/.

Skoruppa:1992:CSM

[Sko92] Nils-Peter Skoruppa. Computations of Siegel modular forms of genus two. Mathematics of Computation, 58(197): 381–398, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sloan:1990:LRP

[SL90] I. H. Sloan and J. N. Lyness. Lattice rules: projection regularity and unique representations. Mathematics of Computation, 54(190):649–660, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smart:1995:STC

[Sma95] N. P. Smart. The solution of triangularly connected decomposable form equations. Mathematics of Computation, 64(210):819–840, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smart:1999:DSS

[Sma99] N. P. Smart. Determining the small solutions to S-unit equations. Mathematics of Computation, 68(228):1687– 1699, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01140-
0&u=/mcom/1999-68-228/.

Smith:1996:MPD

[Smi96] David M. Smith. A multiple-precision division algorithm. Mathematics of Computation, 65(213):157–163, January 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/
journals/mcom/1996-65-213/S0025-
5718-96-00688-6/;   http://www.
ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00688-
6&u=/mcom/1996-65-213/.

Smyth:1991:IOM

[Smy91] C. J. Smyth. Ideal 9th-order multigrades and Letac’s elliptic curve. Mathematics of Computation, 57(196):817– 823, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stynes:1991:ASP

[SO91] Martin Stynes and Eugene O’Riordan.

An analysis of a singularly perturbed two-point boundary value problem using only finite element techniques. Mathematics of Computation, 56(194): 663–675, April 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Solderitsch:1992:QFS

[Sol92] James J. Solderitsch. Quadratic fields with special class groups. Mathematics of Computation, 59(200):633– 638, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Souvignier:1994:IFI

[Sou94] Bernd Souvignier. Irreducible finite integral matrix groups of degree 8 and 10. With microfiche supplement. Mathematics of Computation, 63(207): 335–350, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schwarz:1994:TQN

[SPD94] A. Schwarz, M. Pohst, and F. Diaz y Diaz. A table of quintic number fields. Mathematics of Computation, 63(207): 361–376, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spiess:1990:SII

[Spi90] Jürgen Spieß. Some identities involving harmonic numbers. Mathematics of Computation, 55(192):839– 863, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spira:1994:SZT

[Spi94] Robert Spira. Some zeros of the Titchmarsh counterexample. Mathematics of Computation, 63(208):747– 748, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Srinivasan:1998:CCN

[Sri98] Anitha Srinivasan. Computations of class numbers of real quadratic fields. Mathematics of Computation, 67(223): 1285–1308, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00965-
X&u=/mcom/1998-67-223/.

Saavedra:1991:VFM

[SS91] Patricia Saavedra and L. Ridgway Scott. Variational formulation of a model free-boundary problem. Mathematics of Computation, 57(196):451– 475, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schwab:1994:LBL

[SS94] Christoph Schwab and Manil Suri. Locking and boundary layer effects in the finite element approximation of the Reissner–Mindlin plate model. In Gautschi [Gau94], pages 367–371. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Schwab:1996:VFE

[SS96a] Christoph Schwab and Manil Suri. The p and hp versions of the finite element method for problems with boundary layers. Mathematics of Computation, 65(216):1403– 1429, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00781-
8&u=/mcom/1996-65-216/.

Sen:1996:BRW

[SS96b] Mrinal K. Sen and Paul L. Stoffa.

Book reviews: Wavelets in geophysics, edited by Efi Foufoula-Georgiou and Praveen Kumar. Mathematics of Computation, 65(213): ??, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Sun:1996:UCM

[SS96c] Guang Fu Sun and Martin Stynes. A uniformly convergent method for a singularly perturbed semilinear reaction– diffusion problem with multiple solutions. Mathematics of Computation, 65(215):1085–1109, July 1996. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
96-00753-3&u=/mcom/1996-65-215/.

Shanks:1999:ZAF

[SSW99] Daniel C. Shanks, Patrick J. Sime, and Lawrence C. Washington. Zeros of 2adic L-functions and congruences for class numbers and fundamental units. Mathematics of Computation, 68(227): 1243–1255, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

                                              fn=110&arg1=S0025-5718-99-01046-
                                              7&u=/mcom/1999-68-227/.

Steidl:1991:PAF

[ST91] G. Steidl and M. Tasche. A polynomial approach to fast algorithms for discrete Fourier-cosine and Fourier-sine transforms. Mathematics of Computation, 56(193):281–296, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Strela:1996:WCP

[ST96] V. V. Strela and E. E. Tyrtyshnikov. Which circulant preconditioner is better? Mathematics of Computation, 65(213):137– 150, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00682-
                                              5&u=/mcom/1996-65-213/.

Stenberg:1990:EAS

[Ste90] Rolf Stenberg. Error analysis of some finite element methods for the Stokes problem. Mathematics of Computation, 54(190):495–508, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1992:EAU

[Ste92] G. W. Stewart. Error analysis of QR updating with exponential windowing. Mathematics of Computation, 59(199): 135–140, July 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stetter:1994:IVP

[Ste94a] Hans J. Stetter. Initial value problems for ordinary differential equations: development of ideas, techniques, and implementation. In Gautschi [Gau94], pages 205–224. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Stevenhagen:1994:CNP

[Ste94b] Peter Stevenhagen. Class number parity for the pth cyclotomic field. Mathematics of Computation, 63(208):773– 784, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stenger:1995:CC

[Ste95] Frank Stenger. Collocating convolutions. Mathematics of Computation, 64 (209):211–235, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stenger:1996:BRS

[Ste96a] Frank Stenger. Book reviews: Solving Problems in Scientific Computing Using MAPLE and MATLAB, by Walter Gander and Jírí Hrebícek. Mathematics of Computation, 65(214):880– 882, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Stevenson:1996:FDM

[Ste96b]  Rob Stevenson.  The frequency

decomposition multilevel method: A robust additive hierarchical basis preconditioner. Mathematics of Computation, 65(215):983–997, July 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-96-00740-
5&u=/mcom/1996-65-215/.

Stenger:1997:RDT

[Ste97a] Frank Stenger. Reviews and descriptions of tables and books: 22. Integral equations: Theory and numerical treatment, by Wolfgang Hackbusch. Mathematics of Computation, 66(220): ??, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00843-
0&u=/mcom/1997-66-220/.

Stetter:1997:RDT

[Ste97b] Hans J. Stetter. Reviews and descriptions of tables and books: 19. Lectures on finite precision computations, by Françoise Chaitin-Chatelin and Valérie Frayssé. Mathematics of Computation, 66(220):1751– ??, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00843-
0&u=/mcom/1997-66-220/.

Steiner:1998:CNB

[Ste98a] Ray Steiner. Class number bounds and Catalan’s equation. Mathematics of Computation, 67(223):1317– 1322, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00966-
1&u=/mcom/1998-67-223/.

Stetter:1998:BRC

[Ste98b] Hans J. Stetter. Book reviews: Computational differentiation techniques, applications, and tools, edited by Martin Berg, Christian Bischof, George Corliss, and Andreas Griewank. Mathematics of Computation, 67(224): ??, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00989-
2&u=/mcom/1998-67-224/.

Stenger:1999:BRB

[Ste99a]  Frank Stenger.  Book reviews:

Boundary element method, fundamentals and applications, by Frederico Paris and Jose Canas. Mathematics of Computation, 68(225):??, January 1999. URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-00992-
8&u=/mcom/1999-68-225/.

Stetter:1999:BRN

[Ste99b]  Hans J. Stetter.  Book reviews:

Numerical analysis, an introduction, by Walter Gautschi. Mathematics of Computation, 68(226): 887, April 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01060-
1&u=/mcom/1999-68-226/.

Stewart:1999:BRN

[Ste99c] G. W. Stewart. Book reviews: Numerical linear algebra, by Lloyd N. Trefethen and David Bau, III. Mathematics of Computation, 68(225):??, January 1999. URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-00992-
8&u=/mcom/1999-68-225/.

Strain:1992:FLT

[Str92a] John Strain. A fast Laplace transform based on Laguerre functions. Mathematics of Computation, 58(197):275– 283, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroeker:1992:INU

[Str92b] R. J. Stroeker. Improvement of Nakamula’s upper bound for the absolute discriminant of a sextic number field with two real conjugates. Mathematics of Computation, 59(199):203–211, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Styer:1993:STV

[Sty93] Robert Styer. Small two-variable exponential Diophantine equations. Mathematics of Computation, 60(202):811– 816, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Suli:1992:ACV

[Sül92] Endre Süli. The accuracy of cell vertex finite volume methods on quadrilateral meshes. Mathematics of Computation, 59(200):359–382, October 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Suleiman:1993:SNC

[Sul93] Mohamed Bin Suleiman. Some necessary conditions for convergence of the GBDF methods. Mathematics of Computation, 60(202):635–649, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sun:1995:SOA

[Sun95] Zhi Zhong Sun. A second-order accurate linearized difference scheme for the two-dimensional Cahn–Hilliard equation. Mathematics of Computation, 64 (212):1463–1471, October 1995. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Suri:1990:SCH

[Sur90] Manil Suri. On the stability and convergence of higher-order mixed finite element methods for second-order elliptic problems. Mathematics of Computation, 54(189):1–19, January 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Suri:1997:RCF

[Sur97] Manil Suri. A reduced constraint hp finite element method for shell problems. Mathematics of Computation, 66(217): 15–29, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00806-
5&u=/mcom/1997-66-217/.

Suter:1998:NLP

[Sut98] Ruedi Suter. The number of lattice points in alcoves and the exponents of the finite Weyl groups. Mathematics of Computation, 67(222):751– 758, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00919-
3&u=/mcom/1998-67-222/.

Spigler:1990:NME

[SV90] Renato Spigler and Marco Vianello. A numerical method for evaluating zeros of solutions of second-order linear differential equations. Mathematics of Computation, 55(192):591– 612, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Saranen:1998:FSI

[SV98] J. Saranen and G. Vainikko. Fast solvers of integral and pseudodifferential equations on closed curves. Mathematics of Computation, 67(224):1473– 1491, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00997-
                                              1&u=/mcom/1998-67-224/.

Sloan:1990:CSR

[SW90] Ian H. Sloan and Linda Walsh. A computer search of rank-2 lattice rules for multidimensional quadrature. Mathematics of Computation, 54(189):281– 302, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scheidler:1992:MTN

[SW92] Renate Scheidler and Hugh C. Williams. A method of tabulating the numbertheoretic function g(k). Mathematics of Computation, 59(199):251–257, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Silverman:1993:PAE

[SW93] Robert D. Silverman and Samuel S. Wagstaff, Jr. A practical analysis of the elliptic curve factoring algorithm. Mathematics of Computation, 61(203): 445–462, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schlafly:1994:CCE

[SW94] Aaron Schlafly and Stan Wagon. Carmichael’s conjecture on the Euler function is valid below 10^{10,000,000}. Mathematics of Computation, 63(207):

415–419, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schatz:1995:IMN

[SW95a] A. H. Schatz and L. B. Wahlbin. Interior maximum-norm estimates for finite element methods. part II. Mathematics of Computation, 64(211):907–928, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1995:MSM

[SW95b] Daniel Shanks and Samuel S. Wagstaff, Jr. 48 more solutions of Martin Davis’s quaternary quartic equation. Mathematics of Computation, 64(212):1717– 1731, October 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schatz:1996:SNE

[SW96] Alfred H. Schatz and Jun Ping Wang. Some new error estimates for Ritz–Galerkin methods with minimal regularity assumptions. Mathematics of Computation, 65(213):19– 27, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00649-
7&u=/mcom/1996-65-213/.

Sloan:1997:IRM

[SW97] I. H. Sloan and H. Woźniakowski. An intractability result for multiple integration. Mathematics of Computation, 66(219):1119–1124, July 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
97-00834-X&u=/mcom/1997-66-219/.

Szabo:1998:FEG

[SW98] Sándor Szabó and Coburn Ward. Factoring elementary groups of prime cube order into subsets. Mathematics of Computation, 67(223):1199– 1206, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00929-
                                              6&u=/mcom/1998-67-223/.

Schwab:1999:ETB

[SW99] C. Schwab and W. L. Wendland. On the extraction technique in boundary integral equations. Mathematics of Computation, 68(225):91–122, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01044-
                                              3&u=/mcom/1999-68-225/.

Swann:1993:ULM

[Swa93a] Howard Swann. On the use of Lagrange multipliers in domain decomposition for solving elliptic problems. Mathematics of Computation, 60(201): 49–78, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Swarztrauber:1993:PAC

[Swa93b] Paul N. Swarztrauber. A parallel algorithm for computing the eigenvalues of a symmetric tridiagonal matrix. Mathematics of Computation, 60(202): 651–668, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sauer:1995:MLI

[SX95] Thomas Sauer and Yuan Xu. On multivariate Lagrange interpolation. Mathematics of Computation, 64(211):1147– 1170, July 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scott:1990:FEI

[SZ90] L. Ridgway Scott and Shangyou Zhang.

Finite element interpolation of nonsmooth functions satisfying boundary conditions. Mathematics of Computation, 54(190):483–493, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scott:1992:HDN

[SZ92] L. Ridgway Scott and Shangyou Zhang.

Higher-dimensional nonnested multigrid methods. Mathematics of Computation, 58(198):457–466, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tadmor:1993:TVE

[Tad93] Eitan Tadmor. Total variation and error estimates for spectral viscosity approximations. Mathematics of Computation, 60(201):245–256, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Takeuchi:1994:EZ

[Tak94] Kisao Takeuchi. Evaluation of ζ_{K}(2)

for some totally real algebraic number fields K of degree 9. In Gautschi [Gau94], pages 589–593. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Taya:1996:CIR

[Tay96] Hisao Taya. Computation of ℤ_{3}- invariants of real quadratic fields. Mathematics of Computation, 65(214):779– 784, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00721-
1&u=/mcom/1996-65-214/.

Tourigny:1997:AAG

[TB97] Y. Tourigny and M. J. Baines. Analysis of an algorithm for generating locally optimal meshes for L_{2} approximation by discontinuous piecewise polynomials. Mathematics of Computation, 66(218): 623–650, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00823-
5&u=/mcom/1997-66-218/.

Tzanakis:1991:SST

[Td91] N. Tzanakis and B. M. M. de Weger. Solving a specific Thue–Mahler equation. Mathematics of Computation, 57 (196):799–815, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Teitelbaum:1990:CCR

[Tei90] Jeremy Teitelbaum. The computational complexity of the resolution of plane curve singularities. Mathematics of Computation, 54(190):797–837, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Teitelbaum:1998:EAL

[Tei98]  Jeremy Teitelbaum.  Euclid’s al-

gorithm and the Lanczos method over finite fields. Mathematics of Computation, 67(224):1665–1678, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-98-00973-
9&u=/mcom/1998-67-224/.

Temam:1991:SAN

[Tem91] R. Temam. Stability analysis of the nonlinear Galerkin method. Mathematics of Computation, 57(196):477– 505, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Temme:1992:AII

[Tem92] N. M. Temme. Asymptotic inversion of incomplete gamma functions. Mathematics of Computation, 58(198): 755–764, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Temam:1994:MMP

[Tem94a] Roger Temam. Multiresolution methods for partial differential equations. In Gautschi [Gau94], pages 225–240. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Temme:1994:NAU

[Tem94b] N. M. Temme. Numerical aspects of uniform Airy-type asymptotic expansions. In Gautschi [Gau94], pages 395– 398. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Temme:1997:BRI

[Tem97]  Nico M. Temme.  Book reviews:

Integral transforms and their applications, by L. Debnath. Mathematics of Computation, 66(217):

??, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00793-
X&u=/mcom/1997-66-217/.

Temme:1998:BRT

[Tem98]  Nico M. Temme.  Book reviews:

Tables addenda for Laplace transforms, by R. G. Buschman. Mathematics of Computation, 67(222): 875, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00928-
4&u=/mcom/1998-67-222/.

Teske:1998:SEA

[Tes98] Edlyn Teske. A space efficient algorithm for group structure computation. Mathematics of Computation, 67(224):1637– 1663, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00968-
5&u=/mcom/1998-67-224/.

Tezuka:1994:DDC

[Tez94] Shu Tezuka. The k-dimensional distribution of combined GFSR sequences. Mathematics of Computation, 62(206): 809–817, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tezuka:1993:CFP

[TF93] Shu Tezuka and Masanori Fushimi. Calculation of Fibonacci polynomials for GFSR sequences with low discrepancies. Mathematics of Computation, 60(202):763–770, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.jstor.org/stable/
pdfplus/2153114.pdf.

Tourigny:1994:DSD

[TG94] Yves Tourigny and Michael Grinfeld.

Deciphering singularities by discrete methods. Mathematics of Computation, 62(205):155–169, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thomason:1994:EAF

[Tho94] George Thomason. Erratum: “Zeros of approximate functional approximations” [Math. Comp. 21 (1967), 41–48, MR 38 #2945] by R. Spira. Mathematics of Computation, 63(208): 829, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thomee:1997:BRA

[Tho97] Vidar Thomée. Book reviews: Acta numerica 1995, edited by A. Iserles. Mathematics of Computation, 66 (218):??, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00856-
9&u=/mcom/1997-66-218/.

Tang:1997:VMP

[ThT97] Tao Tang and Zhen huan Teng. Viscosity methods for piecewise smooth solutions to scalar conservation laws. Mathematics of Computation, 66(218): 495–526, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00822-
3&u=/mcom/1997-66-218/.

Tilli:1997:ASH

[Til97] Paolo Tilli. On the asymptotic spectrum of Hermitian block Toeplitz matrices with Toeplitz blocks. Mathematics of Computation, 66(219):1147– 1159, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00840-
5&u=/mcom/1997-66-219/.

Anonymous:1996:BRI

[Tod96] M. J. Todd. Book reviews: Interiorpoint polynomial algorithms in convex programming, by Yurii Nesterov and Arkadii Nemirovskii. Mathematics of Computation, 65(214):884– 886, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Todd:1998:BRP

[Tod98]  Michael J. Todd.  Book reviews:

Primal-dual interior-point methods, by Stephen J. Wright. Mathematics of Computation, 67(222):867– 870, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00928-
4&u=/mcom/1998-67-222/.

Tollis:1997:ZDZ

[Tol97] Emmanuel Tollis. Zeros of Dedekind zeta functions in the critical strip. Mathematics of Computation, 66(219): 1295–1321, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?

fn=110&arg1=S0025-5718-97-00871-
5&u=/mcom/1997-66-219/.

teRiele:1994:NMF

[tR94] H. J. J. te Riele. A new method for finding amicable pairs. In Gautschi [Gau94], pages 577–581. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Tran:1995:OGM

[Tra95] Thanh Tran. The K-operator and the Galerkin method for strongly elliptic equations on smooth curves: Local estimates. Mathematics of Computation, 64(210):501–513, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tsuchiya:1990:NDS

[Tsu90] Takuya Tsuchiya. A note on discrete solutions of the Plateau problem. Mathematics of Computation, 54(189):131– 138, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tang:1994:TSM

[TT94] Tao Tang and Zhen Huan Teng. Timesplitting methods for nonhomogeneous conservation laws. In Gautschi [Gau94], pages 389–393. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Tang:1995:SKE

[TT95] Tao Tang and Zhen Huan Teng. The sharpness of Kuznetsov’s O(\sqrt{Δx})L^{1}- error estimate for monotone difference schemes. Mathematics of Computation, 64(210):581–589, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Trompert:1993:RKM

[TV93] R. A. Trompert and J. G. Verwer. Runge–Kutta methods and local uniform grid refinement. Mathematics of Computation, 60(202):591–616, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Traub:1992:MCA

[TW92] J. F. Traub and H. Woźniakowski. The Monte Carlo algorithm with a pseudorandom generator. Mathematics of Computation, 58(197):323–339, January 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thomee:1994:LTN

[TW94] Vidar Thomée and Lars B. Wahlbin. Long-time numerical solution of a parabolic equation with memory. Mathematics of Computation, 62(206):477– 496, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Urabe:1996:CMI

[Ura96a]  Tohsuke Urabe.  Calculation of

Manin’s invariant for Del Pezzo surfaces. Mathematics of Computation, 65(213):247–258, S15–S23, January 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00681-
                                              3&u=/mcom/1996-65-213/.

Urabe:1996:SCM

[Ura96b] Tohsuke Urabe. Supplement to “Calculation of Manin’s invariant for Del Pezzo surfaces”. Mathematics of Computation, 65(213):S15–S23, January 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vallee:1991:GES

[Val91] Brigitte Vallée. Generation of elements with small modular squares and provably fast integer factoring algorithms. Mathematics of Computation, 56(194): 823–849, April 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

VanTuyl:1994:ACF

[Van94] Andrew H. Van Tuyl. Acceleration of convergence of a family of logarithmically convergent sequences. Mathematics of Computation, 63(207):229– 246, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

VanLoan:1997:RDT

[Van97] Charles Van Loan. Reviews and descriptions of tables and books: 20. Accuracy and stability of numerical algorithms, by Nicholas J. Higham. Mathematics of Computation, 66(220): ??, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00843-
0&u=/mcom/1997-66-220/.

vanWamelen:1999:EGT

[van99]  Paul van Wamelen.  Examples

of genus two CM curves defined over the rationals. Mathematics of Computation, 68(225):307–320, January 1999. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01020-
                                              0&u=/mcom/1999-68-225/.

Vaserstein:1991:SCP

[Vas91] L. N. Vaserstein. Sums of cubes in polynomial rings. Mathematics of Computation, 56(193):349–357, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vassilevski:1992:HCA

[Vas92] P. S. Vassilevski. Hybrid V -cycle algebraic multilevel preconditioners. Mathematics of Computation, 58(198):489– 512, April 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

VanDaalen:1992:VPC

[vBv92] Edwin F. G. van Daalen, Jan Broeze, and Embrecht van Groesen. Variational principles and conservation laws in the derivation of radiation boundary conditions for wave equations. Mathematics of Computation, 58(197):55– 71, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Verlinden:1994:PCA

[VC94] P. Verlinden and R. Cools. Proof of a conjectured asymptotic expansion for the approximation of surface integrals. Mathematics of Computation, 63(208): 717–725, October 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Verfurth:1994:PEE

[Ver94] R. Verfürth. A posteriori error estimates for nonlinear problems. Finite element discretizations of elliptic equations. Mathematics of Computation, 62(206):445–475, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Verfurth:1998:PEE

[Ver98]  R. Verfürth.  A posteriori er-

ror estimates for nonlinear problems. L^{r}(0, T; L^{ρ}(Ω))-error estimates for finite element discretizations of parabolic equations. Mathematics of Computation, 67(224):1335–1360, October 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-01011-
4&u=/mcom/1998-67-224/.

Viallon:1991:CTP

[Via91] Marie-Claude Viallon. Convergence of the two-point upstream weighting scheme. Mathematics of Computation, 57(196):569–584, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vandenHijligenberg:1995:DVF

[vKP95] Nico van den Hijligenberg, Youri Kotchetkov, and Gerhard Post. Deformations of vector fields and Hamiltonian vector fields on the plane. Mathematics of Computation, 64(211):1215–1226, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Voutier:1995:PDL

[Vou95] Paul M. Voutier. Primitive divisors of Lucas and Lehmer sequences. Mathematics of Computation, 64(210):869– 888, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vonMatt:1996:RES

[vS96] Urs von Matt and G. W. Stewart. Rounding errors in solving block Hessenberg systems. Mathematics of Computation, 65(213):115– 135, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00667-
                                              9&u=/mcom/1996-65-213/.

VanDerHouwen:1992:EDI

[vSC92] P. J. van der Houwen, B. P. Sommeijer, and W. Couzy. Embedded diagonally implicit Runge–Kutta algorithms on parallel computers. Mathematics of Computation, 58(197):135– 159, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanderGeer:1992:WFT

[vSv92] Gerard van der Geer, René Schoof, and Marcel van der Vlugt. Weight formulas for ternary Melas codes. Mathematics of Computation, 58(198):781–792, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

VanAssche:1993:QFB

[VV93a] Walter Van Assche and Ingrid Vanherwegen. Quadrature formulas based on rational interpolation. Mathematics of Computation, 61(204):765– 783, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

VanGeemen:1993:CFW

[vv93b] Bert van Geemen and Duco van Straten. The cusp forms of weight 3 on Γ_{2}(2, 4, 8). Mathematics of Computation, 61(204):849–872, October 1993.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vanmaele:1996:VAM

[Vv96] M. Vanmaele and R. van Keer. On a variational approximation method for a class of elliptic eigenvalue problems in composite structures. Mathematics of Computation, 65(215):999– 1017, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00741-
7&u=/mcom/1996-65-215/.

vanWamelen:1999:PGC

[vW99] Paul van Wamelen. Proving that a genus 2 curve has complex multiplication. Mathematics of Computation, 68(228):1663–1677, October 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01101-
1&u=/mcom/1999-68-228/.

Verner:1995:CER

[VZ95] J. H. Verner and M. Zennaro. Continuous explicit Runge–Kutta methods of order 5. Mathematics of Computation, 64(211):1123–1146, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wade:1990:SFD

[Wad90] Bruce A. Wade. Symmetrizable finite difference operators. Mathematics of Computation, 54(190):525–543, April 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wagstaff:1994:PBE

[Wag94] Samuel S. Wagstaff, Jr. The period of the Bell exponential integers modulo a prime. In Gautschi [Gau94], pages 595– 598. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Wagon:1995:QST

[Wag95] Stan Wagon. Quintuples with square triplets. Mathematics of Computation, 64(212):1755–1756, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wagner:1996:CN

[Wag96a] Christian Wagner. Class number 5, 6 and 7. Mathematics of Computation, 65(214):785– 800, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00722-
                                              3&u=/mcom/1996-65-214/.

Wagstaff:1996:BRP

[Wag96b] S. S. Wagstaff, Jr. Book reviews: Prime numbers and computer methods for factorization, by Hans Riesel. Mathematics of Computation, 65(215): ??, July 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-96-00705-
                                              3&u=/mcom/1996-65-215/.

Wagstaff:1996:AFP

[Wag96c] Samuel S. Wagstaff, Jr. Aurifeuillian factorizations and the period of the Bell numbers modulo a prime. Mathematics of Computation, 65(213):383– 391, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00683-
7&u=/mcom/1996-65-213/.

Wagstaff:1998:BRA

[Wag98]  S. S. Wagstaff, Jr.  Book re-

views: Algorithmic number theory, Volume I: Efficient algorithms, by Eric Bach and Jeffrey Shallit. Mathematics of Computation, 67(223): ??, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00982-
X&u=/mcom/1998-67-223/.

Wahlbin:1997:BRN

[Wah97a] Lars B. Wahlbin. Book reviews: Numerical solution of partial differential equations, by K. W. Morton and D. F. Mayers. Mathematics of Computation, 66(218):??, April 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00856-
9&u=/mcom/1997-66-218/.

Wahlbin:1997:RDT

[Wah97b] Lars B. Wahlbin. Reviews and descriptions of tables and books:

27. Number-theoretic function products, by R. G. Buschman. Mathematics of Computation, 66(220): ??, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00843-
0&u=/mcom/1997-66-220/.

Wahlbin:1998:BRP

[Wah98a] Lars B. Wahlbin. Book reviews: Proceedings of the Third International Conference on Spectral and High Order Methods, edited by Andrew V. Ilin and L. Ridgeway Scott. Mathematics of Computation, 67(223): ??, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-98-00982-
                                              X&u=/mcom/1998-67-223/.

Wahlbin:1998:BRM

[Wah98b] Lars B. Wahlbin. Book reviews: The mathematics of surfaces, edited by Glen Mullineux. Mathematics of Computation, 67(223):??, July 1998. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                              http://www.ams.org/jourcgi/jour-
                                              pbprocess?fn=110&arg1=S0025-5718-
                                              98-00982-X&u=/mcom/1998-67-223/.

Walker:1991:IDG

[Wal91] Peter Walker. Infinitely differentiable generalized logarithmic and exponential functions. Mathematics of Computation, 57(196):723–733, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Walkington:1992:CNF

[Wal92] Noel J. Walkington. Convergence of nonconforming finite element approximations to first-order linear hyperbolic equations. Mathematics of Computation, 58(198):671–691, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Walsh:1999:TCS

[Wal99] P. G. Walsh. On two classes of simultaneous Pell equations with no solutions.

Mathematics of Computation, 68(225): 385–388, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01048-
0&u=/mcom/1999-68-225/. See corri-

gendum [Wal21].

Walsh:2021:CTC

[Wal21] P. G. Walsh. Corrigendum to “On two classes of simultaneous Pell equations with no solutions”. Mathematics of Computation, 90(331):2503– 2505, July 2021. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.

ams.org/AMSMathViewer;   https:/
/www.ams.org/journals/mcom/2021-
90-331/S0025-5718-2021-03677-7;
https://www.ams.org/journals/mcom/
2021-90-331/S0025-5718-2021-03677-
7/S0025-5718-2021-03677-7.pdf;
https://www.ams.org/mathscinet/
search/authors.html?mrauthid=230332.

Wan:1990:FMP

[Wan90] Daqing Wan. Factoring multivariate polynomials over large finite fields. Mathematics of Computation, 54(190): 755–770, April 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wang:1991:SEM

[Wan91] Jun Ping Wang. Superconvergence and extrapolation for mixed finite element methods on rectangular domains. Mathematics of Computation, 56(194): 477–503, April 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wan:1997:GIP

[Wan97] Daqing Wan. Generators and irreducible polynomials over finite fields. Mathematics of Computation, 66(219): 1195–1212, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00835-
1&u=/mcom/1997-66-219/.

Wang:1999:CNM

[Wan99] Xinghua Wang. Convergence of Newton’s method and inverse function theorem in Banach space. Mathematics of Computation, 68(225):169– 186, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-00999-
0&u=/mcom/1999-68-225/.

Ward:1996:BRA

[War96] J. D. Ward. Book reviews: Advances in computational mathematics: New Delhi, India, edited by H. P. Dikshit and C. A. Micchelli. Mathematics of Computation, 65(213): 432, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00666-
7&u=/mcom/1996-65-213/.

Ward:1998:BRI

[War98] Joseph D. Ward. Book reviews: An introduction to computational geometry for curves and surfaces, by Alan Davies and Philip Samuels. Mathematics of Computation, 67(223): ??, July 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00982-
X&u=/mcom/1998-67-223/.

Washington:1991:FCQ

[Was91] Lawrence C. Washington. A family of cyclic quartic fields arising from modular curves. Mathematics of Computation, 57(196):763–775, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Watkins:1998:BRA

[Wat98] David S. Watkins. Book reviews:

Applied numerical linear algebra, by James W. Demmel. Mathematics of Computation, 67(224):??, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00989-
2&u=/mcom/1998-67-224/.

Wang:1994:LDB

[WBL94] Xin Wang, Edward K. Blum, and Qingnan Li. Local dynamics and bifurcation consistencies of continuous-time dynamical systems and their numerical discretizations. In Gautschi [Gau94], pages 399–402. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Wang:1993:URP

[WC93] Da Kai Wang and Aaldert Compagner. On the use of reducible polynomials as random number generators. Mathematics of Computation, 60(201):363– 374, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weideman:1994:CIC

[Wei94] J. A. C. Weideman. Computing integrals of the complex error function. In Gautschi [Gau94], pages 403–407. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Weideman:1995:CHT

[Wei95] J. A. C. Weideman. Computing the Hilbert transform on the real line. Mathematics of Computation, 64(210): 745–762, April 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wei:1997:EFS

[Wei97]  Musheng Wei.  Equivalent formu-

lae for the supremum and stability of weighted pseudoinverses. Mathematics of Computation, 66(220):1487– 1508, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00899-
                                              5&u=/mcom/1997-66-220/.

Wendland:1999:MGM

[Wen99] Holger Wendland. Meshless Galerkin methods using radial basis functions. Mathematics of Computation, 68(228): 1521–1531, October 1999. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                              http://www.ams.org/jourcgi/jour-
                                              pbprocess?fn=110&arg1=S0025-5718-
                                              99-01102-3&u=/mcom/1999-68-228/.

Wasilkowski:1992:PAI

[WG92] G. W. Wasilkowski and F. Gao. On the power of adaptive information for functions with singularities. Mathematics of Computation, 58(197):285– 304, January 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1993:RHC

[WH93] Kenneth S. Williams and Kenneth Hardy. A refinement of H. C. Williams’ qth root algorithm. Mathematics of Computation, 61(203):475–483, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wohlmuth:1999:CPE

[WH99] Barbara I. Wohlmuth and Ronald H. W. Hoppe. A comparison of a posteriori error estimators for mixed finite element discretizations by Raviart–Thomas elements. Mathematics of Computation, 68(228):1347– 1378, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01125-
4&u=/mcom/1999-68-228/.

Williams:1993:HWF

[Wil93] H. C. Williams. How was F_{6} factored.

Mathematics of Computation, 61(203): 463–474, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1997:DS

[Wil97] H. C. Williams. Daniel Shanks (1917– 1996). Mathematics of Computation, 66(219):929–934, July 1997. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-
97-00889-2&u=/mcom/1997-66-219/.

Wimp:1998:BRS

[Wim98]  Jet Wimp.  Book reviews: Spe-

cial functions: An introduction to the classical functions of mathematical physics, by Nico M. Temme. Mathematics of Computation, 67(222):872– 874, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00928-
4&u=/mcom/1998-67-222/.

Wei:1999:PEE

[WL99] Dongming Wei and Lew Lefton. A priori L^{ρ} error estimates for Galerkin approximations to porous medium and fast diffusion equations. Mathematics of Computation, 68(227):971– 989, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01021-
2&u=/mcom/1999-68-227/.

Woznicki:1994:SSA

[Woź94] Zbigniew I. Woźnicki. The Sigma-SOR algorithm and the optimal strategy for the utilization of the SOR iterative method. Mathematics of Computation, 62(206):619–644, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Woznicki:1997:C

[Woz97a] Zbigniew Woznicki. Corrigendum. Mathematics of Computation, 66(220): 1769, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00908-
3&u=/mcom/1997-66-220/.

Woznicki:1997:ESS

[Woz97b] Zbigniew Woznicki. Erratum: “The sigma-SOR algorithm and the optimal strategy for the utilization of the SOR iterative method” [math. comp. 62 (1994), no. 206, 619–644, MR 94g:65040]. Mathematics of Computation, 66(220):1769, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wright:1996:BRM

[Wri96] Francis J. Wright. Book reviews:

Maple V by example, by Martha L. Abell and James P. Braselton. Mathematics of Computation, 65(214):881– 884, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00700-
4&u=/mcom/1996-65-214/.

Wright:1997:RDT

[Wri97] Stephen J. Wright. Reviews and descriptions of tables and books:

25. Linear and nonlinear conjugate gradient-related methods, edited by Loyce Adams and J. L. Nazareth. Mathematics of Computation, 66(220): ??, October 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00843-
0&u=/mcom/1997-66-220/.

Williams:1994:FIB

[WS94] H. C. (Hugh Cowie) Williams and J. O.

Shallit. Factoring integers before computers. In Gautschi [Gau94], pages 481– 531. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Wei:1990:NAI

[WT90] Yuan Wei and Tang Tao. The numerical analysis of implicit Runge– Kutta methods for a certain nonlinear integro-differential equation. Mathematics of Computation, 54(189):155– 168, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wu:1995:SSD

[Wu95] Lixin Wu. The semigroup stability of the difference approximations for initial-boundary value problems. Mathematics of Computation, 64(209):71– 88, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wasilkowski:1997:EDM

[WW97] Grzegorz W. Wasilkowski and Henryk Woźniakowski. The exponent of discrepancy is at most 1.4778 · · ·. Mathematics of Computation, 66(219):1125– 1132, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-97-00824-
                                              7&u=/mcom/1997-66-219/.

Xu:1992:PGM

[XC92] Jinchao Xu and Xiao-Chuan Cai. A preconditioned GMRES method for nonsymmetric or indefinite problems. Mathematics of Computation, 59(200): 311–319, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Xu:1991:CCW

[Xu91] Jinchao Xu. Counterexamples concerning a weighted L^{2} projection. Mathematics of Computation, 57(196):563– 568, October 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Xu:1992:GCB

[Xu92] Yuan Xu. Gaussian cubature and bivariate polynomial interpolation. Mathematics of Computation, 59(200):547– 555, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Xu:1994:CIP

[Xu94a] Changsheng Xu. Computing invariant polynomials of p-adic reflection groups. In Gautschi [Gau94], pages 599–602. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Xu:1994:CPQ

[Xu94b] Yuan Xu. A characterization of positive quadrature formulae. Mathematics of Computation, 62(206):703–718, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Xu:1994:RFM

[Xu94c] Yuan Xu. Recurrence formulas for multivariate orthogonal polynomials. Mathematics of Computation, 62(206):687– 702, April 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Xu:1994:QII

[XZ94] Yuesheng Xu and Yunhe Zhao. Quadratures for improper integrals and their applications in integral equations. In Gautschi [Gau94], pages 409–413. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Xu:1996:EMC

[XZ96] Yuesheng Xu and Yunhe Zhao. An extrapolation method for a class of boundary integral equations. Mathematics of Computation, 65(214):587– 610, April 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00723-
5&u=/mcom/1996-65-214/.

Xu:1997:PPS

[XZ97] Jinchao Xu and Sheng Zhang. Preconditioning the Poincaré–Steklov operator by using Green’s function. Mathematics of Computation, 66(217):125– 138, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00799-
0&u=/mcom/1997-66-217/.

Xu:1999:MFE

[XZ99] Jinchao Xu and Ludmil Zikatanov. A monotone finite element scheme for convection-diffusion equations. Mathematics of Computation, 68(228):1429– 1446, October 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-99-01148-
5&u=/mcom/1999-68-228/.

Yamamura:1994:DIA

[Yam94] Ken Yamamura. The determination of the imaginary Abelian number fields with class number one. Mathematics of Computation, 62(206):899–921, April 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yan:1990:CMF

[Yan90] Yi Yan. The collocation method for first-kind boundary integral equations on polygonal regions. Mathematics of Computation, 54(189):139– 154, January 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yang:1995:GMS

[Yan95] Dao Qi Yang. Grid modification for second-order hyperbolic problems. Mathematics of Computation, 64(212): 1495–1509, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yang:1996:WEI

[Yan96a] Huanan Yang. On wavewise entropy inequalities for high-resolution schemes. I: The semidiscrete case. Mathematics of Computation, 65(213):45–67, S1– S13, January 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00668-
0&u=/mcom/1996-65-213/.

Yang:1996:SWE

[Yan96b] Huanan Yang. Supplement to “On wavewise entropy inequalities for highresolution schemes. I: The semidiscrete case”. Mathematics of Computation, 65 (213):S1–S13, January 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yeung:1993:CPT

[YC93] Man-Chung Yeung and Raymond H. Chan. Circulant preconditioners for Toeplitz matrices with piecewise continuous generating functions. Mathematics of Computation, 61(204):701– 718, October 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ye:1991:CAN

[Ye91] Qiang Ye. A convergence analysis for nonsymmetric Lánczos algorithms. Mathematics of Computation, 56(194): 677–691, April 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ye:1994:BFV

[Ye94] Qiang Ye. A breakdown-free variation of the nonsymmetric Lánczos algorithms. Mathematics of Computation, 62(205):179–207, January 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ying:1990:VSU

[Yin90] Lung-An Ying. Viscous splitting for the unbounded problem of the Navier– Stokes equations. Mathematics of Computation, 55(191):89–113, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yin:1994:SCM

[Yin94]  Guang Yan Yin.  Sinc-collocation

method with orthogonalization for singular Poisson-like problems. Mathematics of Computation, 62(205):21– 40, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ying:1998:VMT

[Yin98]  Lung-An Ying.  Vortex method

for two-dimensional Euler equations in bounded domains with boundary correction. Mathematics of Computation, 67(224):1383–1400, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00970-
3&u=/mcom/1998-67-224/.

Young:1998:LPF

[You98] Jeff Young. Large primes and Fermat factors. Mathematics of Computation, 67(224):1735–1738, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00989-
2&u=/mcom/1998-67-224/.

Yeow:1992:ESN

[YR92] Meng Chee Yeow and Gordon F.

Royle. Enumeration of small nonisomorphic 1-rotational twofold triple systems. with microfiche supplement. Mathematics of Computation, 59(200): 609–612, October 1992. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yui:1997:SVW

[YZ97] Noriko Yui and Don Zagier. On the singular values of Weber modular functions. Mathematics of Computation, 66(220):1645–1662, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00854-
5&u=/mcom/1997-66-220/.

Zagier:1993:ANC

[Zag93] D. Zagier. Algebraic numbers close to both 0 and 1. Mathematics of Computation, 61(203):485–491, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zenisek:1995:MAC

[Žen95] Alexander Ženíšek. Maximum-angle condition and triangular finite elements of Hermite type. Mathematics of Computation, 64(211):929–941, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zhou:1993:LSP

[ZF93] Tian Xiao Zhou and Min Fu Feng. A least squares Petrov–Galerkin finite element method for the stationary Navier– Stokes equations. Mathematics of Computation, 60(202):531–543, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zha:1990:RGG

[ZH90] Hongyuan Zha and Per Christian Hansen. Regularization and the general Gauss–Markov linear model. Mathematics of Computation, 55(192):613– 624, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zhang:1990:OONa

[Zha90a]  Shangyou Zhang.  Optimal-order

nonnested multigrid methods for solving finite element equations. I. on quasiuniform meshes. Mathematics of Computation, 55(191):23–36, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zhang:1990:OONb

[Zha90b]  Shangyou Zhang.  Optimal-order

nonnested multigrid methods for solving finite element equations II: On non-quasi-uniform meshes. Mathematics of Computation, 55(192):439– 450, October 1990. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zhang:1993:FDG

[Zha93a]  Nai Ying Zhang.  On fully dis-

crete Galerkin approximations for partial integro-differential equations of parabolic type. Mathematics of Computation, 60(201):133–166, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zhang:1993:SLR

[Zha93b] Zhen Xiang Zhang. A B_{2}-sequence with larger reciprocal sum. Mathematics of Computation, 60(202):835–839, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zhang:1993:PEC

[Zha93c] Zhen Xiang Zhang. On a problem of Erdős concerning primitive sequences. Mathematics of Computation, 60(202): 827–834, April 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zhang:1994:FFS

[Zha94] Zhen Xiang Zhang. Finding finite B_{2}-sequences with larger m − a_{m}^{1/2}. Mathematics of Computation, 63(207): 403–414, July 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zhang:1995:YC

[Zha95a] Ming Yao Zhang. On Yokoi’s conjecture. Mathematics of Computation, 64(212):1675–1685, October 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zhang:1995:OON

[Zha95b]  Shangyou Zhang.  Optimal-order

nonnested multigrid methods for solving finite element equations. III. On degenerate meshes. Mathematics of Computation, 64(209):23–49, January 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zhang:1996:UPR

[Zha96] Zhimin Zhang. Ultraconvergence of the patch recovery technique. Mathematics of Computation, 65(216):1431– 1437, October 1996. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-96-00782-
X&u=/mcom/1996-65-216/.

Zhang:1998:BRC

[Zha98a] Zhimin Zhang. Book reviews: Computational mathematics in engineering and applied science: ODEs, DAEs, and PDEs, by William E. Schiesser. Mathematics of Computation, 67(222): 870–871, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00928-
4&u=/mcom/1998-67-222/.

Zhang:1998:DSP

[Zha98b] Zhimin Zhang. Derivative superconvergent points in finite element solutions of Poisson’s equation for the serendipity and intermediate families — a theoretical justification. Mathematics of Computation, 67(222):541– 552, April 1998. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-98-00942-
9&u=/mcom/1998-67-222/.

Zheludev:1994:SHA

[Zhe94] Valery A. Zheludev. Spline harmonic analysis and wavelet bases. In Gautschi [Gau94], pages 415–419. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Zhou:1997:HAS

[Zho97]  Guohui Zhou.  How accurate is

the streamline diffusion finite element method? Mathematics of Computation, 66(217):31–44, January 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00788-
6&u=/mcom/1997-66-217/.

Zhu:1992:FEA

[Zhu92] Jiang Zhu. Finite element approximation to initial-boundary value problems of the semiconductor device equations with magnetic influence. Mathematics of Computation, 59(199):39–62, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zivkovic:1994:TPBa

[Živ94a] Miodrag Živković. A table of primitive binary polynomials. Mathematics of Computation, 62(205):385– 386, January 1994. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zivkovic:1994:TPBb

[Živ94b] Miodrag Živković. Table of primitive binary polynomials. II. Mathematics of Computation, 63(207):301–306, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ziv:1995:CAE

[Ziv95] Abraham Ziv. Converting approximate error bounds into exact ones. Mathematics of Computation, 64(209):265– 277, January 1995. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ziv:1999:SUR

[Ziv99a] Abraham Ziv. Sharp ULP rounding error bound for the hypotenuse function. Mathematics of Computation, 68(227): 1143–1148, July 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-01103-
                                              5&u=/mcom/1999-68-227/.

Zivkovic:1999:NPF

[Živ99b] Miodrag Živković. The number of primes ∑_{i=1}^{n}(−1)^{n−i}i! is finite. Mathematics of Computation, 68(225):403– 409, January 1999. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                              ams.org/jourcgi/jour-pbprocess?
                                              fn=110&arg1=S0025-5718-99-00990-
                                              4&u=/mcom/1999-68-225/.

Zhang:1997:TDB

[ZZ97] Shangyou Zhang and Zhimin Zhang. Treatments of discontinuity and bubble functions in the multigrid method.

Mathematics of Computation, 66(219): 1055–1072, July 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-97-00853-
3&u=/mcom/1997-66-219/.